### v1

Abstract
In stars that lie on the main sequence in the Hertzsprung Russel diagram, like our sun, hydrogen is fused to helium in a number of nuclear reaction chains and series, such as the proton-proton chain and the carbon-nitrogen-oxygen cycles. Precisely determined thermonuclear rates of these reactions lie at the foundation of the standard solar model.
This review, the third decadal evaluation of the nuclear physics of hydrogen-burning stars, is motivated by the great advances made in recent years by solar neutrino observatories, putting experimental knowledge of the proton-proton chain neutrino fluxes in the few-percent precision range. The basis of the review is a one-week community meeting held in July 2022 in Berkeley, California, and many subsequent digital meetings and exchanges.
Each of the relevant reactions of solar and quiescent stellar hydrogen burning is reviewed here, from both theoretical and experimental perspectives. Recommendations for the state of the art of the astrophysical S-factor and its uncertainty are formulated for each of them.
Several other topics of paramount importance for the solar model are reviewed, as well: recent and future neutrino experiments, electron screening, radiative opacities, and current and upcoming experimental facilities. In addition to reaction-specific recommendations, also general recommendations are formed.
Contents
I Introduction
The present review summarizes the state of our understanding, in the third decade of the 21st century, of the nuclear reactions and decays taking place inside the Sun and other hydrogen-burning stars. It also addresses related issues, including solar neutrino detection, radiative opacities, electron screening of nuclear reactions, and the status of the current facilities for measuring cross sections and opacities.
As was the case for two previous decadal reviews [1998RvMP...70.1265A, 2011RvMP...83..195A], the present review summarizes the progress made over the past decade in advancing the nuclear physics of main-sequence stars and makes recommendations for future work.
I.1 Purpose
Ray Davis’s measurements of the flux of solar neutrinos [1968PhRvL..20.1205D, 2003RvMP...75..985D] showed a significant deficit with respect to the predictions of the standard solar model (SSM) [1966PhRvL..17..398B]. This discrepancy became known as the Solar Neutrino Problem. An intense debate ensued about possible explanations for this discrepancy, summarized in the entertaining review by 1988Natur.334..487B. Suggestions included plausible flaws in the SSM that might produce a cooler solar core, new particle physics such as neutrino oscillations, and changes in the nuclear physics governing He synthesis, such as an unidentified low-energy resonance in the 3He+3He reaction affecting extrapolations of laboratory cross sections to solar energies. By the mid-1990s, with new results from Kamioka II/III and the SAGE/GALLEX experiments indicating a pattern of neutrino fluxes inconsistent with the expected scaling of those fluxes with the solar core temperature, attention increasingly turned to neutrino oscillations and other potential particle physics solutions. Concerns about the nuclear physics of the SSM also evolved, focusing more on the uncertainties that might inhibit extraction of any such “new physics.”
At that time, during a meeting on the Solar Neutrino Problem hosted by the Institute for Nuclear Theory (INT), University of Washington, Seattle, a suggestion was made to convene the community working on solar nuclear reactions, in order to reach consensus on the best current values of cross sections and their uncertainties. The INT hosted the proposed workshop in February, 1997, drawing representatives from almost every experimental group active in this area, as well as many of the theorists who were engaged in solar neutrino physics. The working groups that formed during this meeting worked over the next year to evaluate past work, determining the needed cross sections and their uncertainties. The results of this evaluation, which came to be known as Solar Fusion I or SF-I [1998RvMP...70.1265A], became the standard for use in solar modeling over the next decade.
An update of SF I was launched with the January, 2009, workshop “Solar Fusion Cross Sections II”, hosted again by the INT in Seattle. Initial results from SuperKamiokande, the Sudbury Neutrino Observatory, and Borexino were then in hand, and the conversion of approximately two thirds of solar electron neutrinos into other flavors had been firmly established [2002PhRvL..89a1301A]. Consequently, the motivation for the study was to ensure that solar model predictions would be based on the most current nuclear physics, so that meaningful uncertainties could be placed on neutrino parameters derived from solar neutrino measurements, such as the mixing angle . An additional source of uncertainty had arisen at the time, with the advent of three-dimensional (3D) radiative-hydrodynamic models of the Sun’s atmosphere for determining element abundances from photo-absorption lines [2006NuPhA.777....1A, 2009ARA&A..47..481A]. This improved analysis, though, lowered the inferred metallicities, leading to tension with the Sun’s interior sound speed profile determined from helioseismology [2002RvMP...74.1073C]. The discrepancy between higher metallicity models that accurately reproduce the Sun’s interior sound speed and lower metallicity models based on the most current treatment of the photosphere was termed the “solar composition problem”. While the neutrino fluxes were not known to the precision needed to distinguish between the competing models, it had been shown that the solar core metallicity could be extracted directly from future CN solar neutrino measurements, if the precision of associated nuclear cross sections were improved [2008ApJ...687..678H]. Reflecting the impact of new experimental work and improved theory, the SF II study of 2011RvMP...83..195A significantly revised the SF I cross section for the driving reaction of the CN (or CNO-I) cycle, and updated key pp-chain cross sections.
The present SF-III review began with a four-day workshop in July, 2022, “Solar Fusion Cross Sections III.” The meeting was hosted by Physics Frontier Center N3AS (Network for Neutrinos, Nuclear Astrophysics, and Symmetries)111https://n3as.berkeley.edu/, a consortium of institutions involved in multi-messenger astrophysics, whose central hub is the University of California, Berkeley. Sessions were held at the David Brower Center, Berkeley, and the Physics Department, UC Berkeley, with the assistance of N3AS staff. The workshop was co-hosted by N3AS, by Solar Fusion’s long-term sponsor, the INT, and by the European ChETEC-INFRA Starting Community for Nuclear Astrophysics222https://www.chetec-infra.eu/.
The 47 workshop participants represented most of the leading experimental and theoretical research groups that are active in the field. Workshop participants were organized in nine working groups, who were charged with reviewing and evaluating work completed since SF II, to produce updated recommendations. The structure of this review largely follows that of the working groups. The timing of SF-III was driven in part by planned large-scale asteroseismic surveys of hydrogen-burning stars such as ESA’s PLATO333https://platomission.com/ mission [2014ExA....38..249R], and the expectation that associated data analysis will require large libraries of stellar models in which masses, ages, metallicities, and other parameters are varied. SF-III will provide the best current nuclear astrophysics input for such modeling. This review summarizes that input, based on the recommendations of the nine SF-III working groups.
I.2 Terminology used
The terminology used here follows recent practice, as summarized for example in the NACRE compilation [1999NuPhA.656....3A]. It is briefly described to aid the reader.
Nuclear reactions are designated as
where is the target nucleus, is the bombarding particle, and and are the reaction products, with distinguished as the heavier product. Frequently in nuclear astrophysics, the reaction cross section is re-expressed in terms of the factor , with related to as described below, and with the mass numbers of entrance channel nuclei and .
The relation between the nuclear cross section and astrophysical factor is
with and the charge numbers of the target and projectile, the fine structure constant, the relative velocity of the interacting particles in units of , the reduced mass, and the center-of-mass energy. (In this review we use energy units of MeV unless otherwise stated.) This removes the sharp energy dependence from the cross section associated with point Coulomb scattering. The quantity is known as the Sommerfeld parameter. Consequently, unless there are resonances, varies gently, and at low energies can generally be expressed as a low-order polynomial in .
The rapid rise in the nuclear cross section with energy, reflecting the weakening of the Coulomb barrier due to the dependence of the Sommerfeld parameter, competes with a rapid decrease in the probability of finding interacting particles with the requisite center-of-mass energy , in the high-energy tail of the Maxwell-Boltzmann distribution for the solar plasma. Consequently, the thermonuclear reaction rate for a given stellar temperature takes the form
where is proportional to the Sommerfeld parameter of Eq. (2) and is Boltzmann’s constant. The maximum of the integrand defines the Gamow peak – the most probable energy for interactions – while the range around the peak is frequently termed the Gamow window.
For the solar reactions of interest, the energy of the Gamow peak is well below the height of the Coulomb barrier. The slowly varying can be expanded as a power series around , with the leading term being , and with corrections given by derivatives taken at ,
Most of the results presented in this review will be given in terms of and its derivatives.
For most of the nuclear reactions studied here, the notation is adopted, where and are the mass numbers of the projectile and target nuclei, respectively. A summary of SF-III reactions and recommended astrophysical factors are presented in Table 1.
| Reaction | (MeV b) | (b) | (MeV-1 b) | Section | |
| III | |||||
| see text | IV | ||||
| V | |||||
| – | VI | ||||
| – | – | VII | |||
| – | – | IX | |||
| – | – | LABEL:subsec:WG3:Total_$S$_factors | |||
| LABEL:sec:OtherCNO:CNO1:c12pg | |||||
| LABEL:sec:OtherCNO:CNO1:c13pg | |||||
| LABEL:sec:OtherCNO:CNO2:n15pg | |||||
| LABEL:sec:OtherCNO:CNO1:n15pa | |||||
| LABEL:sec:OtherCNO:CNO2:o16pg | |||||
| – | – | LABEL:sec:OtherCNO:CNO2:o17pg | |||
| – | – | LABEL:sec:OtherCNO:CNO3:o18pg | |||
| – | – | LABEL:sec:OtherCNO:NeNa:ne20pg | |||
| – | – | LABEL:sec:21nepg | |||
| – | – | LABEL:sec:OtherCNO:NeNa:ne22pg | |||
| LABEL:sec:OtherCNO:NeNa:na23pg |
I.3 Scope and structure of the review
In the Sun, the primary mechanism for the conversion of four protons into 4He is the proton-proton chain of Fig. 1, during which two charge-changing weak interactions take place, each converting a proton into a neutron with the emission of a neutrino. The pp-chain consists of three main branches, ppI, ppII, and ppIII. These branches are distinguished by their dependence on the solar core temperature and by the neutrinos they produce, with the pp, 7Be, and 8B neutrinos serving as “tags” for the ppI, ppII, and ppIII chains, respectively. The SSM predicts that the fluxes of pp, 7Be, and 8B neutrinos will have an approximate relative temperature scaling of 1:T:T, where Tc is the solar core temperature [annurev:/content/journals/10.1146/annurev.aa.33.090195.002331]. By the mid-1990s solar neutrino experiments had established that the measured fluxes differed significantly from this expected pattern. This contributed to growing expectations that new neutrino physics might be the solution of the Solar Neutrino Problem. It also was an important motivation for SF I, as errors in the nuclear physics of the pp chains could distort the relationship between the various fluxes.
The pp chains account for 99% of solar hydrogen burning and are the main focus of this review. The remaining 1% is generated through the CN cycle (or CNO-I cycle, see Fig. 2). Motivated in part by Borexino’s recent success in measuring the flux of solar CN neutrinos, this review also considers in more detail than past Solar Fusion reviews the reactions driving hydrogen burning in stars more massive than our Sun.
The review is structured as follows. Section II begins with a summary of past solar neutrino experiments, highlighting some of the open questions that remain. It discusses the second-generation experiments Super-Kamiokande, the Sudbury Neutrino Observatory, and Borexino that helped resolve the solar neutrino problem, as well as new detectors in various stages of development, including Hyper-Kamiokande, SNO+, JUNO, DUNE, and others.
The cross section of the driving reaction of the pp chains, , is too small to be measured and thus must be taken from theory. This important reaction is reviewed in section III. All the other pp-chain reactions and decays have been studied in the laboratory, and used together with theory – which guides fits and extrapolations, and corrects for difference in the atomic environments of solar and terrestrial reactions – to predict solar fusion rates. The 2H(He, 3He(3He,2)4He, and 3HeB reactions are reviewed in Secs. IV, V, and VI, respectively. The minor hep pp-chain branch is reviewed in Section VII, and the electron-capture reactions on pp and 7Be are discussed in Section VIII. The 7Be(,B reaction responsible for the high energy neutrinos measured in Super-Kamiokande is treated in Section IX
While a modest (%) contributor to solar energy production, the CN (or CNO-I) cycle is a potentially important probe of core metallicity as its hydrogen burning is catalyzed by the Sun’s primordial C and N. It is also the primary mode of hydrogen burning in stars more massive than the Sun, with central temperatures above 2K. This review treats the CNO cycles in somewhat more detail than in SF I and SF II. In SF II the cross section for the leading reaction, 14N()15O, was revised downward by a factor of two from the SF I recommended value. The SF-III value is close to that of SF II but with an updated uncertainty evaluation described in Section LABEL:sec:S114.
A current focus of laboratory astrophysics is the higher-temperature pathways for hydrogen burning that operate in massive stars. Many of the contributing reactions have been re-measured. That progress is reviewed in Section LABEL:sec:OtherCNO. Reactions of the NeNa cycle are also treated in this section.
Charged-particle nuclear reactions occurring at energies below the Coulomb barrier are affected by screening, the shielding of the nuclear charge by electrons. Because of their very different atomic environments, laboratory reactions and those occurring in the solar plasma are affected in distinct ways. Section LABEL:sec:Screening describes the current status of efforts to account for these differences, when solar cross sections are extracted from laboratory measurements.
The energy transport inside the radiative zone is affected by the metals that were incorporated into the Sun when it first formed. While metals comprise less than 2% of the Sun by mass, they play an outsized role in determining the opacity. Very few of the needed radiative opacities have been measured in the laboratory, and the conditions under which measurements are made typically are not identical to those in the Sun. Section LABEL:sec:Opacities describes the current state of the art, and discusses connections between opacities and solar composition that have complicated efforts to resolve the solar metallicity problem.
Much of the progress made in constraining the reactions of the pp-chain and CNO-I cycles has been enabled by new laboratories and the facilities they house. The underground LUNA laboratory is a prominent example [2018PrPNP..98...55B]. Section LABEL:sec:Facilities describes the capabilities of LUNA and other laboratories responsible for the measurements reviewed here.
II Solar neutrino observations
Solar neutrinos offer a unique tool box for probing both the fundamental properties of these elusive particles, and their interactions with matter, as well as understanding their source: the fusion reactions that power our Sun. The original motivation for observations of solar neutrinos was precisely the hope to probe solar fusion. The first successful experiment beginning in 1967 at the Homestake mine [1998ApJ...496..505C] offered the surprising result of a neutrino flux suppressed to approximately one third of expectation. The intervening years have seen a tour de force of experimental efforts, leading to the resolution of the Solar Neutrino Problem and the confirmation of neutrino oscillations [1998PhRvL..81.1562F, 2001PhRvL..87g1301A, 2002PhRvL..89a1301A, 2003PhRvL..90b1802E, 2004PhRvL..92r1301A]. These high-precision flux and spectral measurements mapped out the details of solar fusion, probed the structure of the Sun, and demonstrated (together with atmospheric neutrino measurements) that neutrinos have masses and mix.
II.1 Open questions
The majority of solar neutrino fluxes have now been measured, in some cases with a precision equivalent to or better than theoretical predictions. Improved precision in several cases can offer further insight. A more precise measurement of the flux of neutrinos from the CNO cycle (Fig. 2) would test important aspect of the SSM including, in particular, the initial abundance of heavy elements [2007A&A...463..755C, 2011ApJ...743...24S, 2022A&A...667L...2K] and the redistribution of the CNO elements from out-of-equilibrium burning at the onset of the main sequence. Solar neutrino flux measurements can constrain physical processes in stars such as chemical mixing. The fluxes are sensitive to solar/stellar model inputs, including the radiative opacities (see Sect. LABEL:sec:Opacities). Measurements of the 8B and 7Be fluxes can offer a handle on the temperature of the solar core, and other environmental factors. A percent-level measurement of the pp flux would provide a test of the luminosity constraint, further probing solar power generation mechanisms. The one branch of the pp-chain that yet remains undetected, the hep flux, has both the highest energy but also the lowest flux of neutrinos from the Sun. The spectrum extends beyond that of the 8B neutrinos, offering a small window for unequivocal observation. Limits have been placed on this flux [2020PhRvD.102f2006A], but a definitive observation will likely require a large, next-generation detector such as Hyper-Kamiokande (HyperK) [2018arXiv180504163H] or the Deep Underground Neutrino Experiment (DUNE) [2019PhRvL.123m1803C]. Finally, the observation of monochromatic neutrinos produced by electron capture reactions on 13N, 15O and 17F (which we refer to as ecCNO neutrinos) [1990PhRvD..41.2964B, 2004PhRvC..69a5801S, 2015PhLB..742..279V] would provide important new physics. They can be used as probes of the metal content of the solar core. They provide a measure of the electron neutrino survival probability at specific neutrino energies MeV within the broad transition region between vacuum and matter-enhanced oscillations, where currently we have no constraints. This would constitute a new test of the large-mixing-angle (LMA) MSW flavor oscillation paradigm. The detection of this subdominant component of the CNO-cycle is extremely difficult but it could be within reach of future very large ultra-pure liquid scintillator detectors [2015PhLB..742..279V].
There is also opportunity to leverage solar neutrinos to understand the interaction of neutrinos with matter. The effect of MSW oscillations has a significant impact on the observed solar neutrino spectrum: with vacuum oscillation dominating at low energy, below approximately 1 MeV, where the survival probability is roughly one half, while matter effects further suppress the flux to a survival probability of roughly one third above approximately 5 MeV. The transition region between these two regimes offers an extremely sensitive probe of the details of the interactions of neutrinos with matter, including the potential to search for new physics such as sterile neutrinos, or non-standard interactions, by looking for distortions to the expected spectral shape.
The same theory of matter effects predicts a small regeneration of electron neutrinos during the night time, as they propagate through the bulk of the earth. This so-called “day/night effect” has been sought after by both Super-Kamiokande (SuperK) and the Sudbury Neutrino Observatory (SNO) [2014PhRvL.112i1805R, 2002PhRvL..89a1302A], but a significant observation is still limited by statistics. Future data from HyperK or DUNE may be needed to confirm our understanding of this prediction of the MSW effect.
Mild tension lingers between measurements of the mass splitting parameters, between solar neutrino experiments and terrestrial data sensitive to the same parameter, from the KamLAND reactor experiment [Esteban2020TheFO]. Further study of solar neutrinos could allow resolution of this question, potentially offering new insights.
II.1.1 The solar composition problem
Solar photospheric abundances, determined with spectroscopic techniques, are a fundamental input for the construction of SSMs, but also they are used as input to nearly all models in astrophysics, including stellar evolution, protoplanetary disks, and galactic chemical evolution. The development of three dimensional radiation hydrodynamic models of the solar atmosphere, of techniques to study line formation under non-local thermodynamic conditions and the improvement in atomic properties (e.g. transition strengths) have led since 2001 to a complete revision of solar abundances. Initial results based on these more sophisticated methods pointed towards a markedly lower solar metallicity [2009ARA&A..47..481A], in particular regarding CNO elements, than results from the 90s based on older techniques [1993oee..conf...15G, 1998SSRv...85..161G].
Recently, two groups have revisited the solar photospheric abundances using modern methods [2021A&A...653A.141A, 2022A&A...661A.140M]. While 2021A&A...653A.141A obtain results consistent with their previous findings, 2022A&A...661A.140M find an O abundance intermediate between that of Asplund’s group and those from the 90s, in agreement with another 3D-based determination by 2011SoPh..268..255C. Interestingly, 2022A&A...661A.140M find higher C and N abundances and, indirectly, a higher Ne abundance due to the larger Ne to O ratio measured in the solar corona [2018ApJ...855...15Y]. This leads to a combined metal-to-hydrogen ratio that is by chance comparable to those from 1993oee..conf...15G and 1998SSRv...85..161G, albeit with a different mixture of elements.
Considering that uncertainties in element abundances are difficult to quantify, it has become customary to consider two canonical sets of abundances, which we refer to as high metallicity (HZ) and low metallicity (LZ) solar admixtures. See, for example, 2017ApJ...835..202V for SSM reference values. In this context, the new solar abundance determinations by 2021A&A...653A.141A fall into the LZ category, while those from 2022A&A...661A.140M are HZ. Solar models employing the LZ abundances fail to reproduce most helioseismic probes of solar properties. This disagreement constitutes the so-called solar composition problem [2004ApJ...606L..85B, 2005ApJ...618.1049B, 2006ApJ...649..529D] that has defied a complete solution. All proposed modifications to physical processes in SSMs offer, at best, only partial improvements in some helioseismic probes, see e.g. [2005ApJ...627.1049G, 2007A&A...463..755C, 2008PhR...457..217B, 2010ApJ...713.1108G, 2011ApJ...743...24S]. The same conclusions are obtained with SSMs computed with the newest LZ [2021A&A...653A.141A] and HZ [2022A&A...661A.140M] abundances as discussed also there. An alternative possibility is to consider modifications to the physical inputs of SSMs at the level of the constitutive physics, radiative opacities in particular. This is possible because most helioseismic probes depend actually not directly on the solar composition, but on the radiative opacity profile in the solar interior, i.e. on the combination of solar composition and atomic opacities (see § LABEL:sec:Opacities). The same can be said for solar neutrinos from the pp chain. Early work [2005ApJ...621L..85B, 2004ESASP.559..574M] already suggested that a localized increase in opacities could solve or, at least, alleviate the disagreement of low-Z solar models with helioseismology, and 2009A&A...494..205C and 2010ApJ...724...98V showed that a tilted increase in radiative opacities, with a few percent increase in the solar core and a larger (15-20%) increase at the base of the convective envelope could lead to LZ SSMs that would satisfy helioseismic probes equally well as HZ SSMs.
The degeneracy between solar composition and opacities can be broken using CNO solar neutrinos, e.g., following the methodology developed in 2008ApJ...687..678H. Such a study was recently carried out by the Borexino collaboration [2020Natur.587..577B, 2022PhRvL.129y2701A, 2023PhRvD.108j2005B]. The Borexino measurement of the CN-cycle neutrino flux (the 13N and 15O fluxes) were used to determine the C+N core abundance. Results show a tension with LZ metallicity determinations, while being in better agreement with HZ mixtures. While the error budget is presently dominated by the uncertainty of the Borexino CNO neutrino measurement, a significant contributor to the error () is nuclear, due to uncertainties in , , and . The interpretation of future improved CNO neutrino measurements will be impacted, unless these nuclear physics uncertainties are reduced.
Very recently, 2024JHEP...02..064G presented a new global determination of all solar neutrino fluxes using all available experimental data, including the latest phases of Borexino. Results from this global analysis are in line with those from Borexino, although the added 13N and 15O fluxes are about 10% lower than the Borexino result alone. A comparison of the solar fluxes with SSM calculations [B23Fluxes] shows that HZ SSMs are in better agreement with solar neutrino fluxes than the LZ SSMs, pointing toward a C+N solar core abundance consistent with HZ abundances. The discrimination that solar neutrino fluxes can offer between solar compositions is at most, however, of the order of 2. (See Table 2 in 2024JHEP...02..064G).
Helioseismology also offers the potential to determine the total solar metallicity, i.e. without disentangling individual element abundances, rather independently from opacities. This relies on using the so-called adiabatic index , which deviates from 5/3 in regions of partial ionization. The underlying technique has been used widely to determine the helium abundance in the solar envelope and it has also been extended to determine solar metallicity. Attempting the latter is difficult because the imprint of partial ionization of metals is quite subtle. Previous work along this line [2006ApJ...644.1292A] found an overall metallicity consistent with HZ abundances, but new work [2024A&A...681A..57B] claims to favor LZ values. While this method depends very weakly on radiative opacities, the abundance determination is degenerate with the equation of state. Independent confirmation of these results would be desirable. In short, the controversy related to the solar composition is far from being resolved, currently with different indicators showing contradictory results.
II.1.2 The Gallium anomaly
With the aid of very intense radioactive sources of 51Cr and 37Ar, tests have been made of the rate of production of 71Ge by neutrino interactions on 71Ga, the basis of radiochemical measurements of the low-energy solar neutrino flux. Initial experiments showed lower rates than expected, with interesting but inconclusive statistical precision. Very recently the BEST experiment [2022PhRvL.128w2501B, 2022PhRvC.105f5502B, 2018PhRvD..97g3001B] has confirmed this anomaly at more than 4. Many possible explanations have been explored, but for each there are contradictions [2023EPJC...83..578K, 2022JHEP...11..082A, 2023PhLB..84237983G, 2023JHEP...05..143B]. A sterile neutrino explanation is disfavored [2022JHEP...10..164G, 2022EPJC...82..116G], particularly because of conflict with solar neutrino limits (see Fig. 3), as well as cosmological bounds [2021PhRvD.104l3524H].
The sterile neutrino contribution to the solar flux is fundamentally limited by the luminosity constraint, that the Sun’s total energy output is the result of nuclear reactions that produce neutrinos, whether active or sterile. If precise neutral-current data were available across the solar spectrum, a completely model-independent limit on a possible sterile component could be determined. The present neutral-current data at low energies (from Borexino) are imprecise but can be supplemented by much more precise charged-current data together with 3-flavor oscillation physics. Still better limits, at the cost of some model dependence, can be obtained with fits of experimental data to solar models, as shown in Fig. 3 above, as well as Fig. 1 of 2022EPJC...82..116G.
The precisely known rate of electron capture on 71Ge to 71Ga, for which the half-life is 11.43(3) d [1985PhRvC..31..666H], places an important constraint on the neutrino absorption rate [2023PhLB..84237983G, PhysRevC.108.035502]. The allowed matrix element that governs neutrino capture to the ground state of 71Ge can be extracted to a precision of 1% from the electron capture rate. The ground-state transition by itself generates a significant anomaly: inclusion of the excited-state contributions doubles the effect. In Appendix A we describe some of the details of both the 1997 extraction of the allowed (Gamow Teller) strength by 1997PhRvC..56.3391B and the recent update of PhysRevC.108.035502, ELLIOTT2024104082.
PhysRevC.108.035502, ELLIOTT2024104082 find the ground-state cross section for absorbing 51Cr neutrinos is 2.5% lower than Bahcall’s value. This work includes the contributions of weak magnetism and radiative corrections, which are shown to be sub-1% effects. An improved extraction of the excited-state contribution to neutrino absorption was performed using data from forward-angle (p,n) scattering. This led to a slight increase in that contribution relative to 1997PhRvC..56.3391B, thereby reducing the 2.5% difference above by about half, for the total 51Cr and 37Ar capture cross sections.
The procedures followed by PhysRevC.108.035502, ELLIOTT2024104082 could be extended to the solar neutrino 71Ga capture cross section, including the contribution from the high-energy 8B neutrinos. That has not been done, but as noted in Appendix A, could help resolve small discrepancies noted there.
Results from the SAGE and GALLEX/GNO experiments remain part of the solar neutrino database used in various global fits to neutrino parameters. A resolution of the anomaly that emerged from BEST and the four gallium detector calibration experiments is important as it would increase confidence in these data.
II.1.3 The Boron-8 neutrino spectrum
The neutrino spectrum from , extending to approximately 17 MeV, has been the most accessible part of the solar neutrino spectrum [2016PhRvD..94e2010A, 2020PhRvD.102f2006A], playing an important role in disclosing the physics of neutrino oscillation and testing the SSM. Knowing it with precision, particularly at the high energy end, has renewed interest in the context of observing the hep neutrinos [2006ApJ...653.1545A, 2020PhRvD.102f2006A], which could be accessible with future observatories [2022arXiv220212839A].
In SF II it was decided to recommend the spectrum calculated by Winter et al. [2003PhRvL..91y2501W, 2006PhRvC..73b5503W, Win:2007] based on their measurement of the alpha spectrum from . This spectrum showed excellent agreement with an independent experiment [2006PhRvC..73e5802B]. Both radiative and recoil-order corrections were included by Winter et al. The radiative corrections are relatively small, due to a cancellation between the real and virtual contributions [1995PhRvD..52.5362B]. The recoil-order corrections are dominated by the weak magnetism part, which has been deduced from measurements of the analog electromagnetic decays [1995PhRvC..51.2778D].
Since then three additional measurements [2011PhRvC..83f5802K, 2012PhRvL.108p2502R, 2023PhRvC.107c2801L] of the -decay alpha spectrum have been performed. All three find that the peak of the alpha spectrum appears about 20 keV lower than determined by Winter et al. Overall, these new measurements yield differences in the neutrino spectrum of 5% below MeV, where the contribution becomes small compared to the hep component. Figure 4 shows the 8B spectral uncertainties on the scale of the hep spectrum. With resolution effects considered, it will be difficult to extract the hep flux from solar neutrino measurements without further reduction in 8B uncertainties.
In producing a recommendation for the neutrino spectrum we considered the following:
1.
Apparently the uncertainty estimations for the neutrino spectrum in 2012PhRvL.108p2502R are not quite correct. 2023PhRvC.107c2801L show that the uncertainty due only to the weak magnetism part, for which all authors follow similar prescriptions, is larger in the 0-12 MeV range than the overall uncertainty estimated by Roger et al. We estimated the uncertainties due to the weak magnetism term only and found agreement with those of Longfellow et al.
2.
All three of the recent efforts [2011PhRvC..83f5802K, 2012PhRvL.108p2502R, 2023PhRvC.107c2801L] used detectors (Si strip) more complicated than those used by Winter et al. (a Si surface barrier detector). The strip detectors allow for better position resolution and reduced beta summing. However, for detection from sources external to the detectors (as opposed to detection from sources implanted deep into a detector), the complicated pattern of dead layers needs to be measured and taken into consideration for accurate calibrations. Effects of partial charge collection in-between strips are also significant. The effects are different for a calibration source than from the emission from trapped as an ion.
Based on these considerations, we recommend the spectrum of 2023PhRvC.107c2801L, but suggest that conclusions sensitive to the choice of neutrino spectrum, such as evidence for hep neutrinos, should be evaluated using both the Winter et al. and the Longfellow et al. spectra.
Given the importance of the weak magnetism contributions, it would be good to have a new experiment with reduced uncertainties that could be compared directly to the results of 1995PhRvC..51.2778D. Recoil-order effects in the system were most recently measured by 2011PhRvC..83f5501S, making use of the alignment- correlation. Using combinations of their measurements and correlation data from 1980PhRvC..22..738M, they were able to extract some of the recoil-order matrix elements. Calculations of these matrix elements in the symmetry-adapted no-core shell model were recently performed by 2022PhRvL.128t2503S. The comparison to the experimentally derived matrix elements of 1980PhRvC..22..738M, 1995PhRvC..51.2778D, and 2011PhRvC..83f5501S is not straightforward because the experiments quote averages over the whole beta spectrum as opposed to fits for each of the four levels used in the calculation. A more detailed comparison of theory to experiment that takes into account the averaging would be helpful. An additional point of comparison that has been used previously is the beta spectrum from [1996PhRvC..54..411B]. An experiment with improved statistics and systematics that extends over the full range of beta energies would be very useful. The existing data from 1987PhRvC..36..298N is in good agreement with the alpha spectrum from Winter et al. But given the apparent 20 keV difference that the newer alpha-spectrum experiments described above report, it would be good to have a modern beta spectrum measurement specifically designed to address this difference.
II.2 Experimental program on solar neutrinos
A broad range of technology can be used to interrogate solar neutrinos. Radiochemical experiments utilizing the interaction on either chlorine or gallium were the first to identify the mystery that became known as the Solar Neutrino Problem. Large monolithic detectors have achieved great success in real-time measurements – from the water Cherenkov detectors such as Kamiokande, SNO, and SuperK, which could use the unique topology of Cherenkov light to point back to the neutrinos’ origin, to liquid scintillator detectors such as KamLAND and Borexino, whose high light yield and low threshold allow for precision spectroscopy.
II.2.1 Super-Kamiokande and Hyper-Kamiokande
The first real-time detection of solar neutrinos was achieved by the Kamiokande experiment [1989PhRvL..63...16H], which detected the neutrinos via elastic scattering (ES) of electrons, a process in which the outgoing electron’s direction is highly correlated to that of the incoming neutrino. This allowed Kamiokande to directly point back to the neutrinos’ origin: our Sun. The successor experiment, Super-Kamiokande, is a 50-kton volume of pure water, surrounded by over 10,000 photon detectors, which has achieved an unparalleled program of neutrino observations and other physics over its several decades of operation. This program has included the highest-precision measurement of the ES signal from solar neutrino interactions, as well as sensitive searches for the day/night distortions of the spectral shape [2016PhRvD..94e2010A]. Now filled with a gadolinium additive to enhance neutron capture, SuperK’s primary current focus is the search for the Diffuse Supernova Neutrino Background (DSNB) via inverse beta decay [2023ApJ...951L..27H]. It also continues its atmospheric neutrino program: SuperK measurements made 25 years ago demonstrated the neutrino oscillations were responsible for the puzzling zenith-angle-dependence of this flux [PhysRevLett.81.1562].
In parallel to ongoing operation of SuperK, an even larger sister project is under construction. At over 250 kton in total mass, and with improved light collection relative to SuperK, Hyper-Kamiokande (HyperK) will impact statistics-limited searches the involve the higher energy 8B solar neutrinos, including day/night effects, the search for hep neutrinos, and the shape of the 8B neutrino spectrum. After 10 years of operations, HyperK will reach a sensitivity to day/night effects 4 (8) , given the oscillation parameters deduced from reactor (solar) experiments. HyperK will serve as the far detector for JPARC’s long-baseline neutrino program.
II.2.2 Sudbury Neutrino Observatory and SNO+
SNO’s unique use of heavy water as a target medium offered the ability to detect solar neutrinos via two additional interactions besides ES. In the charged current (CC) interaction an electron neutrino interacts with the neutron in the deuteron, producing an electron and a proton. This interaction path is sensitive only to electron-flavor neutrinos at the few-MeV energy scale of solar neutrinos, providing a measurement of the pure flux. SNO also had access to the neutral current (NC) interaction, in which a neutrino of any flavor interacts with the deuteron, breaking it apart into its constituent nucleons. Being equally sensitive to all active flavors, the NC interaction yields a measurement of the total, flavor-blind neutrino flux. It was this capability that allowed SNO to resolve the Solar Neutrino Problem, and demonstrate that the produced in the Sun were changing flavor prior to detection [2001PhRvL..87g1301A, 2002PhRvL..89a1301A, 2004PhRvL..92r1301A].
SNO ceased data taking in 2006. Analysis of the data set has continued, and produced a number of new results, including constraints on non-standard effects such as Lorentz violation and neutrino decay [2018PhRvD..98k2013A, 2019PhRvD..99c2013A], searches for hep neutrinos and the diffuse supernova neutrino background (DSNB) [2020PhRvD.102f2006A], and studies of neutron production from cosmogenic muons and from atmospheric neutrino interactions [2019PhRvD..99k2007A, 2019PhRvD.100k2005A].
After decommissioning, the detector was repurposed for the SNO+ experiment, in which the target material was replaced with a pure organic liquid scintillator (LS): linear alkyl benzene (LAB), loaded with 2.2 g/L of the secondary fluor, PPO. The high light yield of this scintillator, coupled with the location in SNOLAB –at 6-km water equivalent, one of the deepest underground laboratories in the world– and an extremely well-understood detector, allow for a range of high-precision measurement programs. In a preliminary water phase, SNO+ demonstrated detection of neutron captures on hydrogen [PhysRevC.102.014002], an impressive technical achievement in an unloaded water detector; a low background measurement of 8B solar neutrinos [PhysRevD.99.012012]; several searches for invisible modes of nucleon decay [PhysRevD.99.032008, PhysRevD.105.112012]; and detection of antineutrinos from reactors over 200 km away [PhysRevLett.130.091801]. These results demonstrate both the low backgrounds and the technical capabilities of the detector, which will facilitate the future program. In a “partial-fill” stage, in which the upper half of the detector was filled with LS while the lower half still contained water, the first demonstration of event-level direction reconstruction was achieved for 8B neutrinos in a scintillator detector [PhysRevD.109.072002]. This was facilitated by a lower loading of PPO at that time, which results in lower scintillation yield and a slower time profile, effectively enhancing the Cherenkov component, which can be leveraged for directional information.
Now fully filled with LS, SNO+ is the deepest, largest operating LS detector in the world. The future program will include measurements of several solar neutrino fluxes, as well as the 8B neutrino energy spectrum.
II.2.3 Borexino
Water Cherenkov experiments such as Kamiokande, SuperK and SNO provided the first real-time measurement of solar neutrinos but their relatively high energy threshold made them sensitive only to a small fraction of the solar neutrino flux. To study in real time the bulk of solar neutrino emission a different detector technology was required. The Borexino experiment, located deep underground at the INFN Laboratori Nazionali del Gran Sasso, used an organic LS target made of pseudocumene with PPO to detect the ES of solar neutrinos off electrons. The high light yield of the LS made it possible to significantly lower the energy threshold, but given the small signal rate and the lack of a clear signature to separate it from the background (such as the direction indicated by the emission of Cherenkov radiation), the experiment required a long preparatory phase to develop the most advanced techniques to suppress the background, especially the one due to radioactive contamination of the LS itself.
The detector’s extreme radiopurity was key to its success, and over its 14 years of data taking, ending in October 2021, Borexino proved itself capable of covering the entire solar neutrino spectrum. At the beginning of its data taking in 2007 the level of 238U and 232Th contamination in Borexino were lower than and respectively [2008PhLB..658..101B], paving the way for the first measurement of the 7Be sub-MeV solar neutrinos, followed by a low-threshold measurement of the 8B flux and a first indication of the pep neutrinos. After a purification campaign that further reduced the LS contamination, Borexino aimed at improving the accuracy of its results in the measurement of the pp-chain solar neutrinos. The flux of neutrinos produced in the pp fusion process was first measured in [2014Natur.512..383B] and later improved along with all the other fluxes produced throughout the pp-chain, with the only exception of hep neutrinos [2018Natur.562..505B, 2019PhRvD.100h2004A, 2020PhRvD.101f2001A]: the pp flux was determined to an uncertainty of 11%, improving the neutrino-based estimate of solar luminosity, while 7Be neutrinos were measured with an uncertainty of , half that of the SSM prediction. Furthermore, applying a constraint on the flux of CNO neutrinos derived from the SSM predictions, the signal of pep neutrinos was established with a significance larger than for the first time, and the 8B flux was measured with a threshold as low as 3 MeV.
After completing the investigation of the pp-chain neutrinos, Borexino reported the first detection of solar neutrinos produced in the CNO-cycle [2020Natur.587..577B], providing the first direct evidence of the occurrence of that process in stars. Profiting from a larger exposure and a better understanding of the radioactive backgrounds enabled by the unprecedented thermal stability of the detector, Borexino further improved its measurement of the CNO neutrino flux [2022PhRvL.129y2701A, 2023PhRvD.108j2005B], where the new CNO result was used in combination with existing solar neutrino results to probe solar composition.
Borexino’s physics program was not limited to solar neutrinos: its outstanding radiopurity made it a excellent detector for geoneutrinos [2020PhRvD.101a2009A] and for searches for various rare processes. Borexino data were used to constrain exotic properties of neutrinos, such as their magnetic moments [2017PhRvD..96i1103A] and non standard interactions with matter [2020JHEP...02..038B].
II.2.4 JUNO
The Jiangmen Underground Neutrino Observatory (JUNO) is a large liquid scintillator detector currently under construction in an underground laboratory with a vertical overburden of about 650 m (roughly 1800 m water equivalent) in Jiangmen city in Southern China [2015arXiv150807166A]. JUNO is located 52.5 km from two nuclear power plants, a baseline optimized for JUNO’s primary goal, the determination of the neutrino mass ordering [28, 29]. To achieve this, JUNO requires a large target mass (20 kton) and excellent energy resolution, attributes also import in solar neutrino detection. The relatively shallow overburden limits the solar neutrino program, due to cosmogenic activity, but the low threshold and large detector mass may allow measurements of day/night effects, non-standard neutrino interactions affecting the 8B spectral shape, and 7Be neutrinos, as well as providing a new solar determination of . JUNO’s reactor data will constrain to extremely high precision, allowing comparisons between the solar and reactor determinations of this parameter. To the extent that oscillation effects can be treated with greater confidence, the connections between solar neutrino flux measurements and the solar fusion reactions generating those fluxes will be sharpened [2016JPhG...43c0401A].
II.2.5 DUNE
As recommended in the 2023 P5 report [P5Report], the DUNE Collaboration seeks to construct two 10-kton fiducial liquid argon time projection chambers (LArTPC), deep underground in the Homestake mine in South Dakota [2016arXiv160102984A]. A third LArTPC will follow in Phase II, along with other upgrades. The newly developed Sanford Underground Research Facility (SURF) offers 4800-m water equivalent overburden, and will form the far site for a long baseline neutrino program utilizing a high-energy neutrino beam directed from Fermilab to SURF. The DUNE LArTPC detectors may be sensitive to measurements of the high-energy solar neutrinos via CC interactions on argon, which offers good precision on the spectral shape and, thus, the potential to measure day/night effects [2019PhRvL.123m1803C]. The solar neutrino program at DUNE is limited primarily by backgrounds (the detector is optimized for GeV-scale physics) and energy threshold. Multiple technologies are under consideration for a fourth detector module, with the motivation of expanding the physics program, including improved sensitivity to solar neutrinos. Use of underground argon, or alternative technologies such as organic or water-based scintillators, could preserve the long baseline neutrino sensitivity while also opening up a rich program of low-energy physics.
II.2.6 Future prospects
A “hybrid” detector that could utilize both Cherenkov and scintillation light simultaneously could achieve unprecedented levels of particle and event identification and, hence, background rejection [2014arXiv1409.5864A]. The Cherenkov signature offers directional information, while the high light yield scintillation offers precision energy and vertex reconstruction. A full waveform analysis of detected light offers yet more information, based on the pulse shape of the scintillation, which is subject to species-dependent quenching effects, and the impact of the Cherenkov threshold. As a result, both the shape of the waveform and the ratio of the two signals will differ for different particle types. A full-scale detector utilizing novel scintillators along with fast and spectrally-sensitive photon detectors, such as the proposed Theia experiment [2020EPJC...80..416A], or the Jinping detector [2017ChPhC..41b3002B] could achieve percent-level precision on the CNO neutrino flux, as well as improving precision across the suite of solar neutrino measurements, such as the 8B spectral shape, and hep neutrino flux.
As experiments grow larger, and capabilities increase, there are also opportunities to leverage detectors designed for other purposes. Noble liquid detectors built primarily for dark matter, or even long baseline neutrino experiments, may offer sensitivity to solar neutrinos. Noble liquids have high light yields, and are transparent to their own scintillation. Coupled with the high voltage under which such detectors are often operated, the combined ionisation and scintillation signatures can be powerful for discriminating signal from background. The low threshold of these detectors may allow detection of neutrinos from the high-energy tail of the 8B spectrum via coherent elastic scattering off a target nucleus. Solid state detectors built purposely for observing this “CENS” signature have been used to detect neutrinos from stopped pion decay [doi:10.1126/science.aao0990].
III The reaction ()
The cross section for the initial reaction in the pp chain (Figure 1), p+p, is too small to be measured in the laboratory. It must be calculated from the standard theory of weak interactions.
III.1 Introduction and terminology
Near the Gamow peak energy 6 keV for temperatures characteristic of the Sun center, the first and second derivatives of the astrophysical factor at zero energy, and , generate 7% and 0.5% corrections, respectively, in Taylor’s series expansion of around . Higher derivative terms are neglected in this review since they only contribute at the level. The recommended values for and will be discussed in Section III.5. Here we focus on .
At zero relative energy, can be written as [1968ApJ...152L..17B, 1969ApJ...155..501B],
where is the fine-structure constant, MeV is the proton mass, fm-1 is the deuteron binding wave number, being the proton-neutron reduced mass and the deuteron binding energy, and are the Fermi vector and axial-vector weak coupling constants. Finally, is the phase-space factor for the pp reaction with radiative corrections, is the value for superallowed transitions, and is proportional to the transition matrix element connecting the pp and deuteron states.
III.2 Adopted parameters for this review
For the phase-space factor , we use the same value as in SF-II. It comes from the value without radiative corrections, [1969ApJ...155..501B], increased by 1.62% to take into account radiative corrections to the cross section [2003PhRvC..67c5502K]. The main source of uncertainty in arises from neglected diagrams in which the lepton exchanges a weak boson and a photon with different nucleons. These diagrams are estimated to modify by , based on scaling the similar nucleus-dependent correction in superallowed decay [2003PhRvC..67c5502K]. Direct computations of these diagrams were recommended in Solar Fusion II. Here we urge again that this computation be carried out.
For , we use the PDG [2022PTEP.2022h3C01W] value, , whose central value is larger than the 2008 PDG value, , used in SF-II by (or 1.9 ). Naively, this would lead to a increase in the central value of according to Eq. (5). Its effect will be discussed below. For , we take s from the most updated comprehensive analysis of experimental rates with the radiative and Coulomb effects corrected [2020PhRvC.102d5501H]. This value is consistent with s [2009PhRvC..79e5502H] used in SF-II with a larger error.
The dominant uncertainty in comes from the normalized Gamow-Teller (GT) matrix element . Reducing this uncertainty has been the main focus of theoretical work since SF-I. In SF-I, was written as , i.e. the sum of the one- and two-body current matrix elements, and , respectively, with their uncertainties estimated independently. In SF-II, two major steps had contributed to reducing the uncertainty on . The first was a much deeper understanding of the correlation between the uncertainties in and : the overall uncertainty in could be described by a universal parameter that could be fixed by a single measurement. The study of 1998PhRvC..58.1263S demonstrated this phenomenologically in the context of potential-model approaches, while later analysis via effective field theory (EFT) provided a more formal justification [2001PhRvC..63c5501B, 2003PhRvC..67e5206P]. The second step was the use of the precisely known tritium decay rate to fix this universal parameter, as first proposed by 1991PhRvC..44..619C. This has been done in both potential models [1998PhRvC..58.1263S] and in the hybrid EFT approach [2003PhRvC..67e5206P] as explained below.
In SF-II, was determined with three approaches. The first one was the potential model approach. In the most elaborate calculation for the pp fusion process, a comparison of the results for five representative modern potentials designed to accurately reproduce nucleon-nucleon scattering data was carried out [1998PhRvC..58.1263S]. After adjusting the unknown strength of the two-body exchange currents to reproduce , the variation in that otherwise would come from the choice of the phenomenological potential was largely removed. Predictions for five representative high-precision phenomenological potentials fell in a narrow interval . There were additional uncertainties in the three-body potentials and three-body currents in , of the order of , and a 0.5% uncertainty due to effective range parameters for nucleon-nucleon scattering. Hence, the recommended value from the potential model approach was .
The second and third approaches were both based on EFT. The second one was a hybrid EFT (EFT*), which used the current operators derived from EFT in conjunction with the initial and final state wave functions generated by a potential model [2003PhRvC..67e5206P]. For pp fusion, the relevant two-body current contained only one unknown low-energy constant (LEC) which parameterized the contact axial coupling to two nucleons [2003PhRvC..67e5206P]. A weakness of this approach was the mismatch between the operators and wave functions. However, it was argued that the mismatch only happened for short distance physics which could be absorbed by the LECs. Hence, when the ultraviolet cutoff was changed over a physically reasonable range, the residual cutoff dependence of physical observables provided a measure of the model dependence of the EFT* calculation. By combining the error from changing the cutoff in the range of MeV and the 0.4% higher order correction, obtained by multiplying the contribution of the highest calculated order with the small expansion parameter , 2003PhRvC..67e5206P provided the value , in perfect agreement with the one obtained within the phenomenological approach.
The third study was performed with the pionless EFT (EFT) approach. It is a framework applicable to processes with the characteristic momentum much smaller than the pion mass , such that the pion field can be “integrated out” and becomes non-dynamical (1996APS..DNP..BA03K; 1999PhRvL..82..463B; 1999NuPhA.653..386C). In this approach, all nucleon-nucleon interactions and two-body currents are described by point-like contact interactions with a systematic expansion in powers of . For all the deuteron weak breakup processes (e.g. and scattering) and their inverse processes, including the pp fusion, only one two-body current (with coupling ) is needed up to next-to-next-to-leading order (N2LO) [2001PhRvC..63c5501B]. Therefore a single measurement will fix and the rates of all such processes. This feature is shared by the other approaches discussed above. The computation of in EFT was carried out to the second order in the expansion by 2001PhRvC..64d4002K and then to the fifth order by 2001PhLB..520...87B. Constraints on from two-nucleon systems [2002PhLB..549...26B, 2003PhRvC..67b5801C] yielded .
Based on the consistent results of the above three approaches, SF-II recommended . The SF-I value from 1969ApJ...155..501B was not re-computed in SF-II. However, with reaching a accuracy, new calculations of and , together with the full pionful chiral EFT (EFT) computations to remove the unknown systematics of the hybrid EFT, were called for in SF-II. These two challenges have been met in SF-III and will be described below.
III.3 Experimental progress on muon capture of the deuteron
In SF-II it was also recommended to carry out the experimental determination of the muon capture rate on deuteron, as proposed in the MuSun experiment [2003nucl.ex...4019K, 2017APS..DNP.CG007S]. This quantity could be used to constrain without the need to rely on the three-body calculation of . The theoretical calculations for the muon capture rate on deuteron have been carried out in EFT [2005PhRvC..72f1001C] (it is possible to impose neutron energy cut to isolate the low-energy neutron events so EFT is applicable [2003nucl.ex...4019K]), chiral hybrid EFT* [2002PhLB..533...25A], phenomenological potential model [2011PhRvC..83a4002M], and recently in EFT (2012PhRvL.108e2502M, 2018PhRvL.121d9901M; 2018PhRvC..98f5506A; 2023FrP....1049919C; 2023PhRvC.107f5502B; PhysRevC.109.035502). However, the MuSun result is yet to be released. On the other hand, we should notice that the muon capture processes happen at a rather large momentum transfer compared to pp fusion. The momentum transfer dependence of the single nucleon axial coupling constant , with , has been recently studied by 2018RPPh...81i6301H, who have provided an experimental determination for the axial charge radius, given by fm2, where is defined by the relation for small . The % uncertainty on has an impact on the ability of the MuSun experiment alone to directly constrain [2018PhRvC..98f5506A, 2023FrP....1049919C, 2023PhRvC.107f5502B, PhysRevC.109.035502].
In the next few years, lattice Quantum Chromodynamics (QCD) calculations of are expected to reduce the uncertainty by a factor of two or more (see 2023slft.confE.240M for the most recent review of the lattice results). Currently, lattice QCD results are consistent with 2018RPPh...81i6301H on , often with comparable uncertainties. At larger ( GeV2), there is a growing tension between lattice QCD predictions of and the phenomenological determination from older neutrino-deuterium bubble chamber data [2016PhRvD..93k3015M] with the lattice QCD results yielding a 30% larger neutrino-nucleon cross section over a large range of [2022ARNPS..72..205M].
In any case, it is evident that accurate experimental determinations of muon capture rates on the deuteron and other light nuclei, which can be addressed theoretically ab initio approaches, represent fundamental tests for the theoretical approaches themselves, either within EFT or EFT, and might be able to provide the necessary experimental information to fix the unknown parameters of the theory.
III.4 Progress in calculations since SF II
Below we summarize the calculations performed after SF-II using different approaches.
III.4.1 EFT
The pioneering work of 2013PhRvL.110s2503M used the next-to-next-to-next-to-leading-order (N3LO) chiral two-nucleon potential [2003PhRvC..68d1001E, 2011PhR...503....1M] augmented with higher order () two-photon and vacuum-polarization electromagnetic interactions. These corrections reduced by % mainly due to the vacuum-polarization-induced pp wave function distortion. This was consistent with the 0.84% first found in the potential model calculation of 1998PhRvC..58.1263S. This correction was also included in the EFT* calculation of [2003PhRvC..67e5206P]. The relevant LECs were fitted to reproduce the binding energies, magnetic moments, and GT matrix element in to obtain , with the -wave initial state contributing at , which was about the accuracy level of the calculation [2013PhRvL.110s2503M]. However, using EFT at next-to-leading-order (NLO), 2019PhRvC.100b1001A later found that -wave only contributed at the order of . Re-examining the computer programs, 2013PhRvL.110s2503M found an error in the determination of one of the -wave reduced matrix elements (associated with the longitudinal multipole operator). Consequently, 2019PhRvL.123a9901M reported in the Erratum that .
In more recent work, 2016PhLB..760..584A used chiral interactions and consistent currents up to N3LO (called next-to-next-to-leading-order (N2LO) in the original literature since the second order vanished) and developed a robust procedure for the error quantification. In particular, they analyzed a family of 42 interactions [2016PhRvX...6a1019C] with 7 different cutoff values from 450 to 600 MeV. The 26 LECs were fitted to 6 different pools of input data including and scatterings, as well as the binding energies and charge radii of 3H and 3He, the quadrupole moment of 2H, and . This thorough study yielded .
In both of the EFT calculations mentioned above, a widely used relation first proposed by 2009PhRvL.103j2502G that linked the two-body axial current LEC with the LEC from the vertex was employed. However, later 2018PhRvL.121d9901M found that there was a factor missing in this - relation which was then acknowledged in the Erratum of 2019PhRvL.122b9901G. Fortunately, this error was unimportant in muon capture on deuteron [2018PhRvL.121d9901M] and it only affected at the level [2023arXiv230403327A].
In the most recent and comprehensive EFT study, 2023arXiv230403327A compared the above calculations of 2019PhRvL.123a9901M and 2016PhLB..760..584A in detail. In addition to the increase of from using the correct - relation, updating the input parameters to their most recent values increased by in both calculations, mainly due to the increase of from its SF-II value mentioned above. Furthermore, 2019PhRvL.123a9901M received a increase from removing the truncation error of the basis functions which effectively cut off the long distance part of the wave functions, as 2017PhRvC..95c1301A advocated. After these corrections, the difference between 2019PhRvL.123a9901M and 2016PhLB..760..584A was reconciled and the combined result was found to be [2023arXiv230403327A]
2023arXiv230403327A obtained consistent values of using four different EFT models for the nuclear interaction. This value is also in agreement with the result obtained in EFT by 2022arXiv220710176D (see below). In estimating the order-by-order convergence, however, 2023arXiv230403327A used one of the EFT models, which is able to nicely reproduce the deuteron properties already at leading order. Therefore, the EFT error of Eq. (6) is likely to be an underestimate, warranting the enlarged error advocated in Section III.4.4.
Finally, we would like to mention the work of 2022PhRvC.106e5501L, where the power counting of the EFT weak current operator involved in the pp reaction is revisited using renormalization group (RG) invariance as the guideline. In particular, it is argued that the contact two-body axial current proportional to the LEC must appear one order lower than assessed by naive dimensional analysis. Then it can be shown that RG invariance is fulfilled at N2LO. However, the estimate for obtained by 2022PhRvC.106e5501L does not use a value for obtained by fitting , but extracted in order to match the value of obtained in SF-II. Therefore, we will not consider the work of 2022PhRvC.106e5501L in the present evaluation.
III.4.2 EFT
The universal two-body current coupling was determined using in EFT for the first time by 2019PhRvC.100e5502D. This calculation was carried out up to NLO using the dibaryon formulation of 2001NuPhA.694..511B, which partially resumed higher order effective range contributions to improve the convergence. Then, this result was used by 2022arXiv220710176D, with updated input parameters, to obtain
where the electromagnetic correction was not calculated but assumed to be the same as the potential model value, [1998PhRvC..58.1263S]. While this number was not model independent, the model dependence was believed to be well below the assigned error.
The small error assigned to this NLO result has been justified by drawing an analogy from the corresponding electromagnetic processes. Using the same approach, the matrix element at threshold was predicted at NLO within 0.5% to the experimental value after the electromagnetic two-body current was fit to the magnetic moments of 3He and 3H. This indicates that the contribution of the three-body current at N2LO is small in this electromagnetic case. The weak sector is shown to follow the same operator structure and hence provides support for the calculation procedure and the uncertainty estimate.
It is important to note that the reason that the three-body current is an N2LO effect in EFT is related to the non-trivial renormalization of the non-derivative three-body contact interaction, which shows up at LO to absorb the cutoff dependence of Feynman diagrams. The subleading two-derivative three-body contact interaction is expected to show up at N2LO. This interaction, combined with the one-body current, renormalizes the three-body current. Hence, the three-body current should also appear at the same order, N2LO. However, if the non-derivative three-body contact interaction were counted as higher order, such as N3LO, as in EFT (because cutoff independence is not strictly enforced order by order in EFT), then the three-body current would contribute at much higher order. Although the EFT power counting indeed yields good convergence in the expansions, it is unsatisfactory that one can not remove the cutoff dependence at each order of the expansion. In addition, the uncertainty estimate of the EFT still lacks a broad inspection of the specific nuclear EFT potential implementation. As a consequence, we consider the EFT 1.5% error a better estimate for the theoretical uncertainty.
III.4.3 Lattice QCD and Lattice EFT
Ideally lattice QCD would provide a first principles prediction of the pp fusion rate and the GT matrix element of , for both pure QCD and with QED effects incorporated. However, such calculations are quite challenging and not yet available at the required precision. A proof-of-principle calculation was carried out by 2017PhRvL.119f2002S using a background field method to determine both the pp fusion GT matrix element and . This exploratory calculation utilized a single pion mass at the SU(3)-flavor–symmetric point with MeV, a single and relatively coarse lattice spacing of fm, and a single volume. The calculation was performed under the assumption that the two- and three-nucleon systems were deeply bound. Without this assumption, matrix elements computed in the finite volume can be significantly different from those in infinite volume, due to Lellouch-Lüscher factors [2001CMaPh.219...31L] that lead to power-law finite volume corrections. These can range from the few-percent level to [2013PhRvD..88i4507B, 2015PhRvL.115x2001B].
More recent lattice QCD calculations have found that two-nucleon systems at heavy pion masses are in fact not bound (2016JHEP...10..101I; 2019PhRvD..99g4505F; 2021PhRvC.103a4003H; 2021arXiv210810835A). These efforts, which employ interpolating operators more sophisticated than those of 2017PhRvL.119f2002S, suggest there could be large systematic uncertainties affecting the conclusions of 2017PhRvL.119f2002S stemming from misidentification of the spectrum and inaccurate Lellouch-Lüscher factors. In addition, 2021PhRvL.127x2003G found that the two-baryon spectrum may be particularly sensitive to discretization effects, which would also have important implications for the continuum extrapolation of the matrix elements. These issues are discussed in some detail in 2022FBS....63...67T. Finally, results with pion masses MeV are needed for accurate extrapolation to the physical pion mass.
Given these unresolved systematic issues, the result of 2017PhRvL.119f2002S is not included in the present evaluation – even though the extracted value is consistent with our recommended range quoted in Section III.4.4.
The lattice EFT computation of 2015PhLB..741..301R performed the pp fusion calculation by implementing EFT on a spacetime lattice. The purpose of this leading-order study was to demonstrate that lattice EFT could reproduce the infinite volume and continuum result of EFT such that it could be applied to various reactions of astrophysical interest in the future. Therefore, for , this result is considered as a subset of the EFT calculation.
III.4.4 Final Recommendation of
The above discussions show that determinations of from EFT in Eq.(6), EFT in Eq.(7), and lattice QCD (although with unquantified systematics) are all consistent with each other. Furthermore, these values are also consistent with the recommended value of SF-II, provided the central value is increased by 444To know the precise shift requires an explicit calculation. However, if the shift is within the range of 0.8-1.0 , the averaged value and remain the same within the significant digits. to to account for the update. Averaging this value with the EFT value in Eq.(6) and the EFT value in Eq.(7) yields with per degree of freedom to be . This shows that the EFT, EFT and SF-II estimates of are all mutually consistent.
In addition, we would like to advocate adding an additional correlated error to account for any input that would tend to move all results in a coordinated way. For example, from SF-II to SF-III, we experienced the shift due to the update of the PDG value of . It is not inconceivable that or other input parameters or physics could change again by similar amounts: the large PDG inflation factor of 2.7 reflects the tension that continues to exist among measurements [2022PTEP.2022h3C01W]. Therefore, assigning an additional correlated error seems reasonable. This is also in line with the subtleties discussed in Secs.III.4.1 and III.4.2 which call for an enlarged error. Therefore, our final recommended value for is
where we have added the correlated and uncorrelated errors linearly to be conservative.
III.5 Progress in and
Using EFT, 2013PhLB..720..385C computed analytically with all partial waves included up to N2LO. The Fermi matrix element only contributed at the level and was neglected compared with the GT matrix element. The energy dependence of the phase factor of Eq. (5) was the dominant effect in and , the energy dependence of pp scattering was subdominant, while the contribution was much less important in these derivatives than in . Therefore, these derivatives could be predicted more reliably than . Furthermore, the derivatives were computed analytically and were free from errors of fitting to a polynomial. The result was MeV-1 and MeV-2.
In EFT, for keV was fit to polynomials of [2019PhRvL.123a9901M]. Depending on using a quadratic or quartic fit, changed from 12.23 to 10.82 MeV-1 and changed from 178.4 to 317.4 MeV-2. 2016PhLB..760..584A used cubit fit and keV to obtain =10.84(2) MeV-1 and =317.8(13) MeV-2. Recently 2023arXiv230403327A repeated the calculation of 2019PhRvL.123a9901M with the same energy range and cubic fit as 2016PhLB..760..584A, obtaining a consistent result with 2016PhLB..760..584A: MeV-1 and MeV-2. We will take these as the recommended values from EFT.
Although the face values of and from EFT and EFT look quite different, we would like to remark that they actually agree on better than 0.1% below the keV Gamow peak. For massive stars with central temperatures keV, the agreement is better than 0.8% to second order in the derivatives and 0.5% if MeV-3 is included in the EFT result. Hence, we take the average of the EFT and EFT results as the recommended value:
Finally, we comment on the work of 2019PhRvC.100c5805G which performed a study of in a wide energy range. The main focus of this work was to perform a proper energy-dependence analysis of the pp process, in order to reliably extract and . However, the calculation was performed within a phenomenological approach, using a quite simplified model for the nuclear currents (i.e. not including two-body currents, which are well known to be significant) and structure (i.e. neglecting the -wave components in the deuteron wave function). They found results for and compatible with those of Eq. (9). However, since these values were not obtained with state of the art calculations, they have not been considered in the determination of and .
IV The reaction ()
IV.1 Introduction
The 2H(,)3He reaction is the second step in the pp chain (Figure 1). Compared with the reactions mediated by the weak interaction, this reaction occurs much more rapidly. As a result, deuterium is effectively and instantaneously converted to 3He without leaving any observable signature. Therefore, uncertainties in the rate of this reaction do not affect solar energy generation.
However, the 2H(,)3He reaction plays an important role in the development of protostars, because the onset of deuterium burning slows down the protostars contraction and heating, increasing their lifespan. Accurate knowledge of the 2H(,)3He reaction rate, particularly within the few keV range corresponding to the Gamow peak in protostars, is vital for modeling protostellar evolution effectively [1988ApJ...332..804S].
Another astrophysical scenario where the 2H(,)3He reaction plays a key role is Big Bang Nucleosynthesis (BBN), responsible for the production of light elements during the first few minutes of the Universe. Among these elements, deuterium is an excellent indicator of cosmological parameters because its primordial abundance is the most sensitive to the baryon density and critically depends on the radiation density of the early Universe. The reactions involved in the synthesis of deuterium are: production via the well known )2H process and destruction via the 2H(2H,)3He, 2H(2H,)3H and 2H(,)3He reactions [2020JCAP...03..010F].
Since the comprehensive review performed in SF II, there have been both new measurements as well as advances in the theoretical and phenomenological analysis of the 2H(,)3He reaction. The new experimental results have been determined with accelerator-based measurements of the cross section [2019EPJA...55..137T, 2020Natur.587..210M, 2021PhRvC.103d5805T], as well as plasma-based, intertial confinement fusion measurements [2020PhRvC.101d2802Z, 2022FrP....10.4339M]. On the theoretical side, there have been advances in ab initio calculations, where nucleons are the fundamental degrees of freedom interacting among themselves and with the external electromagnetic probe. Finally, Bayesian analysis methods have been used to model the energy dependence of the factor, starting from the ab initio predictions and applying a polynomial approximation to it.
We review these updates and provide recommended values and uncertainties of the factor over the energy range of interest for Solar fusion, based upon a Bayesian averaging of various models.
IV.2 Data sets used for the present review
The 2H(,)3He reaction has a Q value of 5.5 MeV and proceeds through the direct capture mechanism. Different experimental approaches were followed to measure its cross section. Tisma et al. [2019EPJA...55..137T] irradiated deuterated titanium targets with a proton beam and detected the -rays with two high-purity germanium (HPGe) detectors placed at different angles. The final factor is provided at four energies in the 47 - 210 keV range, with approximately 15% uncertainty. More recently, the LUNA Collaboration performed a measurement underground in the Gran Sasso Laboratories, exploiting the six orders of magnitude suppression of the cosmic radiation background [2021FrASS...7..119Z, 2018IJMPA..3343010C]. A windowless deuterium gas target was used and the -rays emitted by the 2H(,)3He reaction were detected by a large HPGe detector at 90∘ with respect to the beam axis. Great care was taken to minimize all sources of systematic uncertainties in the factor data, resulting in high precision and accuracy at the 3% level [2020EPJA...56..144M, 2020Natur.587..210M]. These new results provided stringent constraints on cosmological parameters obtained by comparing the precise primordial deuterium abundance predictions of the standard BBN model with astronomical observations [2018ApJ...855..102C]. A deeper discussion of the LUNA results and their implications can be found in [2021MNRAS.502.2474P, 2021JCAP...04..020P, 2021JCAP...03..046Y, 2021ApJ...923...49M]. Presently a study of the 2H(,)3He angular distribution is ongoing. By studying the shape of the peak produced in the -spectra by the reaction, it is possible to obtain the angular distribution; a dedicated paper containing the results will be published soon. Finally, a new measurement was performed at the Helmholtz-Zentrum Dresden-Rossendorf in the 300-1000 keV energy range [2021PhRvC.103d5805T] using implanted deuterium targets on tantalum backings and two HPGe detectors. The results are affected by very large systematic uncertainties and they show a different trend of the factor at energies above 300 keV with respect to the fit of the LUNA data of 2020Natur.587..210M. To investigate the tension between the LUNA fit and the more recent high-energy data of 2021PhRvC.103d5805T, a new measurement of the 2H(,)3He reaction is planned at the Felsenkeller laboratory in Germany [2019sone.conf..249B].
In addition to the new accelerator-based results, two recent sets of measurements [2020PhRvC.101d2802Z, 2022FrP....10.4339M] have also been performed using the inertial confinement fusion plasma-based platform [2017PhPl...24d1407G, 2022FrP....10.4339M, 2023FrP....1180821G], which has recently begun to be exploited for this type of work, with initial results on other reactions presented in, e.g., [2016PhRvL.117c5002Z, 2017NatPh..13.1227C]. Both of these measurements, which were done at the OMEGA laser facility (see section LABEL:subsec:WG8:Plasma), used laser-driven implosions of spherical plastic-shell capsules filled with H2D2 gas, and measured the emitted -rays using a gas-Cherenkov-detector [2014RScI...85kE124H] that was calibrated applying the technique described in 2019RScI...90l3504Z. The initial experiment [2020PhRvC.101d2802Z] obtained good statistics at an energy of 16 keV by making several repeated measurements, with a final statistical uncertainty of 6% and a systematic one of 17% (dominated by uncertainty in the absolute calibration of the detector). The second experiment [2022FrP....10.4339M] obtained data at three different energies in the 17-37 keV region, with comparable systematic uncertainty, but with larger statistical uncertainty due to fewer repeated measurements. The results obtained on this unique platform agree within error bars with the accelerator-based measurements.
IV.3 Theoretical studies
Nuclear reactions of astrophysical interest in general, and the 2H(,)3He in particular, are of great importance in nuclear theory because the available experimental data can be used to test the adopted theoretical framework. The 2H(,)3He reaction has the great advantage of involving only nuclear systems, and can be addressed with a microscopic ab initio study. This means that the nuclear systems involved in the process are viewed as made up of nucleons, interacting among themselves and with the external electromagnetic probes. Within such an approach, the following ingredients are essential for the calculation: (i) realistic models for the nuclear interactions and currents, possibly rooted in Quantum Chromodynamics (QCD); (ii) a numerical technique able to solve the -body bound and scattering state problem, including the Coulomb interaction without approximation. Such a technique is usually referred to as an ab initio method. The agreement (or disagreement) between ab initio theoretical predictions and experimental data represents a validation (or indicates the necessity of improvement) mostly for ingredient (i), i.e. the models of the nuclear interactions and currents, within the accuracy of the ab initio method, i.e., ingredient (ii). This is why few-nucleon reactions can be used as an “ideal” laboratory, where the ab initio framework can be stringently tested.
The most recent ab initio calculation of the 2H(,)3He reaction is that of 2016PhRvL.116j2501M. Here the pair-correlated Hyperspherical Harmonics ab initio method was used to calculate the initial scattering and final bound state wave functions (see 2008JPhG...35f3101K and 2020FrP.....8...69M for details). The nuclear interaction model adopted to describe the nuclear state in 2016PhRvL.116j2501M consists of a two-nucleon term, the Argonne (AV18) potential [1995PhRvC..51...38W], augmented by a three-nucleon contribution, the Urbana IX (UIX) potential [1995PhRvL..74.4396P]. The AV18 potential can reproduce the large two-nucleon database with a /datum [1995PhRvC..51...38W], while the combination AV18/UIX can describe quite accurately the properties of 3He, the spectra of light p-shell nuclei [2001ARNPS..51...53P], and scattering observables (see for instance 2002PhRvC..65c4002W). The electromagnetic current operator used in 2016PhRvL.116j2501M includes, in addition to the non-relativistic one-body operator, two- and three-body terms required by gauge invariance in a system of interacting particles. These terms were constructed in 2005PhRvC..72a4001M. The model was then tested against various electromagnetic observables, to access the quality of its predictions. As a potential model, however, there is no systematic procedure for assigning uncertainties for observables whose values are unknown.
The results of 2016PhRvL.116j2501M have been found to be about 10% higher than the experimental data of 2020Natur.587..210M. Given the lack of a procedure for quantifying errors in calculations based on phenomenological interactions and currents, it is difficult to access the significance of this discrepancy. This leads us to make several recommendations to the theory community. First, the 2H(,)3He reaction should be studied within the framework of chiral effective field theory, which has reached a degree of accuracy and predictive power comparable to potential-based phenomenology. This approach is formally ab initio and rooted in QCD, and as an operator expansion will provide an estimate of the theoretical uncertainty. Work along this line is currently underway. Second, the angular distribution of the 2H(,)3He capture reaction, and possibly also polarization observables, should be both calculated and measured. This would be valuable even if measurements were limited to higher energies, where they are less difficult. Such measurements would provide a further test of the predictive power of the theory.
IV.4 Phenomenological and Bayesian Analyses
In the energy range of astrophysical interest, the 2H(,)3He reaction does not have any resonance or coupled channels that can give rise to non-trivial energy dependence, and the factor can be modelled by a low-order polynomial in energy [1967ARA&A...5..525F]. Moreover, the recent results from LUNA [2020Natur.587..210M], combined with previous measurements, place stringent constraints on the cross section such that a rigorous analysis can capture both the average value and uncertainty over the range of interest to Solar fusion and Big Bang Nucleosynthesis [2020Natur.587..210M, 2021MNRAS.502.2474P, 2021JCAP...04..020P, 2021JCAP...03..046Y].
Another change since SF-II has been the advent of Bayesian analysis methods to evaluate thermonuclear reaction rates and provide more rigorous uncertainty estimates in a statistical sense [2016ApJ...831..107I]. To describe this approach, we closely follow the analysis of the 2H(,)3He reaction in 2021ApJ...923...49M, and references therein.
Consider an optimization that utilizes Gaussian distributed priors in a constrained Bayesian analysis, such that the augmented (for uncorrelated data) is given by
The first term in the above equation is the standard and the second term derives from the Bayesian constraint on the parameters with prior mean () and width (), which are chosen with some prior knowledge or can be optimized as described below. In the first term, the double sum runs over the datasets () and individual results from each dataset (), with the mean value and stochastic uncertainty of each data point given by and , respectively.
The theoretical model describing the factor data, , is a function of the energy () and a set of parameters () that must be determined. The quoted systematic uncertainties are parameterized by the normalization factors , with a prior of unit normalization and a width characterized by the quoted systematic uncertainty. Within a Bayesian framework, it is straightforward to utilize distribution functions for that are not Gaussian, such as a log-normal or other distributions. The parameters represent a normalization of the model function for a given dataset, which, viewed from a non-Bayesian perspective, can also be interpreted as normalization factors that must be applied to the data to match the “true” underlying distribution.
Finally, are unknown extrinsic uncertainties [2019PhRvC..99a4619D]. If one assumes that a smooth function of energy can accurately describe the data, then the data in a given experimental set may scatter about this presumed “true” value by more than is reflected by the quoted statistical uncertainties. It was suggested in 2019PhRvC..99a4619D that this extra scatter might be explained by some additional source of statistical uncertainty unbeknownst to the experimenter, or by some unknown systematic uncertainty that is different for each data point in the same data set, as opposed to a correlated systematic that affects all data points similarly. In either case, this extrinsic uncertainty can be accommodated in a Bayesian analysis framework by adding the additional uncertainty as a normal-distributed source of noise with a width that is constrained by the data. In 2019PhRvC..99a4619D and 2021ApJ...923...49M, was added as an absolute uncertainty, independent of the energy. For some data, 2022PhRvC.105a4625O suggested adopting instead a relative uncertainty, such that the scale of the extrinsic fluctuations is proportional to the mean value of . This strategy of adding extra extrinsic uncertainty to the data sets can be viewed as an alternative to that of inflating the quoted statistical uncertainties by , which is often used for seemingly incompatible data sets [2022PTEP.2022h3C01W], where is the per degree of freedom. An advantage of the extrinsic uncertainty method is that it uses the observed scatter within a given data set as a measure of the possible size of unreported uncertainties, as opposed to uniformly increasing the uncertainty in all data sets by the same relative amount.
Such a Bayesian analysis of the 2H(,)3He reaction data was performed by 2021ApJ...923...49M, using both a third-order polynomial and the ab initio prediction from 2005PhRvC..72a4001M. As noted in the literature, the energy dependence of 2005PhRvC..72a4001M is more reliable than the absolute normalization. Furthermore, the updated prediction in 2016PhRvL.116j2501M, as noted above, is larger than that of 2005PhRvC..72a4001M. Therefore, in 2021ApJ...923...49M, the ab initio prediction is modeled as
where is the prediction from 2005PhRvC..72a4001M (or alternatively from 2016PhRvL.116j2501M), and and are an unknown scale factor and offset to be determined in the analysis. While was not determined with theoretical uncertainty, model uncertainty is introduced through the parameters and . The resulting mean values and uncertainties of determined from Eq. (11) were found to be comparable. For a given model, after optimizing the posterior parameter distributions, the Baye factor (BF) is proportional to the probability of the model given the data [2017bmad.book.....H]. Therefore, for a fixed data set, the BF can be used as a relative probability of each model, thus enabling a weighted model-averaging procedure. If we assume a uniform likelihood for each model, the expectation value and variance of a quantity is given by
where denotes the expectation of given the model and denotes the probability of the model given the data (), which is given by
with the Baye factor of model .
Similarly, for a given model, in the absence of prior information on the size of an unknown parameter, the optimal width of its prior can be estimated by finding the value of that maximizes the BF, which typically provides a reasonable approximation to marginalizing over the prior width.
| keV | 10 | 20 | 40 | 80 | 91 | 100 | 120 |
| 2.028(51)(9) | 2.644(60)(8) | 3.276(70)(7) | 4.579(94)(5) | 7.31(15)(0) | 8.11(16)(0) | 8.77(18)(0) | 10.24(21)(0) |
IV.5 Summary and Recommendations
| Model | logGBF | weight |
| ab initio, | 1478.7 | 0.585 |
| ab initio, | 1478.3 | 0.414 |
| 3rd order polynomial | 1470.2 | 1.2 |
| 4th order polynomial | 1469.2 | 4.6 |
| 5th order polynomial | 1468.2 | 1.6 |
| 6th order polynomial | 1464.1 | 2.9 |
With such a Bayes model averaging, we can compare and contrast the polynomial parameterizations of 2H(,)3He along with ab initio results predicted by 2005PhRvC..72a4001M, 2016PhRvL.116j2501M, including the extra variance that arises from this set of reasonable models, e.g., using polynomials of different order, as well as the phenomenological model of Eq. (11). Results are listed in Table 3. The first column lists the model (i.e., polynomial or ab initio), the second column the natural logarithm of the Gaussian approximation to the BF, and the third column the corresponding weight in the model averaging. From this model averaging, the resulting prediction for the 2H(,)3He factor at a few representative energies is provided in Table 2, where the first uncertainty arises from the first term in Eq. (IV.4) and the second uncertainty is from the second term, which we denote as model selection uncertainty. The resulting Bayes model average prediction of the factor over the entire kinematic range considered is depicted in Fig. 5, with the gray band representing the 68% coverage probability 555The analysis in Table 3 and Fig. 5 can be reproduced with the code at the git repository https://github.com/nrp-g/leaner.. For this particular reaction, it is interesting to note that the Bayesian analysis strongly favors the phenomenological models of Eq. (11) over the polynomial approximations. One reason for this might be that the scaled ab initio models have only one or two free parameters, and the energy dependence of given in 2005PhRvC..72a4001M, 2016PhRvL.116j2501M is sufficient to accurately describe the various data sets. On the other hand, polynomial approximations are disfavored as they require more parameters to capture the energy dependence. Also of note, in the simplest model, using 2016PhRvL.116j2501M, the scale factor is given by
indicating that the prediction in 2016PhRvL.116j2501M overestimates the 2H(,)3He data by 7.9%, consistent with expectations. In comparison, the third order polynomial fit predicts values of that are higher at and lower at keV as compared to those in Table 2.
V The reaction ()
The 3He(3He,)4He reaction (astrophysical factor ) terminates the pp-I chain. The ratio of its rate to that of the 3He()7Be reaction controls the branching to the pp-II and pp-III chains, so that in SF-I increasing was discussed as a potential solution to the Solar Neutrino Problem. Subsequent experiments, notably a very low energy measurement at the LUNA 50 kV accelerator deep underground in Gran Sasso [1999PhRvL..82.5205B] and a complementary experiment at somewhat higher energies [2004PhRvC..69a5802K] ruled out such an increase in , as summarized in SF-II.
V.1 Shape of the particle spectrum
Since no new absolute measurements of have been reported since SF-II, we consider the same four experiments [1999PhRvL..82.5205B, 1998PhRvC..57.2700J, 2004PhRvC..69a5802K, 1987NuPhA.467..273K]. However, there is new information on the energy spectrum of the emitted protons. This spectrum has recently been measured using Inertial Confinement Fusion (ICF) plasmas for a Gamow peak energy of 165 keV [2017PhRvL.119v2701Z]. The results show significant structure, indicating the presence of a sequential reaction mechanism passing through the unbound ground state of . This spectrum is important for the measurements [1999PhRvL..82.5205B, 1998PhRvC..57.2700J, 2004PhRvC..69a5802K, 1987NuPhA.467..273K] since they determined cross sections by detecting only the protons above an energy threshold.
The efficiency correction by which 1987NuPhA.467..273K account for the threshold is not well documented, but the more recent measurements [1999PhRvL..82.5205B, 1998PhRvC..57.2700J, 2004PhRvC..69a5802K] utilized the genbod event generator [James68-GENBOD]. It employs a simple reaction model without final state interactions, the Pauli principle, or Coulomb effects that are important near the spectrum endpoint. The only angular correlations are those required by energy and momentum conservation, and together these simplifications give simple ellipses for the singles energy distributions of the emitted nuclei. Published proton spectra obtained using accelerator beams do exist (Fig. 2 of 1974PhRvC...9..805D and Fig. 3 of 1987NuPhA.467..273K), and although they are not corrected for instrumental effects, they do not appear to be well described by ellipses.
The 3He(3He,)4He reaction may proceed via several sequential mechanisms, including (with in its ground or first excited state) and (where the di-proton is two correlated protons in a singlet state) [2015PhRvC..92a4003B]. More complicated three-body decay channels, sometimes called direct decays, are also possible. Only a relatively narrow intermediate state (here, the 1 MeV wide 5Li ground state) could produce a peak in the energy spectrum, and precise classification of the reaction mechanism is in general both experimentally and theoretically ambiguous. When coincident detection of reaction products is used [as in 1999PhRvL..82.5205B], possible angular correlations between the reaction products also matter. A reaction through the ground state of would emit the second proton preferentially either along or opposite the direction of the first proton, while di-proton emission would tend to send both protons in the same direction [2015PhRvC..92a4003B].
The solid curves in Fig. 6 show two different R-matrix models of the proton energy spectrum at keV, fitted to the spectrum measured by 2017PhRvL.119v2701Z for MeV. The measured spectrum differs significantly from the elliptical spectrum of genbod. Based on their Fig. 6, the experiment of 1998PhRvC..57.2700J had a detection threshold of about 5 MeV proton energy. The measurements by 1999PhRvL..82.5205B required a coincidence between two detectors, with a detection threshold of 2 MeV proton energy in each detector. Finally, Fig. 19 of 2004PhRvC..69a5802K indicates a detection threshold of about 4 MeV proton energy. We have estimated their sensitivities to the assumed proton spectrum by integrating the curves shown in Fig. 6 above energy thresholds of 2, 4, and 6 MeV. These integrals vary from 3% below to 6% above the result from an elliptical spectrum, depending on the specific threshold and the assumed spectrum.
V.2 Recommendation
Based on these investigations, our recommended fit is the one from SF-II, but with an additional 4% systematic uncertainty. This gives
Only the larger constant term in the uncertainty differs from the previous evaluation.
New measurements of this reaction could provide more accurate absolute cross sections by using lower energy thresholds and improve our understanding of the proton energy spectra and angular correlations. The analysis of new measurements should use Monte Carlo simulations considering a variety of plausible energy spectra and angular correlations, in order to estimate the sensitivity to these effects. We note finally that the proton energy spectrum may depend on , as this has been found to be the case for neutrons from the mirror reaction [2018PhRvL.121d2501G].
VI The reaction ()
The 3He()7Be reaction, or , proceeds via direct capture to the ground and 429 keV first excited states of 7Be. It has been studied experimentally using three main methods: First, by detecting at least two of the three -rays from the reaction and taking the angular correlation with the alpha beam direction into account (prompt- method). Second, by detecting the induced 7Be [2002NuPhA.708....3T] radioactivity (activation method). Third, by counting the 7Be recoils (recoil method).
VI.1 Previous recommendation in SF II
The 3He()7Be factor recommended in SF II was developed in multiple steps. First, a model was selected for the shape of , based on existing nucleon-level calculations [1986NuPhA.460..559K, 2001PhRvC..63e4002N]. A rescaling of the curve was fitted to data, but only at 1.002 MeV center-of-mass to minimize the role of short-distance physics in the models.
In SF I, a possible systematic discrepancy between data from the two methods previously used was discussed. By the time of SF II, in-depth studies from two groups using both activation and prompt methods were available, from LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G] and Seattle [2007PhRvC..76e5801B] groups. These studies did not find any discrepancy between activation and prompt- data in direct comparisons. SF II opted to limit the fitting to the data by the activation and recoil methods. The prompt- data were left out because of their somewhat larger common mode errors, concerns about how well the -ray angular distribution was known, and avoidance of the correlated errors between activation and prompt data from the same experiment.
In SF II, only data published after 1998 were included. This was partly because the newer data are in general better documented than the older data, and partly because the newer experiments were perceived as addressing systematics more quantitatively. Data from each of the retained experiments were used to determine a separate rescaling of the theory curve, and the rescalings were averaged to find a recommended and recommended error. Theory error was estimated by dispersion of the extrapolated among several models, including versions of the main fitting model that were modified to cover the experimentally-allowed range of scattering lengths.
The SF II recommended value was , based on the data by the Weizmann [2004PhRvL..93z2503S], LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G], Seattle [2007PhRvC..76e5801B], and ERNA [2009PhRvL.102w2502D] groups.
VI.2 Theory progress on
Significant theoretical work on this reaction has occurred since SF II, but the basic understanding of its mechanism remains unchanged from the 1960s [1963PhRv..131.2582T]. It is dominated by external direct capture into the two bound states, and most of the dipole strength at low energy arises beyond the range of nuclear interaction. The external capture part of the cross section is determined by the ANCs of the bound states and by scattering phase shifts; near threshold, most of the strength lies at –20 fm [2011PhRvL.106d2502N]. In models with explicit wave functions the shorter-range strength largely cancels out due to effects of nucleon-exchange antisymmetry. All models feature a shallow minimum of near 1.25 MeV, where capture from -waves becomes comparable to that from -waves.
The first fully ab initio calculation [2011PhRvL.106d2502N] appeared just after the SF II analysis concluded; it used the fermionic molecular dynamics (FMD) method and a softened representation of the Argonne interaction [1995PhRvC..51...38W]. It agrees well with both the scale and the energy dependence of the modern factor data, and it is very close to the energy dependence assumed in SF II. Neff’s elastic-scattering phase shifts also agree well with experiment. Notably, Neff found that the dipole strength distribution departs significantly from pure external capture at 3He-4He separations as large as 9 fm, compared with 3-4 fm in potential models.
Another ab initio model [2016PhLB..757..430D, 2019PhRvC.100b4304V] has also appeared, based on the no-core shell model with continuum (NCSMC) method and the chiral interaction of 2003PhRvC..68d1001E, “softened” by the similarity renormalization group (SRG) procedure. Freedom to choose the SRG stopping point allowed exact reproduction of the 7Be breakup energy, which corrects for the main effect on the external capture of omitting the three-nucleon potential. The results agree well with the overall scale of the modern data, but their energy dependence and phase shifts depart from experiment, possibly due to the omission of explicit three-body forces.
An important advance since SF II has been the application of halo effective field theory (halo EFT) methods to astrophysical capture reactions [2002NuPhA.712...37B, 2018EPJA...54...89H, 2020EPJA...56..166P, 2020JPhG...47e4002Z]. EFTs are valid for systems with a natural separation between the momentum scales probed in low-energy experiment (e.g., corresponding to a binding energy or a projectile energy) and the much larger momentum scale where the low-energy degrees of freedom are no longer valid: the word “halo” here refers to nuclei with only shallow bound states. In halo EFT the only degrees of freedom are the initial- and final-state nuclei treated as point particles, plus photons. For , MeV corresponds to the momentum needed to separate a proton from 3He or to excite 4He; the thresholds to separate the two 7Be bound states into 3He and 4He correspond to momenta MeV and 60.9 MeV.
Given a sufficient separation of scales, one constructs a Lagrangian that respects the system’s symmetries and known qualitative features systematically, organized by powers of . This series can be truncated (at leading order or LO, next-to-leading order or NLO, next-to-next-to-leading-order or NNLO, etc.), and the precision of the resulting theory depends on the sizes of the omitted terms. The precision of a calculation and the energy where it breaks down can be estimated by assuming coefficients of the first omitted term to have a “natural” size. Coupling constants of the Lagrangian must be fitted to data, and consistency of the power counting scheme (identification of powers of for the main operator terms of a system) is tested by whether the fitted constants have natural sizes. Low-order coupling constants in a halo EFT can often be identified with familiar quantities like ANCs, scattering lengths, and effective ranges; the description of elastic scattering in halo EFT reproduces the (Coulomb-modified) effective-range expansion [2022JPhG...49d5102P, 2000NuPhA.665..137K, 2008NuPhA.809..171H]. When a separation of scales exists, halo EFT is well-suited to data extrapolation because it avoids the tacit, hard-to-test, and unavoidable prior assumptions present in the short-range parts of models based on explicit wave functions.
Two groups have studied in halo EFT [2018EPJA...54...89H, 2020EPJA...56..166P, 2020JPhG...47e4002Z]. The sizes of the Coulomb interaction and the large -wave scattering length make the correct power counting tricky to establish for this system. The work of 2018EPJA...54...89H and 2020EPJA...56..166P includes a careful examination of possible power-counting schemes and strict adherence to a power counting once established, up to terms of NLO or NNLO, respectively, in two different power-countings. These authors fitted EFTs to factor data, and they examined effects of including or excluding elastic scattering constraints in their fits. When they included scattering phase shifts from 1972NuPhA.195..241B in their fits, they concluded that the large-scattering-length power counting of the NLO theory was favored. This result promotes two-body currents (i.e., contributions not equivalent to external direct capture) to leading order in the theory. The -wave contribution that becomes important above 1 MeV first appears at NLO. In addition to data-fitting errors, these theories have errors estimated to be 10% from EFT truncation.
The work of 2020JPhG...47e4002Z followed a different approach to power counting up through NLO. At MeV, their derived expression for is essentially the same as that from the 2020EPJA...56..166P NLO theory. 2020JPhG...47e4002Z also developed some ad hoc (i.e., not systematically developed) higher-order EFT terms, referred to as partial-N4LO, to test for their impact on the fitting. The additional terms proved not to be required by the data and did not improve the fit; this result was taken to indicate that corrections from omitted terms are not large compared with experimental errors below 2 MeV. Scattering data were not considered in this work apart from very broad priors on scattering length and effective range; the correlated errors in the factor data were taken into account.
Like halo EFT, the phenomenological R-matrix approach avoids a model of nuclear interactions and uses a systematic parameterization to fit data [2010RPPh...73c6301D]. At the time of SF II, had been the subject of very little R-matrix fitting, apparently consisting only of the simple treatment in 2004ADNDT..88..203D that focused on BBN energies. After SF-II an R-matrix analysis using the AZURE2 code [2010PhRvC..81d5805A, azure2-manual] was carried out on both elastic scattering and factor data in conjunction with the Notre Dame experiment [2013PhRvC..87f5804K], and a more elaborate analysis using Monte Carlo sampling to estimate errors was later reported in 2014PhRvC..90c5804D. In the latter analysis both the fitted value and error bars were heavily influenced by the numerous scattering data of 1964NucPh..50..629B. Very recently, the BRICK software package has been constructed to carry out Bayesian parameter estimation for AZURE2 and applied to both capture and elastic scattering at all energies in the 7Be system by 2022FrP....10.8476O. In this work it was shown that markedly different values result from inclusion or not of the older scattering data [1964NucPh..50..629B] alongside the very recent SONIK scattering data of 2024PhRvC.109a5802P.
A small amount of additional theoretical work in more traditional frameworks has appeared since SF-II. This includes fits of potential models to factor data [2018PhRvC..97c5802T, 2021NuPhA100622108T] and some calculations of the RGM type [2017PhRvC..96f4605S, 2019PhRvC..99e4618S].
VI.3 Experimental progress on
Since SF II, five new experiments have been reported: Four from the Madrid and ATOMKI groups by the activation technique, at relatively high center of mass energies [2012PhRvC..86c2801C, 2013NuPhA.908....1B, 2019PhRvC..99e5804S, 2023PhRvC.108b5802T], and one from the Notre Dame group using the prompt- method [2013PhRvC..87f5804K].
Following the approach adopted in SF II, we again only use the recoil and activation data for the fits below. The -ray angular distribution is not known experimentally [2007PhRvC..76e5801B], and the resultant uncertainty increases the common-mode error for the prompt- studies somewhat. (A related experiment has recently concluded at Felsenkeller Dresden. The data suggest a higher than expected anisotropy but are so far only available in the form of a PhD thesis [Turkat23-PhD].) In addition, all of the “modern” works reporting prompt- data except for Notre Dame [2013PhRvC..87f5804K] also include data obtained with other methods, leading to partial correlations between data sets that would complicate fitting.
Following this restriction, our data selection proceeds as follows. First, the four data sets previously used in SF II are carried over here: Weizmann [2004PhRvL..93z2503S], LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G] (only the activation data), Seattle [2007PhRvC..76e5801B] (only the activation data), and ERNA [2009PhRvL.102w2502D].
Two of the archival data sets excluded by the 1998 cutoff date in SF II merit some further discussion here: The data of 1982PhRvL..48.1664O, 1984NuPhA.419..115O consist of two points measured by activation in a 3He gas cell. However, their uncertainties are not separated into statistical and systematic components, making them unsuitable for the data model used below. The work by 1983PhRvC..27...11R consists of one data point that was obtained by averaging two separate activation measurements with a 3He and a 4He gas cell, respectively. We do not use this result because we lack details to verify the background subtraction. We also choose not to use the activation study by 1983ZPhyA.310...91V. In that experiment there were thick entrance foils and the 4He beam was completely stopped inside a high-pressure 3He gas cell. This gave an integrated measurement over a wide energy range, so that the analysis depends strongly on the assumed shape of the curve, and the result was reported only as an extrapolated .
We now consider the new data since SF II. An activation experiment in Madrid [2012PhRvC..86c2801C] used a 3He beam incident on a 4He gas cell, reporting three data points. Another activation experiment was reported by the ATOMKI group, using a 4He beam on 3He gas cells [2013NuPhA.908....1B]. Two higher-energy campaigns at ATOMKI were again performed using the activation method. These latter data are at = 2.5-4.4 MeV [2019PhRvC..99e5804S] and = 4.3-8.3 MeV [2023PhRvC.108b5802T], respectively, above the energy range suitable for halo EFT, and therefore not included.
A detailed study by the Notre Dame group reported 17 data points [2013PhRvC..87f5804K], using the primary -ray from ground state capture and the secondary -ray from the deexcitation of the 429 keV first excited state of 7Be. No activation data are reported in that work, and the -ray detector was placed at just one angle, 90∘. Since only prompt- data are reported, we did not include this data set. In order to test the effects of this decision, we repeated some of our fits (see below, Section VI.4) with the modern prompt- experiments included: the 17 points from 2013PhRvC..87f5804K, 3 prompt- points from LUNA [2007PhRvC..75f5803C] and 8 from Seattle [2007PhRvC..76e5801B]. The extrapolated -factor changed by less than 1%, well within the error bars for the recommended .
Finally, an indirect experiment using the 6Li(3He,)7Be reaction and the ANC technique has recently been reported [2020PhLB..80735606K]. It is left out of our fits due to the additional normalization and theory uncertainties involved in determining an ANC from a transfer experiment, which are larger than the errors in the factor measurements used for the present fits. All the data used in the present fits are summarized in Table 4.
As in SF-II, we model data uncertainties as consisting of a component that is independent for each point and a common-mode component that applies to all data from a given experiment as a multiplicative factor. This separation is well-documented for all of the modern data and is also shown in Table 4. Except at the lowest energies, the common-mode error typically dominates.
| Group and references | Energy | Data points | Common | Rescaling factor | ||||
| range [keV] | used | total | mode uncert. | Higa EFT | Zhang EFT | BRICK | BRICK+S | |
| Weizmann [2004PhRvL..93z2503S] | 4 | 4 | 1.03(2) | 1.02(2) | 1.02(2) | 1.03(2) | ||
| LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G] | 7 | 7 | 1.02(2) | 1.04(2) | 1.05(2) | 1.01(2) | ||
| Seattle [2007PhRvC..76e5801B] | 8 | 8 | 0.96(1) | 0.95(2) | 0.95(2) | 0.98(2) | ||
| ERNA [2009PhRvL.102w2502D] | 27 | 27 | 0.96(3) | 0.94(2) | 0.96(2) | 0.99(2) | ||
| Madrid [2012PhRvC..86c2801C] | 2 | 3 | 0.99(3) | 0.99(3) | 0.97(2) | 0.99(2) | ||
| ATOMKI [2013NuPhA.908....1B] | 4 | 5 | 1.01(3) | 1.00(3) | 1.01(3) | 1.04(3) | ||
VI.4 Data fitting
We base our recommended on fits to the halo EFT and R-matrix parameterizations discussed above. These avoid tacit assumptions present in potential-models and conceptual difficulties involved in combining ab initio constraints with data. The fits presented here differ from the previously published fits of 2020EPJA...56..166P, 2020JPhG...47e4002Z, and 2022FrP....10.8476O mainly in the uniform use of the agreed-upon capture data and uncertainties from Section VI.3 across all fits. We restricted fitting to the MeV range of validity for the NLO halo EFT expressions. Despite uniform handling of capture data, the fits in each framework handle scattering inputs differently for reasons discussed below.
In addition to the total capture cross section and scattering data, the fitted data also include branching ratios for capture into the two 7Be bound states, taken from 2007PhRvC..76e5801B, 2009PhRvL.102w2502D, 2007PhRvC..75f5803C, and 2013PhRvC..87f5804K. These are necessarily from prompt- experiments and suffer from the concerns about angular distribution discussed in Section VI.3. However, their inclusion simplifies the fitting considerably by breaking parameter degeneracies between ground- and excited-state transitions (especially in the fitted ANCs), probably without strong impact on .
In constructing fits we split common-mode and point-to-point errors and “float the norms” of data sets using the cost function
to describe goodness of fit [1994NIMPA.346..306D].
A rescaling factor is fitted for each data set , with its deviation from unity penalized by the common-mode errors given in Table 4. The index sums over all points within a given data set; is a measured cross section, is a predicted cross section, and is the point-to-point error of the th point in data set .
We carried out both frequentist fits that minimize and Bayesian fits based on the posterior probability distribution of a likelihood function computed from . In the Bayesian fits parameters, extrapolated , and their errors were determined by Monte Carlo sampling of the parameter space. The relatively large number of model parameters and the significant fitting degeneracies between some of them make Bayesian analysis a natural choice for finding best fit and confidence intervals for the multi-parameter models applied here. Two of the groups did not produce frequentist fits in which they had confidence.
We performed multiple fits in the NLO halo EFT of 2018EPJA...54...89H and 2020EPJA...56..166P. These fits were carried out both with and without scattering constraints, which mainly impact by removing parameter degeneracies that would otherwise leave the -wave scattering length poorly constrained. The code base for this version of halo EFT incorporates scattering data through phase shifts. These were taken from the partial-wave analysis of 1972NuPhA.195..241B, which at low energy are based mainly on the data of 1964NucPh..50..629B.
One set of fits for the 2018EPJA...54...89H halo EFT was produced by minimization, proceeding in two steps: first LO parameters were fitted to keV data, and then the results were taken as initial values in the search to minimize over all parameters in the NLO theory for the full set of data. These fits gave keV b with the scattering constraint and keV b without (uncertainties being propagated in linear approximation using covariances and partial derivatives). Formally the EFT truncation error from stopping at NLO corresponds to an additional theoretical error that can shift factors by 10%. However, any fitted curve is constrained by low-energy data, which in some sense become effectively renormalization conditions of the field theory; the error on extrapolated should probably be smaller than 10% by an amount that is hard to estimate.
A second set of fits to the 2018EPJA...54...89H halo EFT was carried out using Bayesian methods. Priors for the EFT parameters were developed based on previous experience, and data rescaling factors were incorporated as additional priors. The Bayesian results (including only experimental error) are keV b including the phase shifts by 1972NuPhA.195..241B and keV b excluding them, consistent with the -minimization.
Searches of the parameter space to minimize for the NLO halo EFT of 2020JPhG...47e4002Z ran into difficulties with parameter degeneracy and local minima in the surface; for this formalism we report only Bayesian results. No experimental information about scattering was used, but flat priors on scattering length and effective range were chosen over a 5 range around recent experiment. This fit differs from 2020JPhG...47e4002Z mainly by excluding the Notre Dame data and including the ERNA activation data, and it gives keV b. The error associated with EFT truncation at NLO is estimated in this approach by separately fitting the partial-N4LO theory discussed above. The result suggests that EFT truncation at NLO affects extrapolation from the data to threshold by –3%.
Our R-matrix fits are based mainly on sampling Bayesian posterior probabilities with the BRICK code. We also produced frequentist fits, but we were unable to estimate their errors convincingly. The R-matrix fits used capture data both alone and in combination with the SONIK scattering data [2024PhRvC.109a5802P] – the latter being chosen because of concerns with the 1964NucPh..50..629B data that are discussed in 2022FrP....10.8476O. Since BRICK fits elastic differential cross sections directly, it was not feasible to use the same phase-shift-based scattering constraints as our halo EFT fits. The R-matrix fit that includes scattering data is essentially the “CS” fit of 2022FrP....10.8476O, but with a restriction to only the capture data described in Table 4, only scattering data below 2 MeV center-of-mass, and only the R-matrix parameters relevant below 2 MeV (no level or radiative widths for -wave background poles).
| Fit | ||||
| [keV b] | [MeV-1] | [fm] | [fm-1] | |
| 2018EPJA...54...89H EFT-NLO | 0.554 | -0.56 | 30 | |
| 2020JPhG...47e4002Z EFT-NLO | 0.573 | -0.61 | 46(6) | 29(3) |
| BRICK R-matrix | 0.562 | -0.57 | 24.4(9) | |
| BRICK R-matrix+scattering | 0.531 | -0.47 | 22.6(7) | |
| ANC [2020PhLB..80735606K] | 34 |
Several of the fits are compared in Fig. 7. There it is apparent that they all agree within 5% in the 0–2000 keV energy range, and within 4% in the astrophysically relevant range 0–500 keV. The fit to the 2018EPJA...54...89H NLO EFT lies in the middle of the range, so we adopt it as a reference in the lower panel of the figure. Relative to that fit, the fits to the NLO EFT of 2020JPhG...47e4002Z and the R-matrix fit without scattering rise up as much as 4% higher at 0–200 keV.
The most striking difference among the curves is between the BRICK R-matrix fit including the SONIK scattering data [2024PhRvC.109a5802P] and all the other fits. While the R-matrix fit with scattering finds rescaling factors somewhat closer to unity than the other fits do, its energy dependence in the 0–1000 keV range is qualitatively different, as is especially visible in the lower panel of Fig. 7 where residuals relative to the fitted 2018EPJA...54...89H EFT are shown. The BRICK fit without the scattering data is much closer to all the halo EFT fits, so the recent scattering data apparently have a large impact.
VI.5 Recommended value
Because the fit to the NLO EFT of 2018EPJA...54...89H gives the central result among those attempted here, we adopt it for our recommended . For ease of adoption, it is noted that this recommended curve can be empirically parameterized using the same shape as in SF II, by the numerical equation (following a customary form in the past literature)
Here is the center of mass energy in MeV. Equation (VI.5) reproduces our recommended curve within 0.3% and is only applicable for = 0–1600 keV.
Propagation of uncertainties from measurement to extrapolated is relatively straightforward and unambiguous (at least in the Bayesian fitting). However, the uncertainty due to theoretical approximation is more complicated, and error estimation on extrapolated quantities in EFTs remains an open area of research; estimates from the groups working on are discussed above. We estimate the overall theoretical uncertainty on our recommended factor to be the measured by the 4% dispersion among fitted curves; this is comparable to the more formal estimates.
Our recommendations for the zero energy astrophysical factor and the low-energy slope are then
The given value of applies over = 0-500 keV. Its considerable uncertainty is given by the full span of slopes among the four fits discussed above, which is dominated by the inclusion of elastic scattering in one of the R-matrix fits. The curve is quite straight at low energy in all models; its second derivative is consistent with zero and even difficult to compute for a given fit because of numerical cancellations.
The first error bar for is given by the fit uncertainty, conservatively using the largest value among the various fits, and it encodes the experimental uncertainty. The second error bar reflects the 4% dispersion among theoretical formalisms. The two errors can be added in quadrature for applications. Our recommendation is essentially the same as in SF II, but derived from completely different theoretical approaches that are much more directly connected to the data and much less dependent on model assumptions. It is also based on data over a wider energy range. The dispersion of and among models at the solar Gamow peak is illustrated in Table 5, along with some indication of how fitted scattering lengths and squared ANCs vary among frameworks. (The latter are highly degenerate with short-range parameters in the fitting; in the 2018EPJA...54...89H EFT they are sensitive to small variations in the -wave effective ranges that are fitted instead of ANCs.)
VI.6 Recommendations for future work on
Here we present recommendations for further work that would improve understanding of the 3He()7Be reaction. Our first recommendation for experiment concerns elastic 3He+4He scattering. New experiments over a wide energy range including at least part of the astrophysical range are needed to address the marked impact of the recent SONIK scattering data [2024PhRvC.109a5802P] on fitting, for independent confirmation. Data below 2 MeV would be most useful for consistent fitting of halo EFTs.
Second, new measurements (besides the preliminary data of Turkat23-PhD) are needed for the angular distribution of the emitted -rays. This would relieve concerns about including prompt- data, most notably the Notre Dame [2013PhRvC..87f5804K] that do not include activation measurements, in future fits. These experiments should also be designed to provide new angle-corrected branching ratios for capture into the two bound states.
Our third recommendation is a study of the astrophysical factor over a wide energy range, 50-1500 keV and extending to even lower energies when technically feasible, with small point-to-point errors. Even with a large common-mode error, this would significantly constrain theoretical curve shapes near threshold, and providing an independent check on the influence of scattering data on .
On the theoretical side, further development of ab initio reaction calculations is desirable, improving on points of current difficulty like explicit inclusion of three-body potentials. Calculations with the same potential but by different methods would help to separate computational issues from the impact of potential choice. Further exploration of how ab initio inputs like ANCs and scattering lengths could serve as fitting priors would also be a useful avenue of research.
Authors of future theoretical calculations should consider including -ray angular distributions in their results; until very recently angular distributions were only available from two potential models [1981PhRvC..23...33K, 1963PhRv..131.2582T]. New calculations would inform analysis of experiments, and the angular distribution measurements we recommend would provide tests of the calculations (or fitting constraints).
A source of complication here has been the need for a good meeting point between scattering data and theory. Given the difficulties of constructing a unique phase shift analysis without an underlying theory, and the recent arrival of modern scattering data, theorists should consider including elastic differential cross sections directly as outputs and fitting constraints.
VII The reaction ()
The reaction , also known as the hep reaction, is one of the possible processes taking place after the , and it produces the most energetic neutrinos, with an endpoint energy of 18.8 MeV. It is however the least probable, with its factor, and consequently its rate, even beyond the reach of current experiments. This is due to the fact that the hep reaction is induced by the weak interaction and further suppressed by Coulomb barrier and by the fact that the leading one-body axial (Gamow-Teller) current operator cannot connect the main -state components of the and initial- and final-state wave functions. As a consequence, the one-body axial current non-zero contribution is due to the small components of the and wave functions. Furthermore, other contributions, such as weak magnetism and other one-body corrections to the vector current, and two-body vector and axial currents, and the -waves in the initial state, normally suppressed at solar temperatures, are significantly enhanced. This is further complicated by cancellations between the one-body and two-body axial current contributions.
Due to this large suppression of the contributions that are nominally expected to be the largest, the zero-energy factor, , is unreachable by experiment and only theory can provide an estimate. The most recent studies of this reaction, already evaluated in SF-II, are those of 2000PhRvC..63a5801M and 2003PhRvC..67e5206P. The first one was performed within a phenomenological approach, similar to the calculation of 1998PhRvC..58.1263S for the pp fusion. In this particular case, the and initial and final nuclear wave functions were obtained with the correlated Hyperspherical Harmonics (CHH) variational method [1995FBS....18...25V, 1998PhRvL..81.1580V], using the Argonne (AV18) two-nucleon potential [1995PhRvC..51...38W] augmented by the Urbana IX (UIX) three-nucleon interaction [1995PhRvL..74.4396P]. In this way, the binding energies of 3He and 4He, and the singlet and triplet scattering lengths were in good agreement with experiment. The weak vector and axial transition operators were obtained within a phenomonological approach. In particular, the largest two-body contribution to the axial current arising from the excitation of intermediate -isobar degrees of freedom was included using the transition-correlation operator scheme of 1991PhRvC..44..619C, 1992PhRvC..45.2628S and fixing the nucleon-to- axial coupling constant to reproduce as in the pp case [1998PhRvC..58.1263S].
The second study of 2003PhRvC..67e5206P was performed within the same hybrid EFT approach used for the pp funsion (EFT∗), with the same nuclear weak current operators obtained within EFT as discussed in Section III, the LEC being fixed to reproduce . The initial and final wave functions were derived using the CHH method with the AV18/UIX potential, as in 2000PhRvC..63a5801M.
Following the results of 2000PhRvC..63a5801M and 2003PhRvC..67e5206P, SF-II recommended for the value keV b. Three recommendations were also made: (i) to perform new studies with a broad spectrum of Hamiltonian models, in order to properly access the theoretical uncertainty; (ii) to study other weak reactions for which experiments can test the predictions made by employing the same theoretical ingredients as those used for the hep reaction calculation; (iii) to further understand the relation between the hep reaction and the so-called hen process (). Of these recommended works, only the second one has been performed so far, applying the theoretical framework used for the hep study to the muon capture reactions on light nuclei [2011PhRvC..83a4002M].
Given the lack of further studies on the hep reaction after SF-II, we decide to maintain
as recommended value. Notice that we could have increased by %, in analogy with , due to the updated values of input parameters (mostly ). However, we have decided not to apply such an increase, because it does not derive from a direct calculation, and, most of all, because such an increase lies well within the quoted % uncertainty. However, given the advances in ab initio studies of the last decade and in the development of the EFT framework, we update the recommendations of SF-II as follows: (i) the hep process should be revisited within a fully consistent EFT approach, similarly to what has been done in these years for the pp fusion and other processes, as, for instance, muon captures on light nuclei [2012PhRvL.108e2502M, 2018PhRvL.121d9901M, 2018PhRvC..98f5506A, 2023FrP....1049919C, 2023PhRvC.107f5502B, PhysRevC.109.035502]; (ii) the initial and final wave functions should be calculated with the most accurate ab initio methods, as the (uncorrelated) HH method [2008JPhG...35f3101K, 2020FrP.....8...69M], which has been proven to provide more accurate bound and scattering wave functions than the CHH method, using both local and non-local interactions; (iii) the relation between the hep and the hen process should be understood. The hen process has been studied within EFT and the HH method by 2022PhRvC.105a4001V, but a consistent parallel study of both reactions is still missing.
A relatively recent SNO Collaboration analysis of data from all three of the detector’s running phases yielded a one-sided confidence-interval bound on the hep neutrino flux of 2020PhRvD.102f2006A
a result in agreement with the SSM prediction. The lack of a definite measurement, however, has meant that there is no substantive experimental test of the predicted Various global analyses of solar neutrino data have provided weak evidence for a nonzero flux, e.g., from Bergstrom:2016cbh
which is consistent with the experiment bound given above.
An extraction of the hep flux from experiment will likely require a detailed shape analysis to separate hep neutrinos from the high-energy tail of the 8B spectrum and the low-energy tail of atmospheric neutrinos. Typically the hep spectrum is assumed to have an allowed shape: indeed, in the work reported here spectra are not provided, where -waves, weak magnetism, and similar corrections have been included. Because these corrections are unusually large for hep neutrinos, future studies should provide spectra and evaluate the impact of such corrections on the shape.
VIII Electron capture by and
Electron capture reactions are the sources of lines found in the solar neutrino spectrum. In particular, electron capture on , i.e., the process , also known as the pep reaction, competes with fusion and depends on the same nuclear matrix element. Therefore, the ratio between the pp and pep rates is independent of nuclear physics. Based on this consideration, in SF-II the result of SF-I, which is based on the work of 1969ApJ...155..501B, was multiplied by the radiative corrections calculated by 2003PhRvC..67c5502K, leading to the final result given by
where is the density in units of g/cm3, is the mean molecular weight per free electron, and is the temperature in units of MK. The range of validity is . Given that no new evaluation of the pep rate has been performed since SF-II, the result of Eq. (23) represents also the present recommended value.
Competition between electron and proton capture on 7Be fixes the branching ratio of the ppII and ppIII chains, and thereby the 7Be and 8B neutrino fluxes. Superallowed electron capture on the ground state of 7Be leads to the ground state or the , 478 keV first excited state of 7Li with a measured (terrestrial) branching ratio of 10.44(4)% [2002NuPhA.708....3T]. It is customary to calculate the solar decay rate in terms of the measured terrestrial decay rate, which has the advantage of removing the nuclear physics dependence, but it requires a proper calculation of the densities of electrons at the nucleus. This has usually been done by considering continuum and bound electrons separately, using the Debye-Hückel (DH) approximation pioneered by 1954AuJPh...7..373S for the screening of bound electrons in the solar plasma; the capture rate of continuum electrons is not significantly influenced by electron screening. Furthermore, it is necessary to calculate the atomic probability densities governing the and terrestrial electron capture rates. 2020PhRvL.125c2701F recently measured the ratio of to capture using superconducting tunnel junctions. The recommendation of both SF I and SF II was
which is valid for . This expression is based on the continuum capture rate calculated by 1969ApJ...155..511B, the average ratio of the total capture rate to the continuum rate calculated using three solar models in 1994PhRvD..49.3923B, and the current value of the half-life, 53.22(6) d [2002NuPhA.708....3T]. It agrees within 1% with a density matrix calculation of the rate that makes no assumptions regarding the nature of electronic quantum states in the solar plasma and allows for aspherical fluctuations in the spatial distributions of plasma ions [1997ApJ...490..437G]. The estimated uncertainty of 2% accounts for possible corrections to the Debye-Hückel approximation due to thermal fluctuations in the small number of ions in the Debye sphere, and breakdowns in the adiabatic approximation [1992ApJ...392..320J].
Although calculations have appeared in the literature claiming new plasma effects based on model calculations (see 2013ApJ...764..118S, 2019A&A...623A.126V), we continue to regard the arguments in 2002A&A...383..291B as definitive. That reference presented five distinct derivations demonstrating that the Salpeter formula for screening corrections is valid for pp-chain reactions and solar conditions, up to corrections on the order of a few percent.
IX The reaction ()
IX.1 Introduction
Radiative proton capture on (=3/2-) at solar energies proceeds through direct capture to the ground state of (=2+). This capture occurs predominantly at separations well beyond the range of the strong interaction via transitions from and partial waves, as the amplitude and contributions from higher partial waves are negligible in the energy range of interest [1961NucPh..24...89C]. The importance of this reaction for the determination of the high energy solar neutrino spectrum and the experimental data reviewed by SF I inspired several experiments that, taken together, provided a consistent picture of the energy dependence of the reaction cross section, despite discrepancies in absolute scale. Prior to the review of SF II, most direct measurements were performed using a radioactive target and an intense proton beam. In spite of the great care taken in evaluating the common mode errors (CMEs), a remarkable discrepancy persisted even in the most recent experiments, that was finally handled in SF II by inflating the stated experimental uncertainties. Indirect measurements of the cross section of based on Coulomb dissociation were considered, but not included in the final recommendation of SF II for the zero-energy astrophysical factor , whose uncertainty was dominated by the theoretical contribution. Therefore, our analysis focused on these aspects (discrepancy of experimental data, indirect measurements, theoretical uncertainties), as discussed in the next sections.
IX.2 Experimental Data
IX.2.1 Direct Measurements
The thorough review of published work done in SF II investigated the influence of beam-target overlap, target stoichiometry, beam energy loss, and the backscattering of 8B recoils on CMEs and led to the selection of a homogeneous group of well-documented data sets. Since then, no further information became available on these aspects of direct measurements. The data that formed the basis of the SF II recommendation were those of 2003PhRvC..67f5805B, 2003PhRvL..90b2501B, 1998PhRvL..80..928H, 2001PhRvL..86.3985H, 1999PhLB..462..237H, 2001NuPhA.696..219S, and the BE3 data set of 2003PhRvC..68f5803J. Here we adopt the revised BE3 data presented by 2010PhRvC..81a2801J. In addition to the complete set of radiative capture measurements that were the basis of the SF II recommendation for , we consider the only new experiment reported thereafter of 2022PhLB..82436819B, which used a radioactive ion beam and the recoil mass separator ERNA to detect the 8B recoils. While these new results contribute to the determination of with different, well controlled systematics [2018EPJA...54...92B], the relatively large statistical uncertainty of the measurement limits its impact.
New scattering data have been published. Elastic and inelastic scattering cross sections at relative kinetic energies between 474 keV and 2.74 MeV were measured and wave scattering lengths were inferred by 2019PhRvC..99d5807P. At higher energies between 1.6 and 3.4 MeV, thick target elastic and inelastic scattering excitation function measurements were used to infer the existence of new resonances above 1.8 MeV [2013PhRvC..87e4617M]. These high lying resonances have no direct influence on .
IX.2.2 Indirect Measurements
In SF II, Coulomb dissociation measurements based on the formalism developed by 1986NuPhA.458..188B were not included in the determination of due to disagreements over whether the analyses of the various measurements properly accounted for the contributions of transitions. Despite these reservations, the Coulomb breakup data were found to be in agreement with radiative capture measurements. In order to include these consistent data in formulating our recommendation for , we inflated the common mode errors of the Coulomb breakup measurements to adequately account for the uncertainty in the components assumed in each analysis.
The common mode errors are obtained from a linear sum of those given in the manuscripts and an additional common mode error due to the estimated size of the component in each measurement. This additional error is applicable in one direction only, downward for the 1998EPJA....3..213K and 2006PhRvC..73a5806S analyses of the RIKEN and GSI measurements, respectively, and upward for the 2003PhRvC..68d5802D analysis of the Michigan State University (MSU) measurement, according to whether the component was included or excluded. Though the 2003PhRvC..68d5802D analysis did include a common mode error contribution of 2.5% due to the component (which represented 5% of the total cross section), we have included an additional 2.5% contribution to the MSU experiment. The size of the additional common mode errors for the RIKEN and GSI measurements was based on scaling this 2.5% according to the predicted size of the contribution to the cross section in these measurements relative to that in the MSU measurement, evaluated using first-order perturbation theory and the and factors calculated in a potential model of 8B [1996ZPhyA.356..293B]. Thus the additional single-sided common mode errors added to the GSI, MSU, and RIKEN measurements are -2.1%, 2.5%, and -4.3%, respectively.
IX.3 Theory
The size of the theoretical contribution to the uncertainty in the SF II estimate of was twice that of the experimental contribution. Since then, significant progress has been made. A new potential model calculation was performed by 2019NuPhA.983..175D. The large-scale computational demands of calculations using high-quality nucleon-nucleon (NN) and three-nucleon (3N) interactions beyond =5 had hindered the ab initio approach in the past. Nevertheless progress has been rapid and it has been possible to extend ab initio calculations to , albeit without 3N forces. 2011PhLB..704..379N performed such a calculation using the no-core shell model in the continuum with a nucleon-nucleon interaction derived from chiral EFT at fourth order. The calculation converged with respect to the model space size and reproduces experimental data for the factor despite having virtually no free parameters (though the similarity renormalization group evolution parameter was tuned to reproduce the binding energy of 8B). The final value of =19.4(7) eV b agrees with the SF II recommendation and, as will be shown later, with our analysis. More recently, a new approach was introduced to combine experimental measurements with ab initio predictions, resulting in an “ab initio-informed evaluation” Kravvaris23-PLB that arrived at a value of eV b.
A significant theoretical advance was the application of the formalism of halo EFT developed by 2002NuPhA.712...37B to 7Be(B calculations by 2015PhLB..751..535Z, 2018PhRvC..98c4616Z, and 2022PhRvC.106a4601H. It treats the incoming and 7Be as point-like particles and the 8B final state as a bound state of the two. The calculation and the associated errors depend upon the accuracy with which one is able to describe the incoming scattering states, the final bound state and the relevant electromagnetic currents. The relative kinetic or CM energy range MeV, corresponding to a relative momentum MeV/c, is considered within the domain of applicability of halo EFT. The two cluster physical description of is expected to hold at 3He- relative momenta below a physical cutoff MeV/c set by the threshold (binding momentum) for breaking the 7Be core into these constituents. Further, at energies above the excitation energy of the first excited state, MeV, including the contribution of the excited core is imperative in the EFT.
2015PhLB..751..535Z and 2018PhRvC..98c4616Z include such an excited contribution. 2022PhRvC.106a4601H provides a calculation without the contribution below and one with it above that also includes an contribution to the resonance. The two groups’ halo EFT calculations with the contributions differ in several respects that are elaborated in 2022PhRvC.106a4601H. Differences in the final bound state calculation affect the treatment of the short-range interactions and divergences in the EFTs. The interpretation of the asymptotic normalization constants (ANCs) in terms of elastic scattering parameters is affected. However, the momentum dependence of the capture cross section is not impacted by the ANCs, which are fitted to capture data. Thus there is no effect on the cross section. In the incoming channel, 2022PhRvC.106a4601H use a more general form of the -wave short-ranged interaction involving the core based on the low-energy symmetry that is included only at next-to-next-to-leading order (NNLO). In contrast, 2015PhLB..751..535Z and 2018PhRvC..98c4616Z include a short-ranged interaction at LO that could make a difference in the momentum dependence of the cross section. However, as shown in an order-by-order calculation [2022PhRvC.106a4601H], the short-ranged interaction is a sub-leading effect so the two halo EFTs should have similar low-momentum dependence. The third important difference is in the treatment of two-body currents, which affects EFT error estimates. 2015PhLB..751..535Z and 2018PhRvC..98c4616Z include this at next-to-leading order (NLO) whereas 2022PhRvC.106a4601H estimate it to be an N3LO effect. Bayesian estimates of the two-body current [2015PhLB..751..535Z, 2018PhRvC..98c4616Z] are not strongly constrained by capture data, consistent with this being a higher-order contribution.
Another widely used approach for describing the cross sections of low-energy nuclear reactions is R-matrix theory. In R-matrix theory the 7Be(B cross section can be parametrized in terms of a few parameters representing either “real” or “background” poles. The real poles correspond to the observed states and resonances, while the background terms correspond to the “mean-field” effects that are of the non-resonant type. In the present case for example, a background pole in the 1- channel is needed so as to reproduce the non-resonant part of the capture. The parameters of the real pole corresponding to the bound state are determined by the binding energy and the two ANCs, while the 1+ resonance parameters are determined by its energy and partial decay widths (both proton and radiative). The R-matrix analysis of this reaction carried out by 1995NuPhA.588..693B arrived at a value of eV b; however, it was based on data that were excluded in SF II and here, with the exception of those of 1983PhRvC..28.2222F.
We fit data using R-matrix theory and halo EFTs, with both frequentist and Bayesian approaches, investigating the impact of different choices for the energy range in which data were fitted, and obtained consistent results. Our recommended value and its uncertainty are based on a Bayesian fit of both direct and Coulomb dissociation data taken at energies using the 2022PhRvC.106a4601H halo EFT⋆. This theory provides a good description of capture data over a wide energy range including the contribution from the resonance. The theoretical uncertainties are well understood. Further, the contributions of the and spin channels to the capture cross section are self-consistently parametrized with the elastic scattering information contained in the known ANCs for this system.
The Bayesian formalism naturally incorporates prior information concerning the ANCs in the fits. The upper limit on the energy in the fit was determined to be high enough to include some data from all the modern experiments, and yet low enough that the contribution of the wide resonance at does not exceed of the total cross section at any energy, as estimated by our R-matrix analysis. The fitting procedure is described in the next section, whereas the details, as well as the results of the fits done in different energy ranges and with different theories are reported in the supplemental material.
IX.4 Fitting Procedures
In SF II, it was noted that the different data sets had similar energy dependences. The discrepancies among the data sets were primarily due to different absolute normalizations. We address the normalization issue by introducing scaling factors associated with the CMEs of measurements that arise from the parts of the systematic uncertainties that are independent of energy or only very weakly energy-dependent. This contrasts with the SF II approach, in which the errors were inflated by the factor . The procedure for handling discrepant data sets adopted here is mathematically rigorous and aligns naturally with the method of estimating theoretical uncertainty discussed below. We prefer this approach to the inflation factor method used in SF II, since the absolute scales of the different data sets can be used as constraints for each other, while there is no information available for a reevaluation of the CMEs. The procedure we used, described below, can be applied to both frequentist and Bayesian analyses.
Both minimization (a frequentist approach) and Bayesian posterior probability distribution function evaluation were performed to fit EFTs to the data; in the R-matrix analysis, only the former was used. The likelihood function, defined as the conditional probability for the data given theoretical parameters , enters both fitting procedures. It is expressed as [1994NIMPA.346..306D]
with labeling the different experimental data sets and labeling the individual data points of each experiment. In the second term, we divide theoretical predictions for the data points of the experiment by a scaling factor and consider the combination as our full model for , the th data point of experiment with point-to-point uncertainty .
In the first term, we assign a Gaussian distribution centered at with a width equal to the CME ; for the GSI, MSU and RIKEN data shown in Table LABEL:table:normsBayesian, asymmetric Gaussians are used with depending on the sign of . In our minimization, the definition is derived from equating the full exponents in Eq. 25 to . In the absence of CMEs with , Eq. 25 leads to the usual uncorrelated least-squares minimization formula.
In the Bayesian analysis, the terms in the first exponential are considered as part of the prior distribution that are conditioned on prior knowledge and assumptions such as the CMEs in the scaling factors or estimates of theoretical parameters in the EFT. The posterior distribution for the theoretical parameters is then derived via Bayes’s theorem,
where the likelihood function is defined without the first exponential in Eq. (25). However, since the GSI, MSU and RIKEN data have asymmetric CMEs (see Table LABEL:table:normsBayesian), we draw priors from a uniform distribution spanning 50% to 150% of the central values for these that are then used in the asymmetric normalization exponentials in the likelihood function. This was computationally simpler than generating asymmetric Gaussian priors that span the entire range, which is neither necessary nor physical for a normalization constant.
The constant in Eq. (26), known as the Bayesian evidence, guarantees the correct normalization of the posterior distribution. It is useful in comparisons of different theories but does not affect the estimation of parameters.
IX.5 Data Analysis and Determination
We performed fits using several theoretical expressions over multiple energy ranges: , , and over a “non-resonant” energy range and , in which the contribution from the resonance can be ignored. Fits with and without Coulomb dissociation data were performed that resulted in overlapping determinations within the estimated fitting uncertainties. The different low energy fits included one using the halo EFT from 2018PhRvC..98c4616Z and one using the halo EFT⋆ from 2022PhRvC.106a4601H, both including the excited contribution. The simpler halo EFT expressions, not including an excited component , were also fitted in this low energy region with comparable results. For the fits in the other two energy ranges, halo EFT expressions with excited contributions from both 2018PhRvC..98c4616Z and 2022PhRvC.106a4601H were used. The 2018PhRvC..98c4616Z halo EFT expression, which does not include an contribution, was supplemented with a Breit-Wigner resonance [1988ccna.book.....R] with its energy (around the resonance energy ), width, and proton-decay branching ratio determined by the fits. An R-matrix model was fitted to data in the three energy regions with the AZURE2 code [2010PhRvC..81d5805A], again resulting in an consistent with the halo EFT values. The results from the various fits are included in the supplemental material.
Bayesian fits were performed with halo EFT expressions from both 2020JPhG...47e4002Z and 2018EPJA...54...89H that are compatible with the various fits. The Bayesian analysis had advantages over the fits. The capture proceeds through the and channels. The reaction is known to be peripheral, resulting in a that has a sub-leading dependence on the strong interaction in the initial state [2000PhRvC..62f5803B]. This behavior was shown to persist away from the threshold in halo EFT, in which for capture the strong interaction in the incoming -wave channel only contributes at NNLO.
Therefore, to high order, the capture is only sensitive to the sum of the squares of the ANCs in the and channels. This results in an arbitrariness in the relative contributions from the two spin channels in the fits, without affecting the final determination. R-matrix, EFT, and EFT⋆ fits confirm this behaviour. On the other hand, the ANCs have been extracted experimentally [2003PhRvC..67f2801T, 2006PhRvC..73b5808T] and also theoretically [2011PhRvC..83d1001N, 2018PhRvC..98c4616Z] in ab initio calculations. Although the determination of is insensitive to the relative contributions of the two spin channels, knowledge of the relative contributions of the ANCs is important in the EFT framework. It establishes the hierarchy of different contributions in the perturbative expansion. This is crucial in developing a self-consistent expansion that is necessary to quantify the theoretical errors at any finite order of the perturbation. Prior knowledge of the ANCs and other parameters can be accounted for naturally in the Bayesian framework in which one specifies the prior probabilities . Similar constraints can be included in the fit by modifying the first exponential in Eq. (25) to center the theoretical parameters around known values.
The Bayesian fits involved 6 theoretical parameters and 11 normalization constants for the data sets. The priors for the normalization constants were as described earlier. The priors for the theoretical parameters are based on the underlying EFT assumptions [2022PhRvC.106a4601H] about the expansion in a small ratio where represents the low momentum of infrared physics and the cutoff of the theory. The and scattering lengths fm and fm [2019PhRvC..99d5807P] are taken as inputs. For capture at LO, the only fit parameter, the effective momentum is assumed to scale as and accordingly we assume a uniform prior for between -100 MeV/ and 1.5 MeV/, where the upper limit is set by the physical constraint that the ANC squares have to be positive. At NLO, the ANC2 is a fit parameter whose prior is taken from a normal distribution determined by experiment. At NNLO, we use the ab initio calculation of [2018PhRvC..98c4616Z] to draw the ANC2 from a normal distribution. A ratio of the -wave scattering lengths in the coupled-channel - that is assumed to be in the EFT is drawn from a normal distribution with a mean of 0 and a standard deviation of 10 to cover a wide range. The capture is dominated by the transition and it requires three parameters: a -wave scattering volume , an effective momentum and a two-body current to regulate divergences. 2022PhRvC.106a4601H determined and from the the narrow resonance energy and width. Here, we kept fixed while drawing from a uniform distribution between -150 and -50 MeV. The Bayesian fit gave . The two-body coupling was drawn from a uniform distribution and the fits gave a numerical value consistent with the EFT estimate. We obtain the posterior distribution using a probabilistic integration method called Nested Sampling [Skilling06-BA] implemented in Python [2009MNRAS.398.1601F] that calculates both the posterior and the evidence.
The experimental data and EFT⋆ fits at LO, NLO, and NNLO are shown in Fig. 8. and its derivatives at threshold are given in Table 6. The quality of the fits, the consistency of the numerical values of the fit parameters with EFT assumptions, and the order-by-order improvement in the numerical value of the fitted capture cross section and factor give us confidence in the estimated theoretical uncertainty. We find the posterior probability distribution functions to be symmetric and well described by Gaussians, and as such we report the mean and standard deviation in Table 6. Table LABEL:table:normsBayesian gives the scaling factors for the experimental data sets determined from the NNLO fit along with their common mode errors.
| Theory | () | () | () |
[…]
| Data set | Scaling | Data points | CME |
[…]

### v2

@@
 Precisely determined thermonuclear rates of these reactions lie at the foundation of the standard solar model.
-This review, the third decadal evaluation of the nuclear physics of hydrogen-burning stars, is motivated by the great advances made in recent years by solar neutrino observatories, putting experimental knowledge of the proton-proton chain neutrino fluxes in the few-percent precision range.
+This review, the third decadal evaluation of the nuclear physics of hydrogen-burning stars, is motivated by the great advances made in recent years by solar neutrino observatories, putting experimental knowledge of the pp chain neutrino fluxes in the few-percent precision range.
 The basis of the review is a one-week community meeting held in July 2022 in Berkeley, California, and many subsequent digital meetings and exchanges.
-Each of the relevant reactions of solar and quiescent stellar hydrogen burning is reviewed here, from both theoretical and experimental perspectives.
+The relevant reactions of solar and stellar hydrogen burning are reviewed here, from both theoretical and experimental perspectives.
 Recommendations for the state of the art of the astrophysical S-factor and its uncertainty are formulated for each of them.
@@
 Contents
-I Introduction
-The present review summarizes the state of our understanding, in the third decade of the 21st century, of the nuclear reactions and decays taking place inside the Sun and other hydrogen-burning stars.
+Introduction
+The present review summarizes the state of our understanding, in the third decade of the 21st century, of the nuclear reactions and decays taking place in solar and stellar hydrogen burning.
 It also addresses related issues, including solar neutrino detection, radiative opacities, electron screening of nuclear reactions, and the status of the current facilities for measuring cross sections and opacities.
@@
 Ray Davis’s measurements of the flux of solar neutrinos [1968PhRvL..20.1205D, 2003RvMP...75..985D] showed a significant deficit with respect to the predictions of the standard solar model (SSM) [1966PhRvL..17..398B].
-This discrepancy became known as the Solar Neutrino Problem.
-An intense debate ensued about possible explanations for this discrepancy, summarized in the entertaining review by 1988Natur.334..487B.
-Suggestions included plausible flaws in the SSM that might produce a cooler solar core, new particle physics such as neutrino oscillations, and changes in the nuclear physics governing He synthesis, such as an unidentified low-energy resonance in the 3He+3He reaction affecting extrapolations of laboratory cross sections to solar energies.
-By the mid-1990s, with new results from Kamioka II/III and the SAGE/GALLEX experiments indicating a pattern of neutrino fluxes inconsistent with the expected scaling of those fluxes with the solar core temperature, attention increasingly turned to neutrino oscillations and other potential particle physics solutions.
+This discrepancy became known as the "solar neutrino problem." An intense debate ensued about possible explanations for this discrepancy, summarized in the entertaining review by 1988Natur.334..487B.
+Suggestions included plausible flaws in the SSM that might produce a cooler solar core (e.g.
+1981A&A....96....1S; 1996PhRvL..77.4286C), new particle physics such as neutrino oscillations (1969PhLB...28..493G; 1978PhRvD..17.2369W), changes in the nuclear physics governing He synthesis, such as an unidentified low-energy resonance in the 3He+3He reaction affecting extrapolations of laboratory cross sections to solar energies [1972Natur.238...24F, 1972PhLB...40..602F], or the presence of dark matter particles modifying energy transport in the solar core [1985ApJ...294..663S].
+By the mid-1990s, with new results from Kamioka II/III (1989PhRvL..63...16H; 1996PhRvL..77.1683F) and the SAGE/GALLEX (1991PhRvL..67.3332A; 1992PhLB..285..376A) experiments revealed a pattern of neutrino fluxes inconsistent with the expected scaling of those fluxes with the solar core temperature.
+Attention increasingly turned to neutrino oscillations and other potential particle physics solutions.
 Concerns about the nuclear physics of the SSM also evolved, focusing more on the uncertainties that might inhibit extraction of any such “new physics.”
@@
 The INT hosted the proposed workshop in February, 1997, drawing representatives from almost every experimental group active in this area, as well as many of the theorists who were engaged in solar neutrino physics.
-The working groups that formed during this meeting worked over the next year to evaluate past work, determining the needed cross sections and their uncertainties.
-The results of this evaluation, which came to be known as Solar Fusion I or SF-I [1998RvMP...70.1265A], became the standard for use in solar modeling over the next decade.
-An update of SF I was launched with the January, 2009, workshop “Solar Fusion Cross Sections II”, hosted again by the INT in Seattle.
+The working groups that formed during this meeting worked over the following year to evaluate past work, determining the needed cross sections and their uncertainties.
+The results of this evaluation, which came to be known as Solar Fusion I or 1998RvMP...70.1265A [1998RvMP...70.1265A], became the standard for use in solar modeling over the following decade.
+An update of 1998RvMP...70.1265A was launched with the January, 2009, workshop “Solar Fusion Cross Sections II”, hosted again by the INT in Seattle.
 Initial results from SuperKamiokande, the Sudbury Neutrino Observatory, and Borexino were then in hand, and the conversion of approximately two thirds of solar electron neutrinos into other flavors had been firmly established [2002PhRvL..89a1301A].
@@
 While the neutrino fluxes were not known to the precision needed to distinguish between the competing models, it had been shown that the solar core metallicity could be extracted directly from future CN solar neutrino measurements, if the precision of associated nuclear cross sections were improved [2008ApJ...687..678H].
-Reflecting the impact of new experimental work and improved theory, the SF II study of 2011RvMP...83..195A significantly revised the SF I cross section for the driving reaction of the CN (or CNO-I) cycle, and updated key pp-chain cross sections.
-The present SF-III review began with a four-day workshop in July, 2022, “Solar Fusion Cross Sections III.” The meeting was hosted by Physics Frontier Center N3AS (Network for Neutrinos, Nuclear Astrophysics, and Symmetries)111https://n3as.berkeley.edu/, a consortium of institutions involved in multi-messenger astrophysics, whose central hub is the University of California, Berkeley.
+Reflecting the impact of new experimental work and improved theory, the 2011RvMP...83..195A study of 2011RvMP...83..195A significantly revised the 1998RvMP...70.1265A cross section for the driving reaction of the CNO-I cycle, and updated key pp-chain cross sections.
+The present SF III review began with a four-day workshop in July, 2022, “Solar Fusion Cross Sections III.” The meeting was hosted by Physics Frontier Center N3AS (Network for Neutrinos, Nuclear Astrophysics, and Symmetries)111https://n3as.berkeley.edu/, a consortium of institutions involved in multi-messenger astrophysics, whose central hub is the University of California, Berkeley.
 Sessions were held at the David Brower Center, Berkeley, and the Physics Department, UC Berkeley, with the assistance of N3AS staff.
-The workshop was co-hosted by N3AS, by Solar Fusion’s long-term sponsor, the INT, and by the European ChETEC-INFRA Starting Community for Nuclear Astrophysics222https://www.chetec-infra.eu/.
+The workshop was organized by N3AS, by Solar Fusion’s long-term sponsor, the INT, and by the European ChETEC-INFRA Starting Community for Nuclear Astrophysics222https://www.chetec-infra.eu/.
 The 47 workshop participants represented most of the leading experimental and theoretical research groups that are active in the field.
-Workshop participants were organized in nine working groups, who were charged with reviewing and evaluating work completed since SF II, to produce updated recommendations.
+Workshop participants were organized into nine working groups, who were charged with reviewing and evaluating work completed since 2011RvMP...83..195A, to produce updated recommendations.
 The structure of this review largely follows that of the working groups.
-The timing of SF-III was driven in part by planned large-scale asteroseismic surveys of hydrogen-burning stars such as ESA’s PLATO333https://platomission.com/ mission [2014ExA....38..249R], and the expectation that associated data analysis will require large libraries of stellar models in which masses, ages, metallicities, and other parameters are varied.
-SF-III will provide the best current nuclear astrophysics input for such modeling.
-This review summarizes that input, based on the recommendations of the nine SF-III working groups.
+The timing of SF III was driven in part by planned large-scale asteroseismic surveys of hydrogen-burning stars such as ESA’s PLATO333https://platomission.com/ mission [2014ExA....38..249R], and the expectation that associated data analysis will require large libraries of stellar models in which masses, ages, metallicities, and other parameters are varied.
+SF III will provide the best current nuclear astrophysics input for such modeling.
+This review summarizes that input, based on the recommendations of the nine SF III working groups.
 I.2 Terminology used
@@
 where is the target nucleus, is the bombarding particle, and and are the reaction products, with distinguished as the heavier product.
-Frequently in nuclear astrophysics, the reaction cross section is re-expressed in terms of the factor , with related to as described below, and with the mass numbers of entrance channel nuclei and .
-The relation between the nuclear cross section and astrophysical factor is
+Frequently in nuclear astrophysics, the reaction cross section is re-expressed in terms of the S-factor , with S related to as described below, and with the mass numbers of entrance channel nuclei and .
+The relation between the nuclear cross section and astrophysical S-factor is
 with and the charge numbers of the target and projectile, the fine structure constant, the relative velocity of the interacting particles in units of , the reduced mass, and the center-of-mass energy.
-(In this review we use energy units of MeV unless otherwise stated.) This removes the sharp energy dependence from the cross section associated with point Coulomb scattering.
+(In this review we use energy units of MeV unless otherwise stated.) This removes the sharp energy dependence from the cross section associated with s-wave Coulomb scattering off a point nucleus, thereby accounting for the leading penetration effects associated with the Coulomb barrier.
 The quantity is known as the Sommerfeld parameter.
 Consequently, unless there are resonances, varies gently, and at low energies can generally be expressed as a low-order polynomial in .
-The rapid rise in the nuclear cross section with energy, reflecting the weakening of the Coulomb barrier due to the dependence of the Sommerfeld parameter, competes with a rapid decrease in the probability of finding interacting particles with the requisite center-of-mass energy , in the high-energy tail of the Maxwell-Boltzmann distribution for the solar plasma.
+The S-factor in Eq.
+(2) is defined in terms of the nuclear (or “bare") cross section, while laboratory experiments are performed with targets consisting of neutral atoms.
+As electrons shield the nucleus, lowering the Coulomb barrier, their presence enhances the laboratory cross section.
+The S-factors derived from laboratory data must be corrected for this effect.
+As the energy is lowered, the screening correction becomes increasingly important.
+Thus particular care must be taken when extrapolating laboratory cross sections to lower energies, as Fig.
+4 of 2011RvMP...83..195A illustrates.
+Unless otherwise noted, the S-factors discussed here are those for bare nuclei.
+The rapid rise in the nuclear cross section with energy, reflecting the higher probability of penetrating the Coulomb barrier, competes with a rapid decrease in the probability of finding interacting particles with the requisite center-of-mass energy in the high-energy tail of the Maxwell-Boltzmann distribution.
 Consequently, the thermonuclear reaction rate for a given stellar temperature takes the form
@@
 The maximum of the integrand defines the Gamow peak – the most probable energy for interactions – while the range around the peak is frequently termed the Gamow window.
-For the solar reactions of interest, the energy of the Gamow peak is well below the height of the Coulomb barrier.
+For the solar reactions of interest, the energy at the Gamow peak is well below the height of the Coulomb potential.
 The slowly varying can be expanded as a power series around , with the leading term being , and with corrections given by derivatives taken at ,
@@
 For most of the nuclear reactions studied here, the notation is adopted, where and are the mass numbers of the projectile and target nuclei, respectively.
-A summary of SF-III reactions and recommended astrophysical factors are presented in Table 1.
+A summary of SF III reactions and recommended astrophysical S-factors are presented in Table 1.
+Finally, we provide a short list of abbreviations that are frequently used throughout several sections:
+- ANC
+asymptotic normalization coefficient.
+- BBN
+Big Bang nucleosynthesis.
+- CME
+common-mode error.
+- EFT
+effective field theory, and variants: chiral (EFT), pion-less (EFT), hybrid (EFT*), etc.
+- LO, NLO, NNLO
+leading order, next to (next to) leading order, etc.
+- SSM
+Standard Solar Model.
 | Reaction | (MeV b) | (b) | (MeV-1 b) | Section | |
@@
 | – | – | IX | |||
-| – | – | LABEL:subsec:WG3:Total_$S$_factors | |||
+| – | – | LABEL:subsec:WG3:Total_Sfactors | |||
 | LABEL:sec:OtherCNO:CNO1:c12pg | |||||
@@
 I.3 Scope and structure of the review
-In the Sun, the primary mechanism for the conversion of four protons into 4He is the proton-proton chain of Fig.
+In the Sun, the primary mechanism for the conversion of four protons into 4He is the pp chain of Fig.
 1, during which two charge-changing weak interactions take place, each converting a proton into a neutron with the emission of a neutrino.
-The pp-chain consists of three main branches, ppI, ppII, and ppIII.
+The pp chain consists of three main branches, ppI, ppII, and ppIII.
 These branches are distinguished by their dependence on the solar core temperature and by the neutrinos they produce, with the pp, 7Be, and 8B neutrinos serving as “tags” for the ppI, ppII, and ppIII chains, respectively.
-The SSM predicts that the fluxes of pp, 7Be, and 8B neutrinos will have an approximate relative temperature scaling of 1:T:T, where Tc is the solar core temperature [annurev:/content/journals/10.1146/annurev.aa.33.090195.002331].
-By the mid-1990s solar neutrino experiments had established that the measured fluxes differed significantly from this expected pattern.
+The SSM predicts that the fluxes of pp, 7Be, and 8B neutrinos have an approximate relative temperature scaling of 1:T:T, where Tc is the solar core temperature [annurev:/content/journals/10.1146/annurev.aa.33.090195.002331].
+By the mid-1990s, solar neutrino experiments established that the measured fluxes differed significantly from this expected pattern.
 This contributed to growing expectations that new neutrino physics might be the solution of the solar neutrino problem.
-It also was an important motivation for SF I, as errors in the nuclear physics of the pp chains could distort the relationship between the various fluxes.
-The pp chains account for 99% of solar hydrogen burning and are the main focus of this review.
-The remaining 1% is generated through the CN cycle (or CNO-I cycle, see Fig.
+It also was an important motivation for 1998RvMP...70.1265A, as errors in the nuclear physics of the pp chain could distort the relationship between the various fluxes.
+The pp chain accounts for 99% of solar hydrogen burning and are the main focus of this review.
+The remaining 1% is generated through the CNO-I cycle, see Fig.
 2).
@@
 Section II begins with a summary of past solar neutrino experiments, highlighting some of the open questions that remain.
-It discusses the second-generation experiments Super-Kamiokande, the Sudbury Neutrino Observatory, and Borexino that helped resolve the solar neutrino problem, as well as new detectors in various stages of development, including Hyper-Kamiokande, SNO+, JUNO, DUNE, and others.
-The cross section of the driving reaction of the pp chains, , is too small to be measured and thus must be taken from theory.
+It discusses the second-generation experiments Super-Kamiokande, the Sudbury Neutrino Observatory, and Borexino that contributed in the resolution of the solar neutrino problem.
+It also describes Hyper-Kamiokande, SNO+, JUNO, and DUNE, new detectors in various stages of construction.
+The cross section of the driving reaction of the pp chain, , is too small to be measured and thus must be taken from theory.
 This important reaction is reviewed in section III.
-All the other pp-chain reactions and decays have been studied in the laboratory, and used together with theory – which guides fits and extrapolations, and corrects for difference in the atomic environments of solar and terrestrial reactions – to predict solar fusion rates.
+With the exception of the minor hep branch, all other reactions and decays have been studied in laboratories, with measurements then combined with theory to predict rates under solar conditions.
 The 2H(He, 3He(3He,2)4He, and 3HeB reactions are reviewed in Secs.
 IV, V, and VI, respectively.
-The minor hep pp-chain branch is reviewed in Section VII, and the electron-capture reactions on pp and 7Be are discussed in Section VIII.
-The 7Be(,B reaction responsible for the high energy neutrinos measured in Super-Kamiokande is treated in Section IX
-While a modest (%) contributor to solar energy production, the CN (or CNO-I) cycle is a potentially important probe of core metallicity as its hydrogen burning is catalyzed by the Sun’s primordial C and N.
-It is also the primary mode of hydrogen burning in stars more massive than the Sun, with central temperatures above 2K.
-This review treats the CNO cycles in somewhat more detail than in SF I and SF II.
-In SF II the cross section for the leading reaction, 14N()15O, was revised downward by a factor of two from the SF I recommended value.
-The SF-III value is close to that of SF II but with an updated uncertainty evaluation described in Section LABEL:sec:S114.
-A current focus of laboratory astrophysics is the higher-temperature pathways for hydrogen burning that operate in massive stars.
+The hep reaction is reviewed in Section VII, and the electron-capture reactions on pp and 7Be are discussed in Section VIII.
+The 7Be(,B reaction responsible for the high energy neutrinos measured in Super-Kamiokande is treated in Section IX.
+While a modest contributor to solar energy production, the CNO-I cycle is a potentially important probe of core metallicity as its hydrogen burning is catalyzed by the Sun’s primordial C and N (Section LABEL:sec:OtherCNO).
+A current focus of nuclear astrophysics laboratories is the higher-temperature pathways for hydrogen burning that operate in massive stars.
 Many of the contributing reactions have been re-measured.
 That progress is reviewed in Section LABEL:sec:OtherCNO.
-Reactions of the NeNa cycle are also treated in this section.
-Charged-particle nuclear reactions occurring at energies below the Coulomb barrier are affected by screening, the shielding of the nuclear charge by electrons.
+Charged-particle nuclear reactions occurring at energies below the Coulomb barrier are affected by electron screening, the shielding of the nuclear charge by electrons.
 Because of their very different atomic environments, laboratory reactions and those occurring in the solar plasma are affected in distinct ways.
 Section LABEL:sec:Screening describes the current status of efforts to account for these differences, when solar cross sections are extracted from laboratory measurements.
-The energy transport inside the radiative zone is affected by the metals that were incorporated into the Sun when it first formed.
+The energy transport inside the radiative zone is affected by the metals444In this work, we follow the astrophysical terminology where ”metals” refer to all chemical elements except for the two lightest ones, hydrogen and helium.
+that were incorporated into the Sun when it first formed.
 While metals comprise less than 2% of the Sun by mass, they play an outsized role in determining the opacity.
-Very few of the needed radiative opacities have been measured in the laboratory, and the conditions under which measurements are made typically are not identical to those in the Sun.
+Very few of the needed radiative opacities have been measured, and the conditions under which measurements are made typically are not identical to those in the Sun.
 Section LABEL:sec:Opacities describes the current state of the art, and discusses connections between opacities and solar composition that have complicated efforts to resolve the solar metallicity problem.
-Much of the progress made in constraining the reactions of the pp-chain and CNO-I cycles has been enabled by new laboratories and the facilities they house.
+Much of the progress made in constraining the reactions of the pp chain and CNO-I cycles has been possible thanks to new laboratories and the facilities they host.
 The underground LUNA laboratory is a prominent example [2018PrPNP..98...55B].
-Section LABEL:sec:Facilities describes the capabilities of LUNA and other laboratories responsible for the measurements reviewed here.
-II Solar neutrino observations
+Section LABEL:sec:Facilities describes the capabilities of selected laboratories involved in the measurements reviewed here.
+Solar neutrino observations
 Solar neutrinos offer a unique tool box for probing both the fundamental properties of these elusive particles, and their interactions with matter, as well as understanding their source: the fusion reactions that power our Sun.
@@
 These high-precision flux and spectral measurements mapped out the details of solar fusion, probed the structure of the Sun, and demonstrated (together with atmospheric neutrino measurements) that neutrinos have masses and mix.
+Table 2 lists the pp-chain and CNO cycle neutrino sources, their endpoint ( decay) or line (electron capture) energies , and predicted SSM fluxes.
+| Source | Eν (MeV) | GS98 | AAGS21 | MB22p | units (cm-2s-1) |
+| pp | 0.420 | 5.96 (0.6%) | 6.00 (0.6%) | 5.95 (0.6%) | 1010 |
+| pep | 1.442 | 1.43 (1.1%) | 1.45 (1.1%) | 1.42 (1.3%) | 108 |
+| hep | 18.77 | 7.95 (31%) | 8.16 (31%) | 7.93 (30%) | 103 |
+| 7Be | 0.862(a) 0.384(b) | 4.85 (7.4%) | 4.52 (7.3%) | 4.88 (8.1%) | 109 |
+| 8B | 15 | 5.03 (13%) | 4.31 (13%) | 5.07 (15%) | 106 |
+| 13N | 1.198 | 2.80 (16%) | 2.22 (13%) | 3.10 (15%) | 108 |
+| 15O | 1.732 | 2.07 (18%) | 1.58 (16%) | 2.30 (18%) | 108 |
+| 17F | 1.738 | 5.35 (20%) | 3.40 (16%) | 4.70 (17%) | 106 |
+Notes: (a) 90% and (b) 10% of the 7Be neutrino flux, respectively.
 II.1 Open questions
-The majority of solar neutrino fluxes have now been measured, in some cases with a precision equivalent to or better than theoretical predictions.
-Improved precision in several cases can offer further insight.
-A more precise measurement of the flux of neutrinos from the CNO cycle (Fig.
-2) would test important aspect of the SSM including, in particular, the initial abundance of heavy elements [2007A&A...463..755C, 2011ApJ...743...24S, 2022A&A...667L...2K] and the redistribution of the CNO elements from out-of-equilibrium burning at the onset of the main sequence.
-Solar neutrino flux measurements can constrain physical processes in stars such as chemical mixing.
+The solar neutrino spectrum has been measured for energies above 233 keV, with some fluxes determined to a precision equivalent to or better than theoretical predictions.
+However, improved measurements can offer further insight on several aspects.
+For instance, more precise determinations of the fluxes of neutrinos from the CNO cycle (Fig.
+2) could determine the contemporary abundances of N and O in the solar core [2007A&A...463..755C, 2011ApJ...743...24S, 2022A&A...667L...2K].
+While these abundances differ significantly from the primordial values that are input into the SSM –the primordial abundances are altered by out-of-equilibrium CNO-cycle burning that occurs in the early Sun [Roxburgh85]– this burning redistributes the C, N, and O, but does not alter the net number of such nuclei.
+Consequently, CNO neutrino flux measurements might allow one to experimentally address the solar composition problem discussed below.
+Solar neutrino flux measurements can also constrain physical processes in stars such as chemical mixing.
 The fluxes are sensitive to solar/stellar model inputs, including the radiative opacities (see Sect.
 LABEL:sec:Opacities).
-Measurements of the 8B and 7Be fluxes can offer a handle on the temperature of the solar core, and other environmental factors.
-A percent-level measurement of the pp flux would provide a test of the luminosity constraint, further probing solar power generation mechanisms.
-The one branch of the pp-chain that yet remains undetected, the hep flux, has both the highest energy but also the lowest flux of neutrinos from the Sun.
+Measurements of the neutrino fluxes from 8B and 7Be can offer a handle on the temperature of the solar core, and other environmental factors [1996PhRvD..53.4202B].
+A measurement of the pp flux with percent-level precision would provide a test of the luminosity constraint, further probing solar power generation mechanisms [2002PhRvC..65b5801B, 2021JPhG...48a5201V].
+The one branch of the pp chain that yet remains undetected, the hep flux, has both the highest energy but also the lowest flux of neutrinos from the Sun.
 The spectrum extends beyond that of the 8B neutrinos, offering a small window for unequivocal observation.
 Limits have been placed on this flux [2020PhRvD.102f2006A], but a definitive observation will likely require a large, next-generation detector such as Hyper-Kamiokande (HyperK) [2018arXiv180504163H] or the Deep Underground Neutrino Experiment (DUNE) [2019PhRvL.123m1803C].
-Finally, the observation of monochromatic neutrinos produced by electron capture reactions on 13N, 15O and 17F (which we refer to as ecCNO neutrinos) [1990PhRvD..41.2964B, 2004PhRvC..69a5801S, 2015PhLB..742..279V] would provide important new physics.
-They can be used as probes of the metal content of the solar core.
-They provide a measure of the electron neutrino survival probability at specific neutrino energies MeV within the broad transition region between vacuum and matter-enhanced oscillations, where currently we have no constraints.
+Finally, the observation of monochromatic neutrinos produced by electron capture reactions on 13N, 15O and 17F (which we refer to as ecCNO neutrinos) [1990PhRvD..41.2964B, 2004PhRvC..69a5801S, 2015PhLB..742..279V] could be rewarding in terms of physical implications.
+Indeed, they can be used as probes of the metallicity of the solar core.
+Moreover, they provide a measure of the electron neutrino survival probability at specific neutrino energies MeV within the broad transition region between vacuum and matter-enhanced oscillations, where currently we have no constraints.
 This would constitute a new test of the large-mixing-angle (LMA) MSW flavor oscillation paradigm.
 The detection of this subdominant component of the CNO-cycle is extremely difficult but it could be within reach of future very large ultra-pure liquid scintillator detectors [2015PhLB..742..279V].
-There is also opportunity to leverage solar neutrinos to understand the interaction of neutrinos with matter.
+Solar neutrinos provide an opportunity to understand the interaction of neutrinos with matter.
 The effect of MSW oscillations has a significant impact on the observed solar neutrino spectrum: with vacuum oscillation dominating at low energy, below approximately 1 MeV, where the survival probability is roughly one half, while matter effects further suppress the flux to a survival probability of roughly one third above approximately 5 MeV.
-The transition region between these two regimes offers an extremely sensitive probe of the details of the interactions of neutrinos with matter, including the potential to search for new physics such as sterile neutrinos, or non-standard interactions, by looking for distortions to the expected spectral shape.
-The same theory of matter effects predicts a small regeneration of electron neutrinos during the night time, as they propagate through the bulk of the earth.
+The transition region between these two regimes offers an extremely sensitive probe of the details of the interactions of neutrinos with matter, including the potential to search for new physics such as sterile neutrinos, or non-standard interactions, by looking for distortions to the expected spectral shape [2019arXiv190700991B, 2020JHEP...02..038B, 2023JHEP...08..032C].
+The same theory of matter effects predicts a small regeneration of electron neutrinos during the night time as they propagate through the bulk of the earth.
 This so-called “day/night effect” has been sought after by both Super-Kamiokande (SuperK) and the Sudbury Neutrino Observatory (SNO) [2014PhRvL.112i1805R, 2002PhRvL..89a1302A], but a significant observation is still limited by statistics.
-Future data from HyperK or DUNE may be needed to confirm our understanding of this prediction of the MSW effect.
-Mild tension lingers between measurements of the mass splitting parameters, between solar neutrino experiments and terrestrial data sensitive to the same parameter, from the KamLAND reactor experiment [Esteban2020TheFO].
-Further study of solar neutrinos could allow resolution of this question, potentially offering new insights.
+Future data from HyperKamiokande or DUNE may be needed to confirm our understanding of this prediction of the MSW effect.
+The slight tension between measurements of the mass splitting parameters, in solar neutrino experiments and terrestrial data sensitive to the same parameter, from the KamLAND reactor experiment has now disappeared with the addition of new data on day-night effects and spectral shapes [Esteban2020TheFO, 2018PrPNP.102...48C].
+Further tightening of the oscillation parameters can be expected from JUNO, which will obtain highly precise data on both the solar mass splitting and the mixing angle [2016JPhG...43c0401A].
 II.1.1 The solar composition problem
 Solar photospheric abundances, determined with spectroscopic techniques, are a fundamental input for the construction of SSMs, but also they are used as input to nearly all models in astrophysics, including stellar evolution, protoplanetary disks, and galactic chemical evolution.
-The development of three dimensional radiation hydrodynamic models of the solar atmosphere, of techniques to study line formation under non-local thermodynamic conditions and the improvement in atomic properties (e.g.
-transition strengths) have led since 2001 to a complete revision of solar abundances.
-Initial results based on these more sophisticated methods pointed towards a markedly lower solar metallicity [2009ARA&A..47..481A], in particular regarding CNO elements, than results from the 90s based on older techniques [1993oee..conf...15G, 1998SSRv...85..161G].
+Over the last two decades, the development of three dimensional radiation hydrodynamic models of the solar atmosphere [2009MmSAI..80..711L, 2000A&A...359..729A] and of techniques to study line formation under non-local thermodynamic conditions [2005ARA&A..43..481A], together with improved descriptions of atomic properties (e.g.
+transition strengths 2022A&A...661A.140M, 2021MNRAS.508.2236B), have led to a significant revision of solar abundances.
+Initial results based on these more sophisticated methods favored a markedly lower solar metallicity [2009ARA&A..47..481A], particularly in the CNO elements, than obtained in the 90s from older techniques [1993oee..conf...15G, 1998SSRv...85..161G].
 Recently, two groups have revisited the solar photospheric abundances using modern methods [2021A&A...653A.141A, 2022A&A...661A.140M].
@@
 Such a study was recently carried out by the Borexino collaboration [2020Natur.587..577B, 2022PhRvL.129y2701A, 2023PhRvD.108j2005B].
-The Borexino measurement of the CN-cycle neutrino flux (the 13N and 15O fluxes) were used to determine the C+N core abundance.
+The Borexino measurement of the CNO-I cycle neutrino flux (the 13N and 15O fluxes) were used to determine the C+N core abundance.
 Results show a tension with LZ metallicity determinations, while being in better agreement with HZ mixtures.
@@
 The precisely known rate of electron capture on 71Ge to 71Ga, for which the half-life is 11.43(3) d [1985PhRvC..31..666H], places an important constraint on the neutrino absorption rate [2023PhLB..84237983G, PhysRevC.108.035502].
+The half-life has now been remeasured by two groups, 2024PhRvC.109e5501N and PhysRevC.108.L021602, yielding values of 11.468(8) d and 11.46(4) d, consistent with the previous determination while increasing the precision.
 The allowed matrix element that governs neutrino capture to the ground state of 71Ge can be extracted to a precision of 1% from the electron capture rate.
@@
 Knowing it with precision, particularly at the high energy end, has renewed interest in the context of observing the hep neutrinos [2006ApJ...653.1545A, 2020PhRvD.102f2006A], which could be accessible with future observatories [2022arXiv220212839A].
-In SF II it was decided to recommend the spectrum calculated by Winter et al.
+In 2011RvMP...83..195A it was decided to recommend the spectrum calculated by Winter et al.
 [2003PhRvL..91y2501W, 2006PhRvC..73b5503W, Win:2007] based on their measurement of the alpha spectrum from .
@@
 All three find that the peak of the alpha spectrum appears about 20 keV lower than determined by Winter et al.
-Overall, these new measurements yield differences in the neutrino spectrum of 5% below MeV, where the contribution becomes small compared to the hep component.
+Overall, these new measurements yield differences in the neutrino spectrum of 5% below MeV, the energy above which the contribution becomes small compared to the hep component.
 Figure 4 shows the 8B spectral uncertainties on the scale of the hep spectrum.
@@
 A broad range of technology can be used to interrogate solar neutrinos.
-Radiochemical experiments utilizing the interaction on either chlorine or gallium were the first to identify the mystery that became known as the Solar Neutrino Problem.
+Radiochemical experiments utilizing chlorine and gallium played a crucial role in the early days of the solar neutrino problem.
 Large monolithic detectors have achieved great success in real-time measurements – from the water Cherenkov detectors such as Kamiokande, SNO, and SuperK, which could use the unique topology of Cherenkov light to point back to the neutrinos’ origin, to liquid scintillator detectors such as KamLAND and Borexino, whose high light yield and low threshold allow for precision spectroscopy.
 II.2.1 Super-Kamiokande and Hyper-Kamiokande
-The first real-time detection of solar neutrinos was achieved by the Kamiokande experiment [1989PhRvL..63...16H], which detected the neutrinos via elastic scattering (ES) of electrons, a process in which the outgoing electron’s direction is highly correlated to that of the incoming neutrino.
+The first real-time detection of solar neutrinos was achieved by the Kamiokande experiment [1989PhRvL..63...16H], which detected the neutrinos via elastic scattering of electrons, a process in which the outgoing electron’s direction is highly correlated to that of the incoming neutrino.
 This allowed Kamiokande to directly point back to the neutrinos’ origin: our Sun.
 The successor experiment, Super-Kamiokande, is a 50-kton volume of pure water, surrounded by over 10,000 photon detectors, which has achieved an unparalleled program of neutrino observations and other physics over its several decades of operation.
-This program has included the highest-precision measurement of the ES signal from solar neutrino interactions, as well as sensitive searches for the day/night distortions of the spectral shape [2016PhRvD..94e2010A].
+This program has included the highest-precision measurement of the elastic scattering signal from solar neutrino interactions, as well as sensitive searches for the day/night distortions of the spectral shape [2016PhRvD..94e2010A].
 Now filled with a gadolinium additive to enhance neutron capture, SuperK’s primary current focus is the search for the Diffuse Supernova Neutrino Background (DSNB) via inverse beta decay [2023ApJ...951L..27H].
@@
 II.2.2 Sudbury Neutrino Observatory and SNO+
-SNO’s unique use of heavy water as a target medium offered the ability to detect solar neutrinos via two additional interactions besides ES.
-In the charged current (CC) interaction an electron neutrino interacts with the neutron in the deuteron, producing an electron and a proton.
+SNO’s unique use of heavy water as a target medium offered the ability to detect solar neutrinos via two additional interactions besides elastic scattering.
+In the charged current interaction an electron neutrino interacts with the neutron in the deuteron, producing an electron and a proton.
 This interaction path is sensitive only to electron-flavor neutrinos at the few-MeV energy scale of solar neutrinos, providing a measurement of the pure flux.
-SNO also had access to the neutral current (NC) interaction, in which a neutrino of any flavor interacts with the deuteron, breaking it apart into its constituent nucleons.
-Being equally sensitive to all active flavors, the NC interaction yields a measurement of the total, flavor-blind neutrino flux.
-It was this capability that allowed SNO to resolve the Solar Neutrino Problem, and demonstrate that the produced in the Sun were changing flavor prior to detection [2001PhRvL..87g1301A, 2002PhRvL..89a1301A, 2004PhRvL..92r1301A].
+SNO also had access to the neutral current interaction, in which a neutrino of any flavor interacts with the deuteron, breaking it apart into its constituent nucleons.
+Being equally sensitive to all active flavors, the neutral current interaction yields a measurement of the total, flavor-blind neutrino flux.
+It was this capability that allowed SNO to resolve the solar neutrino problem, demonstrating that most of the s produced by the Sun were arriving at earth in a different flavor state [2001PhRvL..87g1301A, 2002PhRvL..89a1301A, 2004PhRvL..92r1301A].
 SNO ceased data taking in 2006.
-Analysis of the data set has continued, and produced a number of new results, including constraints on non-standard effects such as Lorentz violation and neutrino decay [2018PhRvD..98k2013A, 2019PhRvD..99c2013A], searches for hep neutrinos and the diffuse supernova neutrino background (DSNB) [2020PhRvD.102f2006A], and studies of neutron production from cosmogenic muons and from atmospheric neutrino interactions [2019PhRvD..99k2007A, 2019PhRvD.100k2005A].
-After decommissioning, the detector was repurposed for the SNO+ experiment, in which the target material was replaced with a pure organic liquid scintillator (LS): linear alkyl benzene (LAB), loaded with 2.2 g/L of the secondary fluor, PPO.
+Analysis of the data set has continued, and produced a number of new results, including constraints on non-standard effects such as Lorentz violation and neutrino decay [2018PhRvD..98k2013A, 2019PhRvD..99c2013A], limits on the fluxes of hep neutrinos and the diffuse supernova neutrino background (DSNB) [2020PhRvD.102f2006A], and neutron production from cosmogenic muons and from atmospheric neutrino interactions [2019PhRvD..99k2007A, 2019PhRvD.100k2005A].
+After decommissioning, the detector was repurposed for the SNO+ experiment, in which the target material was replaced with a pure organic liquid scintillator: linear alkyl benzene (LAB), loaded with 2.2 g/L of the secondary fluor, PPO.
 The high light yield of this scintillator, coupled with the location in SNOLAB –at 6-km water equivalent, one of the deepest underground laboratories in the world– and an extremely well-understood detector, allow for a range of high-precision measurement programs.
 In a preliminary water phase, SNO+ demonstrated detection of neutron captures on hydrogen [PhysRevC.102.014002], an impressive technical achievement in an unloaded water detector; a low background measurement of 8B solar neutrinos [PhysRevD.99.012012]; several searches for invisible modes of nucleon decay [PhysRevD.99.032008, PhysRevD.105.112012]; and detection of antineutrinos from reactors over 200 km away [PhysRevLett.130.091801].
-These results demonstrate both the low backgrounds and the technical capabilities of the detector, which will facilitate the future program.
-In a “partial-fill” stage, in which the upper half of the detector was filled with LS while the lower half still contained water, the first demonstration of event-level direction reconstruction was achieved for 8B neutrinos in a scintillator detector [PhysRevD.109.072002].
+The low backgrounds achieved and the detector’s demonstrated technical capabilities set the stage for future measurements.
+In a “partial-fill” stage, in which the upper half of the detector was filled with liquid scintillator while the lower half still contained water, the first demonstration of event-level direction reconstruction was achieved for 8B neutrinos in a scintillator detector [PhysRevD.109.072002].
 This was facilitated by a lower loading of PPO at that time, which results in lower scintillation yield and a slower time profile, effectively enhancing the Cherenkov component, which can be leveraged for directional information.
-Now fully filled with LS, SNO+ is the deepest, largest operating LS detector in the world.
+Now fully filled with liquid scintillator, SNO+ is the deepest, largest operating liquid scintillator detector in the world.
 The future program will include measurements of several solar neutrino fluxes, as well as the 8B neutrino energy spectrum.
@@
 To study in real time the bulk of solar neutrino emission a different detector technology was required.
-The Borexino experiment, located deep underground at the INFN Laboratori Nazionali del Gran Sasso, used an organic LS target made of pseudocumene with PPO to detect the ES of solar neutrinos off electrons.
-The high light yield of the LS made it possible to significantly lower the energy threshold, but given the small signal rate and the lack of a clear signature to separate it from the background (such as the direction indicated by the emission of Cherenkov radiation), the experiment required a long preparatory phase to develop the most advanced techniques to suppress the background, especially the one due to radioactive contamination of the LS itself.
+The Borexino experiment, located deep underground at the INFN Laboratori Nazionali del Gran Sasso, used an organic liquid scintillator target made of pseudocumene with PPO to detect the elastic scattering of solar neutrinos off electrons.
+The high light yield of the liquid scintillator made it possible to significantly lower the energy threshold, but given the small signal rate and the lack of a clear signature to separate it from the background (such as the direction indicated by the emission of Cherenkov radiation), the experiment required a long preparatory phase to develop the most advanced techniques to suppress the background, especially the one due to radioactive contamination of the liquid scintillator itself.
 The detector’s extreme radiopurity was key to its success, and over its 14 years of data taking, ending in October 2021, Borexino proved itself capable of covering the entire solar neutrino spectrum.
 At the beginning of its data taking in 2007 the level of 238U and 232Th contamination in Borexino were lower than and respectively [2008PhLB..658..101B], paving the way for the first measurement of the 7Be sub-MeV solar neutrinos, followed by a low-threshold measurement of the 8B flux and a first indication of the pep neutrinos.
-After a purification campaign that further reduced the LS contamination, Borexino aimed at improving the accuracy of its results in the measurement of the pp-chain solar neutrinos.
-The flux of neutrinos produced in the pp fusion process was first measured in [2014Natur.512..383B] and later improved along with all the other fluxes produced throughout the pp-chain, with the only exception of hep neutrinos [2018Natur.562..505B, 2019PhRvD.100h2004A, 2020PhRvD.101f2001A]: the pp flux was determined to an uncertainty of 11%, improving the neutrino-based estimate of solar luminosity, while 7Be neutrinos were measured with an uncertainty of , half that of the SSM prediction.
-Furthermore, applying a constraint on the flux of CNO neutrinos derived from the SSM predictions, the signal of pep neutrinos was established with a significance larger than for the first time, and the 8B flux was measured with a threshold as low as 3 MeV.
-After completing the investigation of the pp-chain neutrinos, Borexino reported the first detection of solar neutrinos produced in the CNO-cycle [2020Natur.587..577B], providing the first direct evidence of the occurrence of that process in stars.
+After a purification campaign that further reduced the liquid scintillator contamination, Borexino aimed at improving the accuracy of its results in the measurement of the pp-chain solar neutrinos.
+The flux of neutrinos produced in the pp fusion process was first reported in 2014Natur.512..383B and later improved, along with all the other fluxes from the pp chain, apart from the hep neutrinos [2018Natur.562..505B, 2019PhRvD.100h2004A, 2020PhRvD.101f2001A].
+The pp flux was determined to an uncertainty of 11%, improving the neutrino-based estimate of solar luminosity, while 7Be neutrinos were measured with an uncertainty of , half that of the SSM prediction.
+Furthermore, using the SSM to constrain the flux of CNO neutrinos, the pep neutrino signal was established with a significance exceeding for the first time, and the 8B flux was measured with a threshold as low as 3 MeV.
+After completing the investigation of the pp-chain neutrinos, Borexino reported the first detection of solar neutrinos produced in the CNO-cycle [2020Natur.587..577B], demonstrating directly that this mode of hydrogen burning operates in stars.
 Profiting from a larger exposure and a better understanding of the radioactive backgrounds enabled by the unprecedented thermal stability of the detector, Borexino further improved its measurement of the CNO neutrino flux [2022PhRvL.129y2701A, 2023PhRvD.108j2005B], where the new CNO result was used in combination with existing solar neutrino results to probe solar composition.
@@
 The newly developed Sanford Underground Research Facility (SURF) offers 4800-m water equivalent overburden, and will form the far site for a long baseline neutrino program utilizing a high-energy neutrino beam directed from Fermilab to SURF.
-The DUNE LArTPC detectors may be sensitive to measurements of the high-energy solar neutrinos via CC interactions on argon, which offers good precision on the spectral shape and, thus, the potential to measure day/night effects [2019PhRvL.123m1803C].
+The DUNE LArTPC detectors may be sensitive to measurements of the high-energy solar neutrinos via charged current interactions on argon, which offers good precision on the spectral shape and, thus, the potential to measure day/night effects [2019PhRvL.123m1803C].
 The solar neutrino program at DUNE is limited primarily by backgrounds (the detector is optimized for GeV-scale physics) and energy threshold.
@@
 A full-scale detector utilizing novel scintillators along with fast and spectrally-sensitive photon detectors, such as the proposed Theia experiment [2020EPJC...80..416A], or the Jinping detector [2017ChPhC..41b3002B] could achieve percent-level precision on the CNO neutrino flux, as well as improving precision across the suite of solar neutrino measurements, such as the 8B spectral shape, and hep neutrino flux.
-As experiments grow larger, and capabilities increase, there are also opportunities to leverage detectors designed for other purposes.
-Noble liquid detectors built primarily for dark matter, or even long baseline neutrino experiments, may offer sensitivity to solar neutrinos.
-Noble liquids have high light yields, and are transparent to their own scintillation.
-Coupled with the high voltage under which such detectors are often operated, the combined ionisation and scintillation signatures can be powerful for discriminating signal from background.
-The low threshold of these detectors may allow detection of neutrinos from the high-energy tail of the 8B spectrum via coherent elastic scattering off a target nucleus.
-Solid state detectors built purposely for observing this “CENS” signature have been used to detect neutrinos from stopped pion decay [doi:10.1126/science.aao0990].
-III The reaction ()
-The cross section for the initial reaction in the pp chain (Figure 1), p+p, is too small to be measured in the laboratory. It must be calculated from the standard theory of weak interactions.
+As experiments grow larger, and capabilities increase, there are also opportunities to apply detectors designed for other purposes to the detection of solar neutrinos.
+Noble liquid detectors built primarily for dark matter searches and long-baseline neutrino experiments have now reached the size and background levels to be sensitive to solar neutrinos [2024arXiv240802877A].
+The detection of neutrinos from the high-energy tail of the 8B spectrum via coherent elastic neutrino-nucleus scattering (“CENS”) is particularly interesting because it is a purely neutral-current process sensitive to the total flux of active neutrinos (above the experimental threshold, about 8 MeV) from the Sun.
+Only the SNO experiment has had such a capability.
+Initial results reported in [2024arXiv240802877A] are in agreement, with limited statistical precision.
+Solid state detectors built purposely for observing CENS have observed neutrinos from stopped pion decay at the theoretically expected rate [doi:10.1126/science.aao0990].
+The 1HH reaction ()
+The cross section for the initial reaction in the pp chain (Figure 1), , is too small to be measured in the laboratory. It must be calculated from the standard theory of weak interactions.
 III.1 Introduction and terminology
-Near the Gamow peak energy 6 keV for temperatures characteristic of the Sun center, the first and second derivatives of the astrophysical factor at zero energy, and , generate 7% and 0.5% corrections, respectively, in Taylor’s series expansion of around .
+Near the Gamow peak energy 6 keV for temperatures characteristic of the solar core, the first and second derivatives of the astrophysical factor at zero energy, and , generate 7% and 0.5% corrections, respectively, in Taylor’s series expansion of around .
 Higher derivative terms are neglected in this review since they only contribute at the level.
@@
 III.2 Adopted parameters for this review
-For the phase-space factor , we use the same value as in SF-II.
+For the phase-space factor , we use the same value as in 2011RvMP...83..195A.
 It comes from the value without radiative corrections, [1969ApJ...155..501B], increased by 1.62% to take into account radiative corrections to the cross section [2003PhRvC..67c5502K].
@@
 Here we urge again that this computation be carried out.
-For , we use the PDG [2022PTEP.2022h3C01W] value, , whose central value is larger than the 2008 PDG value, , used in SF-II by (or 1.9 ).
+For , we use the value from the Particle Data Group (PDG) compilation [2022PTEP.2022h3C01W], , whose central value is larger than the 2008 PDG value, , used in 2011RvMP...83..195A by (or 1.9 ).
 Naively, this would lead to a increase in the central value of according to Eq.
@@
 For , we take s from the most updated comprehensive analysis of experimental rates with the radiative and Coulomb effects corrected [2020PhRvC.102d5501H].
-This value is consistent with s [2009PhRvC..79e5502H] used in SF-II with a larger error.
+This value is consistent with s [2009PhRvC..79e5502H] used in 2011RvMP...83..195A with a larger error.
 The dominant uncertainty in comes from the normalized Gamow-Teller (GT) matrix element .
-Reducing this uncertainty has been the main focus of theoretical work since SF-I.
-In SF-I, was written as , i.e.
+Reducing this uncertainty has been the main focus of theoretical work since 1998RvMP...70.1265A.
+In 1998RvMP...70.1265A, was written as , i.e.
 the sum of the one- and two-body current matrix elements, and , respectively, with their uncertainties estimated independently.
-In SF-II, two major steps had contributed to reducing the uncertainty on .
+In 2011RvMP...83..195A, two major steps had contributed to reducing the uncertainty on .
 The first was a much deeper understanding of the correlation between the uncertainties in and : the overall uncertainty in could be described by a universal parameter that could be fixed by a single measurement.
@@
 This has been done in both potential models [1998PhRvC..58.1263S] and in the hybrid EFT approach [2003PhRvC..67e5206P] as explained below.
-In SF-II, was determined with three approaches.
+In 2011RvMP...83..195A, was determined with three approaches.
 The first one was the potential model approach.
@@
 Constraints on from two-nucleon systems [2002PhLB..549...26B, 2003PhRvC..67b5801C] yielded .
-Based on the consistent results of the above three approaches, SF-II recommended .
-The SF-I value from 1969ApJ...155..501B was not re-computed in SF-II.
-However, with reaching a accuracy, new calculations of and , together with the full pionful chiral EFT (EFT) computations to remove the unknown systematics of the hybrid EFT, were called for in SF-II.
-These two challenges have been met in SF-III and will be described below.
+Based on the consistent results of the above three approaches, 2011RvMP...83..195A recommended .
+The 1998RvMP...70.1265A value from 1969ApJ...155..501B was not re-computed in 2011RvMP...83..195A.
+However, with reaching a accuracy, new calculations of and , together with the full pionful chiral EFT (EFT) computations to remove the unknown systematics of the hybrid EFT, were called for in 2011RvMP...83..195A.
+These two challenges have been met in SF III and will be described below.
 III.3 Experimental progress on muon capture of the deuteron
-In SF-II it was also recommended to carry out the experimental determination of the muon capture rate on deuteron, as proposed in the MuSun experiment [2003nucl.ex...4019K, 2017APS..DNP.CG007S].
+In 2011RvMP...83..195A it was also recommended to carry out the experimental determination of the muon capture rate on deuteron, as proposed in the MuSun experiment [2003nucl.ex...4019K, 2017APS..DNP.CG007S].
 This quantity could be used to constrain without the need to rely on the three-body calculation of .
@@
 III.4 Progress in calculations since SF II
-Below we summarize the calculations performed after SF-II using different approaches.
+Below we summarize the calculations performed after 2011RvMP...83..195A using different approaches.
 III.4.1 EFT
@@
 In the most recent and comprehensive EFT study, 2023arXiv230403327A compared the above calculations of 2019PhRvL.123a9901M and 2016PhLB..760..584A in detail.
-In addition to the increase of from using the correct - relation, updating the input parameters to their most recent values increased by in both calculations, mainly due to the increase of from its SF-II value mentioned above.
+In addition to the increase of from using the correct - relation, updating the input parameters to their most recent values increased by in both calculations, mainly due to the increase of from its 2011RvMP...83..195A value mentioned above.
 Furthermore, 2019PhRvL.123a9901M received a increase from removing the truncation error of the basis functions which effectively cut off the long distance part of the wave functions, as 2017PhRvC..95c1301A advocated.
@@
 Then it can be shown that RG invariance is fulfilled at N2LO.
-However, the estimate for obtained by 2022PhRvC.106e5501L does not use a value for obtained by fitting , but extracted in order to match the value of obtained in SF-II.
+However, the estimate for obtained by 2022PhRvC.106e5501L does not use a value for obtained by fitting , but extracted in order to match the value of obtained in 2011RvMP...83..195A.
 Therefore, we will not consider the work of 2022PhRvC.106e5501L in the present evaluation.
@@
 The above discussions show that determinations of from EFT in Eq.(6), EFT in Eq.(7), and lattice QCD (although with unquantified systematics) are all consistent with each other.
-Furthermore, these values are also consistent with the recommended value of SF-II, provided the central value is increased by 444To know the precise shift requires an explicit calculation.
+Furthermore, these values are also consistent with the recommended value of 2011RvMP...83..195A, provided the central value is increased by 555To know the precise shift requires an explicit calculation.
 However, if the shift is within the range of 0.8-1.0 , the averaged value and remain the same within the significant digits.
@@
 Averaging this value with the EFT value in Eq.(6) and the EFT value in Eq.(7) yields with per degree of freedom to be .
-This shows that the EFT, EFT and SF-II estimates of are all mutually consistent.
+This shows that the EFT, EFT and 2011RvMP...83..195A estimates of are all mutually consistent.
 In addition, we would like to advocate adding an additional correlated error to account for any input that would tend to move all results in a coordinated way.
-For example, from SF-II to SF-III, we experienced the shift due to the update of the PDG value of .
+For example, from 2011RvMP...83..195A to SF III, we experienced the shift due to the update of the PDG value of .
 It is not inconceivable that or other input parameters or physics could change again by similar amounts: the large PDG inflation factor of 2.7 reflects the tension that continues to exist among measurements [2022PTEP.2022h3C01W].
@@
 However, since these values were not obtained with state of the art calculations, they have not been considered in the determination of and .
-IV The reaction ()
+The reaction ()
 IV.1 Introduction
@@
 Compared with the reactions mediated by the weak interaction, this reaction occurs much more rapidly.
-As a result, deuterium is effectively and instantaneously converted to 3He without leaving any observable signature.
-Therefore, uncertainties in the rate of this reaction do not affect solar energy generation.
+Consequently, on the time scales relevant to solar energy generation, deuterium is effectively converted to 3He instantaneously and thus it is only sensitive to the value of the reaction and not the uncertainties of the rate.
 However, the 2H(,)3He reaction plays an important role in the development of protostars, because the onset of deuterium burning slows down the protostars contraction and heating, increasing their lifespan.
@@
 Another astrophysical scenario where the 2H(,)3He reaction plays a key role is Big Bang Nucleosynthesis (BBN), responsible for the production of light elements during the first few minutes of the Universe.
-Among these elements, deuterium is an excellent indicator of cosmological parameters because its primordial abundance is the most sensitive to the baryon density and critically depends on the radiation density of the early Universe.
+Among these elements, deuterium is an excellent indicator of cosmological parameters because its primordial abundance is the most sensitive to the baryon density and critically depends on the radiation density of the early Universe, see for example a review of BBN [2016RvMP...88a5004C].
 The reactions involved in the synthesis of deuterium are: production via the well known )2H process and destruction via the 2H(2H,)3He, 2H(2H,)3H and 2H(,)3He reactions [2020JCAP...03..010F].
-Since the comprehensive review performed in SF II, there have been both new measurements as well as advances in the theoretical and phenomenological analysis of the 2H(,)3He reaction.
-The new experimental results have been determined with accelerator-based measurements of the cross section [2019EPJA...55..137T, 2020Natur.587..210M, 2021PhRvC.103d5805T], as well as plasma-based, intertial confinement fusion measurements [2020PhRvC.101d2802Z, 2022FrP....10.4339M].
-On the theoretical side, there have been advances in ab initio calculations, where nucleons are the fundamental degrees of freedom interacting among themselves and with the external electromagnetic probe.
-Finally, Bayesian analysis methods have been used to model the energy dependence of the factor, starting from the ab initio predictions and applying a polynomial approximation to it.
-We review these updates and provide recommended values and uncertainties of the factor over the energy range of interest for Solar fusion, based upon a Bayesian averaging of various models.
+Since the comprehensive review performed in 2011RvMP...83..195A, there have been both new measurements as well as advances in the theoretical and phenomenological analysis of the 2H(,)3He reaction.
+The new experimental results have been determined with accelerator-based measurements of the cross section [2019EPJA...55..137T, 2020Natur.587..210M, 2021PhRvC.103d5805T], as well as plasma-based, intertial confinement fusion measurements [2020PhRvC.101d2802Z, 2022FrP....10.4339M] and are reviewed in Section IV.2 below.
+On the theoretical side, there have been advances in ab initio calculations, where nucleons are the fundamental degrees of freedom interacting among themselves and with the external electromagnetic probe, see for example [2023FrP....1129094E].
+Finally, Bayesian analysis methods have been used to model the energy dependence of the S-factor, starting from the ab initio predictions and applying a polynomial approximation to it.
+We review these updates and provide recommended values and uncertainties of the S-factor over the energy range of interest for Solar fusion, based upon a Bayesian averaging of various models.
 IV.2 Data sets used for the present review
-The 2H(,)3He reaction has a Q value of 5.5 MeV and proceeds through the direct capture mechanism.
+The 2H(,)3He reaction has a Q value of 5.5 MeV and proceeds through the radiative capture of a proton on deuterium.
 Different experimental approaches were followed to measure its cross section.
-Tisma et al.
-[2019EPJA...55..137T] irradiated deuterated titanium targets with a proton beam and detected the -rays with two high-purity germanium (HPGe) detectors placed at different angles.
-The final factor is provided at four energies in the 47 - 210 keV range, with approximately 15% uncertainty.
-More recently, the LUNA Collaboration performed a measurement underground in the Gran Sasso Laboratories, exploiting the six orders of magnitude suppression of the cosmic radiation background [2021FrASS...7..119Z, 2018IJMPA..3343010C].
+2019EPJA...55..137T irradiated deuterated titanium targets with a proton beam and detected the -rays with two high-purity germanium (HPGe) detectors placed at different angles.
+The final S-factor is provided at four energies in the 47 - 210 keV range, with approximately 15% statistical uncertainty.
+More recently, the LUNA Collaboration performed a measurement in the energy range of keV at the underground in the Gran Sasso Laboratories, exploiting the six orders of magnitude suppression of the cosmic radiation background [2021FrASS...7..119Z, 2018IJMPA..3343010C].
 A windowless deuterium gas target was used and the -rays emitted by the 2H(,)3He reaction were detected by a large HPGe detector at 90∘ with respect to the beam axis.
-Great care was taken to minimize all sources of systematic uncertainties in the factor data, resulting in high precision and accuracy at the 3% level [2020EPJA...56..144M, 2020Natur.587..210M].
+Great care was taken to minimize all sources of systematic uncertainties in the S-factor overall uncertainty at the 3% level [2020EPJA...56..144M, 2020Natur.587..210M].
 These new results provided stringent constraints on cosmological parameters obtained by comparing the precise primordial deuterium abundance predictions of the standard BBN model with astronomical observations [2018ApJ...855..102C].
-A deeper discussion of the LUNA results and their implications can be found in [2021MNRAS.502.2474P, 2021JCAP...04..020P, 2021JCAP...03..046Y, 2021ApJ...923...49M].
-Presently a study of the 2H(,)3He angular distribution is ongoing.
-By studying the shape of the peak produced in the -spectra by the reaction, it is possible to obtain the angular distribution; a dedicated paper containing the results will be published soon.
-Finally, a new measurement was performed at the Helmholtz-Zentrum Dresden-Rossendorf in the 300-1000 keV energy range [2021PhRvC.103d5805T] using implanted deuterium targets on tantalum backings and two HPGe detectors.
-The results are affected by very large systematic uncertainties and they show a different trend of the factor at energies above 300 keV with respect to the fit of the LUNA data of 2020Natur.587..210M.
-To investigate the tension between the LUNA fit and the more recent high-energy data of 2021PhRvC.103d5805T, a new measurement of the 2H(,)3He reaction is planned at the Felsenkeller laboratory in Germany [2019sone.conf..249B].
-In addition to the new accelerator-based results, two recent sets of measurements [2020PhRvC.101d2802Z, 2022FrP....10.4339M] have also been performed using the inertial confinement fusion plasma-based platform [2017PhPl...24d1407G, 2022FrP....10.4339M, 2023FrP....1180821G], which has recently begun to be exploited for this type of work, with initial results on other reactions presented in, e.g., [2016PhRvL.117c5002Z, 2017NatPh..13.1227C].
-Both of these measurements, which were done at the OMEGA laser facility (see section LABEL:subsec:WG8:Plasma), used laser-driven implosions of spherical plastic-shell capsules filled with H2D2 gas, and measured the emitted -rays using a gas-Cherenkov-detector [2014RScI...85kE124H] that was calibrated applying the technique described in 2019RScI...90l3504Z.
+A deeper discussion of the LUNA results and their implications can be found in [2021MNRAS.502.2474P, 2021JCAP...04..020P, 2021JCAP...03..046Y, 2021ApJ...923...49M] 666After the closing of the present review, the 2020Natur.587..210M data were re-analyzed to extract the experimental 2H()3He -ray angular distribution [2024PhRvC.110c2801S].
+The data confirm the ab initio angular distributions used in the original analysis by 2020Natur.587..210M.
+Finally, a new measurement was performed at the Helmholtz-Zentrum Dresden-Rossendorf in the 300-1000 keV energy range using implanted deuterium targets on tantalum backings and two HPGe detectors [2021PhRvC.103d5805T].
+The resulting S-factors show tension with the analysis of the LUNA results extrapolated to keV [2020Natur.587..210M].
+However, they are affected by large systematic uncertainties.
+A new measurement of the reaction is planned at the Felsenkeller laboratory in Germany [2019sone.conf..249B].
+In addition to the new accelerator-based results, two recent sets of measurements [2020PhRvC.101d2802Z, 2022FrP....10.4339M] have also been performed using the inertial confinement fusion plasma-based platform [2017PhPl...24d1407G, 2022FrP....10.4339M, 2023FrP....1180821G], which has recently begun to be exploited for this type of work [e.g., 2016PhRvL.117c5002Z, 2017NatPh..13.1227C].
+Both measurements, which were performed at the OMEGA laser facility (see section LABEL:subsec:WG8:Plasma), used laser-driven implosions of spherical plastic-shell capsules filled with H2D2 gas, and measured the emitted -rays using a gas-Cherenkov-detector [2014RScI...85kE124H] that was calibrated applying the technique described in 2019RScI...90l3504Z.
 The initial experiment [2020PhRvC.101d2802Z] obtained good statistics at an energy of 16 keV by making several repeated measurements, with a final statistical uncertainty of 6% and a systematic one of 17% (dominated by uncertainty in the absolute calibration of the detector).
@@
 The agreement (or disagreement) between ab initio theoretical predictions and experimental data represents a validation (or indicates the necessity of improvement) mostly for ingredient (i), i.e.
-the models of the nuclear interactions and currents, within the accuracy of the ab initio method, i.e., ingredient (ii).
-This is why few-nucleon reactions can be used as an “ideal” laboratory, where the ab initio framework can be stringently tested.
+are the models of the nuclear interactions and currents accurate within the precision specified by the ab initio method?
+This is why few-nucleon reactions can be used as an “ideal” laboratory, where the ab initio framework can be stringently tested in systems under more theoretical control.
 The most recent ab initio calculation of the 2H(,)3He reaction is that of 2016PhRvL.116j2501M.
@@
 IV.4 Phenomenological and Bayesian Analyses
-In the energy range of astrophysical interest, the 2H(,)3He reaction does not have any resonance or coupled channels that can give rise to non-trivial energy dependence, and the factor can be modelled by a low-order polynomial in energy [1967ARA&A...5..525F].
-Moreover, the recent results from LUNA [2020Natur.587..210M], combined with previous measurements, place stringent constraints on the cross section such that a rigorous analysis can capture both the average value and uncertainty over the range of interest to Solar fusion and Big Bang Nucleosynthesis [2020Natur.587..210M, 2021MNRAS.502.2474P, 2021JCAP...04..020P, 2021JCAP...03..046Y].
-Another change since SF-II has been the advent of Bayesian analysis methods to evaluate thermonuclear reaction rates and provide more rigorous uncertainty estimates in a statistical sense [2016ApJ...831..107I].
-To describe this approach, we closely follow the analysis of the 2H(,)3He reaction in 2021ApJ...923...49M, and references therein.
-Consider an optimization that utilizes Gaussian distributed priors in a constrained Bayesian analysis, such that the augmented (for uncorrelated data) is given by
-The first term in the above equation is the standard and the second term derives from the Bayesian constraint on the parameters with prior mean () and width (), which are chosen with some prior knowledge or can be optimized as described below.
-In the first term, the double sum runs over the datasets () and individual results from each dataset (), with the mean value and stochastic uncertainty of each data point given by and , respectively.
-The theoretical model describing the factor data, , is a function of the energy () and a set of parameters () that must be determined.
-The quoted systematic uncertainties are parameterized by the normalization factors , with a prior of unit normalization and a width characterized by the quoted systematic uncertainty.
-Within a Bayesian framework, it is straightforward to utilize distribution functions for that are not Gaussian, such as a log-normal or other distributions.
-The parameters represent a normalization of the model function for a given dataset, which, viewed from a non-Bayesian perspective, can also be interpreted as normalization factors that must be applied to the data to match the “true” underlying distribution.
-Finally, are unknown extrinsic uncertainties [2019PhRvC..99a4619D].
-If one assumes that a smooth function of energy can accurately describe the data, then the data in a given experimental set may scatter about this presumed “true” value by more than is reflected by the quoted statistical uncertainties.
-It was suggested in 2019PhRvC..99a4619D that this extra scatter might be explained by some additional source of statistical uncertainty unbeknownst to the experimenter, or by some unknown systematic uncertainty that is different for each data point in the same data set, as opposed to a correlated systematic that affects all data points similarly.
-In either case, this extrinsic uncertainty can be accommodated in a Bayesian analysis framework by adding the additional uncertainty as a normal-distributed source of noise with a width that is constrained by the data.
-In 2019PhRvC..99a4619D and 2021ApJ...923...49M, was added as an absolute uncertainty, independent of the energy.
-For some data, 2022PhRvC.105a4625O suggested adopting instead a relative uncertainty, such that the scale of the extrinsic fluctuations is proportional to the mean value of .
-This strategy of adding extra extrinsic uncertainty to the data sets can be viewed as an alternative to that of inflating the quoted statistical uncertainties by , which is often used for seemingly incompatible data sets [2022PTEP.2022h3C01W], where is the per degree of freedom.
-An advantage of the extrinsic uncertainty method is that it uses the observed scatter within a given data set as a measure of the possible size of unreported uncertainties, as opposed to uniformly increasing the uncertainty in all data sets by the same relative amount.
-Such a Bayesian analysis of the 2H(,)3He reaction data was performed by 2021ApJ...923...49M, using both a third-order polynomial and the ab initio prediction from 2005PhRvC..72a4001M.
+In the energy range of astrophysical interest, the 2H(,)3He reaction does not have any resonance or coupled channels that can give rise to non-trivial energy dependence, and the S-factor can be modelled by a low-order polynomial in energy [1967ARA&A...5..525F].
+Moreover, the recent results from LUNA [2020Natur.587..210M], combined with previous measurements, place stringent constraints on the cross section, reducing the uncertainty over the range of interest to Solar fusion and BBN [2002NuPhA.706..203C, 2020Natur.587..210M, 2021MNRAS.502.2474P, 2021JCAP...04..020P, 2021JCAP...03..046Y].
+A significant change since 2011RvMP...83..195A has been the wide adoption of Bayesian analysis methods to evaluate thermonuclear reaction rates and provide more rigorous uncertainty estimates in a statistical sense [2016ApJ...831..107I].
+This approach has been followed in the analysis presented here, which closely follow work on the 2H(,)3He reaction reported in 2021ApJ...923...49M, and references therein.
+The mathematical details of the Bayesian method will be found in Appendix LABEL:AppendixWG4: Here we focus on the application to 2H(,)3He by the authors cited above.
+Both the ab initio prediction from 2005PhRvC..72a4001M as well as a third-order polynomial in the energy were used to constrain the data, where we define an -order polynomial as
 As noted in the literature, the energy dependence of 2005PhRvC..72a4001M is more reliable than the absolute normalization.
 Furthermore, the updated prediction in 2016PhRvL.116j2501M, as noted above, is larger than that of 2005PhRvC..72a4001M.
-Therefore, in 2021ApJ...923...49M, the ab initio prediction is modeled as
-where is the prediction from 2005PhRvC..72a4001M (or alternatively from 2016PhRvL.116j2501M), and and are an unknown scale factor and offset to be determined in the analysis.
+Therefore, 2021ApJ...923...49M modeled the prediction of 2005PhRvC..72a4001M, denoted , as
+where and are an unknown scale factor and offset to be determined in the analysis.
 While was not determined with theoretical uncertainty, model uncertainty is introduced through the parameters and .
 The resulting mean values and uncertainties of determined from Eq.
-(11) were found to be comparable.
-For a given model, after optimizing the posterior parameter distributions, the Baye factor (BF) is proportional to the probability of the model given the data [2017bmad.book.....H].
-Therefore, for a fixed data set, the BF can be used as a relative probability of each model, thus enabling a weighted model-averaging procedure.
-If we assume a uniform likelihood for each model, the expectation value and variance of a quantity is given by
-where denotes the expectation of given the model and denotes the probability of the model given the data (), which is given by
-with the Baye factor of model .
-Similarly, for a given model, in the absence of prior information on the size of an unknown parameter, the optimal width of its prior can be estimated by finding the value of that maximizes the BF, which typically provides a reasonable approximation to marginalizing over the prior width.
-| keV | 10 | 20 | 40 | 80 | 91 | 100 | 120 |
-| 2.028(51)(9) | 2.644(60)(8) | 3.276(70)(7) | 4.579(94)(5) | 7.31(15)(0) | 8.11(16)(0) | 8.77(18)(0) | 10.24(21)(0) |
+(11) were found to be comparable to those with the third-order polynomial.
 IV.5 Summary and Recommendations
@@
 | 6th order polynomial | 1464.1 | 2.9 |
-With such a Bayes model averaging, we can compare and contrast the polynomial parameterizations of 2H(,)3He along with ab initio results predicted by 2005PhRvC..72a4001M, 2016PhRvL.116j2501M, including the extra variance that arises from this set of reasonable models, e.g., using polynomials of different order, as well as the phenomenological model of Eq.
+In this review, we take the further step of applying Bayes Model Averaging to a number of reasonable models that describe the data, which systematically captures additional theoretical “model selection uncertainty”.
+Detailes of the Bayesian analysis and model averaging procdure are presented in Appendix LABEL:AppendixWG4.
+The models we explore include Eq.
+(11) with both and , as done in 2021ApJ...923...49M, as well as setting , and then we consider polynomials in energy of order .777The implementation of the Bayesian analysis presented here is different than that of 2021ApJ...923...49M but it was verified to produce the same results when the same implementation of the systematic uncertainties was made.
+With such a Bayes model averaging, we can quantitatively compare and contrast the polynomial parameterizations of 2H(,)3He along with ab initio results predicted by 2005PhRvC..72a4001M, 2016PhRvL.116j2501M, including the extra variance that arises from this set of reasonable models, e.g., using polynomials of different order, as well as the phenomenological model of Eq.
 (11).
@@
 The first column lists the model (i.e., polynomial or ab initio), the second column the natural logarithm of the Gaussian approximation to the BF, and the third column the corresponding weight in the model averaging.
-From this model averaging, the resulting prediction for the 2H(,)3He factor at a few representative energies is provided in Table 2, where the first uncertainty arises from the first term in Eq.
-(IV.4) and the second uncertainty is from the second term, which we denote as model selection uncertainty.
-The resulting Bayes model average prediction of the factor over the entire kinematic range considered is depicted in Fig.
-5, with the gray band representing the 68% coverage probability 555The analysis in Table 3 and Fig.
-5 can be reproduced with the code at the git repository https://github.com/nrp-g/leaner..
+From this model averaging, the resulting prediction for the 2H(,)3He S-factor at a few representative energies is provided in Table 4, where the first uncertainty arises from the first term in Eq.
+(LABEL:eq:BMA_Var) and the second uncertainty is from the second term, which we denote as model selection uncertainty.
+The resulting Bayes model average prediction of the S-factor over the entire kinematic range considered is depicted in Fig.
+5, with the gray band representing the 68% coverage probability.
 For this particular reaction, it is interesting to note that the Bayesian analysis strongly favors the phenomenological models of Eq.
@@
 On the other hand, polynomial approximations are disfavored as they require more parameters to capture the energy dependence.
-Also of note, in the simplest model, using 2016PhRvL.116j2501M, the scale factor is given by
-indicating that the prediction in 2016PhRvL.116j2501M overestimates the 2H(,)3He data by 7.9%, consistent with expectations.
+Also of note, in the model of Eq.
+(11) with and using 2016PhRvL.116j2501M for , the scale factor is given by
+indicating that the prediction in 2016PhRvL.116j2501M overestimates the 2H(,)3He data by 7.9%, consistent with expectations noted above.
 In comparison, the third order polynomial fit predicts values of that are higher at and lower at keV as compared to those in Table 4.
-V The reaction ()
-The 3He(3He,)4He reaction (astrophysical factor ) terminates the pp-I chain.
-The ratio of its rate to that of the 3He()7Be reaction controls the branching to the pp-II and pp-III chains, so that in SF-I increasing was discussed as a potential solution to the Solar Neutrino Problem.
-Subsequent experiments, notably a very low energy measurement at the LUNA 50 kV accelerator deep underground in Gran Sasso [1999PhRvL..82.5205B] and a complementary experiment at somewhat higher energies [2004PhRvC..69a5802K] ruled out such an increase in , as summarized in SF-II.
+The analysis in Table 3 and Fig.
+5 can be reproduced with the code linked in Table 4.
+The code also can provide a prediction of the mean value and uncertainty of at any energy over the same kinematic range.
+| keV | 10 | 20 | 40 | 80 | 91 | 100 | 120 |
+| 2.028(51)(9) | 2.644(60)(8) | 3.276(70)(7) | 4.579(94)(5) | 7.31(15)(0) | 8.11(16)(0) | 8.77(18)(0) | 10.24(21)(0) |
+The reaction )
+The 3He(3He,)4He reaction terminates the pp-I chain.
+The ratio of its rate to that of the 3He()7Be reaction controls the branching to the pp-II and pp-III chains, so that in 1998RvMP...70.1265A increasing was discussed as a potential solution to the Solar Neutrino Problem.
+Subsequent experiments, notably a very low energy measurement at the LUNA 50 kV accelerator deep underground in Gran Sasso [1999PhRvL..82.5205B] and a complementary experiment at somewhat higher energies [2004PhRvC..69a5802K] ruled out such an increase in , as summarized in 2011RvMP...83..195A.
 V.1 Shape of the particle spectrum
-Since no new absolute measurements of have been reported since SF-II, we consider the same four experiments [1999PhRvL..82.5205B, 1998PhRvC..57.2700J, 2004PhRvC..69a5802K, 1987NuPhA.467..273K].
+Since no new absolute measurements of have been reported since 2011RvMP...83..195A, we consider the same four experiments [1999PhRvL..82.5205B, 1998PhRvC..57.2700J, 2004PhRvC..69a5802K, 1987NuPhA.467..273K].
 However, there is new information on the energy spectrum of the emitted protons.
@@
 The results show significant structure, indicating the presence of a sequential reaction mechanism passing through the unbound ground state of .
-This spectrum is important for the measurements [1999PhRvL..82.5205B, 1998PhRvC..57.2700J, 2004PhRvC..69a5802K, 1987NuPhA.467..273K] since they determined cross sections by detecting only the protons above an energy threshold.
+This spectrum is important for all of the measurements since they determined cross sections by detecting only the protons above an energy threshold.
 The efficiency correction by which 1987NuPhA.467..273K account for the threshold is not well documented, but the more recent measurements [1999PhRvL..82.5205B, 1998PhRvC..57.2700J, 2004PhRvC..69a5802K] utilized the genbod event generator [James68-GENBOD].
@@
 V.2 Recommendation
-Based on these investigations, our recommended fit is the one from SF-II, but with an additional 4% systematic uncertainty. This gives
-Only the larger constant term in the uncertainty differs from the previous evaluation.
+Based on these investigations, our recommended fit is the one from 2011RvMP...83..195A, but with an additional 4% systematic uncertainty typical of the 3% to 6% corrections to the extrapolated proton spectra estimated above to take the uncertain spectral shape into account.
+This gives
+Only the larger constant term in the uncertainty differs from the previous evaluation, and it is due to the spectral shape discussed above.
 New measurements of this reaction could provide more accurate absolute cross sections by using lower energy thresholds and improve our understanding of the proton energy spectra and angular correlations.
@@
 We note finally that the proton energy spectrum may depend on , as this has been found to be the case for neutrons from the mirror reaction [2018PhRvL.121d2501G].
-VI The reaction ()
-The 3He()7Be reaction, or , proceeds via direct capture to the ground and 429 keV first excited states of 7Be.
-It has been studied experimentally using three main methods: First, by detecting at least two of the three -rays from the reaction and taking the angular correlation with the alpha beam direction into account (prompt- method).
+The reaction ()
+The 3He()7Be reaction proceeds via non-resonant capture to the ground and 429 keV first excited states of 7Be.
+It has been studied experimentally using three main methods, with references given in subsections VI.1 and VI.3: First, by detecting at least two of the three -rays from the reaction and taking the angular correlation with the alpha beam direction into account (prompt- method).
 Second, by detecting the induced 7Be [2002NuPhA.708....3T] radioactivity (activation method).
@@
 VI.1 Previous recommendation in SF II
-The 3He()7Be factor recommended in SF II was developed in multiple steps.
+The 3He()7Be S-factor recommended in 2011RvMP...83..195A was developed in multiple steps.
 First, a model was selected for the shape of , based on existing nucleon-level calculations [1986NuPhA.460..559K, 2001PhRvC..63e4002N].
 A rescaling of the curve was fitted to data, but only at 1.002 MeV center-of-mass to minimize the role of short-distance physics in the models.
-In SF I, a possible systematic discrepancy between data from the two methods previously used was discussed.
-By the time of SF II, in-depth studies from two groups using both activation and prompt methods were available, from LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G] and Seattle [2007PhRvC..76e5801B] groups.
+In 1998RvMP...70.1265A, a possible systematic discrepancy between data from the two methods previously used was discussed.
+By the time of 2011RvMP...83..195A, in-depth studies from two groups using both activation and prompt methods were available, from LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G] and Seattle [2007PhRvC..76e5801B] groups.
 These studies did not find any discrepancy between activation and prompt- data in direct comparisons.
-SF II opted to limit the fitting to the data by the activation and recoil methods.
+2011RvMP...83..195A opted to limit the fitting to the data by the activation and recoil methods.
 The prompt- data were left out because of their somewhat larger common mode errors, concerns about how well the -ray angular distribution was known, and avoidance of the correlated errors between activation and prompt data from the same experiment.
-In SF II, only data published after 1998 were included.
-This was partly because the newer data are in general better documented than the older data, and partly because the newer experiments were perceived as addressing systematics more quantitatively.
-Data from each of the retained experiments were used to determine a separate rescaling of the theory curve, and the rescalings were averaged to find a recommended and recommended error.
-Theory error was estimated by dispersion of the extrapolated among several models, including versions of the main fitting model that were modified to cover the experimentally-allowed range of scattering lengths.
-The SF II recommended value was , based on the data by the Weizmann [2004PhRvL..93z2503S], LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G], Seattle [2007PhRvC..76e5801B], and ERNA [2009PhRvL.102w2502D] groups.
+In 2011RvMP...83..195A, only data published after 1998 were included.
+The 2011RvMP...83..195A recommended value was , based on the data by the Weizmann [2004PhRvL..93z2503S], LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G], Seattle [2007PhRvC..76e5801B], and ERNA [2009PhRvL.102w2502D] groups.
 VI.2 Theory progress on
-Significant theoretical work on this reaction has occurred since SF II, but the basic understanding of its mechanism remains unchanged from the 1960s [1963PhRv..131.2582T].
-It is dominated by external direct capture into the two bound states, and most of the dipole strength at low energy arises beyond the range of nuclear interaction.
+Significant theoretical work on this reaction has occurred since 2011RvMP...83..195A, but the basic understanding of its mechanism remains unchanged from the 1960s [1963PhRv..131.2582T].
+It is dominated by external non-resonant capture into the two bound states, and most of the dipole strength at low energy arises beyond the range of nuclear interaction.
 The external capture part of the cross section is determined by the ANCs of the bound states and by scattering phase shifts; near threshold, most of the strength lies at –20 fm [2011PhRvL.106d2502N].
@@
 All models feature a shallow minimum of near 1.25 MeV, where capture from -waves becomes comparable to that from -waves.
-The first fully ab initio calculation [2011PhRvL.106d2502N] appeared just after the SF II analysis concluded; it used the fermionic molecular dynamics (FMD) method and a softened representation of the Argonne interaction [1995PhRvC..51...38W].
-It agrees well with both the scale and the energy dependence of the modern factor data, and it is very close to the energy dependence assumed in SF II.
+The first fully ab initio calculation [2011PhRvL.106d2502N] appeared just after the 2011RvMP...83..195A analysis concluded; it used the fermionic molecular dynamics (FMD) method and a softened representation of the Argonne interaction [1995PhRvC..51...38W].
+It agrees well with both the scale and the energy dependence of the modern S-factor data, and it is very close to the energy dependence assumed in 2011RvMP...83..195A.
 Neff’s elastic-scattering phase shifts also agree well with experiment.
@@
 The results agree well with the overall scale of the modern data, but their energy dependence and phase shifts depart from experiment, possibly due to the omission of explicit three-body forces.
-An important advance since SF II has been the application of halo effective field theory (halo EFT) methods to astrophysical capture reactions [2002NuPhA.712...37B, 2018EPJA...54...89H, 2020EPJA...56..166P, 2020JPhG...47e4002Z].
+An important advance since 2011RvMP...83..195A has been the application of halo effective field theory (halo EFT) methods to astrophysical capture reactions [2002NuPhA.712...37B, 2018EPJA...54...89H, 2020EPJA...56..166P, 2020JPhG...47e4002Z].
 EFTs are valid for systems with a natural separation between the momentum scales probed in low-energy experiment (e.g., corresponding to a binding energy or a projectile energy) and the much larger momentum scale where the low-energy degrees of freedom are no longer valid: the word “halo” here refers to nuclei with only shallow bound states.
@@
 The work of 2018EPJA...54...89H and 2020EPJA...56..166P includes a careful examination of possible power-counting schemes and strict adherence to a power counting once established, up to terms of NLO or NNLO, respectively, in two different power-countings.
-These authors fitted EFTs to factor data, and they examined effects of including or excluding elastic scattering constraints in their fits.
+These authors fitted EFTs to S-factor data, and they examined effects of including or excluding elastic scattering constraints in their fits.
 When they included scattering phase shifts from 1972NuPhA.195..241B in their fits, they concluded that the large-scattering-length power counting of the NLO theory was favored.
-This result promotes two-body currents (i.e., contributions not equivalent to external direct capture) to leading order in the theory.
+This result promotes two-body currents (i.e., contributions not equivalent to external non-resonant capture) to leading order in the theory.
 The -wave contribution that becomes important above 1 MeV first appears at NLO.
@@
 The additional terms proved not to be required by the data and did not improve the fit; this result was taken to indicate that corrections from omitted terms are not large compared with experimental errors below 2 MeV.
-Scattering data were not considered in this work apart from very broad priors on scattering length and effective range; the correlated errors in the factor data were taken into account.
+Scattering data were not considered in this work apart from very broad priors on scattering length and effective range; the correlated errors in the S-factor data were taken into account.
 Like halo EFT, the phenomenological R-matrix approach avoids a model of nuclear interactions and uses a systematic parameterization to fit data [2010RPPh...73c6301D].
-At the time of SF II, had been the subject of very little R-matrix fitting, apparently consisting only of the simple treatment in 2004ADNDT..88..203D that focused on BBN energies.
-After SF-II an R-matrix analysis using the AZURE2 code [2010PhRvC..81d5805A, azure2-manual] was carried out on both elastic scattering and factor data in conjunction with the Notre Dame experiment [2013PhRvC..87f5804K], and a more elaborate analysis using Monte Carlo sampling to estimate errors was later reported in 2014PhRvC..90c5804D.
+At the time of 2011RvMP...83..195A, had been the subject of very little R-matrix fitting, apparently consisting only of the simple treatment in 2004ADNDT..88..203D that focused on BBN energies.
+After 2011RvMP...83..195A an R-matrix analysis using the AZURE2 code [2010PhRvC..81d5805A, azure2-manual] was carried out on both elastic scattering and S-factor data in conjunction with the Notre Dame experiment [2013PhRvC..87f5804K], and a more elaborate analysis using Monte Carlo sampling to estimate errors was later reported in 2014PhRvC..90c5804D.
 In the latter analysis both the fitted value and error bars were heavily influenced by the numerous scattering data of 1964NucPh..50..629B.
@@
 In this work it was shown that markedly different values result from inclusion or not of the older scattering data [1964NucPh..50..629B] alongside the very recent SONIK scattering data of 2024PhRvC.109a5802P.
-A small amount of additional theoretical work in more traditional frameworks has appeared since SF-II.
-This includes fits of potential models to factor data [2018PhRvC..97c5802T, 2021NuPhA100622108T] and some calculations of the RGM type [2017PhRvC..96f4605S, 2019PhRvC..99e4618S].
+A small amount of additional theoretical work in more traditional frameworks has appeared since 2011RvMP...83..195A.
+This includes fits of potential models to S-factor data [2018PhRvC..97c5802T, 2021NuPhA100622108T] and some calculations of the RGM type [2017PhRvC..96f4605S, 2019PhRvC..99e4618S].
 VI.3 Experimental progress on
-Since SF II, five new experiments have been reported: Four from the Madrid and ATOMKI groups by the activation technique, at relatively high center of mass energies [2012PhRvC..86c2801C, 2013NuPhA.908....1B, 2019PhRvC..99e5804S, 2023PhRvC.108b5802T], and one from the Notre Dame group using the prompt- method [2013PhRvC..87f5804K].
-Following the approach adopted in SF II, we again only use the recoil and activation data for the fits below.
+Since 2011RvMP...83..195A, five new experiments have been reported: Four from the Madrid and ATOMKI groups by the activation technique, at relatively high center of mass energies [2012PhRvC..86c2801C, 2013NuPhA.908....1B, 2019PhRvC..99e5804S, 2023PhRvC.108b5802T], and one from the Notre Dame group using the prompt- method [2013PhRvC..87f5804K].
+Following the approach adopted in 2011RvMP...83..195A, we again only use the recoil and activation data for the fits below.
 The -ray angular distribution is not known experimentally [2007PhRvC..76e5801B], and the resultant uncertainty increases the common-mode error for the prompt- studies somewhat.
-(A related experiment has recently concluded at Felsenkeller Dresden.
-The data suggest a higher than expected anisotropy but are so far only available in the form of a PhD thesis [Turkat23-PhD].) In addition, all of the “modern” works reporting prompt- data except for Notre Dame [2013PhRvC..87f5804K] also include data obtained with other methods, leading to partial correlations between data sets that would complicate fitting.
+(An experiment to address previously-raised concerns about angular distribution [2007PhRvC..76e5801B] has recently concluded at Felsenkeller Dresden.
+The data suggest a higher than expected anisotropy but are so far only available in the form of a PhD thesis [Turkat23-PhD].
+If confirmed, they may lead to corrections of a few percent for some of the in-beam experiments where only one angle was instrumented.) In addition, all of the “modern” works reporting prompt- data except for Notre Dame [2013PhRvC..87f5804K] also include data obtained with other methods, leading to partial correlations between data sets that would complicate fitting.
 Following this restriction, our data selection proceeds as follows.
-First, the four data sets previously used in SF II are carried over here: Weizmann [2004PhRvL..93z2503S], LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G] (only the activation data), Seattle [2007PhRvC..76e5801B] (only the activation data), and ERNA [2009PhRvL.102w2502D].
-Two of the archival data sets excluded by the 1998 cutoff date in SF II merit some further discussion here: The data of 1982PhRvL..48.1664O, 1984NuPhA.419..115O consist of two points measured by activation in a 3He gas cell.
-However, their uncertainties are not separated into statistical and systematic components, making them unsuitable for the data model used below.
+First, the four data sets previously used in 2011RvMP...83..195A are carried over here: Weizmann [2004PhRvL..93z2503S], LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G] (only the activation data), Seattle [2007PhRvC..76e5801B] (only the activation data), and ERNA [2009PhRvL.102w2502D].
+Two of the archival data sets excluded by the 1998 cutoff date in 2011RvMP...83..195A merit some further discussion here: The data of 1982PhRvL..48.1664O, 1984NuPhA.419..115O consist of two points measured by activation in a 3He gas cell.
+However, their uncertainties are not separated into statistical and systematic components.
+Including them would have required an uncertain guess on how to divide the error bars, and the fit result would barely change, so we left them out.
 The work by 1983PhRvC..27...11R consists of one data point that was obtained by averaging two separate activation measurements with a 3He and a 4He gas cell, respectively.
-We do not use this result because we lack details to verify the background subtraction.
+We do not use this result because we lack details to verify the background subtraction, an issue that was raised during our meeting by one of its authors.
 We also choose not to use the activation study by 1983ZPhyA.310...91V.
@@
 This gave an integrated measurement over a wide energy range, so that the analysis depends strongly on the assumed shape of the curve, and the result was reported only as an extrapolated .
-We now consider the new data since SF II.
+We now consider the new data since 2011RvMP...83..195A.
 An activation experiment in Madrid [2012PhRvC..86c2801C] used a 3He beam incident on a 4He gas cell, reporting three data points.
@@
 In order to test the effects of this decision, we repeated some of our fits (see below, Section VI.4) with the modern prompt- experiments included: the 17 points from 2013PhRvC..87f5804K, 3 prompt- points from LUNA [2007PhRvC..75f5803C] and 8 from Seattle [2007PhRvC..76e5801B].
-The extrapolated -factor changed by less than 1%, well within the error bars for the recommended .
-Finally, an indirect experiment using the 6Li(3He,)7Be reaction and the ANC technique has recently been reported [2020PhLB..80735606K].
-It is left out of our fits due to the additional normalization and theory uncertainties involved in determining an ANC from a transfer experiment, which are larger than the errors in the factor measurements used for the present fits.
+The extrapolated S-factor changed by less than 1%, well within the error bars for the recommended .
+Finally, an indirect experiment using the 6Li(3He,d)7Be reaction and the ANC technique has recently been reported [2020PhLB..80735606K].
+It is left out of our fits due to the additional normalization and theory uncertainties involved in determining an ANC from a transfer experiment, which are larger than the errors in the S-factor measurements used for the present fits.
 All the data used in the present fits are summarized in Table 5.
-As in SF-II, we model data uncertainties as consisting of a component that is independent for each point and a common-mode component that applies to all data from a given experiment as a multiplicative factor.
+As in 2011RvMP...83..195A, we model data uncertainties as consisting of a component that is independent for each point and a common-mode component that applies to all data from a given experiment as a multiplicative factor.
 This separation is well-documented for all of the modern data and is also shown in Table 5.
@@
 The relatively large number of model parameters and the significant fitting degeneracies between some of them make Bayesian analysis a natural choice for finding best fit and confidence intervals for the multi-parameter models applied here.
-Two of the groups did not produce frequentist fits in which they had confidence.
+Two of the groups ran into serious difficulty with the frequentist fits due to parameter degeneracies and shallow local minima; during this study they were unable to produce frequentist fits in which they had confidence.
 We performed multiple fits in the NLO halo EFT of 2018EPJA...54...89H and 2020EPJA...56..166P.
@@
 These fits gave keV b with the scattering constraint and keV b without (uncertainties being propagated in linear approximation using covariances and partial derivatives).
-Formally the EFT truncation error from stopping at NLO corresponds to an additional theoretical error that can shift factors by 10%.
+Formally the EFT truncation error from stopping at NLO corresponds to an additional theoretical error that can shift S-factors by 10%.
 However, any fitted curve is constrained by low-energy data, which in some sense become effectively renormalization conditions of the field theory; the error on extrapolated should probably be smaller than 10% by an amount that is hard to estimate.
@@
 | BRICK R-matrix+scattering | 0.531 | -0.47 | 22.6(7) | |
-| ANC [2020PhLB..80735606K] | 34 |
+| ANC [2020PhLB..80735606K] |
 Several of the fits are compared in Fig.
@@
 Because the fit to the NLO EFT of 2018EPJA...54...89H gives the central result among those attempted here, we adopt it for our recommended .
-For ease of adoption, it is noted that this recommended curve can be empirically parameterized using the same shape as in SF II, by the numerical equation (following a customary form in the past literature)
+For ease of adoption, it is noted that this recommended curve can be empirically parameterized using the same shape as in 2011RvMP...83..195A, by the numerical equation (following a customary form in the past literature [1987ApJ...319..531K, 2011RvMP...83..195A])
 Here is the center of mass energy in MeV. Equation (VI.5) reproduces our recommended curve within 0.3% and is only applicable for = 0–1600 keV.
@@
 However, the uncertainty due to theoretical approximation is more complicated, and error estimation on extrapolated quantities in EFTs remains an open area of research; estimates from the groups working on are discussed above.
-We estimate the overall theoretical uncertainty on our recommended factor to be the measured by the 4% dispersion among fitted curves; this is comparable to the more formal estimates.
-Our recommendations for the zero energy astrophysical factor and the low-energy slope are then
+We estimate the overall theoretical uncertainty on our recommended S-factor to be the measured by the 4% dispersion among fitted curves; this is comparable to the more formal estimates.
+Our recommendations for the zero energy astrophysical S-factor and the low-energy slope are then
 The given value of applies over = 0-500 keV.
@@
 The two errors can be added in quadrature for applications.
-Our recommendation is essentially the same as in SF II, but derived from completely different theoretical approaches that are much more directly connected to the data and much less dependent on model assumptions.
+Our recommendation is essentially the same as in 2011RvMP...83..195A, but derived from completely different theoretical approaches that are much more directly connected to the data and much less dependent on model assumptions.
 It is also based on data over a wider energy range.
@@
 These experiments should also be designed to provide new angle-corrected branching ratios for capture into the two bound states.
-Our third recommendation is a study of the astrophysical factor over a wide energy range, 50-1500 keV and extending to even lower energies when technically feasible, with small point-to-point errors.
+Our third recommendation is a study of the astrophysical S-factor over a wide energy range, 50-1500 keV and extending to even lower energies when technically feasible, with small point-to-point errors.
 Even with a large common-mode error, this would significantly constrain theoretical curve shapes near threshold, and providing an independent check on the influence of scattering data on .
@@
 Given the difficulties of constructing a unique phase shift analysis without an underlying theory, and the recent arrival of modern scattering data, theorists should consider including elastic differential cross sections directly as outputs and fitting constraints.
-VII The reaction ()
+The hep reaction ()
 The reaction , also known as the hep reaction, is one of the possible processes taking place after the , and it produces the most energetic neutrinos, with an endpoint energy of 18.8 MeV.
-It is however the least probable, with its factor, and consequently its rate, even beyond the reach of current experiments.
+It is however the least probable, with its S-factor, and consequently its rate, even beyond the reach of current experiments.
 This is due to the fact that the hep reaction is induced by the weak interaction and further suppressed by Coulomb barrier and by the fact that the leading one-body axial (Gamow-Teller) current operator cannot connect the main -state components of the and initial- and final-state wave functions.
@@
 This is further complicated by cancellations between the one-body and two-body axial current contributions.
-Due to this large suppression of the contributions that are nominally expected to be the largest, the zero-energy factor, , is unreachable by experiment and only theory can provide an estimate.
-The most recent studies of this reaction, already evaluated in SF-II, are those of 2000PhRvC..63a5801M and 2003PhRvC..67e5206P.
+As a consequence, similarly to the pp fusion, the hep reaction cross section is too small to be measured in laboratory and only theoretical predictions are available.
+The most recent studies of this reaction, already evaluated in 2011RvMP...83..195A, are those of 2000PhRvC..63a5801M and 2003PhRvC..67e5206P.
 The first one was performed within a phenomenological approach, similar to the calculation of 1998PhRvC..58.1263S for the pp fusion.
@@
 The initial and final wave functions were derived using the CHH method with the AV18/UIX potential, as in 2000PhRvC..63a5801M.
-Following the results of 2000PhRvC..63a5801M and 2003PhRvC..67e5206P, SF-II recommended for the value keV b.
+Following the results of 2000PhRvC..63a5801M and 2003PhRvC..67e5206P, 2011RvMP...83..195A recommended for the zero-energy S-factor, , the value keV b.
 Three recommendations were also made: (i) to perform new studies with a broad spectrum of Hamiltonian models, in order to properly access the theoretical uncertainty; (ii) to study other weak reactions for which experiments can test the predictions made by employing the same theoretical ingredients as those used for the hep reaction calculation; (iii) to further understand the relation between the hep reaction and the so-called hen process ().
 Of these recommended works, only the second one has been performed so far, applying the theoretical framework used for the hep study to the muon capture reactions on light nuclei [2011PhRvC..83a4002M].
-Given the lack of further studies on the hep reaction after SF-II, we decide to maintain
+Given the lack of further studies on the hep reaction after 2011RvMP...83..195A, we decide to maintain
 as recommended value.
@@
 However, we have decided not to apply such an increase, because it does not derive from a direct calculation, and, most of all, because such an increase lies well within the quoted % uncertainty.
-However, given the advances in ab initio studies of the last decade and in the development of the EFT framework, we update the recommendations of SF-II as follows: (i) the hep process should be revisited within a fully consistent EFT approach, similarly to what has been done in these years for the pp fusion and other processes, as, for instance, muon captures on light nuclei [2012PhRvL.108e2502M, 2018PhRvL.121d9901M, 2018PhRvC..98f5506A, 2023FrP....1049919C, 2023PhRvC.107f5502B, PhysRevC.109.035502]; (ii) the initial and final wave functions should be calculated with the most accurate ab initio methods, as the (uncorrelated) HH method [2008JPhG...35f3101K, 2020FrP.....8...69M], which has been proven to provide more accurate bound and scattering wave functions than the CHH method, using both local and non-local interactions; (iii) the relation between the hep and the hen process should be understood.
+However, given the advances in ab initio studies of the last decade and in the development of the EFT framework, we update the recommendations of 2011RvMP...83..195A as follows: (i) the hep process should be revisited within a fully consistent EFT approach, similarly to what has been done in these years for the pp fusion and other processes, as, for instance, muon captures on light nuclei [2012PhRvL.108e2502M, 2018PhRvL.121d9901M, 2018PhRvC..98f5506A, 2023FrP....1049919C, 2023PhRvC.107f5502B, PhysRevC.109.035502]; (ii) the initial and final wave functions should be calculated with the most accurate ab initio methods, as the (uncorrelated) HH method [2008JPhG...35f3101K, 2020FrP.....8...69M], which has been proven to provide more accurate bound and scattering wave functions than the CHH method, using both local and non-local interactions; (iii) the relation between the hep and the hen process should be understood.
 The hen process has been studied within EFT and the HH method by 2022PhRvC.105a4001V, but a consistent parallel study of both reactions is still missing.
@@
 Because these corrections are unusually large for hep neutrinos, future studies should provide spectra and evaluate the impact of such corrections on the shape.
-VIII Electron capture by and
+Electron capture by p+p and
 Electron capture reactions are the sources of lines found in the solar neutrino spectrum.
-In particular, electron capture on , i.e., the process , also known as the pep reaction, competes with fusion and depends on the same nuclear matrix element.
+In particular, electron capture on , i.e., the process d, also known as the pep reaction, competes with fusion and depends on the same nuclear matrix element.
 Therefore, the ratio between the pp and pep rates is independent of nuclear physics.
-Based on this consideration, in SF-II the result of SF-I, which is based on the work of 1969ApJ...155..501B, was multiplied by the radiative corrections calculated by 2003PhRvC..67c5502K, leading to the final result given by
+Based on this consideration, in 2011RvMP...83..195A the result of 1998RvMP...70.1265A, which is based on the work of 1969ApJ...155..501B, was multiplied by the radiative corrections calculated by 2003PhRvC..67c5502K, leading to the final result given by
 where is the density in units of g/cm3, is the mean molecular weight per free electron, and is the temperature in units of MK.
 The range of validity is .
-Given that no new evaluation of the pep rate has been performed since SF-II, the result of Eq.
+Given that no new evaluation of the pep rate has been performed since 2011RvMP...83..195A, the result of Eq.
 (20) represents also the present recommended value.
@@
 2020PhRvL.125c2701F recently measured the ratio of to capture using superconducting tunnel junctions.
-The recommendation of both SF I and SF II was
+The recommendation of both 1998RvMP...70.1265A and 2011RvMP...83..195A was
 which is valid for .
@@
 That reference presented five distinct derivations demonstrating that the Salpeter formula for screening corrections is valid for pp-chain reactions and solar conditions, up to corrections on the order of a few percent.
-IX The reaction ()
+The reaction ()
 IX.1 Introduction
-Radiative proton capture on (=3/2-) at solar energies proceeds through direct capture to the ground state of (=2+).
+Radiative proton capture on (=3/2-) at solar energies proceeds through non-resonant capture to the ground state of (=2+).
 This capture occurs predominantly at separations well beyond the range of the strong interaction via transitions from and partial waves, as the amplitude and contributions from higher partial waves are negligible in the energy range of interest [1961NucPh..24...89C].
-The importance of this reaction for the determination of the high energy solar neutrino spectrum and the experimental data reviewed by SF I inspired several experiments that, taken together, provided a consistent picture of the energy dependence of the reaction cross section, despite discrepancies in absolute scale.
-Prior to the review of SF II, most direct measurements were performed using a radioactive target and an intense proton beam.
-In spite of the great care taken in evaluating the common mode errors (CMEs), a remarkable discrepancy persisted even in the most recent experiments, that was finally handled in SF II by inflating the stated experimental uncertainties.
-Indirect measurements of the cross section of based on Coulomb dissociation were considered, but not included in the final recommendation of SF II for the zero-energy astrophysical factor , whose uncertainty was dominated by the theoretical contribution.
+The importance of this reaction for the determination of the high energy solar neutrino spectrum and the experimental data reviewed by 1998RvMP...70.1265A inspired several experiments that, taken together, provided a consistent picture of the energy dependence of the reaction cross section, despite discrepancies in absolute scale.
+Prior to the review of 2011RvMP...83..195A, most direct measurements were performed using a radioactive target and an intense proton beam.
+In spite of the great care taken in evaluating the systematic energy-independent uncertainties, in the following referred to as common-mode errors (CMEs), a remarkable discrepancy persisted even in the most recent experiments, that was finally handled in 2011RvMP...83..195A by inflating the stated experimental uncertainties.
+Indirect measurements of the cross section of B based on Coulomb dissociation were considered, but not included in the final recommendation of 2011RvMP...83..195A for the zero-energy astrophysical S-factor .
+However, despite the experimental issues, the uncertainty on the recommended value of in 2011RvMP...83..195A is dominated by the theoretical contribution.
 Therefore, our analysis focused on these aspects (discrepancy of experimental data, indirect measurements, theoretical uncertainties), as discussed in the next sections.
@@
 IX.2.1 Direct Measurements
-The thorough review of published work done in SF II investigated the influence of beam-target overlap, target stoichiometry, beam energy loss, and the backscattering of 8B recoils on CMEs and led to the selection of a homogeneous group of well-documented data sets.
+The thorough review of published work done in 2011RvMP...83..195A investigated the influence of beam-target overlap, target stoichiometry, beam energy loss, and the backscattering of 8B recoils on CMEs and led to the selection of a homogeneous group of well-documented data sets.
 Since then, no further information became available on these aspects of direct measurements.
-The data that formed the basis of the SF II recommendation were those of 2003PhRvC..67f5805B, 2003PhRvL..90b2501B, 1998PhRvL..80..928H, 2001PhRvL..86.3985H, 1999PhLB..462..237H, 2001NuPhA.696..219S, and the BE3 data set of 2003PhRvC..68f5803J.
-Here we adopt the revised BE3 data presented by 2010PhRvC..81a2801J.
-In addition to the complete set of radiative capture measurements that were the basis of the SF II recommendation for , we consider the only new experiment reported thereafter of 2022PhLB..82436819B, which used a radioactive ion beam and the recoil mass separator ERNA to detect the 8B recoils.
-While these new results contribute to the determination of with different, well controlled systematics [2018EPJA...54...92B], the relatively large statistical uncertainty of the measurement limits its impact.
+The data that formed the basis of the 2011RvMP...83..195A recommendation were those of 2003PhRvC..67f5805B, 2003PhRvL..90b2501B, 1998PhRvL..80..928H, 2001PhRvL..86.3985H, 1999PhLB..462..237H, 2001NuPhA.696..219S, and the BE3 data set of 2003PhRvC..68f5803J.
+We considered the complete set of radiative capture measurements that were the basis of the 2011RvMP...83..195A recommendation for , where we adopted the revised BE3 data presented by 2010PhRvC..81a2801J, which were reanalyzed using a more sophisticated model of the target and more accurate stopping powers.
+In addition, we included in our analysis the only new experiment reported thereafter of 2022PhLB..82436819B, which used a radioactive ion beam and the recoil mass separator ERNA to detect the 8B recoils.
+While these new results contribute to the determination of with different, well controlled systematics [2018EPJA...54...92B], their large statistical uncertainty relative to the other experiments, due to the low counting statistics, limits their impact.
 New scattering data have been published.
@@
 At higher energies between 1.6 and 3.4 MeV, thick target elastic and inelastic scattering excitation function measurements were used to infer the existence of new resonances above 1.8 MeV [2013PhRvC..87e4617M].
-These high lying resonances have no direct influence on .
+These high lying resonances have no direct influence on because of their small partial widths in the channel and/or because they cannot be populated by s- or d-wave capture at low energy.
 IX.2.2 Indirect Measurements
-In SF II, Coulomb dissociation measurements based on the formalism developed by 1986NuPhA.458..188B were not included in the determination of due to disagreements over whether the analyses of the various measurements properly accounted for the contributions of transitions.
-Despite these reservations, the Coulomb breakup data were found to be in agreement with radiative capture measurements.
-In order to include these consistent data in formulating our recommendation for , we inflated the common mode errors of the Coulomb breakup measurements to adequately account for the uncertainty in the components assumed in each analysis.
+In 2011RvMP...83..195A, Coulomb dissociation measurements based on the formalism developed by 1986NuPhA.458..188B were considered [PhysRevLett.83.2910, PhysRevLett.73.2680, KIKUCHI1997261, 1998EPJA....3..213K, PhysRevLett.86.2750, PhysRevC.63.065806, PhysRevLett.90.232501, 2006PhRvC..73a5806S], where in some cases the most recent publications supersede the previously published ones.
+Finally, the results of these experiments were not included in the determination of due to disagreements over whether the analyses of the various measurements properly accounted for the contributions of transitions.
+We disagree with these reservations and find the agreement of these measurements with radiative capture measurements significant.
+However, in order to include these consistent data in formulating our recommendation for , we inflated the CMEs of the Coulomb breakup measurements to adequately account for the uncertainty in the E2 components assumed in each analysis.
 The common mode errors are obtained from a linear sum of those given in the manuscripts and an additional common mode error due to the estimated size of the component in each measurement.
@@
 Though the 2003PhRvC..68d5802D analysis did include a common mode error contribution of 2.5% due to the component (which represented 5% of the total cross section), we have included an additional 2.5% contribution to the MSU experiment.
-The size of the additional common mode errors for the RIKEN and GSI measurements was based on scaling this 2.5% according to the predicted size of the contribution to the cross section in these measurements relative to that in the MSU measurement, evaluated using first-order perturbation theory and the and factors calculated in a potential model of 8B [1996ZPhyA.356..293B].
+The size of the additional common mode errors for the RIKEN and GSI measurements was based on scaling this 2.5% according to the predicted size of the contribution to the cross section in these measurements relative to that in the MSU measurement, evaluated using first-order perturbation theory and the and S-factors calculated in a potential model of 8B [1996ZPhyA.356..293B].
 Thus the additional single-sided common mode errors added to the GSI, MSU, and RIKEN measurements are -2.1%, 2.5%, and -4.3%, respectively.
 IX.3 Theory
-The size of the theoretical contribution to the uncertainty in the SF II estimate of was twice that of the experimental contribution.
+The size of the theoretical contribution to the uncertainty in the 2011RvMP...83..195A estimate of was twice that of the experimental contribution.
 Since then, significant progress has been made.
@@
 The large-scale computational demands of calculations using high-quality nucleon-nucleon (NN) and three-nucleon (3N) interactions beyond =5 had hindered the ab initio approach in the past.
-Nevertheless progress has been rapid and it has been possible to extend ab initio calculations to , albeit without 3N forces.
+Nevertheless progress has been rapid and it has been possible to extend ab initio calculations to B, albeit without 3N forces.
 2011PhLB..704..379N performed such a calculation using the no-core shell model in the continuum with a nucleon-nucleon interaction derived from chiral EFT at fourth order.
-The calculation converged with respect to the model space size and reproduces experimental data for the factor despite having virtually no free parameters (though the similarity renormalization group evolution parameter was tuned to reproduce the binding energy of 8B).
-The final value of =19.4(7) eV b agrees with the SF II recommendation and, as will be shown later, with our analysis.
+The calculation converged with respect to the model space size and reproduces experimental data for the S-factor despite having virtually no free parameters (though the similarity renormalization group evolution parameter was tuned to reproduce the binding energy of 8B).
+The final value of =19.4(7) eV b agrees with the 2011RvMP...83..195A recommendation and, as will be shown later, with our analysis.
 More recently, a new approach was introduced to combine experimental measurements with ab initio predictions, resulting in an “ab initio-informed evaluation” Kravvaris23-PLB that arrived at a value of eV b.
@@
 Thus there is no effect on the cross section.
-In the incoming channel, 2022PhRvC.106a4601H use a more general form of the -wave short-ranged interaction involving the core based on the low-energy symmetry that is included only at next-to-next-to-leading order (NNLO).
+In the incoming channel, 2022PhRvC.106a4601H use a more general form of the -wave short-ranged interaction involving the core based on the low-energy symmetry that is included only at NNLO.
 In contrast, 2015PhLB..751..535Z and 2018PhRvC..98c4616Z include a short-ranged interaction at LO that could make a difference in the momentum dependence of the cross section.
@@
 The third important difference is in the treatment of two-body currents, which affects EFT error estimates.
-2015PhLB..751..535Z and 2018PhRvC..98c4616Z include this at next-to-leading order (NLO) whereas 2022PhRvC.106a4601H estimate it to be an N3LO effect.
+2015PhLB..751..535Z and 2018PhRvC..98c4616Z include this at NLO whereas 2022PhRvC.106a4601H estimate it to be an N3LO effect.
 Bayesian estimates of the two-body current [2015PhLB..751..535Z, 2018PhRvC..98c4616Z] are not strongly constrained by capture data, consistent with this being a higher-order contribution.
@@
 The parameters of the real pole corresponding to the bound state are determined by the binding energy and the two ANCs, while the 1+ resonance parameters are determined by its energy and partial decay widths (both proton and radiative).
-The R-matrix analysis of this reaction carried out by 1995NuPhA.588..693B arrived at a value of eV b; however, it was based on data that were excluded in SF II and here, with the exception of those of 1983PhRvC..28.2222F.
+The R-matrix analysis of this reaction carried out by 1995NuPhA.588..693B arrived at a value of eV b; however, it was based on data that were excluded in 2011RvMP...83..195A and here, with the exception of those of 1983PhRvC..28.2222F.
 We fit data using R-matrix theory and halo EFTs, with both frequentist and Bayesian approaches, investigating the impact of different choices for the energy range in which data were fitted, and obtained consistent results.
@@
 IX.4 Fitting Procedures
-In SF II, it was noted that the different data sets had similar energy dependences.
+In 2011RvMP...83..195A, it was noted that the different data sets had similar energy dependences.
 The discrepancies among the data sets were primarily due to different absolute normalizations.
-We address the normalization issue by introducing scaling factors associated with the CMEs of measurements that arise from the parts of the systematic uncertainties that are independent of energy or only very weakly energy-dependent.
-This contrasts with the SF II approach, in which the errors were inflated by the factor .
+We address the normalization issue by introducing scaling factors associated with the CMEs.
+This contrasts with the 2011RvMP...83..195A approach, in which the errors were inflated by the factor .
 The procedure for handling discrepant data sets adopted here is mathematically rigorous and aligns naturally with the method of estimating theoretical uncertainty discussed below.
-We prefer this approach to the inflation factor method used in SF II, since the absolute scales of the different data sets can be used as constraints for each other, while there is no information available for a reevaluation of the CMEs.
+We prefer this approach to the inflation factor method used in 2011RvMP...83..195A, since the absolute scales of the different data sets can be used as constraints for each other, while there is no information available for a reevaluation of the CMEs.
 The procedure we used, described below, can be applied to both frequentist and Bayesian analyses.
-Both minimization (a frequentist approach) and Bayesian posterior probability distribution function evaluation were performed to fit EFTs to the data; in the R-matrix analysis, only the former was used.
+Although our final recommendation is based on Bayesian analysis alone, both minimization (a frequentist approach) and Bayesian posterior probability distribution function evaluation were performed to fit EFTs to the data; in the R-matrix analysis, only the former was used.
 The likelihood function, defined as the conditional probability for the data given theoretical parameters , enters both fitting procedures.
-It is expressed as [1994NIMPA.346..306D]
+It can be derived from the expression given in 1994NIMPA.346..306D
 with labeling the different experimental data sets and labeling the individual data points of each experiment.
@@
 Bayesian fits were performed with halo EFT expressions from both 2020JPhG...47e4002Z and 2018EPJA...54...89H that are compatible with the various fits.
-The Bayesian analysis had advantages over the fits.
+The Bayesian analysis had advantages over the fits in exploring parameter space and in determining realistic uncertainties [PhysRevLett.122.232502].
 The capture proceeds through the and channels.
@@
 (22) to center the theoretical parameters around known values.
-The Bayesian fits involved 6 theoretical parameters and 11 normalization constants for the data sets.
+The Bayesian fits involved six theoretical parameters and 11 normalization constants for the data sets.
 The priors for the normalization constants were as described earlier.
@@
 and its derivatives at threshold are given in Table 7.
-The quality of the fits, the consistency of the numerical values of the fit parameters with EFT assumptions, and the order-by-order improvement in the numerical value of the fitted capture cross section and factor give us confidence in the estimated theoretical uncertainty.
+The quality of the fits, the consistency of the numerical values of the fit parameters with EFT assumptions, and the order-by-order improvement in the numerical value of the fitted capture cross section and S-factor give us confidence in the estimated theoretical uncertainty.
 We find the posterior probability distribution functions to be symmetric and well described by Gaussians, and as such we report the mean and standard deviation in Table 7.

### v3

@@
-
+SFIII_Supplement.pdf
 Abstract
@@
 Contents
+- I Introduction
+- II Solar neutrino observations
+- III The 1HH reaction ()
+- IV The reaction ()
+- V The reaction )
+- VI The reaction ()
+- VII The hep reaction ()
+- VIII Electron capture by p+p and
+- IX The reaction ()
+- X The reaction (S1 14)
+- XI Other CNO, Ne-Na reactions
+- XII Electron screening of nuclear reactions
+- XIII Radiative opacities
+- XIV Experimental facilities for solar fusion studies
+- XV Closing remarks
+- A Gallium Neutrino Source Cross Sections
+- B Bayesian Methods
 Introduction
@@
 It also addresses related issues, including solar neutrino detection, radiative opacities, electron screening of nuclear reactions, and the status of the current facilities for measuring cross sections and opacities.
-As was the case for two previous decadal reviews [1998RvMP...70.1265A, 2011RvMP...83..195A], the present review summarizes the progress made over the past decade in advancing the nuclear physics of main-sequence stars and makes recommendations for future work.
+As was the case for two previous decadal reviews [11, 12], the present review summarizes the progress made over the past decade in advancing the nuclear physics of main-sequence stars and makes recommendations for future work.
 I.1 Purpose
-Ray Davis’s measurements of the flux of solar neutrinos [1968PhRvL..20.1205D, 2003RvMP...75..985D] showed a significant deficit with respect to the predictions of the standard solar model (SSM) [1966PhRvL..17..398B].
-This discrepancy became known as the "solar neutrino problem." An intense debate ensued about possible explanations for this discrepancy, summarized in the entertaining review by 1988Natur.334..487B.
+Ray Davis’s measurements of the flux of solar neutrinos [225, 224] showed a significant deficit with respect to the predictions of the standard solar model (SSM) [67].
+This discrepancy became known as the "solar neutrino problem." An intense debate ensued about possible explanations for this discrepancy, summarized in the entertaining review by Bahcall et al.
+[75].
 Suggestions included plausible flaws in the SSM that might produce a cooler solar core (e.g.
-1981A&A....96....1S; 1996PhRvL..77.4286C), new particle physics such as neutrino oscillations (1969PhLB...28..493G; 1978PhRvD..17.2369W), changes in the nuclear physics governing He synthesis, such as an unidentified low-energy resonance in the 3He+3He reaction affecting extrapolations of laboratory cross sections to solar energies [1972Natur.238...24F, 1972PhLB...40..602F], or the presence of dark matter particles modifying energy transport in the solar core [1985ApJ...294..663S].
-By the mid-1990s, with new results from Kamioka II/III (1989PhRvL..63...16H; 1996PhRvL..77.1683F) and the SAGE/GALLEX (1991PhRvL..67.3332A; 1992PhLB..285..376A) experiments revealed a pattern of neutrino fluxes inconsistent with the expected scaling of those fluxes with the solar core temperature.
+Schatzman et al.
+573; Cumming and Haxton 207), new particle physics such as neutrino oscillations (Gribov and Pontecorvo 313; Wolfenstein 662), changes in the nuclear physics governing He synthesis, such as an unidentified low-energy resonance in the 3He+3He reaction affecting extrapolations of laboratory cross sections to solar energies [275, 268], or the presence of dark matter particles modifying energy transport in the solar core [602].
+By the mid-1990s, with new results from Kamioka II and III (Hirata et al.
+353; Fukuda et al.
+286) and the SAGE/GALLEX (Abazov et al.
+1; Anselmann et al.
+47) experiments revealed a pattern of neutrino fluxes inconsistent with the expected scaling of those fluxes with the solar core temperature.
 Attention increasingly turned to neutrino oscillations and other potential particle physics solutions.
@@
 The working groups that formed during this meeting worked over the following year to evaluate past work, determining the needed cross sections and their uncertainties.
-The results of this evaluation, which came to be known as Solar Fusion I or 1998RvMP...70.1265A [1998RvMP...70.1265A], became the standard for use in solar modeling over the following decade.
-An update of 1998RvMP...70.1265A was launched with the January, 2009, workshop “Solar Fusion Cross Sections II”, hosted again by the INT in Seattle.
-Initial results from SuperKamiokande, the Sudbury Neutrino Observatory, and Borexino were then in hand, and the conversion of approximately two thirds of solar electron neutrinos into other flavors had been firmly established [2002PhRvL..89a1301A].
+The results of this evaluation, which came to be known as Solar Fusion I or SF I [11], became the standard for use in solar modeling over the following decade.
+An update of SF I was launched with the January, 2009, workshop “Solar Fusion Cross Sections II”, hosted again by the INT in Seattle.
+Initial results from SuperKamiokande, the Sudbury Neutrino Observatory, and Borexino were then in hand, and the conversion of approximately two thirds of solar electron neutrinos into other flavors had been firmly established [23].
 Consequently, the motivation for the study was to ensure that solar model predictions would be based on the most current nuclear physics, so that meaningful uncertainties could be placed on neutrino parameters derived from solar neutrino measurements, such as the mixing angle .
-An additional source of uncertainty had arisen at the time, with the advent of three-dimensional (3D) radiative-hydrodynamic models of the Sun’s atmosphere for determining element abundances from photo-absorption lines [2006NuPhA.777....1A, 2009ARA&A..47..481A].
-This improved analysis, though, lowered the inferred metallicities, leading to tension with the Sun’s interior sound speed profile determined from helioseismology [2002RvMP...74.1073C].
+An additional source of uncertainty had arisen at the time, with the advent of three-dimensional (3D) radiative-hydrodynamic models of the Sun’s atmosphere for determining element abundances from photo-absorption lines [57, 59].
+This improved analysis, though, lowered the inferred metallicities, leading to tension with the Sun’s interior sound speed profile determined from helioseismology [193].
 The discrepancy between higher metallicity models that accurately reproduce the Sun’s interior sound speed and lower metallicity models based on the most current treatment of the photosphere was termed the “solar composition problem”.
-While the neutrino fluxes were not known to the precision needed to distinguish between the competing models, it had been shown that the solar core metallicity could be extracted directly from future CN solar neutrino measurements, if the precision of associated nuclear cross sections were improved [2008ApJ...687..678H].
-Reflecting the impact of new experimental work and improved theory, the 2011RvMP...83..195A study of 2011RvMP...83..195A significantly revised the 1998RvMP...70.1265A cross section for the driving reaction of the CNO-I cycle, and updated key pp-chain cross sections.
+While the neutrino fluxes were not known to the precision needed to distinguish between the competing models, it had been shown that the solar core metallicity could be extracted directly from future carbon-nitrogen (CN) solar neutrino measurements, if the precision of associated nuclear cross sections were improved [340].
+Reflecting the impact of new experimental work and improved theory, the SF II study of Adelberger et al.
+[12] significantly revised the SF I cross section for the driving reaction of the CNO-I cycle, and updated key pp-chain cross sections.
 The present SF III review began with a four-day workshop in July, 2022, “Solar Fusion Cross Sections III.” The meeting was hosted by Physics Frontier Center N3AS (Network for Neutrinos, Nuclear Astrophysics, and Symmetries)111https://n3as.berkeley.edu/, a consortium of institutions involved in multi-messenger astrophysics, whose central hub is the University of California, Berkeley.
@@
 The 47 workshop participants represented most of the leading experimental and theoretical research groups that are active in the field.
-Workshop participants were organized into nine working groups, who were charged with reviewing and evaluating work completed since 2011RvMP...83..195A, to produce updated recommendations.
+Workshop participants were organized into nine working groups, who were charged with reviewing and evaluating work completed since SF II, to produce updated recommendations.
 The structure of this review largely follows that of the working groups.
-The timing of SF III was driven in part by planned large-scale asteroseismic surveys of hydrogen-burning stars such as ESA’s PLATO333https://platomission.com/ mission [2014ExA....38..249R], and the expectation that associated data analysis will require large libraries of stellar models in which masses, ages, metallicities, and other parameters are varied.
+The timing of SF III was driven in part by planned large-scale asteroseismic surveys of hydrogen-burning stars such as ESA’s PLATO333https://platomission.com/ mission [543], and the expectation that associated data analysis will require large libraries of stellar models in which masses, ages, metallicities, and other parameters are varied.
 SF III will provide the best current nuclear astrophysics input for such modeling.
@@
 I.2 Terminology used
-The terminology used here follows recent practice, as summarized for example in the NACRE compilation [1999NuPhA.656....3A]. It is briefly described to aid the reader.
+The terminology used here follows recent practice, as summarized for example in the Nuclear Astrophysics Compilation of Reaction Rates (NACRE) compilation [44].
+It is briefly described to aid the reader.
 Nuclear reactions are designated as
@@
 The quantity is known as the Sommerfeld parameter.
-Consequently, unless there are resonances, varies gently, and at low energies can generally be expressed as a low-order polynomial in .
+Consequently, unless there are resonances, should be a smooth function that at low energies can generally be expressed as a low-order polynomial in .
 The S-factor in Eq.
@@
 Thus particular care must be taken when extrapolating laboratory cross sections to lower energies, as Fig.
-4 of 2011RvMP...83..195A illustrates.
+4 of SF II illustrates.
 Unless otherwise noted, the S-factors discussed here are those for bare nuclei.
@@
 A summary of SF III reactions and recommended astrophysical S-factors are presented in Table 1.
-Finally, we provide a short list of abbreviations that are frequently used throughout several sections:
-- ANC
-asymptotic normalization coefficient.
-- BBN
-Big Bang nucleosynthesis.
-- CME
-common-mode error.
-- EFT
-effective field theory, and variants: chiral (EFT), pion-less (EFT), hybrid (EFT*), etc.
-- LO, NLO, NNLO
-leading order, next to (next to) leading order, etc.
-- SSM
-Standard Solar Model.
+Finally, we provide a short list of abbreviations that are used throughout the review in the List of Symbols and Abbreviations.
 | Reaction | (MeV b) | (b) | (MeV-1 b) | Section | |
@@
 | V | |||||
-| – | VI | ||||
-| – | – | VII | |||
-| – | – | IX | |||
-| – | – | LABEL:subsec:WG3:Total_Sfactors | |||
-| LABEL:sec:OtherCNO:CNO1:c12pg | |||||
-| LABEL:sec:OtherCNO:CNO1:c13pg | |||||
-| LABEL:sec:OtherCNO:CNO2:n15pg | |||||
-| LABEL:sec:OtherCNO:CNO1:n15pa | |||||
-| LABEL:sec:OtherCNO:CNO2:o16pg | |||||
-| – | – | LABEL:sec:OtherCNO:CNO2:o17pg | |||
-| – | – | LABEL:sec:OtherCNO:CNO3:o18pg | |||
-| – | – | LABEL:sec:OtherCNO:NeNa:ne20pg | |||
-| – | – | LABEL:sec:21nepg | |||
-| – | – | LABEL:sec:OtherCNO:NeNa:ne22pg | |||
-| LABEL:sec:OtherCNO:NeNa:na23pg |
+| VI | |||||
+| VII | |||||
+| IX | |||||
+| X.6 | |||||
+| XI.1.1 | |||||
+| XI.1.2 | |||||
+| XI.2.1 | |||||
+| XI.1.3 | |||||
+| XI.2.2 | |||||
+| XI.2.3 | |||||
+| XI.3.1 | |||||
+| XI.4.1 | |||||
+| XI.4.2 | |||||
+| XI.4.3 | |||||
+| XI.4.4 |
 I.3 Scope and structure of the review
@@
 These branches are distinguished by their dependence on the solar core temperature and by the neutrinos they produce, with the pp, 7Be, and 8B neutrinos serving as “tags” for the ppI, ppII, and ppIII chains, respectively.
-The SSM predicts that the fluxes of pp, 7Be, and 8B neutrinos have an approximate relative temperature scaling of 1:T:T, where Tc is the solar core temperature [annurev:/content/journals/10.1146/annurev.aa.33.090195.002331].
+The SSM predicts that the fluxes of pp, 7Be, and 8B neutrinos have an approximate relative temperature scaling of 1:T:T, where Tc is the solar core temperature [339].
 By the mid-1990s, solar neutrino experiments established that the measured fluxes differed significantly from this expected pattern.
 This contributed to growing expectations that new neutrino physics might be the solution of the solar neutrino problem.
-It also was an important motivation for 1998RvMP...70.1265A, as errors in the nuclear physics of the pp chain could distort the relationship between the various fluxes.
+It also was an important motivation for SF I, as errors in the nuclear physics of the pp chain could distort the relationship between the various fluxes.
 The pp chain accounts for 99% of solar hydrogen burning and are the main focus of this review.
@@
 It discusses the second-generation experiments Super-Kamiokande, the Sudbury Neutrino Observatory, and Borexino that contributed in the resolution of the solar neutrino problem.
-It also describes Hyper-Kamiokande, SNO+, JUNO, and DUNE, new detectors in various stages of construction.
+It also describes Hyper-Kamiokande, SNO+, the Jiangmen Underground Neutrino Observatory (JUNO), and the Deep Underground Neutrino Experiment (DUNE), new detectors in various stages of construction.
 The cross section of the driving reaction of the pp chain, , is too small to be measured and thus must be taken from theory.
@@
 The 7Be(,B reaction responsible for the high energy neutrinos measured in Super-Kamiokande is treated in Section IX.
-While a modest contributor to solar energy production, the CNO-I cycle is a potentially important probe of core metallicity as its hydrogen burning is catalyzed by the Sun’s primordial C and N (Section LABEL:sec:OtherCNO).
-A current focus of nuclear astrophysics laboratories is the higher-temperature pathways for hydrogen burning that operate in massive stars.
+The CNO-I cycle is a modest contributor to solar energy generation, accounting for about 1% of the sun’s hydrogen burning.
+It is a potentially important probe of core metallicity as the burning is catalyzed by the Sun’s primordial C and N (Section XI).
+Higher temperature pathways for hydrogen burning are of great importance in more massive stars.
+The associated nuclear reactions are an increasingly important focus of low-energy nuclear astrophysics facilities.
 Many of the contributing reactions have been re-measured.
-That progress is reviewed in Section LABEL:sec:OtherCNO.
+That progress is reviewed in Section XI.
 Charged-particle nuclear reactions occurring at energies below the Coulomb barrier are affected by electron screening, the shielding of the nuclear charge by electrons.
 Because of their very different atomic environments, laboratory reactions and those occurring in the solar plasma are affected in distinct ways.
-Section LABEL:sec:Screening describes the current status of efforts to account for these differences, when solar cross sections are extracted from laboratory measurements.
-The energy transport inside the radiative zone is affected by the metals444In this work, we follow the astrophysical terminology where ”metals” refer to all chemical elements except for the two lightest ones, hydrogen and helium.
+Section XII describes the current status of efforts to account for these differences, when solar cross sections are extracted from laboratory measurements.
+The energy transport inside the radiative zone is affected by the metals444In this review we follow the astrophysical terminology in which refer to all chemical elements except for the two lightest ones, hydrogen and helium.
 that were incorporated into the Sun when it first formed.
@@
 Very few of the needed radiative opacities have been measured, and the conditions under which measurements are made typically are not identical to those in the Sun.
-Section LABEL:sec:Opacities describes the current state of the art, and discusses connections between opacities and solar composition that have complicated efforts to resolve the solar metallicity problem.
+Section XIII describes the current state of the art, and discusses connections between opacities and solar composition that have complicated efforts to resolve the solar metallicity problem.
 Much of the progress made in constraining the reactions of the pp chain and CNO-I cycles has been possible thanks to new laboratories and the facilities they host.
-The underground LUNA laboratory is a prominent example [2018PrPNP..98...55B].
-Section LABEL:sec:Facilities describes the capabilities of selected laboratories involved in the measurements reviewed here.
+The underground LUNA laboratory is a prominent example [146].
+Section XIV describes the capabilities of selected laboratories involved in the measurements reviewed here.
 Solar neutrino observations
@@
 The original motivation for observations of solar neutrinos was precisely the hope to probe solar fusion.
-The first successful experiment beginning in 1967 at the Homestake mine [1998ApJ...496..505C] offered the surprising result of a neutrino flux suppressed to approximately one third of expectation.
-The intervening years have seen a tour de force of experimental efforts, leading to the resolution of the Solar Neutrino Problem and the confirmation of neutrino oscillations [1998PhRvL..81.1562F, 2001PhRvL..87g1301A, 2002PhRvL..89a1301A, 2003PhRvL..90b1802E, 2004PhRvL..92r1301A].
+The first successful experiment beginning in 1967 at the Homestake mine [196] offered the surprising result of a neutrino flux suppressed to approximately one third of expectation.
+The intervening years have seen a tour de force of experimental efforts, leading to the resolution of the Solar Neutrino Problem and the confirmation of neutrino oscillations [284, 25, 23, 253, 26].
 These high-precision flux and spectral measurements mapped out the details of solar fusion, probed the structure of the Sun, and demonstrated (together with atmospheric neutrino measurements) that neutrinos have masses and mix.
 Table 2 lists the pp-chain and CNO cycle neutrino sources, their endpoint ( decay) or line (electron capture) energies , and predicted SSM fluxes.
-| Source | Eν (MeV) | GS98 | AAGS21 | MB22p | units (cm-2s-1) |
+| Source | Eν (MeV) | GS98 | AAG21 | MB22p | units (cm-2s-1) |
 | pp | 0.420 | 5.96 (0.6%) | 6.00 (0.6%) | 5.95 (0.6%) | 1010 |
@@
 For instance, more precise determinations of the fluxes of neutrinos from the CNO cycle (Fig.
-2) could determine the contemporary abundances of N and O in the solar core [2007A&A...463..755C, 2011ApJ...743...24S, 2022A&A...667L...2K].
-While these abundances differ significantly from the primordial values that are input into the SSM –the primordial abundances are altered by out-of-equilibrium CNO-cycle burning that occurs in the early Sun [Roxburgh85]– this burning redistributes the C, N, and O, but does not alter the net number of such nuclei.
+2) could determine the contemporary abundances of N and O in the solar core [179, 587, 418].
+While these abundances differ significantly from the primordial values that are input into the SSM –the primordial abundances are altered by out-of-equilibrium CNO-cycle burning that occurs in the early Sun [560]– this burning redistributes the C, N, and O, but does not alter the net number of such nuclei.
 Consequently, CNO neutrino flux measurements might allow one to experimentally address the solar composition problem discussed below.
@@
 The fluxes are sensitive to solar/stellar model inputs, including the radiative opacities (see Sect.
-LABEL:sec:Opacities).
-Measurements of the neutrino fluxes from 8B and 7Be can offer a handle on the temperature of the solar core, and other environmental factors [1996PhRvD..53.4202B].
-A measurement of the pp flux with percent-level precision would provide a test of the luminosity constraint, further probing solar power generation mechanisms [2002PhRvC..65b5801B, 2021JPhG...48a5201V].
+XIII).
+Measurements of the neutrino fluxes from 8B and 7Be can offer a handle on the temperature of the solar core, and other environmental factors [81].
+A measurement of the pp flux with percent-level precision would provide a test of the luminosity constraint, further probing solar power generation mechanisms [72, 639].
 The one branch of the pp chain that yet remains undetected, the hep flux, has both the highest energy but also the lowest flux of neutrinos from the Sun.
 The spectrum extends beyond that of the 8B neutrinos, offering a small window for unequivocal observation.
-Limits have been placed on this flux [2020PhRvD.102f2006A], but a definitive observation will likely require a large, next-generation detector such as Hyper-Kamiokande (HyperK) [2018arXiv180504163H] or the Deep Underground Neutrino Experiment (DUNE) [2019PhRvL.123m1803C].
-Finally, the observation of monochromatic neutrinos produced by electron capture reactions on 13N, 15O and 17F (which we refer to as ecCNO neutrinos) [1990PhRvD..41.2964B, 2004PhRvC..69a5801S, 2015PhLB..742..279V] could be rewarding in terms of physical implications.
+Limits have been placed on this flux [18], but a definitive observation will likely require a large, next-generation detector such as Hyper-Kamiokande (HyperK) [359] or the Deep Underground Neutrino Experiment (DUNE) [169].
+Finally, the observation of monochromatic neutrinos produced by electron capture reactions on 13N, 15O and 17F (which we refer to as ecCNO neutrinos) [69, 611, 642] could be rewarding in terms of physical implications.
 Indeed, they can be used as probes of the metallicity of the solar core.
 Moreover, they provide a measure of the electron neutrino survival probability at specific neutrino energies MeV within the broad transition region between vacuum and matter-enhanced oscillations, where currently we have no constraints.
-This would constitute a new test of the large-mixing-angle (LMA) MSW flavor oscillation paradigm.
-The detection of this subdominant component of the CNO-cycle is extremely difficult but it could be within reach of future very large ultra-pure liquid scintillator detectors [2015PhLB..742..279V].
+This would constitute a new test of the large-mixing-angle (LMA) Mikheyev-Smirnov-Wolfenstein (MSW) flavor oscillation solution.
+The detection of this subdominant component of the CNO-cycle is extremely difficult but it could be within reach of future very large ultra-pure liquid scintillator detectors [642].
 Solar neutrinos provide an opportunity to understand the interaction of neutrinos with matter.
 The effect of MSW oscillations has a significant impact on the observed solar neutrino spectrum: with vacuum oscillation dominating at low energy, below approximately 1 MeV, where the survival probability is roughly one half, while matter effects further suppress the flux to a survival probability of roughly one third above approximately 5 MeV.
-The transition region between these two regimes offers an extremely sensitive probe of the details of the interactions of neutrinos with matter, including the potential to search for new physics such as sterile neutrinos, or non-standard interactions, by looking for distortions to the expected spectral shape [2019arXiv190700991B, 2020JHEP...02..038B, 2023JHEP...08..032C].
+The transition region between these two regimes offers an extremely sensitive probe of the details of the interactions of neutrinos with matter, including the potential to search for new physics such as sterile neutrinos, or non-standard interactions, by looking for distortions to the expected spectral shape [125, 135, 199].
 The same theory of matter effects predicts a small regeneration of electron neutrinos during the night time as they propagate through the bulk of the earth.
-This so-called “day/night effect” has been sought after by both Super-Kamiokande (SuperK) and the Sudbury Neutrino Observatory (SNO) [2014PhRvL.112i1805R, 2002PhRvL..89a1302A], but a significant observation is still limited by statistics.
+This so-called “day/night effect” has been sought after by both Super-Kamiokande (SuperK) and the Sudbury Neutrino Observatory (SNO) [546, 24], but a significant observation is still limited by statistics.
 Future data from HyperKamiokande or DUNE may be needed to confirm our understanding of this prediction of the MSW effect.
-The slight tension between measurements of the mass splitting parameters, in solar neutrino experiments and terrestrial data sensitive to the same parameter, from the KamLAND reactor experiment has now disappeared with the addition of new data on day-night effects and spectral shapes [Esteban2020TheFO, 2018PrPNP.102...48C].
-Further tightening of the oscillation parameters can be expected from JUNO, which will obtain highly precise data on both the solar mass splitting and the mixing angle [2016JPhG...43c0401A].
+The slight tension between measurements of the mass splitting parameters, in solar neutrino experiments and terrestrial data sensitive to the same parameter, from the KamLAND reactor experiment has now disappeared with the addition of new data on day-night effects and spectral shapes [262, 170].
+Further tightening of the oscillation parameters can be expected from JUNO, which will obtain highly precise data on both the solar mass splitting and the mixing angle [39].
 II.1.1 The solar composition problem
 Solar photospheric abundances, determined with spectroscopic techniques, are a fundamental input for the construction of SSMs, but also they are used as input to nearly all models in astrophysics, including stellar evolution, protoplanetary disks, and galactic chemical evolution.
-Over the last two decades, the development of three dimensional radiation hydrodynamic models of the solar atmosphere [2009MmSAI..80..711L, 2000A&A...359..729A] and of techniques to study line formation under non-local thermodynamic conditions [2005ARA&A..43..481A], together with improved descriptions of atomic properties (e.g.
-transition strengths 2022A&A...661A.140M, 2021MNRAS.508.2236B), have led to a significant revision of solar abundances.
-Initial results based on these more sophisticated methods favored a markedly lower solar metallicity [2009ARA&A..47..481A], particularly in the CNO elements, than obtained in the 90s from older techniques [1993oee..conf...15G, 1998SSRv...85..161G].
-Recently, two groups have revisited the solar photospheric abundances using modern methods [2021A&A...653A.141A, 2022A&A...661A.140M].
-While 2021A&A...653A.141A obtain results consistent with their previous findings, 2022A&A...661A.140M find an O abundance intermediate between that of Asplund’s group and those from the 90s, in agreement with another 3D-based determination by 2011SoPh..268..255C.
-Interestingly, 2022A&A...661A.140M find higher C and N abundances and, indirectly, a higher Ne abundance due to the larger Ne to O ratio measured in the solar corona [2018ApJ...855...15Y].
-This leads to a combined metal-to-hydrogen ratio that is by chance comparable to those from 1993oee..conf...15G and 1998SSRv...85..161G, albeit with a different mixture of elements.
+Over the last two decades, the development of three dimensional radiation hydrodynamic models of the solar atmosphere [450, 60] and of techniques to study line formation under non-local thermodynamic conditions [55], together with improved descriptions of atomic properties (e.g.
+transition strengths Magg et al.
+455, Bergemann et al.
+112), have led to a significant revision of solar abundances.
+Initial results based on these more sophisticated methods favored a markedly lower solar metallicity [59], particularly in the CNO elements, than obtained in the 90s from older techniques [311, 312].
+Recently, two groups have revisited the solar photospheric abundances using modern methods [56, 455].
+While Asplund et al.
+[56] obtain results consistent with their previous findings, Magg et al.
+[455] find an O abundance intermediate between that of Asplund’s group and those from the 90s, in agreement with another 3D-based determination by Caffau et al.
+[168].
+Interestingly, Magg et al.
+[455] find higher C and N abundances and, indirectly, a higher Ne abundance due to the larger Ne to O ratio measured in the solar corona [669].
+This leads to a combined metal-to-hydrogen ratio that is by chance comparable to those from Grevesse and Noels [311] and Grevesse and Sauval [312], albeit with a different mixture of elements.
 Considering that uncertainties in element abundances are difficult to quantify, it has become customary to consider two canonical sets of abundances, which we refer to as high metallicity (HZ) and low metallicity (LZ) solar admixtures.
-See, for example, 2017ApJ...835..202V for SSM reference values.
-In this context, the new solar abundance determinations by 2021A&A...653A.141A fall into the LZ category, while those from 2022A&A...661A.140M are HZ.
+See, for example, Vinyoles et al.
+[643] for SSM reference values.
+In this context, the new solar abundance determinations by Asplund et al.
+[56] fall into the LZ category, while those from Magg et al.
+[455] are HZ.
 Solar models employing the LZ abundances fail to reproduce most helioseismic probes of solar properties.
-This disagreement constitutes the so-called solar composition problem [2004ApJ...606L..85B, 2005ApJ...618.1049B, 2006ApJ...649..529D] that has defied a complete solution.
+This disagreement constitutes the so-called solar composition problem [96, 73, 237] that has defied a complete solution.
 All proposed modifications to physical processes in SSMs offer, at best, only partial improvements in some helioseismic probes, see e.g.
-[2005ApJ...627.1049G, 2007A&A...463..755C, 2008PhR...457..217B, 2010ApJ...713.1108G, 2011ApJ...743...24S].
-The same conclusions are obtained with SSMs computed with the newest LZ [2021A&A...653A.141A] and HZ [2022A&A...661A.140M] abundances as discussed also there.
+[320, 179, 97, 319, 587].
+The same conclusions are obtained with SSMs computed with the newest LZ [56] and HZ [455] abundances as discussed also there.
 An alternative possibility is to consider modifications to the physical inputs of SSMs at the level of the constitutive physics, radiative opacities in particular.
 This is possible because most helioseismic probes depend actually not directly on the solar composition, but on the radiative opacity profile in the solar interior, i.e.
-on the combination of solar composition and atomic opacities (see § LABEL:sec:Opacities).
+on the combination of solar composition and atomic opacities (see § XIII).
 The same can be said for solar neutrinos from the pp chain.
-Early work [2005ApJ...621L..85B, 2004ESASP.559..574M] already suggested that a localized increase in opacities could solve or, at least, alleviate the disagreement of low-Z solar models with helioseismology, and 2009A&A...494..205C and 2010ApJ...724...98V showed that a tilted increase in radiative opacities, with a few percent increase in the solar core and a larger (15-20%) increase at the base of the convective envelope could lead to LZ SSMs that would satisfy helioseismic probes equally well as HZ SSMs.
-The degeneracy between solar composition and opacities can be broken using CNO solar neutrinos, e.g., following the methodology developed in 2008ApJ...687..678H.
-Such a study was recently carried out by the Borexino collaboration [2020Natur.587..577B, 2022PhRvL.129y2701A, 2023PhRvD.108j2005B].
+Early work [80, 483] already suggested that a localized increase in opacities could solve or, at least, alleviate the disagreement of low-Z solar models with helioseismology, and Christensen-Dalsgaard et al.
+[194] and Villante [641] showed that a tilted increase in radiative opacities, with a few percent increase in the solar core and a larger (15-20%) increase at the base of the convective envelope could lead to LZ SSMs that would satisfy helioseismic probes equally well as HZ SSMs.
+The degeneracy between solar composition and opacities can be broken using CNO solar neutrinos, e.g., following the methodology developed in Haxton and Serenelli [340].
+Such a study was recently carried out by the Borexino collaboration [139, 49, 95].
 The Borexino measurement of the CNO-I cycle neutrino flux (the 13N and 15O fluxes) were used to determine the C+N core abundance.
@@
 The interpretation of future improved CNO neutrino measurements will be impacted, unless these nuclear physics uncertainties are reduced.
-Very recently, 2024JHEP...02..064G presented a new global determination of all solar neutrino fluxes using all available experimental data, including the latest phases of Borexino.
+Very recently, Gonzalez-Garcia et al.
+[307] presented a new global determination of all solar neutrino fluxes using all available experimental data, including the latest phases of Borexino.
 Results from this global analysis are in line with those from Borexino, although the added 13N and 15O fluxes are about 10% lower than the Borexino result alone.
-A comparison of the solar fluxes with SSM calculations [B23Fluxes] shows that HZ SSMs are in better agreement with solar neutrino fluxes than the LZ SSMs, pointing toward a C+N solar core abundance consistent with HZ abundances.
+A comparison of the solar fluxes with SSM calculations [345] shows that HZ SSMs are in better agreement with solar neutrino fluxes than the LZ SSMs, pointing toward a C+N solar core abundance consistent with HZ abundances.
 The discrimination that solar neutrino fluxes can offer between solar compositions is at most, however, of the order of 2.
-(See Table 2 in 2024JHEP...02..064G).
+(See Table 2 in Gonzalez-Garcia et al.
+[307]).
 Helioseismology also offers the potential to determine the total solar metallicity, i.e.
@@
 Attempting the latter is difficult because the imprint of partial ionization of metals is quite subtle.
-Previous work along this line [2006ApJ...644.1292A] found an overall metallicity consistent with HZ abundances, but new work [2024A&A...681A..57B] claims to favor LZ values.
+Previous work along this line [48] found an overall metallicity consistent with HZ abundances, but new work [160] claims to favor LZ values.
 While this method depends very weakly on radiative opacities, the abundance determination is degenerate with the equation of state.
@@
 Initial experiments showed lower rates than expected, with interesting but inconclusive statistical precision.
-Very recently the BEST experiment [2022PhRvL.128w2501B, 2022PhRvC.105f5502B, 2018PhRvD..97g3001B] has confirmed this anomaly at more than 4.
-Many possible explanations have been explored, but for each there are contradictions [2023EPJC...83..578K, 2022JHEP...11..082A, 2023PhLB..84237983G, 2023JHEP...05..143B].
-A sterile neutrino explanation is disfavored [2022JHEP...10..164G, 2022EPJC...82..116G], particularly because of conflict with solar neutrino limits (see Fig.
-3), as well as cosmological bounds [2021PhRvD.104l3524H].
+Very recently the BEST experiment [89, 90, 88] has confirmed this anomaly at more than 4.
+Many possible explanations have been explored, but for each there are contradictions [414, 28, 300, 143].
+A sterile neutrino explanation is disfavored [299, 304], particularly because of conflict with solar neutrino limits (see Fig.
+3), as well as cosmological bounds [328].
 The sterile neutrino contribution to the solar flux is fundamentally limited by the luminosity constraint, that the Sun’s total energy output is the result of nuclear reactions that produce neutrinos, whether active or sterile.
@@
 3 above, as well as Fig.
-1 of 2022EPJC...82..116G.
-The precisely known rate of electron capture on 71Ge to 71Ga, for which the half-life is 11.43(3) d [1985PhRvC..31..666H], places an important constraint on the neutrino absorption rate [2023PhLB..84237983G, PhysRevC.108.035502].
-The half-life has now been remeasured by two groups, 2024PhRvC.109e5501N and PhysRevC.108.L021602, yielding values of 11.468(8) d and 11.46(4) d, consistent with the previous determination while increasing the precision.
+1 of Goldhagen et al.
+[304].
+The precisely known rate of electron capture on 71Ge to 71Ga, for which the half-life is 11.43(3) d [334], places an important constraint on the neutrino absorption rate [300, 256].
+The half-life has now been remeasured by two groups, Norman et al.
+[514] and Collar and Yoon [198], yielding values of 11.468(8) d and 11.46(4) d, consistent with the previous determination while increasing the precision.
 The allowed matrix element that governs neutrino capture to the ground state of 71Ge can be extracted to a precision of 1% from the electron capture rate.
 The ground-state transition by itself generates a significant anomaly: inclusion of the excited-state contributions doubles the effect.
-In Appendix A we describe some of the details of both the 1997 extraction of the allowed (Gamow Teller) strength by 1997PhRvC..56.3391B and the recent update of PhysRevC.108.035502, ELLIOTT2024104082.
-PhysRevC.108.035502, ELLIOTT2024104082 find the ground-state cross section for absorbing 51Cr neutrinos is 2.5% lower than Bahcall’s value.
+In Appendix A we describe some of the details of both the 1997 extraction of the allowed (Gamow Teller) strength by Bahcall [71] and the recent update of Elliott et al.
+[256, 255].
+Elliott et al.
+[256, 255] find the ground-state cross section for absorbing 51Cr neutrinos is 2.5% lower than Bahcall’s value.
 This work includes the contributions of weak magnetism and radiative corrections, which are shown to be sub-1% effects.
 An improved extraction of the excited-state contribution to neutrino absorption was performed using data from forward-angle (p,n) scattering.
-This led to a slight increase in that contribution relative to 1997PhRvC..56.3391B, thereby reducing the 2.5% difference above by about half, for the total 51Cr and 37Ar capture cross sections.
-The procedures followed by PhysRevC.108.035502, ELLIOTT2024104082 could be extended to the solar neutrino 71Ga capture cross section, including the contribution from the high-energy 8B neutrinos.
+This led to a slight increase in that contribution relative to Bahcall [71], thereby reducing the 2.5% difference above by about half, for the total 51Cr and 37Ar capture cross sections.
+The procedures followed by Elliott et al.
+[256, 255] could be extended to the solar neutrino 71Ga capture cross section, including the contribution from the high-energy 8B neutrinos.
 That has not been done, but as noted in Appendix A, could help resolve small discrepancies noted there.
@@
 II.1.3 The Boron-8 neutrino spectrum
-The neutrino spectrum from , extending to approximately 17 MeV, has been the most accessible part of the solar neutrino spectrum [2016PhRvD..94e2010A, 2020PhRvD.102f2006A], playing an important role in disclosing the physics of neutrino oscillation and testing the SSM.
-Knowing it with precision, particularly at the high energy end, has renewed interest in the context of observing the hep neutrinos [2006ApJ...653.1545A, 2020PhRvD.102f2006A], which could be accessible with future observatories [2022arXiv220212839A].
-In 2011RvMP...83..195A it was decided to recommend the spectrum calculated by Winter et al.
-[2003PhRvL..91y2501W, 2006PhRvC..73b5503W, Win:2007] based on their measurement of the alpha spectrum from .
-This spectrum showed excellent agreement with an independent experiment [2006PhRvC..73e5802B].
+The neutrino spectrum from , extending to approximately 17 MeV, has been the most accessible part of the solar neutrino spectrum [2, 18], playing an important role in disclosing the physics of neutrino oscillation and testing the SSM.
+Knowing it with precision, particularly at the high energy end, has renewed interest in the context of observing the hep neutrinos [22, 18], which could be accessible with future observatories [53].
+In SF II it was decided to recommend the spectrum calculated by Winter et al.
+[659, 660, 658] based on their measurement of the alpha spectrum from .
+This spectrum showed excellent agreement with an independent experiment [124].
 Both radiative and recoil-order corrections were included by Winter et al.
-The radiative corrections are relatively small, due to a cancellation between the real and virtual contributions [1995PhRvD..52.5362B].
-The recoil-order corrections are dominated by the weak magnetism part, which has been deduced from measurements of the analog electromagnetic decays [1995PhRvC..51.2778D].
-Since then three additional measurements [2011PhRvC..83f5802K, 2012PhRvL.108p2502R, 2023PhRvC.107c2801L] of the -decay alpha spectrum have been performed.
+The radiative corrections are relatively small, due to a cancellation between the real and virtual contributions [98].
+The recoil-order corrections are dominated by the weak magnetism part, which has been deduced from measurements of the analog electromagnetic decays [226].
+Since then three additional measurements [402, 550, 445] of the -decay alpha spectrum have been performed.
 All three find that the peak of the alpha spectrum appears about 20 keV lower than determined by Winter et al.
@@
 1.
-Apparently the uncertainty estimations for the neutrino spectrum in 2012PhRvL.108p2502R are not quite correct.
-2023PhRvC.107c2801L show that the uncertainty due only to the weak magnetism part, for which all authors follow similar prescriptions, is larger in the 0-12 MeV range than the overall uncertainty estimated by Roger et al.
+Apparently the uncertainty estimations for the neutrino spectrum in Roger et al.
+[550] are not quite correct.
+Longfellow et al.
+[445] show that the uncertainty due only to the weak magnetism part, for which all authors follow similar prescriptions, is larger in the 0-12 MeV range than the overall uncertainty estimated by Roger et al.
 We estimated the uncertainties due to the weak magnetism term only and found agreement with those of Longfellow et al.
 2.
-All three of the recent efforts [2011PhRvC..83f5802K, 2012PhRvL.108p2502R, 2023PhRvC.107c2801L] used detectors (Si strip) more complicated than those used by Winter et al.
+All three of the recent efforts [402, 550, 445] used detectors (Si strip) more complicated than those used by Winter et al.
 (a Si surface barrier detector).
@@
 The effects are different for a calibration source than from the emission from trapped as an ion.
-Based on these considerations, we recommend the spectrum of 2023PhRvC.107c2801L, but suggest that conclusions sensitive to the choice of neutrino spectrum, such as evidence for hep neutrinos, should be evaluated using both the Winter et al.
+Based on these considerations, we recommend the spectrum of Longfellow et al.
+[445], but suggest that conclusions sensitive to the choice of neutrino spectrum, such as evidence for hep neutrinos, should be evaluated using both the Winter et al.
 and the Longfellow et al.
 spectra.
-Given the importance of the weak magnetism contributions, it would be good to have a new experiment with reduced uncertainties that could be compared directly to the results of 1995PhRvC..51.2778D.
-Recoil-order effects in the system were most recently measured by 2011PhRvC..83f5501S, making use of the alignment- correlation.
-Using combinations of their measurements and correlation data from 1980PhRvC..22..738M, they were able to extract some of the recoil-order matrix elements.
-Calculations of these matrix elements in the symmetry-adapted no-core shell model were recently performed by 2022PhRvL.128t2503S.
-The comparison to the experimentally derived matrix elements of 1980PhRvC..22..738M, 1995PhRvC..51.2778D, and 2011PhRvC..83f5501S is not straightforward because the experiments quote averages over the whole beta spectrum as opposed to fits for each of the four levels used in the calculation.
+Given the importance of the weak magnetism contributions, it would be good to have a new experiment with reduced uncertainties that could be compared directly to the results of de Braeckeleer et al.
+[226].
+Recoil-order effects in the system were most recently measured by Sumikama et al.
+[613], making use of the alignment- correlation.
+Using combinations of their measurements and correlation data from McKeown et al.
+[472], they were able to extract some of the recoil-order matrix elements.
+Calculations of these matrix elements in the symmetry-adapted no-core shell model were recently performed by Sargsyan et al.
+[569].
+The comparison to the experimentally derived matrix elements of McKeown et al.
+[472], de Braeckeleer et al.
+[226], and Sumikama et al.
+[613] is not straightforward because the experiments quote averages over the whole beta spectrum as opposed to fits for each of the four levels used in the calculation.
 A more detailed comparison of theory to experiment that takes into account the averaging would be helpful.
-An additional point of comparison that has been used previously is the beta spectrum from [1996PhRvC..54..411B].
+An additional point of comparison that has been used previously is the beta spectrum from [76].
 An experiment with improved statistics and systematics that extends over the full range of beta energies would be very useful.
-The existing data from 1987PhRvC..36..298N is in good agreement with the alpha spectrum from Winter et al.
+The existing data from Napolitano et al.
+[505] is in good agreement with the alpha spectrum from Winter et al.
 But given the apparent 20 keV difference that the newer alpha-spectrum experiments described above report, it would be good to have a modern beta spectrum measurement specifically designed to address this difference.
@@
 II.2.1 Super-Kamiokande and Hyper-Kamiokande
-The first real-time detection of solar neutrinos was achieved by the Kamiokande experiment [1989PhRvL..63...16H], which detected the neutrinos via elastic scattering of electrons, a process in which the outgoing electron’s direction is highly correlated to that of the incoming neutrino.
+The first real-time detection of solar neutrinos was achieved by the Kamiokande experiment [353], which detected the neutrinos via elastic scattering of electrons, a process in which the outgoing electron’s direction is highly correlated to that of the incoming neutrino.
 This allowed Kamiokande to directly point back to the neutrinos’ origin: our Sun.
 The successor experiment, Super-Kamiokande, is a 50-kton volume of pure water, surrounded by over 10,000 photon detectors, which has achieved an unparalleled program of neutrino observations and other physics over its several decades of operation.
-This program has included the highest-precision measurement of the elastic scattering signal from solar neutrino interactions, as well as sensitive searches for the day/night distortions of the spectral shape [2016PhRvD..94e2010A].
-Now filled with a gadolinium additive to enhance neutron capture, SuperK’s primary current focus is the search for the Diffuse Supernova Neutrino Background (DSNB) via inverse beta decay [2023ApJ...951L..27H].
-It also continues its atmospheric neutrino program: SuperK measurements made 25 years ago demonstrated the neutrino oscillations were responsible for the puzzling zenith-angle-dependence of this flux [PhysRevLett.81.1562].
+This program has included the highest-precision measurement of the elastic scattering signal from solar neutrino interactions, as well as sensitive searches for the day/night distortions of the spectral shape [2].
+Now filled with a gadolinium additive to enhance neutron capture, SuperK’s primary current focus is the search for the Diffuse Supernova Neutrino Background (DSNB) via inverse beta decay [335].
+It also continues its atmospheric neutrino program: SuperK measurements made 25 years ago demonstrated the neutrino oscillations were responsible for the puzzling zenith-angle-dependence of this flux [285].
 In parallel to ongoing operation of SuperK, an even larger sister project is under construction.
@@
 Being equally sensitive to all active flavors, the neutral current interaction yields a measurement of the total, flavor-blind neutrino flux.
-It was this capability that allowed SNO to resolve the solar neutrino problem, demonstrating that most of the s produced by the Sun were arriving at earth in a different flavor state [2001PhRvL..87g1301A, 2002PhRvL..89a1301A, 2004PhRvL..92r1301A].
+It was this capability that allowed SNO to resolve the solar neutrino problem, demonstrating that most of the s produced by the Sun were arriving at earth in a different flavor state [25, 23, 26].
 SNO ceased data taking in 2006.
-Analysis of the data set has continued, and produced a number of new results, including constraints on non-standard effects such as Lorentz violation and neutrino decay [2018PhRvD..98k2013A, 2019PhRvD..99c2013A], limits on the fluxes of hep neutrinos and the diffuse supernova neutrino background (DSNB) [2020PhRvD.102f2006A], and neutron production from cosmogenic muons and from atmospheric neutrino interactions [2019PhRvD..99k2007A, 2019PhRvD.100k2005A].
+Analysis of the data set has continued, and produced a number of new results, including constraints on non-standard effects such as Lorentz violation and neutrino decay [17, 19], limits on the fluxes of hep neutrinos and the diffuse supernova neutrino background (DSNB) [18], and neutron production from cosmogenic muons and from atmospheric neutrino interactions [21, 20].
 After decommissioning, the detector was repurposed for the SNO+ experiment, in which the target material was replaced with a pure organic liquid scintillator: linear alkyl benzene (LAB), loaded with 2.2 g/L of the secondary fluor, PPO.
 The high light yield of this scintillator, coupled with the location in SNOLAB –at 6-km water equivalent, one of the deepest underground laboratories in the world– and an extremely well-understood detector, allow for a range of high-precision measurement programs.
-In a preliminary water phase, SNO+ demonstrated detection of neutron captures on hydrogen [PhysRevC.102.014002], an impressive technical achievement in an unloaded water detector; a low background measurement of 8B solar neutrinos [PhysRevD.99.012012]; several searches for invisible modes of nucleon decay [PhysRevD.99.032008, PhysRevD.105.112012]; and detection of antineutrinos from reactors over 200 km away [PhysRevLett.130.091801].
+In a preliminary water phase, SNO+ demonstrated detection of neutron captures on hydrogen [42], an impressive technical achievement in an unloaded water detector; a low background measurement of 8B solar neutrinos [41]; several searches for invisible modes of nucleon decay [40, 36]; and detection of antineutrinos from reactors over 200 km away [35].
 The low backgrounds achieved and the detector’s demonstrated technical capabilities set the stage for future measurements.
-In a “partial-fill” stage, in which the upper half of the detector was filled with liquid scintillator while the lower half still contained water, the first demonstration of event-level direction reconstruction was achieved for 8B neutrinos in a scintillator detector [PhysRevD.109.072002].
+In a “partial-fill” stage, in which the upper half of the detector was filled with liquid scintillator while the lower half still contained water, the first demonstration of event-level direction reconstruction was achieved for 8B neutrinos in a scintillator detector [34].
 This was facilitated by a lower loading of PPO at that time, which results in lower scintillation yield and a slower time profile, effectively enhancing the Cherenkov component, which can be leveraged for directional information.
@@
 The detector’s extreme radiopurity was key to its success, and over its 14 years of data taking, ending in October 2021, Borexino proved itself capable of covering the entire solar neutrino spectrum.
-At the beginning of its data taking in 2007 the level of 238U and 232Th contamination in Borexino were lower than and respectively [2008PhLB..658..101B], paving the way for the first measurement of the 7Be sub-MeV solar neutrinos, followed by a low-threshold measurement of the 8B flux and a first indication of the pep neutrinos.
+At the beginning of its data taking in 2007 the level of 238U and 232Th contamination in Borexino were lower than and respectively [137], paving the way for the first measurement of the 7Be sub-MeV solar neutrinos, followed by a low-threshold measurement of the 8B flux and a first indication of the pep neutrinos.
 After a purification campaign that further reduced the liquid scintillator contamination, Borexino aimed at improving the accuracy of its results in the measurement of the pp-chain solar neutrinos.
-The flux of neutrinos produced in the pp fusion process was first reported in 2014Natur.512..383B and later improved, along with all the other fluxes from the pp chain, apart from the hep neutrinos [2018Natur.562..505B, 2019PhRvD.100h2004A, 2020PhRvD.101f2001A].
+The flux of neutrinos produced in the pp fusion process was first reported in BOREXINO Collaboration et al.
+[138] and later improved, along with all the other fluxes from the pp chain, apart from the hep neutrinos [136, 16, 15].
 The pp flux was determined to an uncertainty of 11%, improving the neutrino-based estimate of solar luminosity, while 7Be neutrinos were measured with an uncertainty of , half that of the SSM prediction.
 Furthermore, using the SSM to constrain the flux of CNO neutrinos, the pep neutrino signal was established with a significance exceeding for the first time, and the 8B flux was measured with a threshold as low as 3 MeV.
-After completing the investigation of the pp-chain neutrinos, Borexino reported the first detection of solar neutrinos produced in the CNO-cycle [2020Natur.587..577B], demonstrating directly that this mode of hydrogen burning operates in stars.
-Profiting from a larger exposure and a better understanding of the radioactive backgrounds enabled by the unprecedented thermal stability of the detector, Borexino further improved its measurement of the CNO neutrino flux [2022PhRvL.129y2701A, 2023PhRvD.108j2005B], where the new CNO result was used in combination with existing solar neutrino results to probe solar composition.
-Borexino’s physics program was not limited to solar neutrinos: its outstanding radiopurity made it a excellent detector for geoneutrinos [2020PhRvD.101a2009A] and for searches for various rare processes.
-Borexino data were used to constrain exotic properties of neutrinos, such as their magnetic moments [2017PhRvD..96i1103A] and non standard interactions with matter [2020JHEP...02..038B].
+After completing the investigation of the pp-chain neutrinos, Borexino reported the first detection of solar neutrinos produced in the CNO-cycle [139], demonstrating directly that this mode of hydrogen burning operates in stars.
+Profiting from a larger exposure and a better understanding of the radioactive backgrounds enabled by the unprecedented thermal stability of the detector, Borexino further improved its measurement of the CNO neutrino flux [49, 95], where the new CNO result was used in combination with existing solar neutrino results to probe solar composition.
+Borexino’s physics program was not limited to solar neutrinos: its outstanding radiopurity made it a excellent detector for geoneutrinos [14] and for searches for various rare processes.
+Borexino data were used to constrain exotic properties of neutrinos, such as their magnetic moments [13] and non standard interactions with matter [135].
 II.2.4 JUNO
-The Jiangmen Underground Neutrino Observatory (JUNO) is a large liquid scintillator detector currently under construction in an underground laboratory with a vertical overburden of about 650 m (roughly 1800 m water equivalent) in Jiangmen city in Southern China [2015arXiv150807166A].
+JUNO [10] is a large liquid scintillator detector currently under construction in an underground laboratory with a vertical overburden of about 650 m (roughly 1800 m water equivalent) in Jiangmen city in Southern China.
 JUNO is located 52.5 km from two nuclear power plants, a baseline optimized for JUNO’s primary goal, the determination of the neutrino mass ordering [28, 29].
@@
 JUNO’s reactor data will constrain to extremely high precision, allowing comparisons between the solar and reactor determinations of this parameter.
-To the extent that oscillation effects can be treated with greater confidence, the connections between solar neutrino flux measurements and the solar fusion reactions generating those fluxes will be sharpened [2016JPhG...43c0401A].
+To the extent that oscillation effects can be treated with greater confidence, the connections between solar neutrino flux measurements and the solar fusion reactions generating those fluxes will be sharpened [39].
 II.2.5 DUNE
-As recommended in the 2023 P5 report [P5Report], the DUNE Collaboration seeks to construct two 10-kton fiducial liquid argon time projection chambers (LArTPC), deep underground in the Homestake mine in South Dakota [2016arXiv160102984A].
+As recommended in the 2023 P5 report [522], the DUNE Collaboration seeks to construct two 10-kton fiducial liquid argon time projection chambers (LArTPC), deep underground in the Homestake mine in South Dakota [4].
 A third LArTPC will follow in Phase II, along with other upgrades.
 The newly developed Sanford Underground Research Facility (SURF) offers 4800-m water equivalent overburden, and will form the far site for a long baseline neutrino program utilizing a high-energy neutrino beam directed from Fermilab to SURF.
-The DUNE LArTPC detectors may be sensitive to measurements of the high-energy solar neutrinos via charged current interactions on argon, which offers good precision on the spectral shape and, thus, the potential to measure day/night effects [2019PhRvL.123m1803C].
+The DUNE LArTPC detectors may be sensitive to measurements of the high-energy solar neutrinos via charged current interactions on argon, which offers good precision on the spectral shape and, thus, the potential to measure day/night effects [169].
 The solar neutrino program at DUNE is limited primarily by backgrounds (the detector is optimized for GeV-scale physics) and energy threshold.
@@
 II.2.6 Future prospects
-A “hybrid” detector that could utilize both Cherenkov and scintillation light simultaneously could achieve unprecedented levels of particle and event identification and, hence, background rejection [2014arXiv1409.5864A].
+A “hybrid” detector that could utilize both Cherenkov and scintillation light simultaneously could achieve unprecedented levels of particle and event identification and, hence, background rejection [37].
 The Cherenkov signature offers directional information, while the high light yield scintillation offers precision energy and vertex reconstruction.
@@
 As a result, both the shape of the waveform and the ratio of the two signals will differ for different particle types.
-A full-scale detector utilizing novel scintillators along with fast and spectrally-sensitive photon detectors, such as the proposed Theia experiment [2020EPJC...80..416A], or the Jinping detector [2017ChPhC..41b3002B] could achieve percent-level precision on the CNO neutrino flux, as well as improving precision across the suite of solar neutrino measurements, such as the 8B spectral shape, and hep neutrino flux.
+A full-scale detector utilizing novel scintillators along with fast and spectrally sensitive photon detectors, such as the proposed Theia experiment [54], or the Jinping detector [102] could achieve percent-level precision on the CNO neutrino flux, as well as improving precision across the suite of solar neutrino measurements, such as the 8B spectral shape and hep neutrino flux.
 As experiments grow larger, and capabilities increase, there are also opportunities to apply detectors designed for other purposes to the detection of solar neutrinos.
-Noble liquid detectors built primarily for dark matter searches and long-baseline neutrino experiments have now reached the size and background levels to be sensitive to solar neutrinos [2024arXiv240802877A].
+Noble liquid detectors built primarily for dark matter searches and long-baseline neutrino experiments have now reached the size and background levels to be sensitive to solar neutrinos [50].
 The detection of neutrinos from the high-energy tail of the 8B spectrum via coherent elastic neutrino-nucleus scattering (“CENS”) is particularly interesting because it is a purely neutral-current process sensitive to the total flux of active neutrinos (above the experimental threshold, about 8 MeV) from the Sun.
 Only the SNO experiment has had such a capability.
-Initial results reported in [2024arXiv240802877A] are in agreement, with limited statistical precision.
-Solid state detectors built purposely for observing CENS have observed neutrinos from stopped pion decay at the theoretically expected rate [doi:10.1126/science.aao0990].
+Initial results reported in [50] are in agreement, with limited statistical precision.
+Solid state detectors built purposely for observing CENS have observed neutrinos from stopped pion decay at the theoretically expected rate [29].
 The 1HH reaction ()
@@
 Higher derivative terms are neglected in this review since they only contribute at the level.
-The recommended values for and will be discussed in Section III.5.
+The recommended values for and are discussed in Section III.5.
 Here we focus on .
-At zero relative energy, can be written as [1968ApJ...152L..17B, 1969ApJ...155..501B],
+At zero relative energy, can be written as [77, 78],
 where is the fine-structure constant, MeV is the proton mass, fm-1 is the deuteron binding wave number, being the proton-neutron reduced mass and the deuteron binding energy, and are the Fermi vector and axial-vector weak coupling constants.
@@
 III.2 Adopted parameters for this review
-For the phase-space factor , we use the same value as in 2011RvMP...83..195A.
-It comes from the value without radiative corrections, [1969ApJ...155..501B], increased by 1.62% to take into account radiative corrections to the cross section [2003PhRvC..67c5502K].
+For the phase-space factor , we use the same value as in SF II.
+It comes from the value without radiative corrections, [78], increased by 1.62% to take into account radiative corrections to the cross section [419].
 The main source of uncertainty in arises from neglected diagrams in which the lepton exchanges a weak boson and a photon with different nucleons.
-These diagrams are estimated to modify by , based on scaling the similar nucleus-dependent correction in superallowed decay [2003PhRvC..67c5502K].
+These diagrams are estimated to modify by , based on scaling the similar nucleus-dependent correction in superallowed decay [419].
 Direct computations of these diagrams were recommended in Solar Fusion II.
 Here we urge again that this computation be carried out.
-For , we use the value from the Particle Data Group (PDG) compilation [2022PTEP.2022h3C01W], , whose central value is larger than the 2008 PDG value, , used in 2011RvMP...83..195A by (or 1.9 ).
+For , we use the value from the Particle Data Group (PDG) compilation [666], , whose central value is larger than the 2008 PDG value, , used in SF II by (or 1.9 ).
 Naively, this would lead to a increase in the central value of according to Eq.
@@
 Its effect will be discussed below.
-For , we take s from the most updated comprehensive analysis of experimental rates with the radiative and Coulomb effects corrected [2020PhRvC.102d5501H].
-This value is consistent with s [2009PhRvC..79e5502H] used in 2011RvMP...83..195A with a larger error.
+For , we take s from the most updated comprehensive analysis of experimental rates with the radiative and Coulomb effects corrected [337].
+This value is consistent with s [336] used in SF II with a larger error.
 The dominant uncertainty in comes from the normalized Gamow-Teller (GT) matrix element .
-Reducing this uncertainty has been the main focus of theoretical work since 1998RvMP...70.1265A.
-In 1998RvMP...70.1265A, was written as , i.e.
+Reducing this uncertainty has been the main focus of theoretical work since SF I.
+In SF I, was written as , i.e.
 the sum of the one- and two-body current matrix elements, and , respectively, with their uncertainties estimated independently.
-In 2011RvMP...83..195A, two major steps had contributed to reducing the uncertainty on .
+In SF II, two major steps had contributed to reducing the uncertainty on .
 The first was a much deeper understanding of the correlation between the uncertainties in and : the overall uncertainty in could be described by a universal parameter that could be fixed by a single measurement.
-The study of 1998PhRvC..58.1263S demonstrated this phenomenologically in the context of potential-model approaches, while later analysis via effective field theory (EFT) provided a more formal justification [2001PhRvC..63c5501B, 2003PhRvC..67e5206P].
-The second step was the use of the precisely known tritium decay rate to fix this universal parameter, as first proposed by 1991PhRvC..44..619C.
-This has been done in both potential models [1998PhRvC..58.1263S] and in the hybrid EFT approach [2003PhRvC..67e5206P] as explained below.
-In 2011RvMP...83..195A, was determined with three approaches.
+The study of Schiavilla et al.
+[574] demonstrated this phenomenologically in the context of potential-model approaches, while later analysis via effective field theory (EFT) provided a more formal justification [165, 529].
+The second step was the use of the precisely known tritium decay rate to fix this universal parameter, as first proposed by Carlson et al.
+[171].
+This has been done in both potential models [574] and in the hybrid EFT approach [529] as explained below.
+In SF II, was determined with three approaches.
 The first one was the potential model approach.
-In the most elaborate calculation for the pp fusion process, a comparison of the results for five representative modern potentials designed to accurately reproduce nucleon-nucleon scattering data was carried out [1998PhRvC..58.1263S].
+In the most elaborate calculation for the pp fusion process, a comparison of the results for five representative modern potentials designed to accurately reproduce nucleon-nucleon scattering data was carried out [574].
 After adjusting the unknown strength of the two-body exchange currents to reproduce , the variation in that otherwise would come from the choice of the phenomenological potential was largely removed.
@@
 The second and third approaches were both based on EFT.
-The second one was a hybrid EFT (EFT*), which used the current operators derived from EFT in conjunction with the initial and final state wave functions generated by a potential model [2003PhRvC..67e5206P].
-For pp fusion, the relevant two-body current contained only one unknown low-energy constant (LEC) which parameterized the contact axial coupling to two nucleons [2003PhRvC..67e5206P].
+The second one was a hybrid EFT (EFT*), which used the current operators derived from EFT in conjunction with the initial and final state wave functions generated by a potential model [529].
+For pp fusion, the relevant two-body current contained only one unknown low-energy constant (LEC) which parameterized the contact axial coupling to two nucleons [529].
 A weakness of this approach was the mismatch between the operators and wave functions.
@@
 Hence, when the ultraviolet cutoff was changed over a physically reasonable range, the residual cutoff dependence of physical observables provided a measure of the model dependence of the EFT* calculation.
-By combining the error from changing the cutoff in the range of MeV and the 0.4% higher order correction, obtained by multiplying the contribution of the highest calculated order with the small expansion parameter , 2003PhRvC..67e5206P provided the value , in perfect agreement with the one obtained within the phenomenological approach.
+By combining the error from changing the cutoff in the range of MeV and the 0.4% higher order correction, obtained by multiplying the contribution of the highest calculated order with the small expansion parameter , Park et al.
+[529] provided the value , in perfect agreement with the one obtained within the phenomenological approach.
 The third study was performed with the pionless EFT (EFT) approach.
-It is a framework applicable to processes with the characteristic momentum much smaller than the pion mass , such that the pion field can be “integrated out” and becomes non-dynamical (1996APS..DNP..BA03K; 1999PhRvL..82..463B; 1999NuPhA.653..386C).
+It is a framework applicable to processes with the characteristic momentum much smaller than the pion mass , such that the pion field can be “integrated out” and becomes non-dynamical (389; 106; 191).
 In this approach, all nucleon-nucleon interactions and two-body currents are described by point-like contact interactions with a systematic expansion in powers of .
 For all the deuteron weak breakup processes (e.g.
-and scattering) and their inverse processes, including the pp fusion, only one two-body current (with coupling ) is needed up to next-to-next-to-leading order (N2LO) [2001PhRvC..63c5501B].
+and scattering) and their inverse processes, including the pp fusion, only one two-body current (with coupling ) is needed up to next-to-next-to-leading order (N2LO) [165].
 Therefore a single measurement will fix and the rates of all such processes.
 This feature is shared by the other approaches discussed above.
-The computation of in EFT was carried out to the second order in the expansion by 2001PhRvC..64d4002K and then to the fifth order by 2001PhLB..520...87B.
-Constraints on from two-nucleon systems [2002PhLB..549...26B, 2003PhRvC..67b5801C] yielded .
-Based on the consistent results of the above three approaches, 2011RvMP...83..195A recommended .
-The 1998RvMP...70.1265A value from 1969ApJ...155..501B was not re-computed in 2011RvMP...83..195A.
-However, with reaching a accuracy, new calculations of and , together with the full pionful chiral EFT (EFT) computations to remove the unknown systematics of the hybrid EFT, were called for in 2011RvMP...83..195A.
+The computation of in EFT was carried out to the second order in the expansion by Kong and Ravndal [405] and then to the fifth order by Butler and Chen [164].
+Constraints on from two-nucleon systems [166, 188] yielded .
+Based on the consistent results of the above three approaches, SF II recommended .
+The SF I value from Bahcall and May [78] was not re-computed in SF II.
+However, with reaching a accuracy, new calculations of and , together with the full pionful chiral EFT (EFT) computations to remove the unknown systematics of the hybrid EFT, were called for in SF II.
 These two challenges have been met in SF III and will be described below.
 III.3 Experimental progress on muon capture of the deuteron
-In 2011RvMP...83..195A it was also recommended to carry out the experimental determination of the muon capture rate on deuteron, as proposed in the MuSun experiment [2003nucl.ex...4019K, 2017APS..DNP.CG007S].
+In SF II it was also recommended to carry out the experimental determination of the muon capture rate on deuteron, as proposed in the MuSun experiment [388, 566].
 This quantity could be used to constrain without the need to rely on the three-body calculation of .
-The theoretical calculations for the muon capture rate on deuteron have been carried out in EFT [2005PhRvC..72f1001C] (it is possible to impose neutron energy cut to isolate the low-energy neutron events so EFT is applicable [2003nucl.ex...4019K]), chiral hybrid EFT* [2002PhLB..533...25A], phenomenological potential model [2011PhRvC..83a4002M], and recently in EFT (2012PhRvL.108e2502M, 2018PhRvL.121d9901M; 2018PhRvC..98f5506A; 2023FrP....1049919C; 2023PhRvC.107f5502B; PhysRevC.109.035502).
+The theoretical calculations for the muon capture rate on deuteron have been carried out in EFT [189] (it is possible to impose neutron energy cut to isolate the low-energy neutron events so EFT is applicable [388]), chiral hybrid EFT* [43], phenomenological potential model [463], and recently in EFT (460, 461; 7; 183; 133; 303).
 However, the MuSun result is yet to be released.
 On the other hand, we should notice that the muon capture processes happen at a rather large momentum transfer compared to pp fusion.
-The momentum transfer dependence of the single nucleon axial coupling constant , with , has been recently studied by 2018RPPh...81i6301H, who have provided an experimental determination for the axial charge radius, given by fm2, where is defined by the relation for small .
-The % uncertainty on has an impact on the ability of the MuSun experiment alone to directly constrain [2018PhRvC..98f5506A, 2023FrP....1049919C, 2023PhRvC.107f5502B, PhysRevC.109.035502].
-In the next few years, lattice Quantum Chromodynamics (QCD) calculations of are expected to reduce the uncertainty by a factor of two or more (see 2023slft.confE.240M for the most recent review of the lattice results).
-Currently, lattice QCD results are consistent with 2018RPPh...81i6301H on , often with comparable uncertainties.
-At larger ( GeV2), there is a growing tension between lattice QCD predictions of and the phenomenological determination from older neutrino-deuterium bubble chamber data [2016PhRvD..93k3015M] with the lattice QCD results yielding a 30% larger neutrino-nucleon cross section over a large range of [2022ARNPS..72..205M].
+The momentum transfer dependence of the single nucleon axial coupling constant , with , has been recently studied by Hill et al.
+[352], who have provided an experimental determination for the axial charge radius, given by fm2, where is defined by the relation for small .
+The % uncertainty on has an impact on the ability of the MuSun experiment alone to directly constrain [7, 183, 133, 303].
+In the next few years, lattice Quantum Chromodynamics (QCD) calculations of are expected to reduce the uncertainty by a factor of two or more (see Meyer [476] for the most recent review of the lattice results).
+Currently, lattice QCD results are consistent with Hill et al.
+[352] on , often with comparable uncertainties.
+At larger ( GeV2), there is a growing tension between lattice QCD predictions of and the phenomenological determination from older neutrino-deuterium bubble chamber data [477] with the lattice QCD results yielding a 30% larger neutrino-nucleon cross section over a large range of [478].
 In any case, it is evident that accurate experimental determinations of muon capture rates on the deuteron and other light nuclei, which can be addressed theoretically ab initio approaches, represent fundamental tests for the theoretical approaches themselves, either within EFT or EFT, and might be able to provide the necessary experimental information to fix the unknown parameters of the theory.
 III.4 Progress in calculations since SF II
-Below we summarize the calculations performed after 2011RvMP...83..195A using different approaches.
+Below we summarize the calculations performed after SF II using different approaches.
 III.4.1 EFT
-The pioneering work of 2013PhRvL.110s2503M used the next-to-next-to-next-to-leading-order (N3LO) chiral two-nucleon potential [2003PhRvC..68d1001E, 2011PhR...503....1M] augmented with higher order () two-photon and vacuum-polarization electromagnetic interactions.
+The pioneering work of Marcucci et al.
+[464] used the next-to-next-to-next-to-leading-order (N3LO) chiral two-nucleon potential [260, 453] augmented with higher order () two-photon and vacuum-polarization electromagnetic interactions.
 These corrections reduced by % mainly due to the vacuum-polarization-induced pp wave function distortion.
-This was consistent with the 0.84% first found in the potential model calculation of 1998PhRvC..58.1263S.
-This correction was also included in the EFT* calculation of [2003PhRvC..67e5206P].
-The relevant LECs were fitted to reproduce the binding energies, magnetic moments, and GT matrix element in to obtain , with the -wave initial state contributing at , which was about the accuracy level of the calculation [2013PhRvL.110s2503M].
-However, using EFT at next-to-leading-order (NLO), 2019PhRvC.100b1001A later found that -wave only contributed at the order of .
-Re-examining the computer programs, 2013PhRvL.110s2503M found an error in the determination of one of the -wave reduced matrix elements (associated with the longitudinal multipole operator).
-Consequently, 2019PhRvL.123a9901M reported in the Erratum that .
-In more recent work, 2016PhLB..760..584A used chiral interactions and consistent currents up to N3LO (called next-to-next-to-leading-order (N2LO) in the original literature since the second order vanished) and developed a robust procedure for the error quantification.
-In particular, they analyzed a family of 42 interactions [2016PhRvX...6a1019C] with 7 different cutoff values from 450 to 600 MeV.
-The 26 LECs were fitted to 6 different pools of input data including and scatterings, as well as the binding energies and charge radii of 3H and 3He, the quadrupole moment of 2H, and .
+This was consistent with the 0.84% first found in the potential model calculation of Schiavilla et al.
+[574].
+This correction was also included in the EFT* calculation of [529].
+The relevant LECs were fitted to reproduce the binding energies, magnetic moments, and GT matrix element in to obtain , with the -wave initial state contributing at , which was about the accuracy level of the calculation [464].
+However, using EFT at next-to-leading-order (NLO), Acharya et al.
+[9] later found that -wave only contributed at the order of .
+Re-examining the computer programs, Marcucci et al.
+[464] found an error in the determination of one of the -wave reduced matrix elements (associated with the longitudinal multipole operator).
+Consequently, Marcucci et al.
+[465] reported in the Erratum that .
+In more recent work, Acharya et al.
+[5] used chiral interactions and consistent currents up to N3LO (called next-to-next-to-leading-order (N2LO) in the original literature since the second order vanished) and developed a robust procedure for the error quantification.
+In particular, they analyzed a family of 42 interactions [172] with 7 different cutoff values from 450 to 600 MeV.
+The 26 LECs were fitted to 6 different pools of input data including nucleopn-nucleon () and scatterings, as well as the binding energies and charge radii of 3H and 3He, the quadrupole moment of 2H, and .
 This thorough study yielded .
-In both of the EFT calculations mentioned above, a widely used relation first proposed by 2009PhRvL.103j2502G that linked the two-body axial current LEC with the LEC from the vertex was employed.
-However, later 2018PhRvL.121d9901M found that there was a factor missing in this - relation which was then acknowledged in the Erratum of 2019PhRvL.122b9901G.
-Fortunately, this error was unimportant in muon capture on deuteron [2018PhRvL.121d9901M] and it only affected at the level [2023arXiv230403327A].
-In the most recent and comprehensive EFT study, 2023arXiv230403327A compared the above calculations of 2019PhRvL.123a9901M and 2016PhLB..760..584A in detail.
-In addition to the increase of from using the correct - relation, updating the input parameters to their most recent values increased by in both calculations, mainly due to the increase of from its 2011RvMP...83..195A value mentioned above.
-Furthermore, 2019PhRvL.123a9901M received a increase from removing the truncation error of the basis functions which effectively cut off the long distance part of the wave functions, as 2017PhRvC..95c1301A advocated.
-After these corrections, the difference between 2019PhRvL.123a9901M and 2016PhLB..760..584A was reconciled and the combined result was found to be [2023arXiv230403327A]
-2023arXiv230403327A obtained consistent values of using four different EFT models for the nuclear interaction.
-This value is also in agreement with the result obtained in EFT by 2022arXiv220710176D (see below).
-In estimating the order-by-order convergence, however, 2023arXiv230403327A used one of the EFT models, which is able to nicely reproduce the deuteron properties already at leading order.
+In both of the EFT calculations mentioned above, a widely used relation first proposed by Gazit et al.
+[294] that linked the two-body axial current LEC with the LEC from the vertex was employed.
+However, later Marcucci et al.
+[461] found that there was a factor missing in this - relation which was then acknowledged in the Erratum of Gazit et al.
+[295].
+Fortunately, this error was unimportant in muon capture on deuteron [461] and it only affected at the level [8].
+In the most recent and comprehensive EFT study, Acharya et al.
+[8] compared the above calculations of Marcucci et al.
+[465] and Acharya et al.
+[5] in detail.
+In addition to the increase of from using the correct - relation, updating the input parameters to their most recent values increased by in both calculations, mainly due to the increase of from its SF II value mentioned above.
+Furthermore, Marcucci et al.
+[465] received a increase from removing the truncation error of the basis functions which effectively cut off the long distance part of the wave functions, as Acharya et al.
+[6] advocated.
+After these corrections, the difference between Marcucci et al.
+[465] and Acharya et al.
+[5] was reconciled and the combined result was found to be [8]
+Acharya et al.
+[8] obtained consistent values of using four different EFT models for the nuclear interaction.
+This value is also in agreement with the result obtained in EFT by De-Leon and Gazit [227] (see below).
+In estimating the order-by-order convergence, however, Acharya et al.
+[8] used one of the EFT models, which is able to nicely reproduce the deuteron properties already at leading order.
 Therefore, the EFT error of Eq.
 (6) is likely to be an underestimate, warranting the enlarged error advocated in Section III.4.4.
-Finally, we would like to mention the work of 2022PhRvC.106e5501L, where the power counting of the EFT weak current operator involved in the pp reaction is revisited using renormalization group (RG) invariance as the guideline.
+Finally, we would like to mention the work of Liu et al.
+[438], where the power counting of the EFT weak current operator involved in the pp reaction is revisited using renormalization group (RG) invariance as the guideline.
 In particular, it is argued that the contact two-body axial current proportional to the LEC must appear one order lower than assessed by naive dimensional analysis.
 Then it can be shown that RG invariance is fulfilled at N2LO.
-However, the estimate for obtained by 2022PhRvC.106e5501L does not use a value for obtained by fitting , but extracted in order to match the value of obtained in 2011RvMP...83..195A.
-Therefore, we will not consider the work of 2022PhRvC.106e5501L in the present evaluation.
+However, the estimate for obtained by Liu et al.
+[438] does not use a value for obtained by fitting , but extracted in order to match the value of obtained in SF II.
+Therefore, we will not consider the work of Liu et al.
+[438] in the present evaluation.
 III.4.2 EFT
-The universal two-body current coupling was determined using in EFT for the first time by 2019PhRvC.100e5502D.
-This calculation was carried out up to NLO using the dibaryon formulation of 2001NuPhA.694..511B, which partially resumed higher order effective range contributions to improve the convergence.
-Then, this result was used by 2022arXiv220710176D, with updated input parameters, to obtain
-where the electromagnetic correction was not calculated but assumed to be the same as the potential model value, [1998PhRvC..58.1263S].
+The universal two-body current coupling was determined using in EFT for the first time by De-Leon et al.
+[228].
+This calculation was carried out up to NLO using the dibaryon formulation of Beane and Savage [103], which partially resumed higher order effective range contributions to improve the convergence.
+Then, this result was used by De-Leon and Gazit [227], with updated input parameters, to obtain
+where the electromagnetic correction was not calculated but assumed to be the same as the potential model value, [574].
 While this number was not model independent, the model dependence was believed to be well below the assigned error.
@@
 However, such calculations are quite challenging and not yet available at the required precision.
-A proof-of-principle calculation was carried out by 2017PhRvL.119f2002S using a background field method to determine both the pp fusion GT matrix element and .
+A proof-of-principle calculation was carried out by Savage et al.
+[570] using a background field method to determine both the pp fusion GT matrix element and .
 This exploratory calculation utilized a single pion mass at the SU(3)-flavor–symmetric point with MeV, a single and relatively coarse lattice spacing of fm, and a single volume.
 The calculation was performed under the assumption that the two- and three-nucleon systems were deeply bound.
-Without this assumption, matrix elements computed in the finite volume can be significantly different from those in infinite volume, due to Lellouch-Lüscher factors [2001CMaPh.219...31L] that lead to power-law finite volume corrections.
-These can range from the few-percent level to [2013PhRvD..88i4507B, 2015PhRvL.115x2001B].
-More recent lattice QCD calculations have found that two-nucleon systems at heavy pion masses are in fact not bound (2016JHEP...10..101I; 2019PhRvD..99g4505F; 2021PhRvC.103a4003H; 2021arXiv210810835A).
-These efforts, which employ interpolating operators more sophisticated than those of 2017PhRvL.119f2002S, suggest there could be large systematic uncertainties affecting the conclusions of 2017PhRvL.119f2002S stemming from misidentification of the spectrum and inaccurate Lellouch-Lüscher factors.
-In addition, 2021PhRvL.127x2003G found that the two-baryon spectrum may be particularly sensitive to discretization effects, which would also have important implications for the continuum extrapolation of the matrix elements.
-These issues are discussed in some detail in 2022FBS....63...67T.
+Without this assumption, matrix elements computed in the finite volume can be significantly different from those in infinite volume, due to Lellouch-Lüscher factors [430] that lead to power-law finite volume corrections.
+These can range from the few-percent level to [144, 145].
+More recent lattice QCD calculations have found that two-nucleon systems at heavy pion masses are in fact not bound (377; 279; 355; 38).
+These efforts, which employ interpolating operators more sophisticated than those of Savage et al.
+[570], suggest there could be large systematic uncertainties affecting the conclusions of Savage et al.
+[570] stemming from misidentification of the spectrum and inaccurate Lellouch-Lüscher factors.
+In addition, Green et al.
+[309] found that the two-baryon spectrum may be particularly sensitive to discretization effects, which would also have important implications for the continuum extrapolation of the matrix elements.
+These issues are discussed in some detail in Tews et al.
+[618].
 Finally, results with pion masses MeV are needed for accurate extrapolation to the physical pion mass.
-Given these unresolved systematic issues, the result of 2017PhRvL.119f2002S is not included in the present evaluation – even though the extracted value is consistent with our recommended range quoted in Section III.4.4.
-The lattice EFT computation of 2015PhLB..741..301R performed the pp fusion calculation by implementing EFT on a spacetime lattice.
+Given these unresolved systematic issues, the result of Savage et al.
+[570] is not included in the present evaluation – even though the extracted value is consistent with our recommended range quoted in Section III.4.4.
+The lattice EFT computation of Rupak and Ravi [562] performed the pp fusion calculation by implementing EFT on a spacetime lattice.
 The purpose of this leading-order study was to demonstrate that lattice EFT could reproduce the infinite volume and continuum result of EFT such that it could be applied to various reactions of astrophysical interest in the future.
@@
 The above discussions show that determinations of from EFT in Eq.(6), EFT in Eq.(7), and lattice QCD (although with unquantified systematics) are all consistent with each other.
-Furthermore, these values are also consistent with the recommended value of 2011RvMP...83..195A, provided the central value is increased by 555To know the precise shift requires an explicit calculation.
+Furthermore, these values are also consistent with the recommended value of SF II, provided the central value is increased555To know the precise shift requires an explicit calculation.
 However, if the shift is within the range of 0.8-1.0 , the averaged value and remain the same within the significant digits.
-to to account for the update.
+by to to account for the update.
 Averaging this value with the EFT value in Eq.(6) and the EFT value in Eq.(7) yields with per degree of freedom to be .
-This shows that the EFT, EFT and 2011RvMP...83..195A estimates of are all mutually consistent.
+This shows that the EFT, EFT and SF II estimates of are all mutually consistent.
 In addition, we would like to advocate adding an additional correlated error to account for any input that would tend to move all results in a coordinated way.
-For example, from 2011RvMP...83..195A to SF III, we experienced the shift due to the update of the PDG value of .
-It is not inconceivable that or other input parameters or physics could change again by similar amounts: the large PDG inflation factor of 2.7 reflects the tension that continues to exist among measurements [2022PTEP.2022h3C01W].
+For example, from SF II to SF III, we experienced the shift due to the update of the PDG value of .
+It is not inconceivable that or other input parameters or physics could change again by similar amounts: the large PDG inflation factor of 2.7 reflects the tension that continues to exist among measurements [666].
 Therefore, assigning an additional correlated error seems reasonable.
@@
 III.5 Progress in and
-Using EFT, 2013PhLB..720..385C computed analytically with all partial waves included up to N2LO.
+Using EFT, Chen et al.
+[190] computed analytically with all partial waves included up to N2LO.
 The Fermi matrix element only contributed at the level and was neglected compared with the GT matrix element.
@@
 The result was MeV-1 and MeV-2.
-In EFT, for keV was fit to polynomials of [2019PhRvL.123a9901M].
+In EFT, for keV was fit to polynomials of [465].
 Depending on using a quadratic or quartic fit, changed from 12.23 to 10.82 MeV-1 and changed from 178.4 to 317.4 MeV-2.
-2016PhLB..760..584A used cubit fit and keV to obtain =10.84(2) MeV-1 and =317.8(13) MeV-2.
-Recently 2023arXiv230403327A repeated the calculation of 2019PhRvL.123a9901M with the same energy range and cubic fit as 2016PhLB..760..584A, obtaining a consistent result with 2016PhLB..760..584A: MeV-1 and MeV-2.
+Acharya et al.
+[5] used cubit fit and keV to obtain =10.84(2) MeV-1 and =317.8(13) MeV-2.
+Recently Acharya et al.
+[8] repeated the calculation of Marcucci et al.
+[465] with the same energy range and cubic fit as Acharya et al.
+[5], obtaining a consistent result with Acharya et al.
+[5]: MeV-1 and MeV-2.
 We will take these as the recommended values from EFT.
@@
 Hence, we take the average of the EFT and EFT results as the recommended value:
-Finally, we comment on the work of 2019PhRvC.100c5805G which performed a study of in a wide energy range.
+Finally, we comment on the work of Gaspard et al.
+[290] which performed a study of in a wide energy range.
 The main focus of this work was to perform a proper energy-dependence analysis of the pp process, in order to reliably extract and .
@@
 However, the 2H(,)3He reaction plays an important role in the development of protostars, because the onset of deuterium burning slows down the protostars contraction and heating, increasing their lifespan.
-Accurate knowledge of the 2H(,)3He reaction rate, particularly within the few keV range corresponding to the Gamow peak in protostars, is vital for modeling protostellar evolution effectively [1988ApJ...332..804S].
+Accurate knowledge of the 2H(,)3He reaction rate, particularly within the few keV range corresponding to the Gamow peak in protostars, is vital for modeling protostellar evolution effectively [607].
 Another astrophysical scenario where the 2H(,)3He reaction plays a key role is big bang nucleosynthesis (BBN), responsible for the production of light elements during the first few minutes of the Universe.
-Among these elements, deuterium is an excellent indicator of cosmological parameters because its primordial abundance is the most sensitive to the baryon density and critically depends on the radiation density of the early Universe, see for example a review of BBN [2016RvMP...88a5004C].
-The reactions involved in the synthesis of deuterium are: production via the well known )2H process and destruction via the 2H(2H,)3He, 2H(2H,)3H and 2H(,)3He reactions [2020JCAP...03..010F].
-Since the comprehensive review performed in 2011RvMP...83..195A, there have been both new measurements as well as advances in the theoretical and phenomenological analysis of the 2H(,)3He reaction.
-The new experimental results have been determined with accelerator-based measurements of the cross section [2019EPJA...55..137T, 2020Natur.587..210M, 2021PhRvC.103d5805T], as well as plasma-based, intertial confinement fusion measurements [2020PhRvC.101d2802Z, 2022FrP....10.4339M] and are reviewed in Section IV.2 below.
-On the theoretical side, there have been advances in ab initio calculations, where nucleons are the fundamental degrees of freedom interacting among themselves and with the external electromagnetic probe, see for example [2023FrP....1129094E].
+Among these elements, deuterium is an excellent indicator of cosmological parameters because its primordial abundance is the most sensitive to the baryon density and critically depends on the radiation density of the early Universe, see for example a review of BBN [209].
+The reactions involved in the synthesis of deuterium are: production via the well known )2H process and destruction via the 2H(2H,)3He, 2H(2H,)3H and 2H(,)3He reactions [269].
+Since the comprehensive review performed in SF II, there have been both new measurements and advances in the theoretical and phenomenological analysis of the 2H(,)3He reaction.
+The new experimental results have been determined with accelerator-based measurements of the cross section [622, 490, 635], as well as plasma-based, intertial confinement fusion measurements [682, 481] and are reviewed in Section IV.2 below.
+On the theoretical side, there have been advances in ab initio calculations, where nucleons are the fundamental degrees of freedom interacting among themselves and with the external electromagnetic probe, see for example [254].
 Finally, Bayesian analysis methods have been used to model the energy dependence of the S-factor, starting from the ab initio predictions and applying a polynomial approximation to it.
@@
 Different experimental approaches were followed to measure its cross section.
-2019EPJA...55..137T irradiated deuterated titanium targets with a proton beam and detected the -rays with two high-purity germanium (HPGe) detectors placed at different angles.
+Tišma et al.
+[622] irradiated deuterated titanium targets with a proton beam and detected the -rays with two high-purity germanium (HPGe) detectors placed at different angles.
 The final S-factor is provided at four energies in the 47 - 210 keV range, with approximately 15% statistical uncertainty.
-More recently, the LUNA Collaboration performed a measurement in the energy range of keV at the underground in the Gran Sasso Laboratories, exploiting the six orders of magnitude suppression of the cosmic radiation background [2021FrASS...7..119Z, 2018IJMPA..3343010C].
+More recently, the LUNA Collaboration performed a measurement in the energy range of keV at the underground in the Gran Sasso Laboratories, exploiting the six orders of magnitude suppression of the cosmic radiation background [265, 182].
 A windowless deuterium gas target was used and the -rays emitted by the 2H(,)3He reaction were detected by a large HPGe detector at 90∘ with respect to the beam axis.
-Great care was taken to minimize all sources of systematic uncertainties in the S-factor overall uncertainty at the 3% level [2020EPJA...56..144M, 2020Natur.587..210M].
-These new results provided stringent constraints on cosmological parameters obtained by comparing the precise primordial deuterium abundance predictions of the standard BBN model with astronomical observations [2018ApJ...855..102C].
-A deeper discussion of the LUNA results and their implications can be found in [2021MNRAS.502.2474P, 2021JCAP...04..020P, 2021JCAP...03..046Y, 2021ApJ...923...49M] 666After the closing of the present review, the 2020Natur.587..210M data were re-analyzed to extract the experimental 2H()3He -ray angular distribution [2024PhRvC.110c2801S].
-The data confirm the ab initio angular distributions used in the original analysis by 2020Natur.587..210M.
-Finally, a new measurement was performed at the Helmholtz-Zentrum Dresden-Rossendorf in the 300-1000 keV energy range using implanted deuterium targets on tantalum backings and two HPGe detectors [2021PhRvC.103d5805T].
-The resulting S-factors show tension with the analysis of the LUNA results extrapolated to keV [2020Natur.587..210M].
+Great care was taken to minimize all sources of systematic uncertainties in the S-factor overall uncertainty at the 3% level [489, 490].
+These new results provided stringent constraints on cosmological parameters obtained by comparing the precise primordial deuterium abundance predictions of the standard BBN model with astronomical observations [201].
+A deeper discussion of the LUNA results and their implications666After the closing of the present review, the Mossa et al.
+[490] data were re-analyzed to extract the experimental 2H()3He -ray angular distribution [610].
+The data confirm the ab initio angular distributions used in the original analysis by Mossa et al.
+[490].
+can be found in [535, 534, 668, 488]
+Finally, a new measurement was performed at the Helmholtz-Zentrum Dresden-Rossendorf in the 300-1000 keV energy range using implanted deuterium targets on tantalum backings and two HPGe detectors [635].
+The resulting S-factors show tension with the analysis of the LUNA results extrapolated to keV [490].
 However, they are affected by large systematic uncertainties.
-A new measurement of the reaction is planned at the Felsenkeller laboratory in Germany [2019sone.conf..249B].
-In addition to the new accelerator-based results, two recent sets of measurements [2020PhRvC.101d2802Z, 2022FrP....10.4339M] have also been performed using the inertial confinement fusion plasma-based platform [2017PhPl...24d1407G, 2022FrP....10.4339M, 2023FrP....1180821G], which has recently begun to be exploited for this type of work [e.g., 2016PhRvL.117c5002Z, 2017NatPh..13.1227C].
-Both measurements, which were performed at the OMEGA laser facility (see section LABEL:subsec:WG8:Plasma), used laser-driven implosions of spherical plastic-shell capsules filled with H2D2 gas, and measured the emitted -rays using a gas-Cherenkov-detector [2014RScI...85kE124H] that was calibrated applying the technique described in 2019RScI...90l3504Z.
-The initial experiment [2020PhRvC.101d2802Z] obtained good statistics at an energy of 16 keV by making several repeated measurements, with a final statistical uncertainty of 6% and a systematic one of 17% (dominated by uncertainty in the absolute calibration of the detector).
-The second experiment [2022FrP....10.4339M] obtained data at three different energies in the 17-37 keV region, with comparable systematic uncertainty, but with larger statistical uncertainty due to fewer repeated measurements.
+A new measurement of the reaction is planned at the Felsenkeller laboratory in Germany [111].
+In addition to the new accelerator-based results, two recent sets of measurements [682, 481] have also been performed using the inertial confinement fusion plasma-based platform [293, 481, 292], which has recently begun to be exploited for this type of work [e.g., 681, 177].
+Both measurements, which were performed at the OMEGA laser facility (see section XIV.4), used laser-driven implosions of spherical plastic-shell capsules filled with H2D2 gas, and measured the emitted -rays using a gas-Cherenkov-detector [346] that was calibrated applying the technique described in Zylstra et al.
+[683].
+The initial experiment [682] obtained good statistics at an energy of 16 keV by making several repeated measurements, with a final statistical uncertainty of 6% and a systematic one of 17% (dominated by uncertainty in the absolute calibration of the detector).
+The second experiment [481] obtained data at three different energies in the 17-37 keV region, with comparable systematic uncertainty, but with larger statistical uncertainty due to fewer repeated measurements.
 The results obtained on this unique platform agree within error bars with the accelerator-based measurements.
@@
 This is why few-nucleon reactions can be used as an “ideal” laboratory, where the ab initio framework can be stringently tested in systems under more theoretical control.
-The most recent ab initio calculation of the 2H(,)3He reaction is that of 2016PhRvL.116j2501M.
-Here the pair-correlated Hyperspherical Harmonics ab initio method was used to calculate the initial scattering and final bound state wave functions (see 2008JPhG...35f3101K and 2020FrP.....8...69M for details).
-The nuclear interaction model adopted to describe the nuclear state in 2016PhRvL.116j2501M consists of a two-nucleon term, the Argonne (AV18) potential [1995PhRvC..51...38W], augmented by a three-nucleon contribution, the Urbana IX (UIX) potential [1995PhRvL..74.4396P].
-The AV18 potential can reproduce the large two-nucleon database with a /datum [1995PhRvC..51...38W], while the combination AV18/UIX can describe quite accurately the properties of 3He, the spectra of light p-shell nuclei [2001ARNPS..51...53P], and scattering observables (see for instance 2002PhRvC..65c4002W).
-The electromagnetic current operator used in 2016PhRvL.116j2501M includes, in addition to the non-relativistic one-body operator, two- and three-body terms required by gauge invariance in a system of interacting particles.
-These terms were constructed in 2005PhRvC..72a4001M.
+The most recent ab initio calculation of the 2H(,)3He reaction is that of Marcucci et al.
+[462].
+Here the pair-correlated Hyperspherical Harmonics ab initio method was used to calculate the initial scattering and final bound state wave functions (see Kievsky et al.
+[394] and Marcucci et al.
+[459] for details).
+The nuclear interaction model adopted to describe the nuclear state in Marcucci et al.
+[462] consists of a two-nucleon term, the Argonne (AV18) potential [661], augmented by a three-nucleon contribution, the Urbana IX (UIX) potential [540].
+The AV18 potential can reproduce the large two-nucleon database with a /datum [661], while the combination AV18/UIX can describe quite accurately the properties of 3He, the spectra of light p-shell nuclei [533], and scattering observables (see for instance Wood et al.
+[664]).
+The electromagnetic current operator used in Marcucci et al.
+[462] includes, in addition to the non-relativistic one-body operator, two- and three-body terms required by gauge invariance in a system of interacting particles.
+These terms were constructed in Marcucci et al.
+[467].
 The model was then tested against various electromagnetic observables, to access the quality of its predictions.
 As a potential model, however, there is no systematic procedure for assigning uncertainties for observables whose values are unknown.
-The results of 2016PhRvL.116j2501M have been found to be about 10% higher than the experimental data of 2020Natur.587..210M.
+The results of Marcucci et al.
+[462] have been found to be about 10% higher than the experimental data of Mossa et al.
+[490].
 Given the lack of a procedure for quantifying errors in calculations based on phenomenological interactions and currents, it is difficult to access the significance of this discrepancy.
@@
 IV.4 Phenomenological and Bayesian Analyses
-In the energy range of astrophysical interest, the 2H(,)3He reaction does not have any resonance or coupled channels that can give rise to non-trivial energy dependence, and the S-factor can be modelled by a low-order polynomial in energy [1967ARA&A...5..525F].
-Moreover, the recent results from LUNA [2020Natur.587..210M], combined with previous measurements, place stringent constraints on the cross section, reducing the uncertainty over the range of interest to Solar fusion and BBN [2002NuPhA.706..203C, 2020Natur.587..210M, 2021MNRAS.502.2474P, 2021JCAP...04..020P, 2021JCAP...03..046Y].
-A significant change since 2011RvMP...83..195A has been the wide adoption of Bayesian analysis methods to evaluate thermonuclear reaction rates and provide more rigorous uncertainty estimates in a statistical sense [2016ApJ...831..107I].
-This approach has been followed in the analysis presented here, which closely follow work on the 2H(,)3He reaction reported in 2021ApJ...923...49M, and references therein.
-The mathematical details of the Bayesian method will be found in Appendix LABEL:AppendixWG4: Here we focus on the application to 2H(,)3He by the authors cited above.
-Both the ab initio prediction from 2005PhRvC..72a4001M as well as a third-order polynomial in the energy were used to constrain the data, where we define an -order polynomial as
-As noted in the literature, the energy dependence of 2005PhRvC..72a4001M is more reliable than the absolute normalization.
-Furthermore, the updated prediction in 2016PhRvL.116j2501M, as noted above, is larger than that of 2005PhRvC..72a4001M.
-Therefore, 2021ApJ...923...49M modeled the prediction of 2005PhRvC..72a4001M, denoted , as
+In the energy range of astrophysical interest, the 2H(,)3He reaction does not have any resonance or coupled channels that can give rise to non-trivial energy dependence, and the S-factor can be modelled by a low-order polynomial in energy [276].
+Moreover, the recent results from LUNA [490], combined with previous measurements, place stringent constraints on the cross section, reducing the uncertainty over the range of interest to Solar fusion and BBN [176, 490, 535, 534, 668].
+A significant change since SF II has been the wide adoption of Bayesian analysis methods to evaluate thermonuclear reaction rates and provide more rigorous uncertainty estimates in a statistical sense [367].
+This approach has been followed in the analysis presented here, which closely follow work on the 2H(,)3He reaction reported in Moscoso et al.
+[488], and references therein.
+The mathematical details of the Bayesian method will be found in Appendix B: Here we focus on the application to 2H(,)3He by the authors cited above.
+Both the ab initio prediction from Marcucci et al. [467] as well as a third-order polynomial in the energy were used to constrain the data, where we define an -order polynomial as
+As noted in the literature, the energy dependence of Marcucci et al.
+[467] is more reliable than the absolute normalization.
+Furthermore, the updated prediction in Marcucci et al.
+[462], as noted above, is larger than that of Marcucci et al.
+[467].
+Therefore, Moscoso et al.
+[488] modeled the prediction of Marcucci et al.
+[467], denoted , as
 where and are an unknown scale factor and offset to be determined in the analysis.
@@
 In this review, we take the further step of applying Bayes Model Averaging to a number of reasonable models that describe the data, which systematically captures additional theoretical “model selection uncertainty”.
-Detailes of the Bayesian analysis and model averaging procdure are presented in Appendix LABEL:AppendixWG4.
+Detailes of the Bayesian analysis and model averaging procdure are presented in Appendix B.
 The models we explore include Eq.
-(11) with both and , as done in 2021ApJ...923...49M, as well as setting , and then we consider polynomials in energy of order .777The implementation of the Bayesian analysis presented here is different than that of 2021ApJ...923...49M but it was verified to produce the same results when the same implementation of the systematic uncertainties was made.
-With such a Bayes model averaging, we can quantitatively compare and contrast the polynomial parameterizations of 2H(,)3He along with ab initio results predicted by 2005PhRvC..72a4001M, 2016PhRvL.116j2501M, including the extra variance that arises from this set of reasonable models, e.g., using polynomials of different order, as well as the phenomenological model of Eq.
+(11) with both and , as done in Moscoso et al.
+[488], as well as setting , and then we consider polynomials in energy of order777The implementation of the Bayesian analysis presented here is different than that of Moscoso et al.
+[488] but it was verified to produce the same results when the same implementation of the systematic uncertainties was made.
+.
+With such a Bayes model averaging, we can quantitatively compare and contrast the polynomial parameterizations of 2H(,)3He along with ab initio results predicted by Marcucci et al.
+[467, 462], including the extra variance that arises from this set of reasonable models, e.g., using polynomials of different order, as well as the phenomenological model of Eq.
 (11).
 Results are listed in Table 3.
-The first column lists the model (i.e., polynomial or ab initio), the second column the natural logarithm of the Gaussian approximation to the BF, and the third column the corresponding weight in the model averaging.
+The first column lists the model (i.e., polynomial or ab initio), the second column the natural logarithm of the Gaussian approximation to the Bayes factor (BF), and the third column the corresponding weight in the model averaging.
 From this model averaging, the resulting prediction for the 2H(,)3He S-factor at a few representative energies is provided in Table 4, where the first uncertainty arises from the first term in Eq.
-(LABEL:eq:BMA_Var) and the second uncertainty is from the second term, which we denote as model selection uncertainty.
+(B) and the second uncertainty is from the second term, which we denote as model selection uncertainty.
 The resulting Bayes model average prediction of the S-factor over the entire kinematic range considered is depicted in Fig.
@@
 (11) over the polynomial approximations.
-One reason for this might be that the scaled ab initio models have only one or two free parameters, and the energy dependence of given in 2005PhRvC..72a4001M, 2016PhRvL.116j2501M is sufficient to accurately describe the various data sets.
+One reason for this might be that the scaled ab initio models have only one or two free parameters, and the energy dependence of given in Marcucci et al.
+[467, 462] is sufficient to accurately describe the various data sets.
 On the other hand, polynomial approximations are disfavored as they require more parameters to capture the energy dependence.
 Also of note, in the model of Eq.
-(11) with and using 2016PhRvL.116j2501M for , the scale factor is given by
-indicating that the prediction in 2016PhRvL.116j2501M overestimates the 2H(,)3He data by 7.9%, consistent with expectations noted above.
+(11) with and using Marcucci et al.
+[462] for , the scale factor is given by
+indicating that the prediction in Marcucci et al.
+[462] overestimates the 2H(,)3He data by 7.9%, consistent with expectations noted above.
 In comparison, the third order polynomial fit predicts values of that are higher at and lower at keV as compared to those in Table 4.
@@
 The code also can provide a prediction of the mean value and uncertainty of at any energy over the same kinematic range.
-| keV | 10 | 20 | 40 | 80 | 91 | 100 | 120 |
+| keV | keV | keV | keV | keV | keV | keV | keV |
 | 2.028(51)(9) | 2.644(60)(8) | 3.276(70)(7) | 4.579(94)(5) | 7.31(15)(0) | 8.11(16)(0) | 8.77(18)(0) | 10.24(21)(0) |
@@
 The 3He(3He,)4He reaction terminates the pp-I chain.
-The ratio of its rate to that of the 3He()7Be reaction controls the branching to the pp-II and pp-III chains, so that in 1998RvMP...70.1265A increasing was discussed as a potential solution to the Solar Neutrino Problem.
-Subsequent experiments, notably a very low energy measurement at the LUNA 50 kV accelerator deep underground in Gran Sasso [1999PhRvL..82.5205B] and a complementary experiment at somewhat higher energies [2004PhRvC..69a5802K] ruled out such an increase in , as summarized in 2011RvMP...83..195A.
+The ratio of its rate to that of the 3He()7Be reaction controls the branching to the pp-II and pp-III chains, so that in SF I increasing was discussed as a potential solution to the Solar Neutrino Problem.
+Subsequent experiments, notably a very low energy measurement at the LUNA 50 kV accelerator deep underground in Gran Sasso [132] and a complementary experiment at somewhat higher energies [417] ruled out such an increase in , as summarized in SF II.
 V.1 Shape of the particle spectrum
-Since no new absolute measurements of have been reported since 2011RvMP...83..195A, we consider the same four experiments [1999PhRvL..82.5205B, 1998PhRvC..57.2700J, 2004PhRvC..69a5802K, 1987NuPhA.467..273K].
+Since no new absolute measurements of have been reported since SF II, we consider the same four experiments [132, 384, 417, 408].
 However, there is new information on the energy spectrum of the emitted protons.
-This spectrum has recently been measured using Inertial Confinement Fusion (ICF) plasmas for a Gamow peak energy of 165 keV [2017PhRvL.119v2701Z].
+This spectrum has recently been measured using inertial confinement fusion plasmas for a Gamow peak energy of 165 keV [680].
 The results show significant structure, indicating the presence of a sequential reaction mechanism passing through the unbound ground state of .
 This spectrum is important for all of the measurements since they determined cross sections by detecting only the protons above an energy threshold.
-The efficiency correction by which 1987NuPhA.467..273K account for the threshold is not well documented, but the more recent measurements [1999PhRvL..82.5205B, 1998PhRvC..57.2700J, 2004PhRvC..69a5802K] utilized the genbod event generator [James68-GENBOD].
+The efficiency correction by which Krauss et al.
+[408] account for the threshold is not well documented, but the more recent measurements [132, 384, 417] utilized the genbod event generator [380].
 It employs a simple reaction model without final state interactions, the Pauli principle, or Coulomb effects that are important near the spectrum endpoint.
@@
 Published proton spectra obtained using accelerator beams do exist (Fig.
-2 of 1974PhRvC...9..805D and Fig.
-3 of 1987NuPhA.467..273K), and although they are not corrected for instrumental effects, they do not appear to be well described by ellipses.
-The 3He(3He,)4He reaction may proceed via several sequential mechanisms, including (with in its ground or first excited state) and (where the di-proton is two correlated protons in a singlet state) [2015PhRvC..92a4003B].
+2 of Dwarakanath [251] and Fig.
+3 of Krauss et al.
+[408]), and although they are not corrected for instrumental effects, they do not appear to be well described by ellipses.
+The 3He(3He,)4He reaction may proceed via several sequential mechanisms, including (with in its ground or first excited state) and (where the di-proton is two correlated protons in a singlet state) [151].
 More complicated three-body decay channels, sometimes called direct decays, are also possible.
 Only a relatively narrow intermediate state (here, the 1 MeV wide 5Li ground state) could produce a peak in the energy spectrum, and precise classification of the reaction mechanism is in general both experimentally and theoretically ambiguous.
-When coincident detection of reaction products is used [as in 1999PhRvL..82.5205B], possible angular correlations between the reaction products also matter.
-A reaction through the ground state of would emit the second proton preferentially either along or opposite the direction of the first proton, while di-proton emission would tend to send both protons in the same direction [2015PhRvC..92a4003B].
+When coincident detection of reaction products is used [as in Bonetti et al.
+[132]], possible angular correlations between the reaction products also matter.
+A reaction through the ground state of would emit the second proton preferentially either along or opposite the direction of the first proton, while di-proton emission would tend to send both protons in the same direction [151].
 The solid curves in Fig.
-6 show two different R-matrix models of the proton energy spectrum at keV, fitted to the spectrum measured by 2017PhRvL.119v2701Z for MeV.
+6 show two different R-matrix models of the proton energy spectrum at keV, fitted to the spectrum measured by Zylstra et al.
+[680] for MeV.
 The measured spectrum differs significantly from the elliptical spectrum of genbod.
 Based on their Fig.
-6, the experiment of 1998PhRvC..57.2700J had a detection threshold of about 5 MeV proton energy.
-The measurements by 1999PhRvL..82.5205B required a coincidence between two detectors, with a detection threshold of 2 MeV proton energy in each detector.
+6, the experiment of Junker et al.
+[384] had a detection threshold of about 5 MeV proton energy.
+The measurements by Bonetti et al.
+[132] required a coincidence between two detectors, with a detection threshold of 2 MeV proton energy in each detector.
 Finally, Fig.
-19 of 2004PhRvC..69a5802K indicates a detection threshold of about 4 MeV proton energy.
+19 of Kudomi et al.
+[417] indicates a detection threshold of about 4 MeV proton energy.
 We have estimated their sensitivities to the assumed proton spectrum by integrating the curves shown in Fig.
@@
 V.2 Recommendation
-Based on these investigations, our recommended fit is the one from 2011RvMP...83..195A, but with an additional 4% systematic uncertainty typical of the 3% to 6% corrections to the extrapolated proton spectra estimated above to take the uncertain spectral shape into account.
+Based on these investigations, our recommended fit is the one from SF II, but with an additional 4% systematic uncertainty typical of the 3% to 6% corrections to the extrapolated proton spectra estimated above to take the uncertain spectral shape into account.
 This gives
@@
 The analysis of new measurements should use Monte Carlo simulations considering a variety of plausible energy spectra and angular correlations, in order to estimate the sensitivity to these effects.
-We note finally that the proton energy spectrum may depend on , as this has been found to be the case for neutrons from the mirror reaction [2018PhRvL.121d2501G].
+We note finally that the proton energy spectrum may depend on , as this has been found to be the case for neutrons from the mirror reaction [291].
 The reaction ()
@@
 It has been studied experimentally using three main methods, with references given in subsections VI.1 and VI.3: First, by detecting at least two of the three -rays from the reaction and taking the angular correlation with the alpha beam direction into account (prompt- method).
-Second, by detecting the induced 7Be [2002NuPhA.708....3T] radioactivity (activation method).
+Second, by detecting the induced 7Be [620] radioactivity (activation method).
 Third, by counting the 7Be recoils (recoil method).
 VI.1 Previous recommendation in SF II
-The 3He()7Be S-factor recommended in 2011RvMP...83..195A was developed in multiple steps.
-First, a model was selected for the shape of , based on existing nucleon-level calculations [1986NuPhA.460..559K, 2001PhRvC..63e4002N].
+The 3He()7Be S-factor recommended in SF II was developed in multiple steps.
+First, a model was selected for the shape of , based on existing nucleon-level calculations [386, 512].
 A rescaling of the curve was fitted to data, but only at 1.002 MeV center-of-mass to minimize the role of short-distance physics in the models.
-In 1998RvMP...70.1265A, a possible systematic discrepancy between data from the two methods previously used was discussed.
-By the time of 2011RvMP...83..195A, in-depth studies from two groups using both activation and prompt methods were available, from LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G] and Seattle [2007PhRvC..76e5801B] groups.
+In SF I, a possible systematic discrepancy between data from the two methods previously used was discussed.
+By the time of SF II, in-depth studies from two groups using both activation and prompt methods were available, from LUNA [109, 200, 321] and Seattle [149] groups.
 These studies did not find any discrepancy between activation and prompt- data in direct comparisons.
-2011RvMP...83..195A opted to limit the fitting to the data by the activation and recoil methods.
+SF II opted to limit the fitting to the data by the activation and recoil methods.
 The prompt- data were left out because of their somewhat larger common mode errors, concerns about how well the -ray angular distribution was known, and avoidance of the correlated errors between activation and prompt data from the same experiment.
-In 2011RvMP...83..195A, only data published after 1998 were included.
-The 2011RvMP...83..195A recommended value was , based on the data by the Weizmann [2004PhRvL..93z2503S], LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G], Seattle [2007PhRvC..76e5801B], and ERNA [2009PhRvL.102w2502D] groups.
 VI.2 Theory progress on
-Significant theoretical work on this reaction has occurred since 2011RvMP...83..195A, but the basic understanding of its mechanism remains unchanged from the 1960s [1963PhRv..131.2582T].
+Significant theoretical work on this reaction has occurred since SF II, but the basic understanding of its mechanism remains unchanged from the 1960s [623].
 It is dominated by external non-resonant capture into the two bound states, and most of the dipole strength at low energy arises beyond the range of nuclear interaction.
-The external capture part of the cross section is determined by the ANCs of the bound states and by scattering phase shifts; near threshold, most of the strength lies at –20 fm [2011PhRvL.106d2502N].
+The external capture part of the cross section is determined by the ANCs of the bound states and by scattering phase shifts; near threshold, most of the strength lies at –20 fm [507].
 In models with explicit wave functions the shorter-range strength largely cancels out due to effects of nucleon-exchange antisymmetry.
 All models feature a shallow minimum of near 1.25 MeV, where capture from -waves becomes comparable to that from -waves.
-The first fully ab initio calculation [2011PhRvL.106d2502N] appeared just after the 2011RvMP...83..195A analysis concluded; it used the fermionic molecular dynamics (FMD) method and a softened representation of the Argonne interaction [1995PhRvC..51...38W].
-It agrees well with both the scale and the energy dependence of the modern S-factor data, and it is very close to the energy dependence assumed in 2011RvMP...83..195A.
+The first fully ab initio calculation [507] appeared just after the SF II analysis concluded; it used the fermionic molecular dynamics (FMD) method and a softened representation of the Argonne interaction [661].
+It agrees well with both the scale and the energy dependence of the modern S-factor data, and it is very close to the energy dependence assumed in SF II.
 Neff’s elastic-scattering phase shifts also agree well with experiment.
 Notably, Neff found that the dipole strength distribution departs significantly from pure external capture at 3He-4He separations as large as 9 fm, compared with 3-4 fm in potential models.
-Another ab initio model [2016PhLB..757..430D, 2019PhRvC.100b4304V] has also appeared, based on the no-core shell model with continuum (NCSMC) method and the chiral interaction of 2003PhRvC..68d1001E, “softened” by the similarity renormalization group (SRG) procedure.
+Another ab initio model [248, 649] has also appeared, based on the no-core shell model with continuum (NCSMC) method and the chiral interaction of Entem and Machleidt [260], “softened” by the similarity renormalization group (SRG) procedure.
 Freedom to choose the SRG stopping point allowed exact reproduction of the 7Be breakup energy, which corrects for the main effect on the external capture of omitting the three-nucleon potential.
 The results agree well with the overall scale of the modern data, but their energy dependence and phase shifts depart from experiment, possibly due to the omission of explicit three-body forces.
-An important advance since 2011RvMP...83..195A has been the application of halo effective field theory (halo EFT) methods to astrophysical capture reactions [2002NuPhA.712...37B, 2018EPJA...54...89H, 2020EPJA...56..166P, 2020JPhG...47e4002Z].
+An important advance since SF II has been the application of halo effective field theory (halo EFT) methods to astrophysical capture reactions [121, 350, 539, 677].
 EFTs are valid for systems with a natural separation between the momentum scales probed in low-energy experiment (e.g., corresponding to a binding energy or a projectile energy) and the much larger momentum scale where the low-energy degrees of freedom are no longer valid: the word “halo” here refers to nuclei with only shallow bound states.
@@
 Coupling constants of the Lagrangian must be fitted to data, and consistency of the power counting scheme (identification of powers of for the main operator terms of a system) is tested by whether the fitted constants have natural sizes.
-Low-order coupling constants in a halo EFT can often be identified with familiar quantities like ANCs, scattering lengths, and effective ranges; the description of elastic scattering in halo EFT reproduces the (Coulomb-modified) effective-range expansion [2022JPhG...49d5102P, 2000NuPhA.665..137K, 2008NuPhA.809..171H].
+Low-order coupling constants in a halo EFT can often be identified with familiar quantities like ANCs, scattering lengths, and effective ranges; the description of elastic scattering in halo EFT reproduces the (Coulomb-modified) effective-range expansion [536, 404, 348].
 When a separation of scales exists, halo EFT is well-suited to data extrapolation because it avoids the tacit, hard-to-test, and unavoidable prior assumptions present in the short-range parts of models based on explicit wave functions.
-Two groups have studied in halo EFT [2018EPJA...54...89H, 2020EPJA...56..166P, 2020JPhG...47e4002Z].
+Two groups have studied in halo EFT [350, 539, 677].
 The sizes of the Coulomb interaction and the large -wave scattering length make the correct power counting tricky to establish for this system.
-The work of 2018EPJA...54...89H and 2020EPJA...56..166P includes a careful examination of possible power-counting schemes and strict adherence to a power counting once established, up to terms of NLO or NNLO, respectively, in two different power-countings.
+The work of Higa et al.
+[350] and Premarathna and Rupak [539] includes a careful examination of possible power-counting schemes and strict adherence to a power counting once established, up to terms of NLO or NNLO, respectively, in two different power-countings.
 These authors fitted EFTs to S-factor data, and they examined effects of including or excluding elastic scattering constraints in their fits.
-When they included scattering phase shifts from 1972NuPhA.195..241B in their fits, they concluded that the large-scattering-length power counting of the NLO theory was favored.
+When they included scattering phase shifts from Boykin et al.
+[140] in their fits, they concluded that the large-scattering-length power counting of the NLO theory was favored.
 This result promotes two-body currents (i.e., contributions not equivalent to external non-resonant capture) to leading order in the theory.
@@
 In addition to data-fitting errors, these theories have errors estimated to be 10% from EFT truncation.
-The work of 2020JPhG...47e4002Z followed a different approach to power counting up through NLO.
-At MeV, their derived expression for is essentially the same as that from the 2020EPJA...56..166P NLO theory.
-2020JPhG...47e4002Z also developed some ad hoc (i.e., not systematically developed) higher-order EFT terms, referred to as partial-N4LO, to test for their impact on the fitting.
+The work of Zhang et al.
+[677] followed a different approach to power counting up through NLO.
+At MeV, their derived expression for is essentially the same as that from the Premarathna and Rupak [539] NLO theory.
+Zhang et al.
+[677] also developed some ad hoc (i.e., not systematically developed) higher-order EFT terms, referred to as partial-N4LO, to test for their impact on the fitting.
 The additional terms proved not to be required by the data and did not improve the fit; this result was taken to indicate that corrections from omitted terms are not large compared with experimental errors below 2 MeV.
 Scattering data were not considered in this work apart from very broad priors on scattering length and effective range; the correlated errors in the S-factor data were taken into account.
-Like halo EFT, the phenomenological R-matrix approach avoids a model of nuclear interactions and uses a systematic parameterization to fit data [2010RPPh...73c6301D].
-At the time of 2011RvMP...83..195A, had been the subject of very little R-matrix fitting, apparently consisting only of the simple treatment in 2004ADNDT..88..203D that focused on BBN energies.
-After 2011RvMP...83..195A an R-matrix analysis using the AZURE2 code [2010PhRvC..81d5805A, azure2-manual] was carried out on both elastic scattering and S-factor data in conjunction with the Notre Dame experiment [2013PhRvC..87f5804K], and a more elaborate analysis using Monte Carlo sampling to estimate errors was later reported in 2014PhRvC..90c5804D.
-In the latter analysis both the fitted value and error bars were heavily influenced by the numerous scattering data of 1964NucPh..50..629B.
-Very recently, the BRICK software package has been constructed to carry out Bayesian parameter estimation for AZURE2 and applied to both capture and elastic scattering at all energies in the 7Be system by 2022FrP....10.8476O.
-In this work it was shown that markedly different values result from inclusion or not of the older scattering data [1964NucPh..50..629B] alongside the very recent SONIK scattering data of 2024PhRvC.109a5802P.
-A small amount of additional theoretical work in more traditional frameworks has appeared since 2011RvMP...83..195A.
-This includes fits of potential models to S-factor data [2018PhRvC..97c5802T, 2021NuPhA100622108T] and some calculations of the RGM type [2017PhRvC..96f4605S, 2019PhRvC..99e4618S].
+Like halo EFT, the phenomenological R-matrix approach avoids a model of nuclear interactions and uses a systematic parameterization to fit data [242].
+At the time of SF II, had been the subject of very little R-matrix fitting, apparently consisting only of the simple treatment in Descouvemont et al.
+[241] that focused on BBN energies.
+After SF II an R-matrix analysis using the AZURE2 code [62, 638] was carried out on both elastic scattering and S-factor data in conjunction with the Notre Dame experiment [407], and a more elaborate analysis using Monte Carlo sampling to estimate errors was later reported in deBoer et al.
+[234].
+In the latter analysis both the fitted value and error bars were heavily influenced by the numerous scattering data of Barnard et al.
+[94].
+Very recently, the BRICK software package has been constructed to carry out Bayesian parameter estimation for AZURE2 and applied to both capture and elastic scattering at all energies in the 7Be system by Odell et al.
+[516].
+In this work it was shown that markedly different values result from inclusion or not of the older scattering data [94] alongside the very recent SONIK scattering data of Paneru et al.
+[526].
 VI.3 Experimental progress on
-Since 2011RvMP...83..195A, five new experiments have been reported: Four from the Madrid and ATOMKI groups by the activation technique, at relatively high center of mass energies [2012PhRvC..86c2801C, 2013NuPhA.908....1B, 2019PhRvC..99e5804S, 2023PhRvC.108b5802T], and one from the Notre Dame group using the prompt- method [2013PhRvC..87f5804K].
-Following the approach adopted in 2011RvMP...83..195A, we again only use the recoil and activation data for the fits below.
-The -ray angular distribution is not known experimentally [2007PhRvC..76e5801B], and the resultant uncertainty increases the common-mode error for the prompt- studies somewhat.
-(An experiment to address previously-raised concerns about angular distribution [2007PhRvC..76e5801B] has recently concluded at Felsenkeller Dresden.
-The data suggest a higher than expected anisotropy but are so far only available in the form of a PhD thesis [Turkat23-PhD].
-If confirmed, they may lead to corrections of a few percent for some of the in-beam experiments where only one angle was instrumented.) In addition, all of the “modern” works reporting prompt- data except for Notre Dame [2013PhRvC..87f5804K] also include data obtained with other methods, leading to partial correlations between data sets that would complicate fitting.
+Since SF II, five new experiments have been reported: Four from the Madrid and ATOMKI groups by the activation technique, at relatively high center of mass energies [173, 134, 616, 624], and one from the Notre Dame group using the prompt- method [407].
+Following the approach adopted in SF II, we again only use the recoil and activation data for the fits below.
+The -ray angular distribution is not known experimentally [149], and the resultant uncertainty increases the common-mode error for the prompt- studies somewhat.
+(An experiment to address previously-raised concerns about angular distribution [149] has recently concluded at Felsenkeller Dresden.
+The data suggest a higher than expected anisotropy but are so far only available in the form of a PhD thesis [633].
+If confirmed, they may lead to corrections of a few percent for some of the in-beam experiments where only one angle was instrumented.) In addition, all of the “modern” works reporting prompt- data except for Notre Dame [407] also include data obtained with other methods, leading to partial correlations between data sets that would complicate fitting.
 Following this restriction, our data selection proceeds as follows.
-First, the four data sets previously used in 2011RvMP...83..195A are carried over here: Weizmann [2004PhRvL..93z2503S], LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G] (only the activation data), Seattle [2007PhRvC..76e5801B] (only the activation data), and ERNA [2009PhRvL.102w2502D].
-Two of the archival data sets excluded by the 1998 cutoff date in 2011RvMP...83..195A merit some further discussion here: The data of 1982PhRvL..48.1664O, 1984NuPhA.419..115O consist of two points measured by activation in a 3He gas cell.
+First, the four data sets previously used in SF II are carried over here: Weizmann [596], LUNA [109, 200, 321] (only the activation data), Seattle [149] (only the activation data), and ERNA [243].
+Two of the archival data sets excluded by the 1998 cutoff date in SF II merit some further discussion here: The data of Osborne et al.
+[520, 521] consist of two points measured by activation in a 3He gas cell.
 However, their uncertainties are not separated into statistical and systematic components.
 Including them would have required an uncertain guess on how to divide the error bars, and the fit result would barely change, so we left them out.
-The work by 1983PhRvC..27...11R consists of one data point that was obtained by averaging two separate activation measurements with a 3He and a 4He gas cell, respectively.
+The work by Robertson et al.
+[548] consists of one data point that was obtained by averaging two separate activation measurements with a 3He and a 4He gas cell, respectively.
 We do not use this result because we lack details to verify the background subtraction, an issue that was raised during our meeting by one of its authors.
-We also choose not to use the activation study by 1983ZPhyA.310...91V.
+We also choose not to use the activation study by Volk et al.
+[648].
 In that experiment there were thick entrance foils and the 4He beam was completely stopped inside a high-pressure 3He gas cell.
 This gave an integrated measurement over a wide energy range, so that the analysis depends strongly on the assumed shape of the curve, and the result was reported only as an extrapolated .
-We now consider the new data since 2011RvMP...83..195A.
-An activation experiment in Madrid [2012PhRvC..86c2801C] used a 3He beam incident on a 4He gas cell, reporting three data points.
-Another activation experiment was reported by the ATOMKI group, using a 4He beam on 3He gas cells [2013NuPhA.908....1B].
+We now consider the new data since SF II.
+An activation experiment in Madrid [173] used a 3He beam incident on a 4He gas cell, reporting three data points.
+Another activation experiment was reported by the ATOMKI group, using a 4He beam on 3He gas cells [134].
 Two higher-energy campaigns at ATOMKI were again performed using the activation method.
-These latter data are at = 2.5-4.4 MeV [2019PhRvC..99e5804S] and = 4.3-8.3 MeV [2023PhRvC.108b5802T], respectively, above the energy range suitable for halo EFT, and therefore not included.
-A detailed study by the Notre Dame group reported 17 data points [2013PhRvC..87f5804K], using the primary -ray from ground state capture and the secondary -ray from the deexcitation of the 429 keV first excited state of 7Be.
+These latter data are at = 2.5-4.4 MeV [616] and = 4.3-8.3 MeV [624], respectively, above the energy range suitable for halo EFT, and therefore not included.
+A detailed study by the Notre Dame group reported 17 data points [407], using the primary -ray from ground state capture and the secondary -ray from the deexcitation of the 429 keV first excited state of 7Be.
 No activation data are reported in that work, and the -ray detector was placed at just one angle, 90∘.
 Since only prompt- data are reported, we did not include this data set.
-In order to test the effects of this decision, we repeated some of our fits (see below, Section VI.4) with the modern prompt- experiments included: the 17 points from 2013PhRvC..87f5804K, 3 prompt- points from LUNA [2007PhRvC..75f5803C] and 8 from Seattle [2007PhRvC..76e5801B].
+In order to test the effects of this decision, we repeated some of our fits (see below, Section VI.4) with the modern prompt- experiments included: the 17 points from Kontos et al.
+[407], 3 prompt- points from LUNA [200] and 8 from Seattle [149].
 The extrapolated S-factor changed by less than 1%, well within the error bars for the recommended .
-Finally, an indirect experiment using the 6Li(3He,d)7Be reaction and the ANC technique has recently been reported [2020PhLB..80735606K].
+Finally, an indirect experiment using the 6Li(3He,d)7Be reaction and the ANC technique has recently been reported [403].
 It is left out of our fits due to the additional normalization and theory uncertainties involved in determining an ANC from a transfer experiment, which are larger than the errors in the S-factor measurements used for the present fits.
 All the data used in the present fits are summarized in Table 5.
-As in 2011RvMP...83..195A, we model data uncertainties as consisting of a component that is independent for each point and a common-mode component that applies to all data from a given experiment as a multiplicative factor.
+As in SF II, we model data uncertainties as consisting of a component that is independent for each point and a common-mode component that applies to all data from a given experiment as a multiplicative factor.
 This separation is well-documented for all of the modern data and is also shown in Table 5.
@@
 | range [keV] | used | total | mode uncert. | Higa EFT | Zhang EFT | BRICK | BRICK+S | |
-| Weizmann [2004PhRvL..93z2503S] | 4 | 4 | 1.03(2) | 1.02(2) | 1.02(2) | 1.03(2) | ||
-| LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G] | 7 | 7 | 1.02(2) | 1.04(2) | 1.05(2) | 1.01(2) | ||
-| Seattle [2007PhRvC..76e5801B] | 8 | 8 | 0.96(1) | 0.95(2) | 0.95(2) | 0.98(2) | ||
-| ERNA [2009PhRvL.102w2502D] | 27 | 27 | 0.96(3) | 0.94(2) | 0.96(2) | 0.99(2) | ||
-| Madrid [2012PhRvC..86c2801C] | 2 | 3 | 0.99(3) | 0.99(3) | 0.97(2) | 0.99(2) | ||
-| ATOMKI [2013NuPhA.908....1B] | 4 | 5 | 1.01(3) | 1.00(3) | 1.01(3) | 1.04(3) | ||
+| Weizmann [596] | 4 | 4 | 1.03(2) | 1.02(2) | 1.02(2) | 1.03(2) | ||
+| LUNA [109, 200, 321] | 7 | 7 | 1.02(2) | 1.04(2) | 1.05(2) | 1.01(2) | ||
+| Seattle [149] | 8 | 8 | 0.96(1) | 0.95(2) | 0.95(2) | 0.98(2) | ||
+| ERNA [243] | 22 | 51 | 0.96(3) | 0.94(2) | 0.96(2) | 0.99(2) | ||
+| Madrid [173] | 1 | 3 | 0.99(3) | 0.99(3) | 0.97(2) | 0.99(2) | ||
+| ATOMKI [134] | 2 | 5 | 1.01(3) | 1.00(3) | 1.01(3) | 1.04(3) | ||
 VI.4 Data fitting
@@
 These avoid tacit assumptions present in potential-models and conceptual difficulties involved in combining ab initio constraints with data.
-The fits presented here differ from the previously published fits of 2020EPJA...56..166P, 2020JPhG...47e4002Z, and 2022FrP....10.8476O mainly in the uniform use of the agreed-upon capture data and uncertainties from Section VI.3 across all fits.
+The fits presented here differ from the previously published fits of Premarathna and Rupak [539], Zhang et al.
+[677], and Odell et al.
+[516] mainly in the uniform use of the agreed-upon capture data and uncertainties from Section VI.3 across all fits.
 We restricted fitting to the MeV range of validity for the NLO halo EFT expressions.
 Despite uniform handling of capture data, the fits in each framework handle scattering inputs differently for reasons discussed below.
-In addition to the total capture cross section and scattering data, the fitted data also include branching ratios for capture into the two 7Be bound states, taken from 2007PhRvC..76e5801B, 2009PhRvL.102w2502D, 2007PhRvC..75f5803C, and 2013PhRvC..87f5804K.
+In addition to the total capture cross section and scattering data, the fitted data also include branching ratios for capture into the two 7Be bound states, taken from Brown et al.
+[149], Di Leva et al.
+[243], Confortola et al.
+[200], and Kontos et al.
+[407].
 These are necessarily from prompt- experiments and suffer from the concerns about angular distribution discussed in Section VI.3.
@@
 In constructing fits we split common-mode and point-to-point errors and “float the norms” of data sets using the cost function
-to describe goodness of fit [1994NIMPA.346..306D].
+to describe goodness of fit [215].
 A rescaling factor is fitted for each data set , with its deviation from unity penalized by the common-mode errors given in Table 5.
@@
 Two of the groups ran into serious difficulty with the frequentist fits due to parameter degeneracies and shallow local minima; during this study they were unable to produce frequentist fits in which they had confidence.
-We performed multiple fits in the NLO halo EFT of 2018EPJA...54...89H and 2020EPJA...56..166P.
+We performed multiple fits in the NLO halo EFT of Higa et al.
+[350] and Premarathna and Rupak [539].
 These fits were carried out both with and without scattering constraints, which mainly impact by removing parameter degeneracies that would otherwise leave the -wave scattering length poorly constrained.
 The code base for this version of halo EFT incorporates scattering data through phase shifts.
-These were taken from the partial-wave analysis of 1972NuPhA.195..241B, which at low energy are based mainly on the data of 1964NucPh..50..629B.
-One set of fits for the 2018EPJA...54...89H halo EFT was produced by minimization, proceeding in two steps: first LO parameters were fitted to keV data, and then the results were taken as initial values in the search to minimize over all parameters in the NLO theory for the full set of data.
+These were taken from the partial-wave analysis of Boykin et al.
+[140], which at low energy are based mainly on the data of Barnard et al.
+[94].
+One set of fits for the Higa et al.
+[350] halo EFT was produced by minimization, proceeding in two steps: first LO parameters were fitted to keV data, and then the results were taken as initial values in the search to minimize over all parameters in the NLO theory for the full set of data.
 These fits gave keV b with the scattering constraint and keV b without (uncertainties being propagated in linear approximation using covariances and partial derivatives).
@@
 However, any fitted curve is constrained by low-energy data, which in some sense become effectively renormalization conditions of the field theory; the error on extrapolated should probably be smaller than 10% by an amount that is hard to estimate.
-A second set of fits to the 2018EPJA...54...89H halo EFT was carried out using Bayesian methods.
+A second set of fits to the Higa et al.
+[350] halo EFT was carried out using Bayesian methods.
 Priors for the EFT parameters were developed based on previous experience, and data rescaling factors were incorporated as additional priors.
-The Bayesian results (including only experimental error) are keV b including the phase shifts by 1972NuPhA.195..241B and keV b excluding them, consistent with the -minimization.
-Searches of the parameter space to minimize for the NLO halo EFT of 2020JPhG...47e4002Z ran into difficulties with parameter degeneracy and local minima in the surface; for this formalism we report only Bayesian results.
+The Bayesian results (including only experimental error) are keV b including the phase shifts by Boykin et al.
+[140] and keV b excluding them, consistent with the -minimization.
+Searches of the parameter space to minimize for the NLO halo EFT of Zhang et al.
+[677] ran into difficulties with parameter degeneracy and local minima in the surface; for this formalism we report only Bayesian results.
 No experimental information about scattering was used, but flat priors on scattering length and effective range were chosen over a 5 range around recent experiment.
-This fit differs from 2020JPhG...47e4002Z mainly by excluding the Notre Dame data and including the ERNA activation data, and it gives keV b.
+This fit differs from Zhang et al.
+[677] mainly by excluding the Notre Dame data and including the ERNA activation data, and it gives keV b.
 The error associated with EFT truncation at NLO is estimated in this approach by separately fitting the partial-N4LO theory discussed above.
@@
 We also produced frequentist fits, but we were unable to estimate their errors convincingly.
-The R-matrix fits used capture data both alone and in combination with the SONIK scattering data [2024PhRvC.109a5802P] – the latter being chosen because of concerns with the 1964NucPh..50..629B data that are discussed in 2022FrP....10.8476O.
+The R-matrix fits used capture data both alone and in combination with the SONIK scattering data [526] – the latter being chosen because of concerns with the Barnard et al.
+[94] data that are discussed in Odell et al.
+[516].
 Since BRICK fits elastic differential cross sections directly, it was not feasible to use the same phase-shift-based scattering constraints as our halo EFT fits.
-The R-matrix fit that includes scattering data is essentially the “CS” fit of 2022FrP....10.8476O, but with a restriction to only the capture data described in Table 5, only scattering data below 2 MeV center-of-mass, and only the R-matrix parameters relevant below 2 MeV (no level or radiative widths for -wave background poles).
+The R-matrix fit that includes scattering data is essentially the “CS” fit of Odell et al.
+[516], but with a restriction to only the capture data described in Table 5, only scattering data below 2 MeV center-of-mass, and only the R-matrix parameters relevant below 2 MeV (no level or radiative widths for -wave background poles).
 | Fit | ||||
 | [keV b] | [MeV-1] | [fm] | [fm-1] | |
-| 2018EPJA...54...89H EFT-NLO | 0.554 | -0.56 | 30 | |
-| 2020JPhG...47e4002Z EFT-NLO | 0.573 | -0.61 | 46(6) | 29(3) |
+| Higa et al. [350] EFT-NLO | 0.554 | -0.56 | 30 | |
+| Zhang et al. [677] EFT-NLO | 0.573 | -0.61 | 46(6) | 29(3) |
 | BRICK R-matrix | 0.562 | -0.57 | 24.4(9) | |
 | BRICK R-matrix+scattering | 0.531 | -0.47 | 22.6(7) | |
-| ANC [2020PhLB..80735606K] |
+| ANC [403] |
 Several of the fits are compared in Fig.
@@
 There it is apparent that they all agree within 5% in the 0–2000 keV energy range, and within 4% in the astrophysically relevant range 0–500 keV.
-The fit to the 2018EPJA...54...89H NLO EFT lies in the middle of the range, so we adopt it as a reference in the lower panel of the figure.
-Relative to that fit, the fits to the NLO EFT of 2020JPhG...47e4002Z and the R-matrix fit without scattering rise up as much as 4% higher at 0–200 keV.
-The most striking difference among the curves is between the BRICK R-matrix fit including the SONIK scattering data [2024PhRvC.109a5802P] and all the other fits.
+The fit to the Higa et al.
+[350] NLO EFT lies in the middle of the range, so we adopt it as a reference in the lower panel of the figure.
+Relative to that fit, the fits to the NLO EFT of Zhang et al.
+[677] and the R-matrix fit without scattering rise up as much as 4% higher at 0–200 keV.
+The most striking difference among the curves is between the BRICK R-matrix fit including the SONIK scattering data [526] and all the other fits.
 While the R-matrix fit with scattering finds rescaling factors somewhat closer to unity than the other fits do, its energy dependence in the 0–1000 keV range is qualitatively different, as is especially visible in the lower panel of Fig.
-7 where residuals relative to the fitted 2018EPJA...54...89H EFT are shown.
+7 where residuals relative to the fitted Higa et al.
+[350] EFT are shown.
 The BRICK fit without the scattering data is much closer to all the halo EFT fits, so the recent scattering data apparently have a large impact.
 VI.5 Recommended value
-Because the fit to the NLO EFT of 2018EPJA...54...89H gives the central result among those attempted here, we adopt it for our recommended .
-For ease of adoption, it is noted that this recommended curve can be empirically parameterized using the same shape as in 2011RvMP...83..195A, by the numerical equation (following a customary form in the past literature [1987ApJ...319..531K, 2011RvMP...83..195A])
+Because the fit to the NLO EFT of Higa et al.
+[350] gives the central result among those attempted here, we adopt it for our recommended .
+For ease of adoption, it is noted that this recommended curve can be empirically parameterized using the same shape as in SF II, by the numerical equation (following a customary form in the past literature [387, 12])
 Here is the center of mass energy in MeV. Equation (VI.5) reproduces our recommended curve within 0.3% and is only applicable for = 0–1600 keV.
@@
 The two errors can be added in quadrature for applications.
-Our recommendation is essentially the same as in 2011RvMP...83..195A, but derived from completely different theoretical approaches that are much more directly connected to the data and much less dependent on model assumptions.
+Our recommendation is essentially the same as in SF II, but derived from completely different theoretical approaches that are much more directly connected to the data and much less dependent on model assumptions.
 It is also based on data over a wider energy range.
 The dispersion of and among models at the solar Gamow peak is illustrated in Table 6, along with some indication of how fitted scattering lengths and squared ANCs vary among frameworks.
-(The latter are highly degenerate with short-range parameters in the fitting; in the 2018EPJA...54...89H EFT they are sensitive to small variations in the -wave effective ranges that are fitted instead of ANCs.)
+(The latter are highly degenerate with short-range parameters in the fitting; in the Higa et al.
+[350] EFT they are sensitive to small variations in the -wave effective ranges that are fitted instead of ANCs.)
 VI.6 Recommendations for future work on
@@
 Our first recommendation for experiment concerns elastic 3He+4He scattering.
-New experiments over a wide energy range including at least part of the astrophysical range are needed to address the marked impact of the recent SONIK scattering data [2024PhRvC.109a5802P] on fitting, for independent confirmation.
+New experiments over a wide energy range including at least part of the astrophysical range are needed to address the marked impact of the recent SONIK scattering data [526] on fitting, for independent confirmation.
 Data below 2 MeV would be most useful for consistent fitting of halo EFTs.
-Second, new measurements (besides the preliminary data of Turkat23-PhD) are needed for the angular distribution of the emitted -rays.
-This would relieve concerns about including prompt- data, most notably the Notre Dame [2013PhRvC..87f5804K] that do not include activation measurements, in future fits.
+Second, new measurements (besides the preliminary data of Turkat [633]) are needed for the angular distribution of the emitted -rays.
+This would relieve concerns about including prompt- data, most notably the Notre Dame [407] that do not include activation measurements, in future fits.
 These experiments should also be designed to provide new angle-corrected branching ratios for capture into the two bound states.
@@
 Further exploration of how ab initio inputs like ANCs and scattering lengths could serve as fitting priors would also be a useful avenue of research.
-Authors of future theoretical calculations should consider including -ray angular distributions in their results; until very recently angular distributions were only available from two potential models [1981PhRvC..23...33K, 1963PhRv..131.2582T].
+Authors of future theoretical calculations should consider including -ray angular distributions in their results; until very recently angular distributions were only available from two potential models [398, 623].
 New calculations would inform analysis of experiments, and the angular distribution measurements we recommend would provide tests of the calculations (or fitting constraints).
@@
 Given the difficulties of constructing a unique phase shift analysis without an underlying theory, and the recent arrival of modern scattering data, theorists should consider including elastic differential cross sections directly as outputs and fitting constraints.
+In addition, efforts should be made to extend EFT fits to higher energies, in order to take into account the 3 MeV resonance.
 The hep reaction ()
@@
 As a consequence, similarly to the pp fusion, the hep reaction cross section is too small to be measured in laboratory and only theoretical predictions are available.
-The most recent studies of this reaction, already evaluated in 2011RvMP...83..195A, are those of 2000PhRvC..63a5801M and 2003PhRvC..67e5206P.
-The first one was performed within a phenomenological approach, similar to the calculation of 1998PhRvC..58.1263S for the pp fusion.
-In this particular case, the and initial and final nuclear wave functions were obtained with the correlated Hyperspherical Harmonics (CHH) variational method [1995FBS....18...25V, 1998PhRvL..81.1580V], using the Argonne (AV18) two-nucleon potential [1995PhRvC..51...38W] augmented by the Urbana IX (UIX) three-nucleon interaction [1995PhRvL..74.4396P].
+The most recent studies of this reaction, already evaluated in SF II, are those of Marcucci et al.
+[466] and Park et al.
+[529].
+The first one was performed within a phenomenological approach, similar to the calculation of Schiavilla et al.
+[574] for the pp fusion.
+In this particular case, the and initial and final nuclear wave functions were obtained with the correlated Hyperspherical Harmonics (CHH) variational method [645, 646], using the Argonne (AV18) two-nucleon potential [661] augmented by the Urbana IX (UIX) three-nucleon interaction [540].
 In this way, the binding energies of 3He and 4He, and the singlet and triplet scattering lengths were in good agreement with experiment.
 The weak vector and axial transition operators were obtained within a phenomonological approach.
-In particular, the largest two-body contribution to the axial current arising from the excitation of intermediate -isobar degrees of freedom was included using the transition-correlation operator scheme of 1991PhRvC..44..619C, 1992PhRvC..45.2628S and fixing the nucleon-to- axial coupling constant to reproduce as in the pp case [1998PhRvC..58.1263S].
-The second study of 2003PhRvC..67e5206P was performed within the same hybrid EFT approach used for the pp funsion (EFT∗), with the same nuclear weak current operators obtained within EFT as discussed in Section III, the LEC being fixed to reproduce .
-The initial and final wave functions were derived using the CHH method with the AV18/UIX potential, as in 2000PhRvC..63a5801M.
-Following the results of 2000PhRvC..63a5801M and 2003PhRvC..67e5206P, 2011RvMP...83..195A recommended for the zero-energy S-factor, , the value keV b.
+In particular, the largest two-body contribution to the axial current arising from the excitation of intermediate -isobar degrees of freedom was included using the transition-correlation operator scheme of Carlson et al.
+[171], Schiavilla et al.
+[575] and fixing the nucleon-to- axial coupling constant to reproduce as in the pp case [574].
+The second study of Park et al.
+[529] was performed within the same hybrid EFT approach used for the pp funsion (EFT∗), with the same nuclear weak current operators obtained within EFT as discussed in Section III, the LEC being fixed to reproduce .
+The initial and final wave functions were derived using the CHH method with the AV18/UIX potential, as in Marcucci et al.
+[466].
+Following the results of Marcucci et al.
+[466] and Park et al.
+[529], SF II recommended for the zero-energy S-factor, , the value keV b.
 Three recommendations were also made: (i) to perform new studies with a broad spectrum of Hamiltonian models, in order to properly access the theoretical uncertainty; (ii) to study other weak reactions for which experiments can test the predictions made by employing the same theoretical ingredients as those used for the hep reaction calculation; (iii) to further understand the relation between the hep reaction and the so-called hen process ().
-Of these recommended works, only the second one has been performed so far, applying the theoretical framework used for the hep study to the muon capture reactions on light nuclei [2011PhRvC..83a4002M].
-Given the lack of further studies on the hep reaction after 2011RvMP...83..195A, we decide to maintain
+Of these recommended works, only the second one has been performed so far, applying the theoretical framework used for the hep study to the muon capture reactions on light nuclei [463].
+Given the lack of further studies on the hep reaction after SF II, we decide to maintain
 as recommended value.
@@
 However, we have decided not to apply such an increase, because it does not derive from a direct calculation, and, most of all, because such an increase lies well within the quoted % uncertainty.
-However, given the advances in ab initio studies of the last decade and in the development of the EFT framework, we update the recommendations of 2011RvMP...83..195A as follows: (i) the hep process should be revisited within a fully consistent EFT approach, similarly to what has been done in these years for the pp fusion and other processes, as, for instance, muon captures on light nuclei [2012PhRvL.108e2502M, 2018PhRvL.121d9901M, 2018PhRvC..98f5506A, 2023FrP....1049919C, 2023PhRvC.107f5502B, PhysRevC.109.035502]; (ii) the initial and final wave functions should be calculated with the most accurate ab initio methods, as the (uncorrelated) HH method [2008JPhG...35f3101K, 2020FrP.....8...69M], which has been proven to provide more accurate bound and scattering wave functions than the CHH method, using both local and non-local interactions; (iii) the relation between the hep and the hen process should be understood.
-The hen process has been studied within EFT and the HH method by 2022PhRvC.105a4001V, but a consistent parallel study of both reactions is still missing.
-A relatively recent SNO Collaboration analysis of data from all three of the detector’s running phases yielded a one-sided confidence-interval bound on the hep neutrino flux of 2020PhRvD.102f2006A
+However, given the advances in ab initio studies of the last decade and in the development of the EFT framework, we update the recommendations of SF II as follows: (i) the hep process should be revisited within a fully consistent EFT approach, similarly to what has been done in these years for the pp fusion and other processes, as, for instance, muon captures on light nuclei [460, 461, 7, 183, 133, 303]; (ii) the initial and final wave functions should be calculated with the most accurate ab initio methods, as the uncorrelated hyperspherical harmonic (HH) method [394, 459], which has been proven to provide more accurate bound and scattering wave functions than the CHH method, using both local and non-local interactions; (iii) the relation between the hep and the hen process should be understood.
+The hen process has been studied within EFT and the HH method by Viviani et al.
+[644], but a consistent parallel study of both reactions is still missing.
+A relatively recent SNO Collaboration analysis of data from all three of the detector’s running phases yielded a one-sided confidence-interval bound on the hep neutrino flux of Aharmim et al. [18]
 a result in agreement with the SSM prediction.
-The lack of a definite measurement, however, has meant that there is no substantive experimental test of the predicted Various global analyses of solar neutrino data have provided weak evidence for a nonzero flux, e.g., from Bergstrom:2016cbh
+The lack of a definite measurement, however, has meant that there is no substantive experimental test of the predicted Various global analyses of solar neutrino data have provided weak evidence for a nonzero flux, e.g., from Bergstrom et al.
+[113]
 which is consistent with the experiment bound given above.
@@
 Therefore, the ratio between the pp and pep rates is independent of nuclear physics.
-Based on this consideration, in 2011RvMP...83..195A the result of 1998RvMP...70.1265A, which is based on the work of 1969ApJ...155..501B, was multiplied by the radiative corrections calculated by 2003PhRvC..67c5502K, leading to the final result given by
-where is the density in units of g/cm3, is the mean molecular weight per free electron, and is the temperature in units of MK.
+Based on this consideration, in SF II the result of SF I, which is based on the work of Bahcall and May [78], was multiplied by the radiative corrections calculated by Kurylov et al.
+[419], leading to the final result given by
+where is the density in units of g/cm3, is the mean molecular weight per free electron, and is the temperature in units of megakelvins.
 The range of validity is .
-Given that no new evaluation of the pep rate has been performed since 2011RvMP...83..195A, the result of Eq.
+Given that no new evaluation of the pep rate has been performed since SF II, the result of Eq.
 (20) represents also the present recommended value.
 Competition between electron and proton capture on 7Be fixes the branching ratio of the ppII and ppIII chains, and thereby the 7Be and 8B neutrino fluxes.
-Superallowed electron capture on the ground state of 7Be leads to the ground state or the , 478 keV first excited state of 7Li with a measured (terrestrial) branching ratio of 10.44(4)% [2002NuPhA.708....3T].
+Superallowed electron capture on the ground state of 7Be leads to the ground state or the , 478 keV first excited state of 7Li with a measured (terrestrial) branching ratio of 10.44(4)% [620].
 It is customary to calculate the solar decay rate in terms of the measured terrestrial decay rate, which has the advantage of removing the nuclear physics dependence, but it requires a proper calculation of the densities of electrons at the nucleus.
-This has usually been done by considering continuum and bound electrons separately, using the Debye-Hückel (DH) approximation pioneered by 1954AuJPh...7..373S for the screening of bound electrons in the solar plasma; the capture rate of continuum electrons is not significantly influenced by electron screening.
+This has usually been done by considering continuum and bound electrons separately, using the Debye-Hückel approximation pioneered by Salpeter [564] for the screening of bound electrons in the solar plasma; the capture rate of continuum electrons is not significantly influenced by electron screening.
 Furthermore, it is necessary to calculate the atomic probability densities governing the and terrestrial electron capture rates.
-2020PhRvL.125c2701F recently measured the ratio of to capture using superconducting tunnel junctions.
-The recommendation of both 1998RvMP...70.1265A and 2011RvMP...83..195A was
+Fretwell et al.
+[282] recently measured the ratio of to capture using superconducting tunnel junctions.
+The recommendation of both SF I and SF II was
 which is valid for .
-This expression is based on the continuum capture rate calculated by 1969ApJ...155..511B, the average ratio of the total capture rate to the continuum rate calculated using three solar models in 1994PhRvD..49.3923B, and the current value of the half-life, 53.22(6) d [2002NuPhA.708....3T].
-It agrees within 1% with a density matrix calculation of the rate that makes no assumptions regarding the nature of electronic quantum states in the solar plasma and allows for aspherical fluctuations in the spatial distributions of plasma ions [1997ApJ...490..437G].
-The estimated uncertainty of 2% accounts for possible corrections to the Debye-Hückel approximation due to thermal fluctuations in the small number of ions in the Debye sphere, and breakdowns in the adiabatic approximation [1992ApJ...392..320J].
-Although calculations have appeared in the literature claiming new plasma effects based on model calculations (see 2013ApJ...764..118S, 2019A&A...623A.126V), we continue to regard the arguments in 2002A&A...383..291B as definitive.
+This expression is based on the continuum capture rate calculated by Bahcall and Moeller [79], the average ratio of the total capture rate to the continuum rate calculated using three solar models in Bahcall [70], and the current value of the half-life, 53.22(6) d [620].
+It agrees within 1% with a density matrix calculation of the rate that makes no assumptions regarding the nature of electronic quantum states in the solar plasma and allows for aspherical fluctuations in the spatial distributions of plasma ions [316].
+The estimated uncertainty of 2% accounts for possible corrections to the Debye-Hückel approximation due to thermal fluctuations in the small number of ions in the Debye sphere, and breakdowns in the adiabatic approximation [381].
+Although calculations have appeared in the literature claiming new plasma effects based on model calculations (see Simonucci et al.
+[594], Vescovi et al.
+[640]), we continue to regard the arguments in Bahcall et al.
+[74] as definitive.
 That reference presented five distinct derivations demonstrating that the Salpeter formula for screening corrections is valid for pp-chain reactions and solar conditions, up to corrections on the order of a few percent.
@@
 Radiative proton capture on (=3/2-) at solar energies proceeds through non-resonant capture to the ground state of (=2+).
-This capture occurs predominantly at separations well beyond the range of the strong interaction via transitions from and partial waves, as the amplitude and contributions from higher partial waves are negligible in the energy range of interest [1961NucPh..24...89C].
-The importance of this reaction for the determination of the high energy solar neutrino spectrum and the experimental data reviewed by 1998RvMP...70.1265A inspired several experiments that, taken together, provided a consistent picture of the energy dependence of the reaction cross section, despite discrepancies in absolute scale.
-Prior to the review of 2011RvMP...83..195A, most direct measurements were performed using a radioactive target and an intense proton beam.
-In spite of the great care taken in evaluating the systematic energy-independent uncertainties, in the following referred to as common-mode errors (CMEs), a remarkable discrepancy persisted even in the most recent experiments, that was finally handled in 2011RvMP...83..195A by inflating the stated experimental uncertainties.
-Indirect measurements of the cross section of B based on Coulomb dissociation were considered, but not included in the final recommendation of 2011RvMP...83..195A for the zero-energy astrophysical S-factor .
-However, despite the experimental issues, the uncertainty on the recommended value of in 2011RvMP...83..195A is dominated by the theoretical contribution.
+This capture occurs predominantly at separations well beyond the range of the strong interaction via transitions from and partial waves, as the amplitude and contributions from higher partial waves are negligible in the energy range of interest [195].
+The importance of this reaction for the determination of the high energy solar neutrino spectrum and the experimental data reviewed by SF I inspired several experiments that, taken together, provided a consistent picture of the energy dependence of the reaction cross section, despite discrepancies in absolute scale.
+Prior to the review of SF II, most direct measurements were performed using a radioactive target and an intense proton beam.
+In spite of the great care taken in evaluating the systematic energy-independent uncertainties, in the following referred to as common-mode errors (CMEs), a remarkable discrepancy persisted even in the most recent experiments, that was finally handled in SF II by inflating the stated experimental uncertainties.
+Indirect measurements of the cross section of B based on Coulomb dissociation were considered, but not included in the final recommendation of SF II for the zero-energy astrophysical S-factor .
+However, despite the experimental issues, the uncertainty on the recommended value of in SF II is dominated by the theoretical contribution.
 Therefore, our analysis focused on these aspects (discrepancy of experimental data, indirect measurements, theoretical uncertainties), as discussed in the next sections.
@@
 IX.2.1 Direct Measurements
-The thorough review of published work done in 2011RvMP...83..195A investigated the influence of beam-target overlap, target stoichiometry, beam energy loss, and the backscattering of 8B recoils on CMEs and led to the selection of a homogeneous group of well-documented data sets.
+The thorough review of published work done in SF II investigated the influence of beam-target overlap, target stoichiometry, beam energy loss, and the backscattering of 8B recoils on CMEs and led to the selection of a homogeneous group of well-documented data sets.
 Since then, no further information became available on these aspects of direct measurements.
-The data that formed the basis of the 2011RvMP...83..195A recommendation were those of 2003PhRvC..67f5805B, 2003PhRvL..90b2501B, 1998PhRvL..80..928H, 2001PhRvL..86.3985H, 1999PhLB..462..237H, 2001NuPhA.696..219S, and the BE3 data set of 2003PhRvC..68f5803J.
-We considered the complete set of radiative capture measurements that were the basis of the 2011RvMP...83..195A recommendation for , where we adopted the revised BE3 data presented by 2010PhRvC..81a2801J, which were reanalyzed using a more sophisticated model of the target and more accurate stopping powers.
-In addition, we included in our analysis the only new experiment reported thereafter of 2022PhLB..82436819B, which used a radioactive ion beam and the recoil mass separator ERNA to detect the 8B recoils.
-While these new results contribute to the determination of with different, well controlled systematics [2018EPJA...54...92B], their large statistical uncertainty relative to the other experiments, due to the low counting statistics, limits their impact.
+The data that formed the basis of the SF II recommendation were those of Baby et al.
+[63], Baby et al.
+[64], Hammache et al.
+[332], Hammache et al.
+[333], Hass [338], Strieder et al.
+[612], and the BE3 data set of Junghans et al.
+[382].
+We considered the complete set of radiative capture measurements that were the basis of the SF II recommendation for , where we adopted the revised BE3 data presented by Junghans et al.
+[383], which were reanalyzed using a more sophisticated model of the target and more accurate stopping powers.
+In addition, we included in our analysis the only new experiment reported thereafter of Buompane et al.
+[162], which used a radioactive ion beam and the recoil mass separator ERNA to detect the 8B recoils.
+While these new results contribute to the determination of with different, well controlled systematics [161], their large statistical uncertainty relative to the other experiments, due to the low counting statistics, limits their impact.
 New scattering data have been published.
-Elastic and inelastic scattering cross sections at relative kinetic energies between 474 keV and 2.74 MeV were measured and wave scattering lengths were inferred by 2019PhRvC..99d5807P.
-At higher energies between 1.6 and 3.4 MeV, thick target elastic and inelastic scattering excitation function measurements were used to infer the existence of new resonances above 1.8 MeV [2013PhRvC..87e4617M].
+Elastic and inelastic scattering cross sections at relative kinetic energies between 474 keV and 2.74 MeV were measured and wave scattering lengths were inferred by Paneru et al.
+[527].
+At higher energies between 1.6 and 3.4 MeV, thick target elastic and inelastic scattering excitation function measurements were used to infer the existence of new resonances above 1.8 MeV [479].
 These high lying resonances have no direct influence on because of their small partial widths in the channel and/or because they cannot be populated by s- or d-wave capture at low energy.
 IX.2.2 Indirect Measurements
-In 2011RvMP...83..195A, Coulomb dissociation measurements based on the formalism developed by 1986NuPhA.458..188B were considered [PhysRevLett.83.2910, PhysRevLett.73.2680, KIKUCHI1997261, 1998EPJA....3..213K, PhysRevLett.86.2750, PhysRevC.63.065806, PhysRevLett.90.232501, 2006PhRvC..73a5806S], where in some cases the most recent publications supersede the previously published ones.
+In SF II, Coulomb dissociation measurements based on the formalism developed by Baur et al.
+[99] were considered [379, 491, 396, 395, 220, 221, 578, 579], where in some cases the most recent publications supersede the previously published ones.
 Finally, the results of these experiments were not included in the determination of due to disagreements over whether the analyses of the various measurements properly accounted for the contributions of transitions.
@@
 The common mode errors are obtained from a linear sum of those given in the manuscripts and an additional common mode error due to the estimated size of the component in each measurement.
-This additional error is applicable in one direction only, downward for the 1998EPJA....3..213K and 2006PhRvC..73a5806S analyses of the RIKEN and GSI measurements, respectively, and upward for the 2003PhRvC..68d5802D analysis of the Michigan State University (MSU) measurement, according to whether the component was included or excluded.
-Though the 2003PhRvC..68d5802D analysis did include a common mode error contribution of 2.5% due to the component (which represented 5% of the total cross section), we have included an additional 2.5% contribution to the MSU experiment.
-The size of the additional common mode errors for the RIKEN and GSI measurements was based on scaling this 2.5% according to the predicted size of the contribution to the cross section in these measurements relative to that in the MSU measurement, evaluated using first-order perturbation theory and the and S-factors calculated in a potential model of 8B [1996ZPhyA.356..293B].
+This additional error is applicable in one direction only, downward for the Kikuchi et al.
+[395] and Schümann et al.
+[579] analyses of the RIKEN and GSI measurements, respectively, and upward for the Davids and Typel [222] analysis of the Michigan State University (MSU) measurement, according to whether the component was included or excluded.
+Though the Davids and Typel [222] analysis did include a common mode error contribution of 2.5% due to the component (which represented 5% of the total cross section), we have included an additional 2.5% contribution to the MSU experiment.
+The size of the additional common mode errors for the RIKEN and GSI measurements was based on scaling this 2.5% according to the predicted size of the contribution to the cross section in these measurements relative to that in the MSU measurement, evaluated using first-order perturbation theory and the and S-factors calculated in a potential model of 8B [118].
 Thus the additional single-sided common mode errors added to the GSI, MSU, and RIKEN measurements are -2.1%, 2.5%, and -4.3%, respectively.
 IX.3 Theory
-The size of the theoretical contribution to the uncertainty in the 2011RvMP...83..195A estimate of was twice that of the experimental contribution.
+The size of the theoretical contribution to the uncertainty in the SF II estimate of was twice that of the experimental contribution.
 Since then, significant progress has been made.
-A new potential model calculation was performed by 2019NuPhA.983..175D.
+A new potential model calculation was performed by Dubovichenko et al.
+[249].
 The large-scale computational demands of calculations using high-quality nucleon-nucleon (NN) and three-nucleon (3N) interactions beyond =5 had hindered the ab initio approach in the past.
 Nevertheless progress has been rapid and it has been possible to extend ab initio calculations to B, albeit without 3N forces.
-2011PhLB..704..379N performed such a calculation using the no-core shell model in the continuum with a nucleon-nucleon interaction derived from chiral EFT at fourth order.
+Navrátil et al.
+[506] performed such a calculation using the no-core shell model in the continuum with a nucleon-nucleon interaction derived from chiral EFT at fourth order.
 The calculation converged with respect to the model space size and reproduces experimental data for the S-factor despite having virtually no free parameters (though the similarity renormalization group evolution parameter was tuned to reproduce the binding energy of 8B).
-The final value of =19.4(7) eV b agrees with the 2011RvMP...83..195A recommendation and, as will be shown later, with our analysis.
-More recently, a new approach was introduced to combine experimental measurements with ab initio predictions, resulting in an “ab initio-informed evaluation” Kravvaris23-PLB that arrived at a value of eV b.
-A significant theoretical advance was the application of the formalism of halo EFT developed by 2002NuPhA.712...37B to 7Be(B calculations by 2015PhLB..751..535Z, 2018PhRvC..98c4616Z, and 2022PhRvC.106a4601H.
+The final value of =19.4(7) eV b agrees with the SF II recommendation and, as will be shown later, with our analysis.
+More recently, a new approach was introduced to combine experimental measurements with ab initio predictions, resulting in an “ab initio-informed evaluation” Kravvaris et al.
+[409] that arrived at a value of eV b.
+A significant theoretical advance was the application of the formalism of halo EFT developed by Bertulani et al.
+[121] to 7Be(B calculations by Zhang et al.
+[675], Zhang et al.
+[676], and Higa et al.
+[349].
 It treats the incoming and 7Be as point-like particles and the 8B final state as a bound state of the two.
 The calculation and the associated errors depend upon the accuracy with which one is able to describe the incoming scattering states, the final bound state and the relevant electromagnetic currents.
-The relative kinetic or CM energy range MeV, corresponding to a relative momentum MeV/c, is considered within the domain of applicability of halo EFT.
+The relative kinetic or c.m.
+energy range MeV, corresponding to a relative momentum MeV/c, is considered within the domain of applicability of halo EFT.
 The two cluster physical description of is expected to hold at 3He- relative momenta below a physical cutoff MeV/c set by the threshold (binding momentum) for breaking the 7Be core into these constituents.
 Further, at energies above the excitation energy of the first excited state, MeV, including the contribution of the excited core is imperative in the EFT.
-2015PhLB..751..535Z and 2018PhRvC..98c4616Z include such an excited contribution.
-2022PhRvC.106a4601H provides a calculation without the contribution below and one with it above that also includes an contribution to the resonance.
-The two groups’ halo EFT calculations with the contributions differ in several respects that are elaborated in 2022PhRvC.106a4601H.
+Zhang et al.
+[675] and Zhang et al.
+[676] include such an excited contribution.
+Higa et al.
+[349] provides a calculation without the contribution below and one with it above that also includes an contribution to the resonance.
+The two groups’ halo EFT calculations with the contributions differ in several respects that are elaborated in Higa et al.
+[349].
 Differences in the final bound state calculation affect the treatment of the short-range interactions and divergences in the EFTs.
@@
 Thus there is no effect on the cross section.
-In the incoming channel, 2022PhRvC.106a4601H use a more general form of the -wave short-ranged interaction involving the core based on the low-energy symmetry that is included only at NNLO.
-In contrast, 2015PhLB..751..535Z and 2018PhRvC..98c4616Z include a short-ranged interaction at LO that could make a difference in the momentum dependence of the cross section.
-However, as shown in an order-by-order calculation [2022PhRvC.106a4601H], the short-ranged interaction is a sub-leading effect so the two halo EFTs should have similar low-momentum dependence.
+In the incoming channel, Higa et al.
+[349] use a more general form of the -wave short-ranged interaction involving the core based on the low-energy symmetry that is included only at NNLO.
+In contrast, Zhang et al.
+[675] and Zhang et al.
+[676] include a short-ranged interaction at LO that could make a difference in the momentum dependence of the cross section.
+However, as shown in an order-by-order calculation [349], the short-ranged interaction is a sub-leading effect so the two halo EFTs should have similar low-momentum dependence.
 The third important difference is in the treatment of two-body currents, which affects EFT error estimates.
-2015PhLB..751..535Z and 2018PhRvC..98c4616Z include this at NLO whereas 2022PhRvC.106a4601H estimate it to be an N3LO effect.
-Bayesian estimates of the two-body current [2015PhLB..751..535Z, 2018PhRvC..98c4616Z] are not strongly constrained by capture data, consistent with this being a higher-order contribution.
+Zhang et al.
+[675] and Zhang et al.
+[676] include this at NLO whereas Higa et al.
+[349] estimate it to be an N3LO effect.
+Bayesian estimates of the two-body current [675, 676] are not strongly constrained by capture data, consistent with this being a higher-order contribution.
 Another widely used approach for describing the cross sections of low-energy nuclear reactions is R-matrix theory.
@@
 The parameters of the real pole corresponding to the bound state are determined by the binding energy and the two ANCs, while the 1+ resonance parameters are determined by its energy and partial decay widths (both proton and radiative).
-The R-matrix analysis of this reaction carried out by 1995NuPhA.588..693B arrived at a value of eV b; however, it was based on data that were excluded in 2011RvMP...83..195A and here, with the exception of those of 1983PhRvC..28.2222F.
+The R-matrix analysis of this reaction carried out by Barker [91] arrived at a value of eV b; however, it was based on data that were excluded in SF II and here, with the exception of those of Filippone et al.
+[270].
 We fit data using R-matrix theory and halo EFTs, with both frequentist and Bayesian approaches, investigating the impact of different choices for the energy range in which data were fitted, and obtained consistent results.
-Our recommended value and its uncertainty are based on a Bayesian fit of both direct and Coulomb dissociation data taken at energies using the 2022PhRvC.106a4601H halo EFT⋆.
+Our recommended value and its uncertainty are based on a Bayesian fit of both direct and Coulomb dissociation data taken at energies using the Higa et al.
+[349] halo EFT⋆.
 This theory provides a good description of capture data over a wide energy range including the contribution from the resonance.
@@
 IX.4 Fitting Procedures
-In 2011RvMP...83..195A, it was noted that the different data sets had similar energy dependences.
+In SF II, it was noted that the different data sets had similar energy dependences.
 The discrepancies among the data sets were primarily due to different absolute normalizations.
 We address the normalization issue by introducing scaling factors associated with the CMEs.
-This contrasts with the 2011RvMP...83..195A approach, in which the errors were inflated by the factor .
+This contrasts with the SF II approach, in which the errors were inflated by the factor .
 The procedure for handling discrepant data sets adopted here is mathematically rigorous and aligns naturally with the method of estimating theoretical uncertainty discussed below.
-We prefer this approach to the inflation factor method used in 2011RvMP...83..195A, since the absolute scales of the different data sets can be used as constraints for each other, while there is no information available for a reevaluation of the CMEs.
+We prefer this approach to the inflation factor method used in SF II, since the absolute scales of the different data sets can be used as constraints for each other, while there is no information available for a reevaluation of the CMEs.
 The procedure we used, described below, can be applied to both frequentist and Bayesian analyses.
@@
 The likelihood function, defined as the conditional probability for the data given theoretical parameters , enters both fitting procedures.
-It can be derived from the expression given in 1994NIMPA.346..306D
+It can be derived from the expression given in D’Agostini [215]
 with labeling the different experimental data sets and labeling the individual data points of each experiment.
 In the second term, we divide theoretical predictions for the data points of the experiment by a scaling factor and consider the combination as our full model for , the th data point of experiment with point-to-point uncertainty .
-In the first term, we assign a Gaussian distribution centered at with a width equal to the CME ; for the GSI, MSU and RIKEN data shown in Table LABEL:table:normsBayesian, asymmetric Gaussians are used with depending on the sign of .
+In the first term, we assign a Gaussian distribution centered at with a width equal to the CME ; for the GSI, MSU and RIKEN data shown in Table 8, asymmetric Gaussians are used with depending on the sign of .
 In our minimization, the definition is derived from equating the full exponents in Eq.
@@
 (22).
-However, since the GSI, MSU and RIKEN data have asymmetric CMEs (see Table LABEL:table:normsBayesian), we draw priors from a uniform distribution spanning 50% to 150% of the central values for these that are then used in the asymmetric normalization exponentials in the likelihood function.
+However, since the GSI, MSU and RIKEN data have asymmetric CMEs, we draw priors from a uniform distribution spanning 50% to 150% of the central values for these that are then used in the asymmetric normalization exponentials in the likelihood function.
 This was computationally simpler than generating asymmetric Gaussian priors that span the entire range, which is neither necessary nor physical for a normalization constant.
@@
 Fits with and without Coulomb dissociation data were performed that resulted in overlapping determinations within the estimated fitting uncertainties.
-The different low energy fits included one using the halo EFT from 2018PhRvC..98c4616Z and one using the halo EFT⋆ from 2022PhRvC.106a4601H, both including the excited contribution.
+The different low energy fits included one using the halo EFT from Zhang et al.
+[676] and one using the halo EFT⋆ from Higa et al.
+[349], both including the excited contribution.
 The simpler halo EFT expressions, not including an excited component , were also fitted in this low energy region with comparable results.
-For the fits in the other two energy ranges, halo EFT expressions with excited contributions from both 2018PhRvC..98c4616Z and 2022PhRvC.106a4601H were used.
-The 2018PhRvC..98c4616Z halo EFT expression, which does not include an contribution, was supplemented with a Breit-Wigner resonance [1988ccna.book.....R] with its energy (around the resonance energy ), width, and proton-decay branching ratio determined by the fits.
-An R-matrix model was fitted to data in the three energy regions with the AZURE2 code [2010PhRvC..81d5805A], again resulting in an consistent with the halo EFT values.
+For the fits in the other two energy ranges, halo EFT expressions with excited contributions from both Zhang et al.
+[676] and Higa et al.
+[349] were used.
+The Zhang et al.
+[676] halo EFT expression, which does not include an contribution, was supplemented with a Breit-Wigner resonance [557] with its energy (around the resonance energy ), width, and proton-decay branching ratio determined by the fits.
+An R-matrix model was fitted to data in the three energy regions with the AZURE2 code [62], again resulting in an consistent with the halo EFT values.
 The results from the various fits are included in the supplemental material.
-Bayesian fits were performed with halo EFT expressions from both 2020JPhG...47e4002Z and 2018EPJA...54...89H that are compatible with the various fits.
-The Bayesian analysis had advantages over the fits in exploring parameter space and in determining realistic uncertainties [PhysRevLett.122.232502].
+Bayesian fits were performed with halo EFT expressions from both Zhang et al.
+[677] and Higa et al.
+[350] that are compatible with the various fits.
+The Bayesian analysis had advantages over the fits in exploring parameter space and in determining realistic uncertainties [399].
 The capture proceeds through the and channels.
-The reaction is known to be peripheral, resulting in a that has a sub-leading dependence on the strong interaction in the initial state [2000PhRvC..62f5803B].
+The reaction is known to be peripheral, resulting in a that has a sub-leading dependence on the strong interaction in the initial state [100].
 This behavior was shown to persist away from the threshold in halo EFT, in which for capture the strong interaction in the incoming -wave channel only contributes at NNLO.
@@
 R-matrix, EFT, and EFT⋆ fits confirm this behaviour.
-On the other hand, the ANCs have been extracted experimentally [2003PhRvC..67f2801T, 2006PhRvC..73b5808T] and also theoretically [2011PhRvC..83d1001N, 2018PhRvC..98c4616Z] in ab initio calculations.
+On the other hand, the ANCs have been extracted experimentally [625, 617] and also theoretically [513, 676] in ab initio calculations.
 Although the determination of is insensitive to the relative contributions of the two spin channels, knowledge of the relative contributions of the ANCs is important in the EFT framework.
@@
 The priors for the normalization constants were as described earlier.
-The priors for the theoretical parameters are based on the underlying EFT assumptions [2022PhRvC.106a4601H] about the expansion in a small ratio where represents the low momentum of infrared physics and the cutoff of the theory.
-The and scattering lengths fm and fm [2019PhRvC..99d5807P] are taken as inputs.
+The priors for the theoretical parameters are based on the underlying EFT assumptions [349] about the expansion in a small ratio where represents the low momentum of infrared physics and the cutoff of the theory.
+The and scattering lengths fm and fm [527] are taken as inputs.
 For capture at LO, the only fit parameter, the effective momentum is assumed to scale as and accordingly we assume a uniform prior for between -100 MeV/ and 1.5 MeV/, where the upper limit is set by the physical constraint that the ANC squares have to be positive.
 At NLO, the ANC2 is a fit parameter whose prior is taken from a normal distribution determined by experiment.
-At NNLO, we use the ab initio calculation of [2018PhRvC..98c4616Z] to draw the ANC2 from a normal distribution.
+At NNLO, we use the ab initio calculation of [676] to draw the ANC2 from a normal distribution.
 A ratio of the -wave scattering lengths in the coupled-channel - that is assumed to be in the EFT is drawn from a normal distribution with a mean of 0 and a standard deviation of 10 to cover a wide range.
 The capture is dominated by the transition and it requires three parameters: a -wave scattering volume , an effective momentum and a two-body current to regulate divergences.
-2022PhRvC.106a4601H determined and from the the narrow resonance energy and width.
+Higa et al.
+[349] determined and from the the narrow resonance energy and width.
 Here, we kept fixed while drawing from a uniform distribution between -150 and -50 MeV.
@@
 The two-body coupling was drawn from a uniform distribution and the fits gave a numerical value consistent with the EFT estimate.
-We obtain the posterior distribution using a probabilistic integration method called Nested Sampling [Skilling06-BA] implemented in Python [2009MNRAS.398.1601F] that calculates both the posterior and the evidence.
+We obtain the posterior distribution using a probabilistic integration method called Nested Sampling [597] implemented in Python [264] that calculates both the posterior and the evidence.
 The experimental data and EFT⋆ fits at LO, NLO, and NNLO are shown in Fig.
@@
 We find the posterior probability distribution functions to be symmetric and well described by Gaussians, and as such we report the mean and standard deviation in Table 7.
-Table LABEL:table:normsBayesian gives the scaling factors for the experimental data sets determined from the NNLO fit along with their common mode errors.
+Table 8 gives the scaling factors for the experimental data sets determined from the NNLO fit along with their common mode errors.
 | Theory | () | () | () |
@@
 […]
-[…]
+IX.6 Theoretical Uncertainty
+The theoretical uncertainties in halo EFT are estimated from the perturbative expansion in [349].
+is associated with the physics at low momenta of interest: relative momentum , binding momentum , inverse Bohr radius , excitation momentum and resonance momentum .
+The cutoff scale of is estimated as described above.
+Except for a narrow momentum range around the resonance, capture is dominated by transitions from - and -waves that preserve the channel spin.
+Thus the threshold theoretical estimates of are based on capture.
+The ANC for the final bound channel is known to be about 5 times larger than the bound channel, so capture in the channel dominates.
+The LO contribution in the expansion comes only from -wave Coulomb interactions in the channel.
+The weakly bound 8B state enables peripheral capture without initial state strong interactions.
+At NLO, -wave Coulomb interactions in the channel and -wave Coulomb interactions in the channel are added.
+Strong interactions in the initial and waves of the channel contribute at NNLO.
+Perturbatively, the -wave scattering lengths , contribute at this order.
+Though is much larger than , the channel is one order higher in the perturbation, making these contributions comparable.
+We also see mixing in the coupled - channels due to the excited core contribution, which is parameterized by scattering lengths.
+The mixing requires the ANC in the capture cross section.
+The theory error is estimated as 30% at LO, 10% at NLO, and 3% at NNLO from an estimated .
+This is validated by the relative sizes of the cross section at different orders of the perturbation, and also by the sizes of the fitted parameters that are consistent with theoretical estimates.
+The uncertainty from the higher order contributions is estimated as follows.
+Two-body currents make a relative contribution with photon energy .
+In the channel, so this is a N3LO contribution.
+In the channel, the larger makes this a contribution; however, this spin channel is suppressed by one order in the perturbation.
+Thus two-body currents constitute a N3LO 3% uncertainty; -wave effective range corrections would also constitute a N3LO uncertainty.
+Initial state -wave interactions (suppressed by two powers of momentum) in the channel also constitute a N3LO uncertainty.
+The uncertainty in the experimental determination of the -wave scattering lengths is included in the 3% N3LO uncertainty.
+IX.7 Conclusions and Recommendation for
+R-matrix and EFT fits of direct and Coulomb dissociation data below 1250 keV, where the influence of resonances other than the can be neglected, were carried out.
+Many different theories and energy ranges were considered and consistent results were obtained.
+Our recommendation was obtained using the halo EFT of Higa et al.
+[349] that includes an excited 7Be component.
+This theory was selected due to the generality and rigor of its treatment of short-range interactions in the initial and final states and of two-body currents, as well as its inclusion of transitions and robust uncertainty quantification.
+The resulting recommendation is eV b, where the first error is due to fitting and the second due to neglected higher-order contributions; for the purposes of solar modeling, we intend that the two errors be added together in quadrature and recommend eV b.
+The recommended has a notably smaller uncertainty than the SF II value for two primary reasons.
+Significant theoretical developments in halo EFT with a NNLO calculation have reduced the theoretical uncertainty by a factor of two.
+The use of floating data normalization factors allows discrepant data sets to be brought into agreement for low-energy extrapolation without inflating the measurement uncertainties.
+In the Bayesian fits, the normalization factors are assigned priors, as is done for some of the theoretical parameters.
+Frequentist fits presented in the supplemental material include these in the definition of the cost function.
+Normalization factors different from unity increase the cost, but the overall cost is lowered by bringing precise measurements from different data sets into agreement.
+This is effective because the number of scale factors is much less than the number of high precision data points.
+Due to the small uncertainties and large number of data points, the measurement of Junghans et al.
+[383] exerts a strong influence on the fit.
+Planned measurements, such as those recently approved at and 1.2 MeV using a 7Be beam and the DRAGON recoil separator at TRIUMF, must have comparably small statistical and systematic uncertainties in order to compete effectively.
+The reaction (S1 14)
+The 14NO reaction proceeds through a number of bound-state transitions, but the total capture cross section is dominated by the capture to the 6.79 MeV state (3/2+) in 15O.
+(In this chapter the reported level energies are taken from the most recent compilation [27]).
+The contribution of this transition to the total reaction cross section at solar energies is % [12, 62].
+The second strongest transition with a contribution of % is the capture to the ground state of 15O, while the capture to the 6.18 MeV state (3/2-) is responsible for about % of the total cross section.
+The remaining % is attributed to the capture to the other bound states (i.e.
+5.18 MeV, 1/2+; 5.24 MeV, 5/2+; 6.86 MeV, 5/2+ and 7.28 MeV, 7/2+ ).
+In SF II a lower value, with a greatly reduced uncertainty, for of 1.66(12) keV b was recommended compared to that of SF I, 3.5 keV b.
+This was the result of the following: the new implementation of the more rigorous R-matrix technique [45] over the previously used Breit-Wigner analysis, higher precision low-energy capture data sets [273, 561, 373, 110, 431, 469] over the yields presented in Schröder et al.
+[577], transfer measurements of proton ANCs for the bound states [116, 492], and constraints on the lifetime of the 6.79 MeV bound state [117, 580].
+Since SF II, several new experiments and R-matrix calculations have been completed, but our understanding of the ground state transition still remains incomplete.
+X.1 Current Status and Results
+Since SF II, several new measurements have been reported for the capture reaction [434, 216, 650, 324, 589, 325, 280], the life-time of the 6.79 MeV state [288, 589, 281], the proton bound state ANCs [51], as well as a measurement of the low energy 14NN scattering cross section [230].
+These and past works now include several R-matrix fits [45, 273, 561, 373, 434, 650, 280], resulting in a range of different approaches to the fitting, which gives insight into the systematic uncertainties of the model.
+The recomended partial S(0) are discussed in sec.
+X.5.
+The 6.79 MeV transition, being thre strongest, has seen a great deal of experimental attention.
+Its ease of measurement and relatively well understood reaction contributions result in a small uncertainty in the low energy extrapolation of the S-factor of 3%.
+The ground state and 6.18 MeV transitions have uncertainties of 40%, owing mainly to the complexity of the underlying reaction contributions and the much smaller cross section.
+X.2 Absolute strength of the 259 keV resonance
+The = 259 keV resonance in 14NO is often used as a reference for normalization in many experiments because it is a narrow resonance, has a precisely known strength, and has -ray transitions that span a wide energy range.
+The absolute strength of the resonance has been determined in three works [216, 324, 589] since SF II.
+Based largely on the work of Daigle et al.
+[216], a new resonance strength of meV is recommended, consistent with the previous value but with higher precision thanks to the new data sets.
+For the averaging we adopted the approach followed by SF II.
+The common systematic uncertainty of the nitrogen stopping power of 2.9% [678] was excluded from the weighted mean calculation, and summed in quadrature with the weighted mean uncertainty to obtain the final uncertainty.
+Additional to this, the stopping power and branching ratio corrected resonance strengths from [216] were used in the calculations instead of the originally quoted values.
+The resonance strength determination is summarised in Table 9.
+For the branching ratios, those of Daigle et al.
+[216] are recommended.
+| Becker et al. [105]888As in [216] | 13.7 1.0 |
+| Runkle et al.
+[561]111Categories for the dominant rate contribution at the central temperature of the Sun: (“A”) Narrow resonances.
+(“B”) Broad resonant/non-resonant contributions.
+We define a “narrow resonance” by the criterion that its full contribution to the reaction rate can be calculated from the resonance energy and strength alone.
+| 12.4 0.9 |
+| Imbriani et al. [373] | 12.9 0.9 |
+| Bemmerer et al. [110] | 12.8 0.6 |
+| Daigle et al. [216] | 12.6 0.6 |
+| Gyürky et al. [324] | 13.4 0.8 |
+| Sharma et al. [589] | 12.8 0.9 |
+| SF III recommended value | 12.9 0.5 |
+X.3 Indirect Studies
+There are several different indirect techniques that have been used to help constrain the extrapolation of the low energy cross section.
+A new transfer study of proton ANCs has been reported by Artemov et al.
+[51].
+Since the 6.79 MeV level also acts a subthreshold resonance in the ground state transition, its lifetime is also needed.
+New lifetime measurements have been made by Galinski et al.
+[288], Sharma et al.
+[589], and Frentz et al.
+[281] in addition to the previously avaliable data by [117] and Schürmann et al.
+[580].
+Becasue the lifetime is in a range that is very difficult to access experimentally ( fs) these works still report values either with very large uncertainties or upper limits only.
+Scattering cross sections can also be very helpful in constraining both bound and unbound level parameters for the R-matrix fit to the capture data.
+A consistent set of scattering measurements has been made by deBoer et al.
+[230] superseding the unutilisable data for 14N showing a large amount of discrepancy [318].
+In the new work only yield ratios are reported in order to reduce the uncertainties associated with target degradation.
+This scattering data give added constraint to the broad, high energy, 3/2+ resonance at 2.2 MeV.
+X.4 R-matrix considerations
+The phenomenological R-matrix has been the primary model used to fit and extrapolate cross section data, following the analysis of Angulo and Descouvemont [45].
+Many of the investigations have used the same multichannel R-matrix code, AZURE2, which was introduced by Azuma et al.
+[62].
+The model has proven to be very popular because it is very flexible in its level structure description, yet at the same time retains enough physical constraints that it can be used for extrapolations into unobserved regions of the cross section.
+Of course this flexibility also allows for additional complications like over fitting, mistaking experimental resolution effects for inconsistencies between data sets or the need for additional reaction components [654], to give just a few common examples.
+In the R-matrix framework, the cross section is described by including a number of reaction components, which includes resonances, subthreshold states, and background levels.
+For radiative capture reactions, an additional component is introduced, external capture [354, 93, 45], which is used to model the non-resonant capture process.
+Fig.
+9 shows an example multichannel AZURE2 R-matrix fit to scattering data and the three strongest transitions of the 14NO reaction: 6.79 MeV, ground state, and 6.17 MeV.
+Using the alternative R-matrix parameterization of Brune [150], the resonance components are characterized by observable energies and partial widths, while subthreshold resonances have their strength quantified by the state’s asymptotic normalization coefficient (ANC).
+The strength of the external capture contributions are likewise given by ANCs.
+The strongest transition, to the 6.79 MeV state, is dominated by external capture at solar fusion energies.
+As pointed out in SF II, discrepancies among the data sets over the higher energy range make it unclear whether an additional contribution is needed from a background level.
+However, as Wagner et al.
+[650] showed, even when a background level is introduced, it has a very small effect on the low energy extrapolation owing to the dominance of the external capture component.
+The next strongest transition is to the ground state, where the R-matrix description of the low energy cross section is rather complicated.
+It is thought to have significant competing and interfering contributions from the 6.79 MeV subthreshold state, the narrow = 259 keV resonance, external capture, a background level, and the broad, higher energy, resonance at = 2.2 MeV.
+Despite all of these different contributions, which gives the R-matrix fit a considerable amount of flexibility for this transition, a satisfactory fit to the ground state transition data remains elusive [434, 650, 280].
+Finally, the 6.18 MeV transition is dominated by the narrow = 259 keV resonance and external capture contributions at low energy.
+A contribution from the 6.79 MeV subthreshold state may also be significant.
+X.5 Partial S-factors
+In this compilation a different approach is adopted to determine the value for the two strongest transitions than in SF II.
+Because several are already available in recent publications, no new R-matrix calculation was performed.
+Instead the reported values from the original publications [561, 373, 434, 650, 280] are considered.
+In Runkle et al.
+[561], Li et al.
+[434], and Frentz et al.
+[280], the data sets are normalized to the strength of the 259 keV resonance.
+Therefore the reported S(0) values were renormalized by the ratios of the resonance strengths used in each work and the present recommended one.
+In constrast, Imbriani et al.
+[373] and Wagner et al.
+[650] established an absolute normalization, independent of the 259 keV resonance strength.
+X.5.1 Ground State Transition
+The ground state transition cross section of the 14NO reaction is very challenging to measure accurately and to high precision because the cross section is very low and most experiments perform measurements with high efficiency -ray detectors in very close geometry.
+This increases the yield for these measurements but also greatly increases the effect of -ray summing into the ground state transition, with the largest contribution from the 6.79 MeV transition at low energies.
+In addition, the cross section on the high energy side of the strong, low energy, 259 keV resonance is rather inconsistent between different measurements, likely a result of nitrogen diffusion into the thick target backing material.
+In SF II, the R-matrix fit included the data of Schröder et al.
+[577], Imbriani et al.
+[373], Runkle et al.
+[561] and Marta et al.
+[469], below 2 MeV.
+A good description of the data was obtained over most of this energy region, with the exception of a consistent over estimation of the experimental data from 300 to 500 keV.
+The fit there favored the higher cross sections of Marta et al.
+[469], consistently overestimating the data of [561] throughout this energy range.
+It should also be noted that the data of Schröder et al.
+[577] were clerified to be yields that were uncorrected for experimental effects.
+Cross sections were deduced for these data in SF II but where never pubished.
+Since SF II, there have been three new data sets reporting partial cross sections [434, 650, 280], which all used the AZURE2 R-matrix code to derive .
+These values are considered here as an indication of the uncertainty in the low energy extrapolation.
+Li et al.
+[434] concentrated on differential cross section measurements between 0.5 and 3.5 MeV, while both Wagner et al.
+[650] and Frentz et al.
+[280] focused on bridging the gap between the measurements of Li et al.
+[434] and the lower energy LUNA [373] and Laboratory for Experimental Nuclear Astrophysics (LENA) measurements [561].
+The new data and accompanying R-matrix fits have repeatedly confirmed a discrepancy between the R-matrix fit and the experimental data for the ground state transition, especially when the high energy data of Schröder et al.
+[577] and Li et al.
+[434] are included in the fit.
+On the R-matrix side, this could point to a missing reaction component (i.e.
+missing resonance or subthreshold state) or an incorrect spin-parity assignement.
+On the data side, the close geometry setups of most experiments result in large summing corrections.
+Inaccurate corrections for angular distributions could also play a role.
+Because of this unresolved inconsistency between data and model, a larger uncertainty has been attributed to this transition compared to SF II (see Table 10).
+X.5.2 6.79 MeV Transition
+In contrast to the ground state, repeated measurements of the 6.79 MeV transition continue to produce consistent results, especially at very low energies.
+The measurements of Wagner et al.
+[650] do show a more rapidly increasing cross section at high energies than those of previous works, but are consistent in the low energy region.
+In addition, the ANC determined from the normalization of the external capture cross section used to model the non-resonant capture in the R-matrix fit yields consistent results with those determined from transfer studies [492, 116].
+Because of this continued consistency across measurements, the value adopted here is the weighted average of the reported values, where the uncertainty has been further reduced from that of SF II.
+X.5.3 Other Transitions
+For the three weaker transitions reported in SF II, there were no new experimental data.
+Here the values of have been determined by scaling those from SF II by the ratio of the 259 keV resonance strengths adopted here and from SF II.
+| Transition | (keV b) | Reference | |
+| tr 0 | 0.30 0.11 | 37% | Present |
+| tr 6.79 | 1.17 0.03 | 2.9% | Present |
+| tr 6.17 | 0.13 0.05 | 38% | SF II |
+| tr 5.18 | 0.010 0.003 | 30% | SF II |
+| tr(5.24) 0 | 0.068 0.020 | 30% | SF II |
+| R-matrix sum | 1.68 0.13999The uncertaity is the sum of the partial transition uncertainties. | 7.6% | |
+| Additional syst. uncert.101010From the normalisation to the = 259 keV resonance strength. | 3.5% | ||
+| Total | 1.68 0.14 | 8.4% |
+X.6 Total S-factor and conclusions
+Figure 9 shows the low energy S-factors and R-matrix fits to the strongest capture transitions and the scattering yields at a representative angle and Table 10 summarizes the uncertainties estimated for the zero energy S-factor, for each of the significant capture transitions.
+Their sum gives the adopted total S(0) value.
+The absolute uncertainty of the R-matrix sum is the quadratic sum of the absolute uncertatinties of the partial transitions, while the recomended uncertainty conservatively includes an additional 3.5% relative uncertainty from the resonance strength (see Table 9) to which most of the datasets were normalised.
+As described in Section X.5.1, the uncertainty in the ground state transition has increased to 37% over the 19% estimate of SF II, owing to repeated confirmation of a tension between different experimental data sets and the R-matrix description of the cross section.
+The uncertainty in the 6.79 MeV transition decreased from 4.0% to 2.9% due to the consistency of the various measurements for this transition.
+In addition, the new total capture cross section of Gyürky et al.
+[325], performed using the activation technique over the energy range from 0.55 to 1.4 MeV, is larger than the sum of the partial R-matrix cross sections by 25%, a significant margin compared to the systematic uncertainties of the different data sets.
+However, it should be emphasized that for several of the transitions the only data available are those of Schröder et al.
+[577], which are yields not cross sections and often give only upper limits.
+For example, the capture to the 7.56 MeV level in this energy range may be significant.
+Given the above considerations, a somewhat larger uncertainty of 8.4% has been estimated for , compared to the 7% uncertainty recommended in SF II.
+As one of the main sources of uncertainty remains the ground state transition, far geometry angular distribution measurements are recommended extending the differential measurements of Li et al.
+[434] to lower energies.
+These measurements will both reduce the effect of summing and provide more information on the underlying reaction mechanisms at these low energies.
+Further, as the data of Schröder et al.
+[577] have significant uncorrected experimental effects, new measurements covering the same wide energy range should be undertaken.
+Such measurements are planned at both the LUNA and JUNA underground facilities.
+At very low energies, the LUNA and LENA measurements of the ground state transition by Imbriani et al.
+[373] and Runkle et al.
+[561], respectively, remain the only measurements.
+The discrepancy between the two measurements and the tension between the R-matrix fits make new measurements in this range also of high priority.
+Despite several additional attempts since SF II, the lifetime of the 6.79 MeV transition still remains unmeasured.
+This remains a high priority, as having a well constrained value will greatly improve our understanding of the low energy behaviour of the ground state transition cross section.
+Finally, we recommend that the cross section to the 6.18 MeV and 7.56 MeV final states be measured at high energies in order to address the tension between the measured total cross section and the sum of the partial cross sections.
+Other CNO, Ne-Na reactions
+In the following, we provide information on the proton-induced reactions involving the target nuclei 12,13C, 15N, 16,17,18O, 19F, 20,21,22Ne, and 23Na.
+Unless otherwise noted, we will focus on the low-energy region, i.e., below 150 keV in the center of mass.
+The reactions on lighter target nuclei up to and including 16O are of direct relevance for solar neutrinos; the other reactions play a role in hydrogen burning at somewhat higher temperatures, including asymptotic giant branch (AGB) stars.
+Table 11 lists, where applicable, the S-factor at zero energy, the Gamow peak location, and references.
+Table 12 provides recommended values of the S-factor and corresponding uncertainty on a grid of seven energy values.
+These results are displayed in Fig.
+10, together with their polynomial fits.
+The fit coefficients are listed in Table 1.
+Table 13 gives values of recommended strengths of narrow resonances.
+A detailed discussion of each reaction is provided below.
+XI.1 Reactions of the CNO-I cycle
+XI.1.1
+The low-energy reaction cross section is influenced by the tail of a broad resonance and the non-resonant capture to the ground state of 13N.
+The results of six new direct measurements have been published since SF II [509, 322, 599, 206, 598, 393].
+The first experiment by Nesaraja et al. [509] reports the rection cross section relative to that of the 13C()14N reactions at keV to be .
+Two of the newest experiments were performed at ATOMKI [322, 206].
+The first employed the activation method in a wide laboratory bombarding energy range, keV, to provide a dataset with an absolute normalization having independent uncertatinties, from that of the older data sets obtained using in-beam -ray spectroscopy.
+The main focus of the second experiment was the precise determination of the energies and widths of the two low-energy resonances using in-beam -ray spectroscopy with detectors located at several angles.
+The fourth experiment, by Skowronski et al.
+[599], was performed at the Felsenkeller underground accelerator facility in Dresden, Germany, using in-beam -ray spectroscopy.
+The focus was on determining the parameters of the lowest-lying resonance employing a detector in close geometry with high solid-angle coverage.
+The data covered the energy range of keV, and allowed for an extrapolation to solar and stellar energies using R-matrix theory.
+The absolute normalization of this data set is about 25% lower than the previous results.
+The low-energy slope of the S-factor also changed, resulting in an extrapolated value of keVb in the solar Gamow peak.
+The fifth experiment, by Kettner et al.
+[393], was performed at Bochum and Notre Dame, and focused on the second resonance in the reaction.
+Differential cross sections at and , thick target yields, and angular distributions at selected energies were measured in the energy range of keV.
+A comprehensive R-matrix fit of the available data was also performed, resulting in keVb.
+The sixth experiment, by the LUNA collaboration [598], obtained data in the range of keV, reaching the lowest bombarding energies to date.
+Both the activation technique and in-beam -ray spectroscopy were applied, with a consistent absolute normalization of the data, despite the different systematic uncertainties introduced by the two methods.
+Their S-factor in the solar Gamow peak, keVb, is in good agreement with Kettner et al.
+[393], and is compatible within 2 with Skowronski et al.
+[599].
+Apart from the direct measurements discussed above, the ground state ANC was determined in a recent particle transfer experiment [52].
+That work also reported a fit of the radiative capture data available at the time.
+The R-matrix fits to the data of Skowronski et al.
+[599, 598], Kettner et al.
+[393] give ANC values consistent with the result of the transfer work.
+Thanks to the high consistency of all recent datasets, the extrapolation of the cross section to solar energies has now improved uncertainty.
+Owing to the self-consistency between the used methods and the precise data coovering the lowest energies to date we adopt the S-factor recommended by Skowronski et al.
+[598] (see Tables 11 and 12).
+XI.1.2
+This reaction proceeds to the ground state and five excited states in 14N.
+The low-energy cross section is influenced by the tail of a broad resonance at laboratory bombarding energy of keV.
+Previously, only a single work reported cross sections for all transitions [400], which was the basis of the S-factor recommendations in both SF I and SF II.
+A comprehensive study by the LUNA collaboration [598] provided data in the energy range of keV.
+One of their data sets was obtained with HPGe detectors to study partial cross sections.
+Their second data set was obtained with a -ray calorimeter to measure total reaction cross sections.
+The results from both data sets agree in the absolute cross section scale, despite major differences in the systematic uncertainties of the two methods.
+The new results are about 30% lower than previous results of King et al.
+[400].
+Despite this scale difference the calculated S-factor ratio for the 12C()13N and 13C()14N reactions at keV is , in perfect agreement with the experimental reported one of by Nesaraja et al.
+[509].
+XI.1.3
+Following SF II, the experimental and phenomenological work on the 15NC reaction can be divided into three groups addressing different aspects: (i) new () data [235]; (ii) new () data [375, 374, 545, 323, 231]; and (iii) R-matrix analysis of existing data sets [233, 152].
+Other studies have been conducted for laboratory bombarding energies above 400 keV.
+In Reinhardt et al.
+[545], the strength of a resonance at 403 keV (, Ex keV) is tabulated, and the value is in agreement with previous determinations.
+In Gyürky et al.
+[323], the thick-target yield for the keV resonance is shown, but no further details are reported.
+New cross section data for the () reaction down to about 900 keV center-of-mass energy are displayed in deBoer et al.
+[231].
+They used the data for a comprehensive R-matrix fit, and reported parameters of resonances above about 1500 keV center-of-mass energy.
+Imbriani et al.
+[375, 374] measured the 15N()12C reaction down to about 140 keV laboratory bombarding energy, and showed that the S-factor of the () channel is orders of magnitude lower compared to the ) channel.
+Finally, deBoer et al.
+[233] contains a detailed analysis of all available channels populating the 16O compound nucleus, expanding the results presented in deBoer et al.
+[235].
+In this paper, a wide range of experimental data were considered leading to a single consistent R-matrix fit, accurately describing the broad level structure of 16O below keV.
+The resulting fit was used to extract an improved determination of the low-energy S-factor for the 15N()12C reaction.
+In addition, Brune and deBoer [152] have developed the partial-wave formalism for calculating the angular distributions of secondary -rays following particle emission, providing the framework needed for future studies that can directly utilize differential measurements of the 15NC reaction.
+The available evidence confirms that the () channel dominates the total S-factor.
+The global R-matrix analysis in deBoer et al.
+[233] provides results that slightly exceed the experimental data of Redder et al.
+[544] and also estimates a slightly larger value, keVb, than previous works (see Table VII of deBoer et al.
+[233]), calling for additional low-energy measurements.
+The extrapolated S-factor in deBoer et al.
+[233] agrees with the findings in La Cognata et al.
+[420], which is the most recent indirect study, determining the S-factor down to 20 keV center-of-mass energy.
+The global R-matrix analysis suggests a renormalization of the La Cognata et al.
+[420] S-factor by 6%, which is within the quoted uncertainty.
+Based on the information provided above, we adopt the S-factor of La Cognata et al.
+[420], keVb, which was also adopted in SF II.
+XI.2 Reactions of the NO, or CNO-II, cycle
+XI.2.1
+The low-energy 15NO cross section is dominated by its ground state transition [428, 375].
+This transition’s cross section is enhanced by the constructively interfering tails of two broad resonances at laboratory proton energies of = 338 and 1028 keV, which correspond to 1- levels in the 16O compound system.
+In addition to ground-state proton and -ray de-excitation, these levels also decay strongly through the and channels.
+For the 338 keV resonance, [233].
+Because of interference effects, this ratio does not directly translate to the rate of passage from the CNO cycles to the NeNa cycle.
+Currently, the favored framework for the analysis and extrapolation of the available data is the phenomenological R-matrix.
+Here, the main components that are used to model the low-energy cross section of the 15NO reaction are the two broad 1- resonances, external capture, and a 1- background level.
+The strength of the external capture can be characterized by the ground state proton ANC of 16O, which has been experimentally determined via a proton transfer reaction [493].
+Gamma-ray de-excitation to the ground state has also been observed through a 2+ level in this energy region in the 12C+ reactions, but has been shown to be negligibly weak for 15NO [233].
+The cross section measurements used in SF II were those of Hebbard [342], Rolfs and Rodney [554], and Bemmerer et al.
+[108].
+While the then newly measured data of Bemmerer et al.
+[108] represented a significant improvement in accuracy over Rolfs and Rodney [554] and in precision over Hebbard [342], they only cover a limited energy range.
+For SF II, the two-level R-matrix fit of Mukhamedzhanov et al.
+[493] was adopted, which simultaneously fit 15NC data, giving = 36 6 keVb for the 15NO reaction (see also Barker [92]).
+The fit shows that the extrapolation performed by Rolfs and Rodney [554], yielding 64 6 keVb, overestimated S(0).
+Following SF II, a comprehensive measurement campaign was carried out at LUNA and the University of Notre Dame [428], covering a laboratory energy range from = 130 to 1800 keV.
+In that work, a three-level R-matrix fit was used, but the ground state proton ANC of Mukhamedzhanov et al.
+[493] was not considered, resulting in a somewhat larger extrapolated value of keVb than that adopted in SF II.
+Mukhamedzhanov et al.
+[495] subsequently showed that a three-level R-matrix fit of equal quality could be obtained with the ANC, and reported an extrapolated value of that ranged from 33 to 40 keVb.
+Caciolli et al.
+[167] expanded the energy range of Bemmerer et al.
+[108] by measuring the cross section up to = 370 keV.
+Their experiment scanned over the top of the lowest-energy broad resonance at = 312 keV.
+The data were found to be consistent with the ground state data of Leblanc et al.
+[428].
+deBoer et al.
+[233] performed a comprehensive R-matrix fit that included not only the 15N+ data, but also a wide range of 12C+ data over a similar excitation energy range (see also Section XI.1.3).
+The value from this analysis, 40 3 keVb, was slightly above that of Leblanc et al.
+[428] and agreed with the upper limit of Mukhamedzhanov et al.
+[495].
+However, it is lower than the recent potential model extrapolation of = 45 keVb found in the NACRE II compilation [667].
+The difference between the two fits is likely due to the fact that NACRE II disregarded the higher-energy part of the Leblanc et al.
+[428] data, which maps the interference pattern between the two 1- resonances more accurately than older data sets.
+In order to reproduce experimental data in that region, a 1- background level was added in the R-matrix fits of Leblanc et al.
+[428] and later works.
+Here we adopt the value of 40 3 keVb from deBoer et al. [233] as it considers the most comprehensive set of available data.
+While low-energy future measurements are recommended, experiments at high energy are highly desirable to better understand the reaction mechanisms responsible for the interference pattern that is observed between the two well known 1- resonances, which can not be completely reproduced unless some background contribution is also included.
+XI.2.2
+Below a laboratory energy of about keV, i.e., the energy of the lowest-lying resonance [621], the 16O()17F reaction is a prime example of the non-resonant direct radiative capture process.
+This reaction has been measured many times using a variety of techniques, including activation, prompt -ray detection, and inverse-kinematics experiments.
+The capture can proceed to the ground or first-excited state in 17F, with the latter being the dominant transition.
+In SF II, a value of keVb with a a % uncertainty was adopted.
+This result was obtained by normalizing the microscopic-model calculations of Baye et al.
+[101] to the data of Rolfs [552] and Morlock et al.
+[487].
+A recent statistically rigorous evaluation of the S-factor data of Hester et al.
+[347], Chow et al.
+[192], Becker et al.
+[105], Morlock et al.
+[487] is presented in Iliadis et al.
+[372].
+Their analysis, including a discussion of experimental uncertainties, used a combined fit of the transitions to the ground and first-excited states in 17F, and their sum, and was performed using a Bayesian model.
+The physical model was a single-particle model employing a Woods-Saxon potential for generating the radial bound-state wave function.
+The fit had three adjustable parameters: the radius parameter and diffuseness of the Woods-Saxon potential, and the ANCs for scaling the theoretical non-resonant capture S-factors.
+It was also found that a poor fit was obtained when using spectroscopic factors instead as scaling parameters.
+As the 16O()17F reaction at low bombarding energies is peripheral, the analysis of the S-factor in terms of ANCs greatly reduced the sensitivity of the fit to the single-particle potential parameters.
+For the ANCs of the ground and first-excited state transitions, they found values of fm-1 (%) and fm-1 (%), respectively.
+These results agree with those measured in the 16O(3He,d)17F reaction by Gagliardi et al.
+[287].
+It is not surprising that the uncertainties (%) in the latter reference are significantly larger than for the values quoted above, because the distorted wave Born approximation (DWBA) analysis of Gagliardi et al.
+[287] was subject to ambiguities in the choice of optical model potentials for the incoming and outgoing channels.
+The total S-factor at zero energy reported by Iliadis et al.
+[372] was keVb (%).
+The uncertainty is significantly smaller compared to the results published by Iliadis et al.
+[368], which were derived from an analysis that was not statistically rigorous.
+We adopt the results of Iliadis et al.
+[372] in this work.
+Their S-factor is presented in Table 12.
+We endorse the two recommendations made by Iliadis et al.
+[372].
+First, they advocated for a new low-energy measurement of the 16O()17F reaction, specifically at center-of-mass energies below keV, which have so far only been reached in the experiment of Hester et al.
+[347], albeit with a relatively large systematic uncertainty of %.
+Second, it would be interesting to compare their results to those from a future R-matrix analysis of the same data sets.
+XI.2.3
+In SF II, only the non-resonant component of the 17O()18F reaction cross section was discussed, which dominates the total cross section at solar temperatures.
+At the time, the results of Chafa et al.
+[186] were recommended.
+Since the publication of SF II, several new S-factor measurements have been reported, both at low and high energies [510, 406, 327, 583, 244, 159, 326].
+Despite differences in the techniques exploited and energy ranges covered, the derived values agree across different experiments, both for the non-resonant capture and the total S-factor, where the latter quantity includes the low-energy tails of two broad resonances ( 557 and 677 keV).
+Here we recommend the total S-factor from the most recent analysis of [326].
+The 17O()18F reaction cross section was measured in the laboratory bombarding energy range of 500-1800 keV with a typical total uncertainty of 10% using the activation method.
+A comparison with other independent results, when possible, at plab keV resulted in good agreement with Hager et al.
+[327] and Kontos et al.
+[406], while a 20% difference is observed with respect to Newton et al.
+[510].
+An R-matrix analysis was performed considering all the low energy data, resulting in a value of tot = (4.71.0 keVb111111In the publication of Gyürky et al.
+[326] the uncertainty is referred to as statistical only.
+However, it represents, in fact, the total uncertainty (Gy.
+Gyürky, priv.
+comm.).
+At temperatures near MK, typical for hydrogen-shell burning in AGB stars, the 17O()18F reaction rate is dominated by the 65 keV resonance, which was not considered in SF II.
+No direct measurements are reported for the resonance strength, , which is presently estimated from experimental values of the partial and total widths.
+Partial widths of eV and eV were measured using the 14N+ reaction [456, 114].
+The most uncertain quantity is the proton partial width, , which is estimated from the measured strength of the corresponding resonance in the () channel.
+Our best estimate for the () strength is eV (see Table 13 and the following section for details), which leads to = eV.
+Hence, our recommended value is eV (see Table 13).
+This value is consistent with the strength recommended by Palmerini et al.
+[525], though we have inflated the error to account for the uncertain status of the proton width.121212A direct measurement of the 65 keV resonance strength [297], published after the SF III literature review was completed, supports the proton width from [155].
+Moreover, the 65 keV (1-) resonance is expected to interfere with the 2.0 keV (1-) subthreshold resonance.
+We recommend the treatment of the interference contribution described in Buckner et al.
+[159].
+At typical temperatures of classical novae, the main contributor to the 17O()18F reaction rate is the = 183.9 keV resonance, which was first observed by Fox et al.
+[277].
+It was subsequently remeasured by Chafa et al.
+[186], Scott et al.
+[583], Di Leva et al.
+[244], Buckner et al.
+[159].
+Here we recommend the weighted mean adopted in Buckner et al.
+[159], (1.770.09)10-6 eV.131313The article by Gesuè et al.
+[297] reports an bare value of eV for the 64.5 keV resonance strength.
+XI.2.4
+The 17O()14N reaction rate is determined by the contributions of narrow resonances.
+At solar energies, the dominant contributions originate from resonances at E 2 and 65 keV.
+At typical temperatures of AGB stars and classical novae, the main contributor to the rate is a resonance at E 183.9 keV.
+Concerning the subthreshold resonance, no new results have been published since SF II.
+The E 65-keV resonance has been studied recently both in a direct experiment ( eV; Bruno et al.
+155) and with the Trojan Horse indirect Method ( eV; Sergi et al.
+588).
+The resonance strength reported by Bruno et al.
+[155] deviates by a factor of 2.5 from that of Sergi et al.
+[588] and a factor of 2 from the pre-SF II value of Fox et al.
+[278], which includes a revised analysis of the data from Blackmon et al.
+[126].
+Because the values disagree by more than their uncertainties, we recommend the value eV, with the expanded uncertainty encompassing all values in the literature.
+XI.3 Reactions of the CNO-III cycle
+XI.3.1
+Important components for this reaction channel are a keV resonance, for which only indirect data exist [652, 187, 426], the resonances at keV and keV and the non-resonant capture process.
+Reported strengths of the kev resonance vary by one order of magnitude between [448] and eV [652], depending on the assumed proton widths.
+We recommend a value of eV, based on the results of Champagne and Pitt [187].
+A comprehensive measurement was done by Wiescher et al.
+[652], covering the 80-2200 keV range at multiple angles and measuring branching ratios, the direct component (extrapolated from high-energy data), and strengths down to the keV resonance.
+The strength of the latter resonance, eV, has been remeasured multiple times with consistent results [652, 105, 647, 240, 122, 528].
+The strength of the keV resonance was recently controversial [157, 274], but a measurement by Best et al.
+[122] found it to be of insignificant strength ( eV).
+The non-resonant component was determined in Wiescher et al.
+at an energy of keV and extrapolated to a value of keVb.
+Buckner et al.
+normalized a non-resonant capture model calculation to the experimental value at keV, resulting in keVb.
+No uncertainties were given for these two S-factors.
+Best et al.
+[122] extracted the non-resonant component from a cross section fit at center-of-mass energies between 85 and 150 keV, resulting in ) keVb.
+Here we adopt this value, but further investigation of the low-energy cross section is necessary to resolve these discrepancies.
+XI.3.2
+The low-energy trend of the 18O()15N cross section is determined by three resonances at 20, 90 and 143 keV (see Table 13).
+The most recent determination of the strength of the 20 keV resonance comes from a Trojan horse method (THM) experiment [426].
+Their result was already recommended in SF II and we maintain the same recommendation.
+The resonance at 90 keV was investigated in the past with direct [448] and indirect [426] techniques, finding consistent results.
+Nevertheless, a new direct measurement reported a resonance strength one order of magnitude higher than the previously adopted value [153].
+Since no clear explanation can be found for such a discrepancy, here we adopt a weighted average of the three resonance strengths, but we inflate the uncertainty so that it embraces all three values.
+New experiments to solve this discrepancy are advised.
+The resonance at 143 keV is well known, and different experiments provided consistent results [448, 104, 156].
+Here we adopt a weighted average of Becker et al.
+[104] and Bruno et al.
+[156], which have independent systematic uncertainties (while no information on individual contributions to the total uncertainty is given in Lorenz-Wirzba et al.
+[448]).
+The non-resonant differential cross section has recently been measured down to Ec.m.
+= 55 keV by Bruno et al.
+[153].
+At overlapping energies, the S-factor is consistent with the previous measurement by Lorenz-Wirzba et al.
+[448].
+Bruno et al.
+[153] found a broad structure at keV that they could only explain assuming the existence of a so-far unobserved resonance at E keV.
+Further investigations are needed to shed light on this supposed resonance.
+XI.3.3
+The 19FNe reaction has been challenging to access experimentally because it has a much smaller cross section than the competing 19FO reaction.
+The cross section of both reactions is dominated by a mixture of narrow and broad resonances corresponding to mostly unnatural parity states in 20Ne (mainly = 1+ and 2-).
+At very low energies, the cross section may be enhanced by a near threshold state at 11 keV and subthreshold and non-resonant capture contributions may also contribute significantly, but have not yet been well characterized [232], as few particle transfer studies have been published [123, 401].
+The first low-energy cross section measurement was that of Couture et al.
+[203], who measured only the 19FNe transition.
+While this was likely a strong transition, there can certainly be other transitions with sizable contributions.
+Couture et al.
+[203] covered the center-of-mass energy range from 200 to 760 keV, where three strong resonances were observed at 323.9, 564 and 634 keV.
+Only an upper limit was reported for an expected resonance at keV.
+Also, no evidence of either a near-threshold state, subthreshold state, or non-resonant capture has been observed.
+An investigation of the -ray branching ratios of the keV resonance was undertaken by Williams et al.
+[655], using the DRAGON recoil separator at TRIUMF.
+Recoil detection provided an effective method of discriminating the large background from the 19FO reaction.
+It was found that the primary transitions to both the ground and the 4967 keV state dominate over the primary transition to the 1633 keV first-excited-state.
+Finally, a new direct measurement of the low-energy cross section has been reported by Zhang et al.
+[671], which overlaps with the previous low-energy data of Couture et al.
+[203].
+A new resonance was discovered at = 225.2 keV in the 19FNe cross section, attributable to a 3- state that corresponds to a weak resonance observed in the 19FO data of Spyrou et al.
+[606] at the same energy.
+This low-energy resonance increases the 19FNe reaction rate by about an order of magnitude at 0.1 GK.
+Like Williams et al.
+[655], Zhang et al.
+[671] also observed a sizeable ground state transition for the keV resonance.
+As in the 19FO reaction (see Section XI.3.4), an upper limit was also estimated for the proposed = 11 keV resonance.
+XI.3.4
+The 19F()16O reaction was not included in SF II.
+NACRE [44] pointed out that the total astrophysical S-factor was dominated by the non-resonant channel, though underscoring the large uncertainties () affecting the reaction rate at low temperatures.
+After NACRE, the situation greatly changed following the publication of a large number of papers addressing the measurement of the channel [421, 423, 376, 442, 443], of the channel [606, 267, 247, 672, 674, 532].
+In addition, several authors attempted to reach a consensus astrophysical factor carrying out refined analyses using, for instance, the phenomenological R-matrix [341, 441, 444, 673, 232], and carried out new theoretical calculations of the astrophysical S-factor [563].
+Regarding the measurement of the astrophysical factor for the channel, populating the 16O ground state, indirect (using the THM [627]) and direct methods were used, leading to consistent results.
+While the direct measurement [443] could reach about 170 keV in the center-of-mass, the THM measurements [421, 423, 376] could reach 5 keV, providing an extrapolated keVb.
+The most striking feature of both direct and indirect measurements is the evidence of resonances populating the keV energy region, at odds with the NACRE extrapolation.
+Until the work of Zhang et al.
+[674], direct measurements of the channel stopped at energies well above the Gamow energy.
+Zhang et al.
+[674] reached energies as low as keV in the center of mass, still above the Gamow window for solar fusion, using JUNA.
+Although the lowest energy point is affected by a large uncertainty (87%) and no correction for the electron screening effect is reported (for instance, Assenbaum et al.
+[61] estimated a 20% enhancement at 70 keV center-of-mass energy for proton-induced reactions on oxygen), the R-matrix analysis the authors carried out indicates the presence of a resonance at keV, corresponding to a 20Ne excitation energy of 12.855 MeV.
+This level could significantly increase the S factor, yielding S(0) MeVb.
+If this result is confirmed, the channel would dominate the low-energy total S-factor for 19F()16O at center-of-mass energies below 50 keV.
+Alternatively, if the keV resonance is excluded from the fit, the deduced value would be S(0) MeVb.
+The presence of new data sets triggered the publication of review papers aiming to provide the best fit of the available data.
+In particular, Lombardo et al.
+[444] carried out a comprehensive R-matrix analysis spanning an energy range from about keV to keV.
+The R-matrix fit predicts a S(0) MeVb for the channel, with a conservative uncertainty of about 20%, lower than the value from Indelicato et al.
+[376].
+An additional interesting result from this work is that the channel, corresponding to the population of the 16O first excited state, is negligible with respect to the one below about keV, yet if the state at 13095 keV is assumed to contribute, as supported by the 16O()16O∗ data, the channel may significantly contribute to the total astrophysical factor above about keV, calling for new data at low energies.
+The most comprehensive R-matrix analysis, however not including the results in Zhang et al.
+[674], is the one carried out in deBoer et al.
+[232].
+For the channel, the results are consistent with Lombardo et al.
+[444].
+For the channel, the conclusions are that the channel is the major one and, as shown by Zhang et al.
+[674], low- and sub-threshold resonances can enhance the astrophysical factor to dominate over the channel at astrophysical energies.
+Finally, the result of the R-matrix analysis for the channel supports a flat S-factor devoid of resonances, in contrast with La Cognata et al.
+[421, 423], Indelicato et al.
+[376], Lombardo et al.
+[442, 443], in agreement with the analysis of Herndl et al.
+[344].
+However, this is based on the data from Lorenz-Wirzba [447], which were excluded from the NACRE compilation due to a likely large underestimate of reported absolute cross sections, by about a factor of 2.
+In summary, while the situation for the and channels is well constrained, the contribution of higher 16O excited states and, in particular, of the channel, needs further studies to confirm the occurrence of the keV resonance.
+This represents, at present, the largest source of uncertainty.
+Therefore, the recommended reaction rate lower limit is conservatively set by the Indelicato et al.
+[376] one, while the upper limit is given by Zhang et al.
+[673], assuming the existence of the keV resonance.
+XI.4 Reactions of the neon-sodium (NeNa) cycle
+XI.4.1
+The low energy S-factor of the 20NeNa reaction is dominated by the high energy tail of a subthreshold resonance only a few keV below the reaction threshold.
+In addition, the non-resonant capture into two additional states is relevant and provides the largest contribution in the energy range where data are available.
+All three transitions have been measured over a wide range of energies by two direct experiments [555, 451], and the results can be used to arrive at S-factor recommendations.
+However, no consistent parametrizations have been published that could be used to compare the results.
+S-factor data from these two experiments were extracted either from their accompanying supplemental materials or from digitized forms of the published figures.
+The results were normalized to the strength of the keV resonance [ eV; 451].
+The resonant capture into the tail of the subthreshold resonance ( keV; Firestone 272) can be described by the Breit-Wigner equation, as shown in Rolfs and Rodney [557], Iliadis [364].
+To account, additionally, for the non-resonant capture to the ground state, a constant value of keVb, as recommended by the indirect experiment of Mukhamedzhanov et al.
+[494], was included in the fit to the data of Rolfs et al.
+[555], Lyons et al.
+[451].
+For the two direct-capture transitions, the energy dependence of the S-factor was calculated with the program DIRCAP [365], using a Woods-Saxon potential with a diffuseness of fm to achieve S-factor magnitudes comparable to the experimental data.
+These S-factors were fit individually to the experimental data and then summed to the total S-factor.
+The results are listed in Table 12.
+Uncertainties from the individual contributions were added in quadrature.
+However, significant and unquantifiable uncertainties likely exist because the results of the two experiments differ systematically both in magnitude and energy dependence [451].
+This may be caused by experimental or data-analysis errors in either experiment, or in both.
+In addition, the experiments at low energies only measured in a small angular range and relied on sparse angular distribution measurements at higher energies to infer angular correlation effects.
+Also, the experiments did not reach sufficiently low energies to conclusively determine the potential parameters to be used in the DIRCAP calculation.
+Consequently, this will contribute to the S-factor uncertainty when the energy range of stellar burning, depending on the temperature, is located far below the experimental range.
+A new measurement of the 20NeNa direct-capture cross section down to keV has recently been published by the LUNA collaboration [471]. The potential-model results agree with the new experimental data.
+XI.4.2
+At temperatures below 0.1 GK, the reaction rate is dominated by three narrow resonances at 16.6, 94.6 and 126 keV.
+The 16.6 and 94.6 keV resonances have never been observed directly in the () channel [308].
+Upper limits on their strengths were derived from a 21Ne(d,Na experiment [508, 370].
+Results are reported in Table 13.
+The non-resonant capture S-factor at zero energy was estimated by Rolfs et al.
+[555] using single-particle spectroscopic factors of the bound states in 22Na.
+According to the authors, the result, S(0) 20 keV b, is uncertain by a factor of three.
+XI.4.3
+The 22NeNa reaction used to be the most uncertain of the NeNa cycle, because of a number of poorly-known low-lying resonances.
+At temperatures below 30 MK, the reaction rate is dominated by a resonance at 36 keV.
+Its strength has been determined from a proton-transfer experiment [330, 371].
+Two near-threshold resonances are expected at 5 and 28 keV, but the existing upper limits on their strengths are so low that they do not contribute significantly to the reaction rate [330].
+Two resonances, at 68 and 100 keV, were reported by Powers et al.
+[537], generating a great deal of discussion in the literature.
+Subsequent searches included direct experiments [180, 238, 266], proton transfer reaction [330] and inelastic scattering [175].
+No evidence for the resonances emerged from these searches.
+The latest direct upper limits on their strength suggest that, even if the two resonances exist, their contribution to the mean (recommended) thermonuclear reaction rate is negligible [266].
+A resonance doublet dominates the reaction rate at temperatures around 0.1 GK.
+The strength of the stronger resonance in the doublet was measured both with direct [181, 392, 432, 656] and proton-transfer experiments (data from Powers et al.
+[537], Hale et al.
+[330] re-analyzed in Santra et al.
+[568]).
+The value adopted in Table 13 is an average of all literature values mentioned above.
+The strength of the weaker resonance in the doublet was inferred from proton-transfer data [568].
+At 0.2 GK, the reaction rate is mainly determined by a resonance at 181 keV.
+Its strength was measured in four independent experiments [181, 392, 266, 656].
+The strength recommended in Table 13 is a weighted average of all four values.
+The most intense higher-lying resonances, relevant to nova and supernova explosions, have been recently addressed in Longland et al.
+[446], Depalo et al.
+[239], Kelly et al.
+[392], Williams et al.
+[656].
+The non-resonant capture cross section has been measured by different groups [392, 266, 656].
+The low-energy trend of the ground state capture is described by a constant value of keVb, plus the contribution of a broad subthreshold state at 130 keV (Ex 8664 keV) [266].
+XI.4.4
+The 23NaMg reaction has seen several recent studies [167, 130, 468, 131].
+The cross section is dominated by three narrow resonances over the lowest experimentally measured energy range, located at = 133, 240.6, and 295.9 keV.
+Below the lowest known resonance, non-resonant components are thought to dominate.
+Several indirect measurements have been made to characterize energies and proton spectroscopic factors of bound and unbound states in 24Mg.
+The measurement of Hale et al.
+[331] used the 23NaHe,d)24Mg reaction to investigate unbound states (11690 12520 keV) near the proton separation energy ( 11693 keV).
+For lower-energy bound states that were not measured, the spectroscopic factors of Garrett et al.
+[289] where adopted.
+A direct capture model was used to calculate the non-resonant component of the cross section.
+This was followed by an investigation of the lowest energy resonance at 133 keV by Cesaratto et al.
+[185], where an upper limit for the resonance strength was determined.
+A subsequent study at the LUNA facility [130] was able to make a first measurement of the resonance strength, 1.46 eV.
+A recent 23NaHe,d)24Mg transfer study at the Triangle Universities Nuclear Laboratory (TUNL) revised the energy of the 133 keV resonance to 133(3) keV [468].
+Direct cross section measurements of the 23NaMg reaction have been made at the University of Notre Dame at energies between = 500 and 1000 keV by Boeltzig et al.
+[131].
+Two, broad, strong, resonances were observed in several -ray transitions at 841 and 981 keV whose low energy tails contribute in the low energy region.
+Note that the energy uncertainties presented here reflect the uncertainty in the beam calibration of the accelerator, which was neglected in Boeltzig et al.
+[131].
+Further, measurements of the 10740 keV primary transition in 24Mg, predicted to have the largest non-resonant capture component based on the spectroscopic factors of Garrett et al.
+[289], observed a trend in the off-resonance cross section that was highly suggestive of non-resonant capture.
+Given the level of consistency observed between the non-resonant capture S-factors calculated from the spectroscopic factors of Garrett et al.
+[289] and the external capture S-factors from the R-matrix analysis, an uncertainty in the non-resonant S-factor of 50% was recommended.
+We have adopted the non-resonant component of Boeltzig et al.
+[131] in this work, which, like all previous calculations, relies heavily on the spectroscopic factors of Garrett et al.
+[289].
+XI.4.5
+The 23NaNe reaction rate is dominated by narrow resonances, with a total of 52 resonances in the center-of-mass energy interval keV.
+After NACRE, the state-of-the-art tabulation of the resonance strengths is the one in Iliadis et al.
+[371].
+It is mostly based on the (3He,d) proton transfer measurement by Hale et al.
+[331], supplying both the excitation energies and spectroscopic factors of the near-threshold states.
+In detail, below keV, all the resonance strengths are taken from Table VI of Hale et al.
+[331], with the exception of the keV resonance strength [468], for which the upper limit directly determined by Rowland et al.
+[558], Iliadis et al.
+[369] is adopted, equal to eV.
+Above keV, the Hale et al.
+[331] strengths are rescaled to the one of the keV resonance determined by Rowland et al.
+[559].
+This result is considered as reference value since it provides the most recent directly-measured strength.
+Besides the above-mentioned works, two more should be considered.
+One is Iliadis et al.
+[369], which is not used in Iliadis et al.
+[371] because the same upper limit for the keV resonance as in Rowland et al.
+[558] is reported.
+The other work is Cesaratto et al.
+[185], which used the upper limit on the () resonance strength and the upper limit on the branching ratio (equal to 0.13 at the 95% CL), to provide an upper limit on the strength of the keV resonance equal to eV.
+This implies a negligible contribution to the reaction rate and confirms the conclusions drawn in Iliadis et al.
+[371].
+We note that below K, the resonances dominating the total 23Na(p,)24Mg S-factor differ from those determining the total 23Na(p,)20Ne S-factor.
+Consequently, the S-factors of these two competing reactions are not correlated.
+| Reaction | Category111Categories for the dominant rate contribution at the central temperature of the Sun: (“A”) Narrow resonances.
+(“B”) Broad resonant/non-resonant contributions.
+We define a “narrow resonance” by the criterion that its full contribution to the reaction rate can be calculated from the resonance energy and strength alone.
+| 222Central energy location (center-of-mass system) of Gamow peak at the solar core.
+| Most recent reference | Comments | |
+| (keVb) | (keV) | ||||
+| 12C(p,)13N | B | 25 | Skowronski et al. [598] | SF II: keVb | |
+| 13C(p,)14N | B | 25 | Skowronski et al. [598] | SF II: keVb | |
+| 15N(p,)16O | B | 28 | deBoer et al. [233] | SF II: keVb | |
+| 15N(p,)12C | B | 28 | La Cognata et al. [420] | SF II: keVb | |
+| 16O(p,)17F | B | 30 | Iliadis et al. [372] | SF II: keVb | |
+| 17O(p,)18F | B | 31 | Gyürky et al. [326] | SF II: keVb | |
+| 17O(p,)14N | A | 31 | Bruno et al. [155] | near-threshold resonances | |
+| 18O(p,)19F | B | 31 | Pantaleo et al. [528] | also in literature: S(0) = 7.1 and 15.7 keVb | |
+| 18O(p,)15N | B | 31 | Bruno et al. [153] | low-energy resonances | |
+| 19F(p,)20Ne | A | 33 | Zhang et al. [671] | possible resonance near threshold | |
+| 19F(p,)16O | B | 33 | Zhang et al. [673] | possible resonance near threshold | |
+| 20Ne(p,)21Na | B | 36 | present work | ||
+| 21Ne(p,)22Na | A | 20 | 36 | Rolfs et al. [555] | factor 3 uncertainty in |
+| 22Ne(p,)23Na | A | 36 | Ferraro et al. [266] | sub-threshold resonance at keV | |
+| 23Na(p,)24Mg | B | 18 9 | 38 | Boeltzig et al. [131] | |
+| 23Na(p,)20Ne | A | 38 | Iliadis et al. [371] |
+| Reaction | (keVb) (% uncertainty) | ||||||
+| 5 keV | 10 keV | 20 keV | 40 keV | 60 keV | 100 keV | 140 keV | |
+| 12C(p,)13N111From Skowronski et al. [598]. | 1.46(4.1%) | 1.47(4.1%) | 1.51(4.1%) | 1.58(4.1%) | 1.67(4.1%) | 1.89(4.0%) | 2.19(4.1%) |
+| 13C(p,)14N111From Skowronski et al. [598]. | 6.16(6.0%) | 6.22(6.0%) | 6.34(6.1%) | 6.60(6.1%) | 6.89(6.1%) | 7.58(6.1%) | 8.47(6.1%) |
+| 15N(p,)16O | 40.8(7.5%) | 41.7(7.5%) | 43.0(7.5%) | 45.8(7.5%) | 50.0(7.5%) | 59.0 (7.5%) | 73.4(7.5%) |
+| 15N(p,)12C | 7.7(6.8%) | 8.0(6.8%) | 8.5(6.8%) | 1.00(6.8%) | 1.18(6.8%) | 1.66(6.8%) | 2.53(6.8%) |
+| 16O(p,)17F222From Iliadis et al. [372]. | 10.64(4.0%) | 10.38(4.0%) | 9.92(4.0%) | 9.15(4.0%) | 8.56(4.0%) | 7.67(4.0%) | 7.05(4.0%) |
+| 20Ne(p,)21Na333From present work. | 2586(8%) | 1365(8%) | 593 (7%) | 226(7%) | 128(7%) | 69(6%) | 45(6%) |
+| 23Na(p,)24Mg | 18(50%) | 18(50%) | 18(50%) | 18(50%) | 18(50%) | 18(50%) | 18(50%) |
+| Reaction | (keV) | (eV) 222The resonance strength, in the center-of-mass system, for the reactions listed here is defined as , with , , and the spins of the resonance, target, and projectile, respectively, and , , and the proton partial width, -ray or -particle partial width, and total resonance width, respectively.
+| Comments/references |
+| 17O(p,)18F | 65 | see text for details | |
+| 183 | () | see text for details | |
+| 17O(p,)14N | 65 | see text for details | |
+| 183.9 | Chafa et al. [186], Bruno et al. [156] | ||
+| 18O(p,)19F | 20 | Champagne and Pitt [187] | |
+| 90 | Best et al. [122] | ||
+| 143 | see text for details | ||
+| 18O(p,)15N | 20 | La Cognata et al. [426] | |
+| 90 | see text for details | ||
+| 143 | see text for details | ||
+| 19F(p,)20Ne | 11.0 | 8.510-29 | Betts et al. [123], Kious [401] |
+| 212.7 | 4.210-6 | Zhang et al. [671], Spyrou et al. [606] | |
+| 225.2 | (4.19 0.33)10-5 | Zhang et al. [671], Spyrou et al. [606] | |
+| 323.9 | (3.16 0.33)10-3 | Couture et al. [203], Williams et al. [655], Zhang et al. [671] | |
+| 19F(p,)16O | 11.0 | 10-29 | Zhang et al. [673] |
+| 212.7 | Spyrou et al. [606], Zhang et al. [673] | ||
+| 225.2 | Spyrou et al. [606], Zhang et al. [673] | ||
+| 323.9 | Zhang et al. [673, 671] | ||
+| 21Ne(p,)22Na | 16.6 | 6.2 | Iliadis et al. [370] |
+| 94.6 | 6.4 | Iliadis et al. [370] | |
+| 126 | (3.75 0.75) | Görres et al. [308] | |
+| 22Ne(p,)23Na | 5 | 2.1 | Hale et al. [330] |
+| 28 | 3.2 | Hale et al. [330] | |
+| 35 | (3.1 1.2) | Iliadis et al. [371] | |
+| 68 | 6 | Ferraro et al. [266] | |
+| 100 | 7 | Ferraro et al. [266] | |
+| 150 | (3.9 0.9) | Santra et al. [568] | |
+| 151 | (1.97 0.12) | see text for details | |
+| 181 | (2.39 0.12) | see text for details | |
+| 23Na(p,)24Mg | 133 | 1.4610-9 | Boeltzig et al. [130], Marshall et al. [468] |
+| 240.6 | 5.3(1)10-4 | Switkowski et al. [614], Keinonen et al. [391] | |
+| 295.90 | 1.05(19)10-1 | Switkowski et al. [614], Endt et al. [257] | |
+| 23Na(p,)20Ne | 133 | 1.5 | Rowland et al. [558], Iliadis et al. [369], Marshall et al. [468] |
+| 240.6 | 0.1 | Hale et al. [330] | |
+| 295.90 | 1.03(26)10-2 | Iliadis et al. [371] |
+Electron screening of nuclear reactions
+Nuclear fusion reactions measured in the laboratory and those occurring in the solar core are both affected by the electronic environments in which they take place.
+The effects are, however, different requiring special care in deriving the appropriate solar reaction rates.
+As long as these rates cannot be measured under solar plasma conditions in the laboratory, the usual two-step strategy is to remove the laboratory screening effects from the data to obtain the bare nuclear cross section , which is then modified to incorporate the solar plasma screening modifications.
+Thus, one must address the differential effects of screening for the solar cross section and the laboratory results, .
+XII.1 Screening in laboratory experiments
+Screening diminishes the Coulomb barriers that retard interactions between bare nuclei, thus enhancing cross sections, with the effects becoming more pronounced with decreasing center-of-mass energy.
+Nuclear astrophysics has advanced rapidly over the last two decades with the deployment of underground accelerators and other low-background techniques [147] that have enabled low-energy, low-counting-rate experiments.
+In particular, laboratory measurements of the pp-chain reactions 2H()3He and 3He(3He,2He have been made at energies corresponding to the solar Gamow peak.
+A treatment of the distorting effects of screening then becomes quite important in extracting the bare S-factor .
+The bare cross section is given by
+with , where is the reduced mass and the center-of-mass energy.
+The effects of screening can be incorporated into the Gamow penetration factor through an energy shift, , where the screening potential has the net effect of lowering the Coulomb barrier.
+At the lowest accessible energies in laboratory measurements, the separation of target and projectile during tunnelling is much smaller than the atomic radius: thus the electrons, originally bound to the target (assuming a projectile beam of bare nuclei), are attracted by the joint charge of target and projectile during the fusion.
+The screening potential can then be replaced by a constant screening energy .
+This picture was confirmed in simulations of the electron dynamics in low-energy fusion reactions [591].
+After expanding the argument of the exponential in Eq.
+24 to first order in one obtains the following relation between the laboratory and bare S-factors [61]
+This result neglects a small normalization correction associated with the effects on the matching of the external wave function to that in the Coulomb region.
+The dependence of the exponent is responsible for the growing importance of the screening correction at low energies.
+Although the screening potential in principle can be calculated from the electron charge distribution, in most applications it is treated as a free constant parameter, determined along with and its derivatives from a fit to data.
+Figure 4 of SF II shows the results for , where the fit yielded eV.
+The exponential dependence of the screening correction leads to a rapid increase in relative to , reaching 40%, at the lowest data point measured.
+This example also underscores the potential for screening to complicate the extraction of for energies relevant to the solar core, even in cases where measurements are restricted to higher energy than the Gamow window.
+In principle one can test the adequacy of Eq.
+25 by doing laboratory experiments at low energies – even if the reactions studied are not directly relevant to nuclear astrophysics.
+Such studies began more than 30 years ago [61], using gaseous targets to limit energy loss effects and other systematics.
+Subsequent studies have varied greatly in their configurations, e.g., targets ranging from atomic and molecular gases to metals,141414In this section, the word “metals” is used with the traditional meaning from chemistry, not the astrophysical one.
+thin and thick targets, and use of direct and inverse kinematics [31].
+The experiments can be very challenging at the energies needed, subject to backgrounds from both intrinsic activities in detectors and external sources such as cosmic rays.
+The values of extracted from experiment using gaseous targets have often exceeded the adiabatic limit, defined in atomic physics as the difference between the electron binding energies of the separate atoms in the entrance channel and that of the composite atom [556, 553, 258, 259, 46, 538, 310, 33].
+This in turn has generated unease about the reliability of the values for extracted from laboratory measurements for use in astrophysics [85].
+Several suggestions have been made to resolve this apparent discrepancy.
+For example, an error in the estimated stopping power could result in a smaller reaction effective energy, so that the experimentally deduced screening would be larger than the true value [427, 86].
+Indeed such deviations have been found between theoretical calculations of stopping powers and the tabulation which is traditionally being used in the analysis of low-energy fusion data [120, 119, 305].
+An accurate experimental determination of stopping powers for gaseous hydrogen and helium targets would be very desirable.
+Corrections associated with the non-uniform distribution of the electron cloud around the nucleus [174] and the effects of electron-electron interactions [591] have also been suggested as potential generators of unidentified systematic error.
+For molecular targets the screening energy shows a significant dependence on the scattering angle, being smallest if the projectile passes the spectator nucleus before fusion [590].
+Furthermore, the experimental deduction of screening energies requires assumptions about the bare nuclear cross sections, which is another source of uncertainty.
+This was demonstrated by Tumino et al.
+[630, 631], who, after constraining fits by the addition of higher energy data from a Trojan Horse measurement, obtained a lower value of .
+Clusterization and polarization effects have also been proposed to effect the deduced screening energies in specific low-energy reactions [603, 604]
+For resonances, both the energy and width can be affected by screening (564; 679).
+Particular care has to be taken for narrow, low-energy resonances where the assumption of a constant screening energy is not valid and the radial dependence of the screening potential has to be considered which leads to a significant reduction of screening on the resonance width (366).
+Particularly large screening energies were observed in deuteron- and proton-induced fusion reactions if the projectile was implanted into metals (211, 542, 541; 390; 213; 208, 436).
+While the d+d fusion data obtained with a gaseous target [310] were compatible with the adiabatic limit ( eV, 142, 141), screening energies obtained for various metal hosts varied among themselves and could reach values of several 100 eV.
+As the adiabatic limit derived for the fusion on individual atoms in a gas, does not apply to reactions with targets implanted in a host material, an appropriate model has been developed [212, 357] which agrees nicely with recent screening energies obtained under high-vacuum conditions [214, 210].
+Huke et al.
+[357] also identified oxygen and carbon contamination as the source for some particularly large deduced screening energies.
+We note that the studies of d+d low-energy fusion were motivated by the quest to enhance nuclear fusion rates by environmental effects [210].
+The potential discrepancy between observed and theoretical screening energies would become astrophysically irrelevant if it were possible to measure reaction rates directly under solar plasma conditions.
+This may prove possible using high-intensity lasers.
+First measurements of light-ion reactions in plasma have been performed, though under plasma conditions for which screening effects are expected to be negligible [87, 681, 682, 178].
+A direct measurement of bare nuclear cross section, thus avoiding screening effects, is proposed for storage ring experiments, in which a stored beam of ions can collide with a transverse beam [301].
+XII.2 Screening in the solar core
+The energy-generating reactions that occur in the solar core involve nuclei that are almost completely ionized.
+Consequently the atomic environment differs substantially from that of the terrestrial experiments in which these same reactions are measured.
+Once is determined from these laboratory measurements, the screening effects of the plasma must be folded in to yield .
+Adopting the notation of SF II, the correction factor is defined as
+In the weak screening approximation [564], the ion-ion Coulomb potential is screened on a length scale given by the Debye radius
+yielding
+We refer the reader to SF II for the functional dependence of on temperature, density, and the mass fractions , and for discussions of corrections for electron degeneracy and incomplete ionization.
+SF II discusses the conditions under which the weak screening approximation is valid: under solar core conditions this requires , a condition satisfied by pp-chain and CNO bi-cycle reactions.
+Yet corrections are expected at some level, as the inter-ion potential is not completely negligible, in all collisions, compared to the relative kinetic energy of the ions [564, 565, 378, 174, 517, 271, 74, 148].
+Dynamic corrections –nonadiabatic effects that arise when the velocities of reacting nuclei momentarily exceed typical plasma velocities– have been a point of some controversy that both SF I and SF II addressed.
+However, the absence of such dynamic corrections even for large Gamow energies has been shown to be a consequence of the nearly exact thermodynamic equilibrium of the solar plasma [148, 315, 317].
+Specifically, Bahcall et al.
+[74] found that corrections to the Salpeter formula under solar conditions would be at the few per cent level, and pointed to specific errors in several papers that had come to contrary conclusions.
+There have been a few recent papers advocating changes in solar rates due to dynamical corrections, including Vescovi et al.
+[640], Wood et al.
+[665], Simonucci et al.
+[594], Mussack [496], Mussack and Däppen [497], and Mao et al.
+[458].
+The approaches employed are based on modeling the plasma, and in general the authors do not relate their conclusions to earlier work, particularly Bahcall et al.
+[74], making them difficult to evaluate.
+The essential point of Bahcall et al.
+[74] is the conceptual simplifications that result from recognizing that screening in the solar plasma can be formulated as a problem in equilibrium statistical mechanics.
+This approach removes all need to classified rates as fast or slow relevant to some plasma timescale, as there is no time in equilibrium statistical mechanics.
+Consequently we continue to regard Bahcall et al.
+[74] as the most realistic estimate of corrections to the Salpeter formula.
+Radiative opacities
+XIII.1 Introduction
+Opacity is a measure of the photon absorption of matter and is an essential quantity for understanding radiative heat transfer in the Sun.
+Radiative heat transfer occurs through the absorption and emission processes taking place within the material that the radiation traverses.
+In local thermodynamic equilibrium (LTE)151515Local thermodynamic equilibrium can be defined as the situation where the thermodynamic properties of a microscopic volume of matter are the same as their thermodynamic equilibrium values corresponding to the local electron temperature and density.
+The assumption of LTE is commonly used in stellar modeling, and the discussion of opacities in this section will be limited only to LTE conditions.
+In some circumstances (such as solar coronal modeling), non-LTE conditions should be given careful consideration, but this will not be discussed in this paper., the emission and absorption are straightforwardly related through the Planck function.
+In this scenario, determining the absorption coefficients or opacity, denoted as , provides a complete description of how radiation is transported through the material.
+For example, in the equilibrium diffusion limit, the radiation heat flux is directly related to the Rosseland-mean opacity (RMO) through the following equation [356]
+where represents temperature, denotes its gradient, is the mass density, and is the Stefan-Boltzmann constant.
+The quantity is the solar Rosseland mean opacity (RMO).
+Thus, any error in this quantity introduces error in the radiative heat transport and the simulated solar evolution.
+Quantifying the error in is very challenging because is a weighted mean of a complex solar mixture opacity:
+where is a weighting function related to the temperature derivative of the Planck function :
+where denotes the photon frequency, the Planck constant, the Boltzmann constant, and the speed of light.
+Moreover, the frequency-resolved solar opacity is a sum of elemental opacities weighted by their mass fractions ():
+Equation 30 reveals that the accuracy of is especially important over the spectral range where is high and is low (due to the inverse weighting).
+Equation 32 shows that accuracy in requires good knowledge of both element opacities and mass fractions .
+Figure 11 shows the total monochromatic opacity of the Sun at three different solar radii, highlighting the variations in (gray), (black), and some of the leading individual element contributions (red, green, blue) to the RMO determination (Eq.
+30), as a function of radius.
+This figure was made assuming the GS98 solar abundances and using Opacity Library (OPLIB) opacities [197], and is included primarily for illustrative purposes.
+As shown in the figure, the solar opacity varies considerably in both magnitude (note the log axes) and shape across different solar radii.
+Moreover, each element contributes distinctively to the overall monochromatic opacity, with either a dominant continuous opacity spectrum, or strong lines that arise from bound-bound transitions.
+If there is an error in the calculated solar opacity , it would stem from inaccuracies in the calculated element opacities and/or abundances.
+Additionally, the error is expected to be a complex function of photon energy and would vary at each radius due to changes in temperature and density.
+As a result, it is not possible to provide a single factor for uncertainty or correction.
+It is crucial to assess the accuracy of the intricate spectral features depicted in Fig.
+11 in order to better understand the nature of the errors.
+These complex spectral features arise from microscopic physics at the atomic level – that is, the absorption of radiation through quantum mechanical processes.
+If one knows the atomic energy levels and transition probabilities of all the relevant atomic states (of each atom or atomic ion of each material in the Sun), then the opacity of that material may be computed.161616We note that the opacity of molecules and dust, or grains, can be very important in some circumstances [356]; however they play a minor role in stellar modeling and so will not be discussed here.
+Since atomic structure and transition probabilities can be predicted (under various approximations), starting from the Schrödinger (or Dirac) equation, this seems like an achievable goal.
+However, complications quickly arise from a multitude of factors, such as: how to deal with the infinity of atomic states that may be populated in each atom or atomic ion; how to account for the perturbations caused by the plasma environment on the atomic states; and how to handle the large quantities of data that arise with such calculations, particularly for multi-electron atoms and ions.
+In solar modeling, one requires the opacity of all relevant elements (around 30) at all (plasma) conditions that are found at all radii throughout the solar evolution.
+Thus, the opacity must be tabulated as a function of species, plasma temperature, plasma density, and photon energy.
+Nevertheless, steady progress has been made in determining the opacity of solar mixtures over the last five decades.
+In the 1980s, insufficient accuracy in stellar-envelope opacity was hypothesized through multiple astrophysical puzzles [593], and opacity models were subsequently improved by incorporating more detailed atomic physics and a more complete set of states [362, 584, 115, 454].
+The improved accuracy was confirmed through a few benchmark experiments [531, 605, 657, 223] and improved agreement with some astrophysical observations [363, 218, 219].
+However, the calculated solar opacities are not experimentally validated at most radii, and in particular more deeply into the Sun where the conditions (temperature and density) are more extreme.
+In the last decade, experimental methods have steadily improved, in concert with the availability of powerful x-ray sources, such as the Sandia Z facility.
+Monochromatic opacities were successfully measured at multiple conditions relevant to solar convection-base conditions [83, 82].
+Systematic experimental studies performed on chromium, iron, and nickel helped narrow down hypotheses for the discrepancies [502].
+All of these experimental results provided essential clues for testing various approximations for solar interior opacity calculations; however, the reported model-data disagreements also raised significant controversy.
+This led to various theoretical investigations and the development of independent experimental methods [530, 518].
+Ongoing theoretical and experimental investigations are primarily focused on understanding the origins of reported disagreements, as a first step toward reaching consensus.
+Our aim here is to provide a status report for the field, summarizing past work on solar opacity calculations and experiments, while identifying some of the open questions that are the focus of current work.
+XIII.2 Opacity models
+Over the years, a wide range of approximations has been used for the various components of opacity calculations.
+To help navigate these approximations, it may be useful to briefly summarize the key components of a frequency-dependent opacity.
+Opacities are often broken down into various components.
+The first of these is “bound-bound” absorption (a photoexcitation transition between two levels of an atomic ion), which manifests as strong line features in the opacity spectrum (such as the black ‘spikes’ in Fig.
+11).
+The “bound-free” absorption (photoionization of the ground or excited level of an atomic ion) is normally a continuous (wavelength-dependent) contribution (such as the smooth (red) hydrogen opacity contribution in the middle and lower panels of Fig.
+11).
+The “free-free” absorption (also known as inverse Bremsstrahlung) is where a free electron absorbs a photon, and finally “scattering” (the scattering of a photon by a free electron) contributions can be important at high photon energies.
+Other contributions may be important in molecules or other material.
+Equation 30 shows that due to the harmonic mean inherent in the RMO, the most important photon energy range is when the spectral opacity is low and when the weighting function is high.
+In the solar case, at a given solar radius, it is important to know which elements dominate the opacity at such a region, which will of course depend on the elemental abundance and its contribution at that distance, as illustrated in Fig.
+11.
+The determination of opacities usually starts from one of two fundamental approaches.
+The first approach, which was used in the first opacity computations [356], starts from a “mean ion” model, where a (fictitious) “average” atom or ion of a given element is conceived, with fractional occupation numbers, which is consistent with the given temperature and density of the species under consideration.
+The fractional occupation numbers, usually determined through a self-consistent iterative approach, inform the real (physical) ion stages of importance at such conditions.
+The average atom can then be unfolded through various procedures to the physical ion stages of relevance, and the opacity of such ions can be determined.
+In this model the plasma effects (in particular the perturbation of the average atom by the fields produced by the surrounding plasma ions and electrons) can be included in a natural way, and this approach is often most useful at high densities, where plasma perturbations on the atoms or ions are most important.
+The infinity of atomic states are naturally truncated by plasma effects that remove bound states from consideration.
+The old Inferno model [435] and the more recent STAR tables [411, 412] are based on this method.
+We note that very recent work [298] has considerably improved some of the approximations used in related approaches.
+The second approach is often referred to as detailed configuration (or term) accounting, where one determines all possible ion stages of the species and then determines all possible ground and (multiply) excited states of each ion, each with a corresponding population that is condition dependent.
+This method is often more spectroscopically accurate than the mean ion model, because it considers the structure of each ion separately, but can require enormous computational resources for complex species.
+The structure of each ion is determined from solution of the Schrödinger (or Dirac) equation, which can then also produce transition probabilities.
+Within this procedure, many computational difficulties must be overcome to produce acceptably accurate data for all ion stages of all relevant elements.
+A few examples are given for illustration.
+Within the solution of the Schrödinger equation it is desirable to include “configuration-interaction” (CI), a two-body term in the Hamiltonian that accounts for electron-electron interactions [204], which results in a quite accurate description of atomic energy levels.
+However, the inclusion of this term for atomic ions with large numbers of configurations can quickly result in a computationally intractable problem.
+Therefore, approximations in which CI is included in only a limited manner are often employed.
+Furthermore, careful consideration has to be given to the number of states that should be included in the calculations.
+A compromise is always necessary between including as many states as possible versus the computational resources available for the calculation.
+Convergence checks are necessary to ensure that sufficient states are included.
+However, the rate of convergence of the opacity with respect to the number of states included will depend on the plasma conditions of interest, so this in itself is not necessarily a straightforward matter.
+The computer codes that are used to generate the opacity must also be able to efficiently process large amounts of atomic data.
+Similarly, when considering photoexcitation and photoionization of atomic ions, it is desirable to include coupling of the bound electrons to the continuum electron(s).
+Often termed “close-coupling”, this may be accomplished through techniques such as the R-matrix approach [163, 584] or other related methods.
+However, such approaches are also computationally intensive, and can be severely limited in the number of terms that can be included in the resulting close-coupling expansions.
+It is also not clear how to include plasma perturbations within such approaches.
+As a result, perturbative approaches, such as distorted-wave methods [204, 65, 567], are often employed, which are much more computationally tractable, and of sufficient accuracy for mid- and highly-ionized ions.
+In concert with an atomic model, the thermodynamics of the system must be considered.
+That is, an equation-of-state (EOS) description of the material is required so that the thermodynamic properties and atomic state populations can be determined for a given material temperature and density.
+In the mean ion approach the thermodynamic quantities arise naturally through the description of the atom/ion in a plasma.
+In a Detailed Configuration Accounting approach, a common approach to EOS calculations starts from a chemical picture based on a minimization of the chemical free energy [551, 358, 217, 329].
+Such a model has to treat a wide range of physical conditions, ranging from ideal gas through high density conditions where pressure ionization may rapidly strip ions of all their electrons.
+The number of atomic states retained in the partition function is a key quantity in determining if the EOS is complete at the conditions of interest.
+This assessment of completeness will of course strongly depend on the plasma conditions.
+The final important aspect of opacity models that we wish to briefly discuss concerns line broadening.
+Photoabsorption lines in a plasma become broadened due to a number of factors, such as natural and Doppler broadening.
+Collisional broadening (often known as Stark broadening) is often a dominant broadening mechanism.
+This is broadening due to the plasma microfields caused by the motion of electrons and ions in a plasma.
+As before, line broadening is an important part of any opacity model, but any treatment of this effect has to be tractable for all ions and all conditions encountered when building an opacity table.
+While quite accurate line broadening models exist for one- and two-electron systems [429, 608, 306], the treatment of broadening for multi-electron systems often requires a number of approximations [245, 246].
+XIII.3 Production of tables
+Some of the earliest determinations of opacity for stellar physics started by the pioneering work of Seaton and collaborators in the UK [584, 115].
+This development led to the “Opacity Project” (OP), which in time produced a set of opacity tables for stellar modeling that are still in common use today [65] At around the same time, OPAL (OPacity Astrophysical Library) opacity tables were published by Lawrence Livermore National Laboratory (LLNL) [362].
+Both of these sets of tables were widely adopted by the stellar modeling community.
+Some early theoretical opacity work was performed at Los Alamos National Laboratory (LANL) [356], starting from pioneering work by Mayer and later by Arthur Cox in the 1960s.
+In the 1990s, the LANL work resulted in “OPLIB” (OPacity LIBrary) tables produced from the LEDCOP code [454]; in the mid 2010s the modern ATOMIC code was used to produce a new generation of opacity tables that improved the computation of opacities in a number of ways [197].
+All of these tables are available electronically in various formats.
+Other important opacity tables have been produced in France (the OPAS and SCO-RCG opacity efforts) [128, 482, 524], Israel (using the STAR code) [411, 412], and elsewhere.
+Differences amonmg these tables and their underlying opacity models have been the focus of dedicated workshops and conferences, for example Mendoza et al.
+[475].
+Many of these tables (primarily OP and OPAL) were quickly adopted by the solar modeling community and were useful in solving the Cepheid variability puzzle [593], where updated opacities helped resolve a previous disagreement between observation and modeling of Cepheid variables.
+Detailed comparisons between these two sets of tables have been made on numerous occasions, both on comparisons of individual opacities [65] and on how solar models that use these tables compare [237].
+Overall, the agreement between the various opacity tables was deemed to be “satisfactory” (within 5 %) for the accepted solar models of the late 1990s [454].
+This paradigm was shaken when new determinations of the elemental solar abundances were reported [58], which implied that solar models disagreed with helioseismology observations.
+It was suggested [586] that changes in the opacity of a few species (mainly Fe, Ne, O) could reconcile the solar models with the observations.
+Since then, intensive work has been performed in exploring some of the approximations made in the older opacity tables.
+The OP effort has focused on new and larger R-matrix calculations for important ions of Fe [504, 236], primarily Fe17+.
+At Los Alamos, the modern ATOMIC code was used to create a new generation of opacity tables for H through Zn [197].
+Although significant improvements in many aspects of the calculations were incorporated, the resulting opacity tables did not significantly improve the comparison of solar modeling with helioseismology.
+The study of Krief et al.
+[410] showed that one possible way to produce larger opacities under solar conditions would be if line broadening effects were significantly increased – by several orders of magnitude.
+Opacity increases of 10% near the solar convective zone and 2% in the solar core have been attributed to ionic correlations [413].
+There is need for further investigation of such topics.
+XIII.4 Experimental testing of calculated iron opacity
+Errors in calculated solar opacity can arise from inaccuracies in both the abundance and calculated element opacities in a complex way.
+Therefore, it is important to investigate the accuracy of abundance and elemental opacities separately.
+The accuracy of elemental opacities can be best evaluated by examining the frequency-resolved (monochromatic) opacity.
+Since each spectral feature of the element opacity often relies on different physics and approximations, disagreements between model predictions and experimental data on a frequency-resolved basis can help identify which approximations are invalid and guide refinements in opacity theory.
+Here, we provide an overview of the fundamentals as well as previous investigations into measurements of solar-interior opacity.
+A more general and tutorial discussion of the experimental methods and a historical overview can be found elsewhere [84].
+The material opacity is related to its transmission by the following equation:
+where is the backlight spectrum, is the backlight transmitted through the sample, is a sum of the sample plasma emission and other backgrounds, and is the areal density of the opacity sample.
+Thus, sample opacity can be experimentally determined by heating the sample and accurately measuring the following quantities: the backlight with and without the heated sample (, ), the plasma self-emission and background (), and the sample areal density ().
+Several challenging criteria must be met if opacity measurements are to be both interpretable and useful [Perry 1996, Bailey 2009].
+First, the elements, conditions, and spectral ranges of interest must be identified.
+Second, a macroscopic opacity sample must be uniformly heated to the desired conditions to achieve local thermodynamic equilibrium.
+Third, the heated sample must be backlit with a bright and spectrally smooth radiation, , to accurately determine frequency-resolved absorption.
+It is critical that the backlight is significantly brighter than the sample plasma emission or other background signals, .
+Fourth, the absorption spectrum must be recorded with spectrometers that provide sufficient signals and spectral resolving power for the opacity study.
+Fifth, the condition of the heated sample (, , ) must be diagnosed independently of the opacity in question.
+Simulated conditions are not appropriate for this purpose since their accuracy depends heavily on the accuracy of the opacity in question.
+Finally, the opacity spectrum and its uncertainty must be accurately determined from the measured , , , and and their uncertainties.
+After the revision of solar abundance in 2005 [58], solar models and helioseismology disagreed, and the accuracy of solar-interior opacity was called into question.
+This disagreement was primarily due to a significant reduction in solar opacity resulting from the reduced metallicity, and it was the greatest at the base of the solar convection zone (hereafter convection-zone base or CZB).
+Since then, solar abundance has been continuously revised, with the latest abundances determined by two groups being Z/X=0.0225 [56] and Z/X=0.0187 [455], which still disagree with each other, leaving significant uncertainty in the solar abundance.
+For determining the accuracy of calculated solar opacity, the accuracy of calculated element opacities must also be experimentally scrutinized.
+Since Fe and O are the two dominant sources of opacity at the CZB, their opacities must be experimentally tested at CZB conditions (i.e., , ).
+Unfortunately, previous experimental approaches were not suitable for this purpose because their backlighters were not bright enough to mitigate the bright self-emission produced at these conditions.
+Over the past two decades, experimental methods have been refined using one of the brightest x-ray sources, Z-pinch.
+In 2007, iron opacities were successfully measured at 150 eV and , providing support for the accuracy of calculated Fe opacities under these specific conditions.
+In 2015, iron opacities were measured at larger temperatures and densities.
+This study revealed a significant discrepancy between the calculated and measured opacities as the temperature and density approached the CZB conditions, as shown in Fig.
+12.
+This is a significant concern since the experimental density was still approximately three times lower than the actual density at the CZB.
+Opacity models employ approximations to account for density effects such as line broadening and pressure ionization.
+If these approximations are incorrect, the disagreement between calculated and measured opacities could be even more pronounced at the densities found at the CZB.
+The disagreement observed in the iron opacity was complex, as depicted in Fig.
+12.
+The modeled opacities were lower in the quasi-continuum region at short wavelengths ( Å), narrower in bound-bound lines, and deeper in the opacity valleys known as windows.
+These discrepancies are likely caused by various factors, but it is difficult to disentangle different sources of discrepancies based solely on the disagreement in iron opacity alone.
+In 2019, a comprehensive study was conducted to measure opacities for Cr, Fe, and Ni at solar interior temperatures [502].
+This systematic study helped narrow down hypotheses on the sources of disagreement by observing how the discrepancies change as a function of atomic number.
+The measurements revealed that measured opacities for all three elements had narrower line features than the model predictions, suggesting potentially inaccurate density effects in opacity models.
+Disagreement in the opacity windows was observed for Cr and Fe, while Ni did not show significant disagreement.
+At the experimental conditions used in the study, Cr and Fe were in an open L-shell configuration, whereas Ni was closer to a closed L-shell configuration.
+It is known that the calculations of both population and equation of state (EOS) are more challenging in open L-shell configurations.
+Therefore, the observed trend in window disagreement may suggest a connection to the difficulties associated with EOS calculations in such configurations.
+One intriguing finding from this systematic study was the disagreement observed in the quasi-continuum region.
+Unlike the discrepancies observed in bound-bound line-width and window opacities, no clear trend was observed in the quasi-continuum disagreement.
+The model and data agree well for Cr and Ni, but not for Fe, at high temperatures.
+Two possible explanations have been proposed for this unexpected behavior.
+The first hypothesis suggests the presence of missing physics in the opacity models that becomes significant under the conditions encountered by high-temperature Fe-opacity experiments.
+The second hypothesis proposes that the opacity experiments or analyses are flawed, but only for Fe at high temperatures.
+Until opacity experiments and theories are reconciled, neither possibility can be definitively ruled out.
+Despite numerous investigations, the discrepancies between experimental and theoretical iron opacities remain unresolved [82, 500, 499, 498, 501, 503, 197, 504, 502, 127, 397, 457].
+On the experimental side, measurements of sample spatial gradients have been conducted and found to be negligible [499].
+Numerical tests have also shown that temporal gradients, self-emission, and the presence of tamping material do not have significant impact on the results [500, 501, 503].
+Furthermore, the uncertainties in temperature and density diagnostics resulting from the choice of spectral model have been determined, and it has been found that these uncertainties are too small to account for the reported discrepancies [498].
+Additionally, background measurements have also been performed and found to be negligible [250].
+The persistent differences between experiment [82, 502] and opacity models has led to a sustained effort to reconcile these discrepancies.
+The experimental efforts have been previously discussed.
+On the theoretical side, tests of some of the approximations used in the opacity tables have been conducted (197, 504, 127, 236), particularly in assessing the convergence of opacity models with respect to the number of configurations included in the calculations.
+However, no significant change in the final opacity has been reported from such studies.
+Several groups have postulated new physical effects that could produce more opacity contributions than in previous works.
+Of particular note is the possibility of two-photon opacity contributions, where an atom or ion could simultaneously absorb two photons (possibly of different frequency) [486, 485].
+All current opacity tables omit such contributions, as they are believed to be a minor contribution to the overall opacity.
+Preliminary studies by More and co-workers [486, 485] have suggested this could be a factor in the comparison of theory with the Z measurements, but other studies [360, 415, 416, 523] find that such a contribution should be small.
+Another intriguing study has reported that transient spatial localization may result in increased opacities [437, 670].
+This is a density-dependent effect where the plasma effects localize the continuum wavefunctions near the absorbing ion, potentially resulting in increased broadening and bound-free cross sections.
+While the work of [670] suggests that this effect is significant, other studies cast doubt on this claim [361].
+Opacity theorists and experimentalists continue to collaborate closely to address discrepancies and uncertainties in the calculated solar RMO.
+They carefully examine each other’s work, seeking to resolve the discrepancies and accurately quantify the uncertainty associated with calculated solar opacity.
+XIII.5 Future opacity work
+The continuing disagreement between the Z Fe experiments [82, 502] and theoretical models is by far the biggest concern of the opacity community.
+New theoretical efforts, while intriguing, have not yet resolved these differences.
+New experimental results, particularly the opacity-on-NIF (National Ignition Facility) campaign [530, 518], are eagerly anticipated to shed some light on the current impasse.
+New approaches are also being developed for opacity measurements using the APOLLON laser at LULI (France) and on the GEKKO XII laser (Japan) [283].
+There has been a suggestion for measuring heat conductivity under thermodynamic conditions resembling those found deeper in the radiative zone, using hydro-carbon foams doped to imitate the solar composition [413].
+Such progress will have implications for solar modeling, as the combination of opacity and elemental abundances is key to resolving the significant discrepancy between helioseismology and solar models.
+As previously noted, the abundance of solar elements is also the subject of renewed scrutiny [56, 455].
+Experimental facilities for solar fusion studies
+Improved experimental facilities have led to significant progress in nuclear astrophysics research in recent years, and future instrumental developments will likely continue to push boundaries in the next decade.
+Key goals driving technological developments include, among others, tests of the weak interactions and of solar properties that make use of high precision solar neutrino measurements, helioseismological mappings of the core metallicity, and detailed solar modeling.
+There is also a host of related open questions in different areas, such as Big Bang nucleosynthesis, red giant evolution, the evolution of supernova progenitors, and a variety of transient explosive phenomena in astrophysics where a quantitative understanding of the nuclear physics is essential.
+Due to the small cross sections of low-energy charged-particle reactions, experiments must be designed to achieve signal rates significantly lower than the background rates from cosmic rays, natural radioactivity in the laboratory, and induced activity from beam interactions with target impurities.
+In the following, we outline current experiences and future facilities—either in development or planned—that will pave the way for possible major breakthroughs in the coming years.
+XIV.1 Above-ground facilities
+Above-ground facilities always play a crucial role in investigating nuclear reactions of astrophysical interest.
+Various methods are applied to minimize background.
+A prevalent approach involves employing passive shielding around the detection area, typically utilizing a layered configuration of lead, copper, and polyethylene.
+This setup helps to decrease unwanted signals and neutron background in detectors with relatively small capacities.
+However, there are additional techniques that can be utilized to further diminish background interference, enabling measurements at energy levels closer to those pertinent to astrophysics.
+XIV.1.1 LENA
+The Laboratory for Experimental Nuclear Astrophysics (LENA) is part of the Triangle Universities Nuclear Laboratory, located on the campus of Duke University, in North Carolina, US.
+LENA features two accelerators: a 230-keV ECR accelerator that produces the world’s most intense low-energy proton beams [202] and a 2-MV Singletron from High Voltage Engineering Europa B.V.
+[592].
+These machines deliver ion beams with intensities ranging up to 20 mA DC proton current at lower energies and 2 mA at several MeV—along with advanced beam pulsing capabilities.
+Each accelerator has dedicated transport systems and control mechanisms for simultaneous operation.
+A key component of LENA is the -ray coincidence spectrometer, which allows measurements with sensitivities comparable to underground facilities [158].
+The detector setup includes a 130 p-type HPGe detector, HPGe clover detectors, and a 16-segment NaI(Tl) annulus, along with the APEX detector featuring 24 position-sensitive NaI(Tl) bars for enhanced measurements.
+Recently, LENA added a cosmic-ray veto system with nine plastic scintillators to reduce cosmic-ray muon background, conveniently positioned due to its wheeled assembly.
+XIV.1.2 NSL
+The Nuclear Science Laboratory (NSL) at the University of Notre Dame, hosts a number of low energy accelerators and the TriSOL radioactive beam facility to maintain a rigorous experimental program in nuclear astrophysics, nuclear structure physics,and fundamental symmetries as well as in a broad range of nuclear physics and nuclear chemistry applications.
+The 5MV single ended pelletron provides high intensity proton, helium, and A40 heavy ion beams for nuclear astrophysics related experiments in forward kinematics using solid and gas target technologies and in inverse kinematic using the St.
+George recoil separator.
+The 10 MV FN pelletron tandem is the main machine for nuclear structure physics experiments, but also serves the accelerator mass spectrometry program on long-lived radioactivities in terrestrial and meteoritic samples.
+The FN is also used for indirect measurements of reactions of astrophysical interest and serves as a driver for the production of light radioactive isotopes beams (A40) for nuclear astrophysics and fundamental symmetry studies.
+XIV.2 Underground Facilities
+Underground facilities are essential to push the boundary of direct measurements towards the lowest energies of astrophysical interest (10s-100s keV).
+In particular, capture-reaction measurements leading to neutron- or -ray emissions are severely hampered in surface laboratories because of the overwhelming background associated with cosmic rays.
+This background can be suppressed by many orders of magnitude by exploiting the natural shielding provided by the rock overburden in underground sites.
+The improvements possible with this strategy have been demonstrated by pioneering work at the Laboratory for Underground Nuclear Astrophysics (LUNA) at INFN-LNGS, first with a 50-kV accelerator (LUNA I) and then with a 400-kV one (LUNA II) still in operation today [30].
+Since the last Solar Fusion review, a new accelerator has been installed at INFN-LNGS, and other underground laboratories have become operational in China (JUNA), the US [Compact Accelerator System for Performing Astrophysical Research (CASPAR)], and Germany (Felsenkeller).
+XIV.2.1 Bellotti Ion Beam Facility
+The INFN-LNGS has recently expanded its accelerator capabilities with the installation of a new 3.5 MV SingletronTM machine designed and set up by High Voltage Engineering Europe (HVEE) [585], see Fig.
+13.
+The 3.5 MV machine is equipped with two independent beam lines which can be operated with solid and gas target systems.
+Acceptance tests at HVEE demonstrated that the machine can deliver intense proton, helium and carbon beams (1, 0.5 and 0.15 mA, respectively) with well defined energy resolution (0.01% of a teravolt) and stability (0.001%h-1 of a teravolt) (Di Leva, 2020 - private comm.).
+The 1.4km of rock overburden at LNGS, equivalent to 3000 m.w.e.
+in a flat site, affords significant background reductions for -ray, neutron-, and charged-particles detection (see [30] for a recent review).
+A first experimental proposal presented by the LUNA-Collaboration focuses on measurements of the reactions 14N(p,)15O, 12C+12C, 13C(,n)16O and 22Ne(,n)25Mg, the latter being in the context of the ERC Starting Grant SHADES.
+The 3.5MV accelerator is now part of the Bellotti Ion Beam Facility [385] and open to external users.171717 https://www.lngs.infn.it/en/pagine/bellotti-facility-en
+XIV.2.2 JUNA
+JUNA is part of the China Jinping Underground Laboratory, established on the site of hydropower plants in the Jinping mountain, Sichuan, China [439].
+The laboratory, located near the middle of a traffic tunnel, is shielded by 2400 m of mainly marble overburden (6720 m.w.e.).
+In December 2020, the JUNA collaboration installed a 400 kV accelerator capable of delivering H+ and He+ beams with intensities of up to 10 particle mA, and a He2+ beam with an intensity of up to 1 particle mA.
+Some important results have already been published.
+For example, the lowest energy study to date of the reaction advocates a breakout from the CNO cycle as a possible route for calcium production in population III stars [671].
+XIV.2.3 CASPAR
+The Compact Accelerator System for Performing Astrophysical Research (CASPAR) laboratory is the only US-based deep underground accelerator and is operated by a collaboration of the University of Notre Dame and the South Dakota School of Mines and Technology [547].
+The accelerator system has been fully operational since 2018 and is located 1480 m below surface at the Sanford Underground Research Facility (SURF)181818See http://www.sanfordlab.org in Lead, South Dakota, formerly the Homestake gold mine.
+The rock overburden results in a 4300 m.w.e shielding effect, significantly decreasing cosmic-ray induced background with a muon flux level of cm-2 s-1.
+The residual neutron flux consists of primarily low-energy (10 MeV) neutrons generated by () reactions induced by the decay of naturally occurring uranium and thorium radionuclides in the surrounding rock, and is generally of the order of 10-6 neutrons cm-2 s-1 [473, 474].
+The CASPAR accelerator is a 1 MV Van de Graff style JN accelerator with a 150 to 1100 kV operational range, well suited for overlap with higher energy measurements.
+The accelerator provides proton and beams with up to A on target.
+The scientific program continues to explore stellar neutron sources and expands the present studies into the magnesium range probing -capture reactions on 24Mg, 25Mg, and 26Mg isotopes.
+A new program has been initiated to explore the endpoint of nova nucleosynthesis, studying proton capture reactions in the Ar to Fe range.
+CASPAR is well suited for these measurements, but will be complemented by a new low energy machine, presently under development at Notre Dame.
+XIV.2.4 Felsenkeller
+The Felsenkeller laboratory is hosted within a network of tunnels, excavated into hornblende monzonite rock, previously used as a cool storage place for the homonymous nearby brewery.
+Jointly funded and built by TU Dresden and HZDR, the laboratory operates a 5 MV Pelletron-type accelerator built by NEC.
+With the addition of a RF ion source in the high-voltage terminal, the accelerator can be operated in tandem mode or as a single-ended machine and is capable of delivering proton, -particle, and carbon beams with currents up to several tens of A.
+Unlike the other sites mentioned in this section, Felsenkeller is a shallow underground laboratory, with a rock overburden of only 140 m.w.e.191919See Bemmerer et al.
+[107] for a more complete description of this laboratory.
+Such a depth is sufficient to shield all components of cosmic-ray induced radiation, except muons, for which further active shielding is generally required [449].
+Detailed studies of muon, neutron, and backgrounds have been carried out [449, 615, 314, 634], and the 12C()13N reaction has been studied [599].
+The further scientific program of Felsenkeller foresees the study of several reactions of astrophysical interest, including 3He(,)7Be and 12C(,)16O.
+XIV.3 Indirect methods
+Indirect methods are also pivotal in nuclear astrophysics due to the challenges associated with directly measuring reactions at extremely low energies.
+They can have systematic uncertainties that are different from those of direct measurements, and they provide supplementary information that can constrain R-matrix and other models used in the extrapolation of data from direct measurements.
+For the present review, it is worth further exploring here the ANC (asymptotic normalization coefficient) and THM (Trojan Horse) methods.
+For a detailed description of the methods, refer to [626, 629].
+The ANC method determines the zero-energy cross-section for radiative-capture reactions by exploiting their peripheral nature.
+This means the reaction depends mostly on the long-distance behavior of the wave function.
+The ANC is extracted from transfer reactions using the DWBA, with uncertainties primarily arising from the optical model description, still much smaller than those affecting spectroscopic factors.
+While the method is focused to zero energy, it complements direct measurements and can be applied to loosely bound nuclei.
+The method has been applied to the 3He(,)7Be (section VI.C).
+The THM is a powerful technique for indirectly measuring astrophysical S(E) factors of reactions involving charged particles.
+By studying a related reaction with a spectator particle, the THM allows for the extraction of the desired cross section without the need for extrapolation.
+The THM relies on the assumption of quasifree kinematics, where the spectator particle has a minimal effect on the reaction of interest.
+The method requires careful selection of beam energy and momentum transfer to ensure that the Coulomb barrier is overcome and the spectator particle remains relatively undisturbed.
+Since Solar Fusion II, significant advancements have been made, focused on improving the model’s accuracy and assessing its systematic uncertainties [626, 629].
+For reactions dominated by broad resonances, the modified R-matrix approach [424, 628] allows for the incorporation of half-off-energy-shell and energy resolution effects within a well-established framework.
+This method enables multi-channel descriptions of reactions, such as in the 12C+12C fusion studies [632], and includes a DWBA-based normalization procedure that does not rely on direct data [425].
+The modified R-matrix framework has been applied to the 15N(p,)12C (section XI.A.3), 19F(p,)16O (section XI.C.4), and 23Na(p,)20Ne (section XI.D.5) reactions.
+For narrow resonance reactions, a simplified approach has been introduced [422] to deduce resonance strengths.
+This method reduces systematic errors from normalization and theory to the percent level through multi-resonance normalization and covariance in error propagation.
+The narrow resonance approach has been used in analyzing the 17O(p,)18F (section XI.B.3), 17O(p,)14N (section XI.B.4), and 18O(p,)15N (section XI.C.2) reactions.
+XIV.4 Plasma facilities
+XIV.4.1 NIF/OMEGA
+Several efforts have recently been devoted to studying nuclear reactions in a plasma environment reminiscent of stellar conditions [293] using the two large laser facilities: OMEGA at the Laboratory for Laser Energetics, University of Rochester [129], and the National Ignition Facility (NIF) at Lawrence Livermore National Laboratory [663].
+Both facilities use high-power lasers to implode spherical capsules containing reactants of interest to high temperatures (of order 1-20 keV) and densities (up to 103 g cm-3 in the extreme case of implosions with a fuel ice layer [3]).
+The OMEGA laser has the capability of delivering 30 kJ of laser energy divided between 60 laser beams to the target; the NIF can deliver up to 2.0 MJ using 192 laser beams, and can thus be used to implode larger amounts of materials to more extreme conditions compared to OMEGA.
+Nuclear experiments at these two facilities are enabled by an extensive suite of nuclear diagnostics, originally developed to do inertial confinement fusion experiments [184].
+Initial results have been obtained on the 2H(,)3He reaction [682, 481], the 3He(3H, )6Li reaction [681], and the 3H(3H,2n) reaction [177].
+Proton spectra have also been measured for the 3He(3He,2p) reaction [680].
+In addition to low-Z reaction nuclear S-factor studies, the NIF and OMEGA facilities are also promising for studies of plasma effects on nuclear reactions, including screening [32, 178]; for studies of charged-particle-induced reactions [653]; and, thanks to the high neutron fluxes achievable (up to neutrons cm-2 s-1 [3]), for studying reactions on excited states [619].
+However, there are challenges remaining to be addressed in fully developing this new platform for nuclear experiments, including the impact on the results of rapid gradients in space and time [205].
+XIV.4.2 PANDORA
+PANDORA (Fig.
+14) is a device conceived for multidisciplinary studies, including many of astrophysical interest [470].
+Two of its main objectives are: (a) to perform the first measurements of decays in plasmas of astrophysical relevance, in order to verify the results obtained in storage rings with 187Re (a lifetime reduction by 9 orders of magnitude), and (b) to measure the opacities of plasmas of astrophysical interest (kilonova ejecta).
+PANDORA will mainly consist of three subsystems:
+an innovative superconducting magnetic plasma trap, able to produce and confine plasmas with electron-ion density up to 1013 cm-3 and electron temperature of keV;
+an advanced plasma multi-diagnostic system, consisting of a set of non-invasive diagnostic tools capable of operating simultaneously, for non-intrusive monitoring of the thermodynamic plasma properties and parameters;
+an array of 14 HPGe (High-purity Germanium) detectors for -ray spectroscopy, surrounding the plasma trap.
+The next development steps will involve: preliminary numerical simulation studies to assess the possibility of measuring opacities in astrophysical plasmas (relevant for kilonova ejecta); study of expected abundances and constraints in AGB stars for some (key) nuclides of interest; experimental investigation of magnetic confinement and turbulence in plasma, using an existing trap at ATOMKI-Debrecen; definition of the analysis algorithm for X-ray imaging and spatially resolved spectroscopy.
+Initial physics cases to be investigated include 176Lu (a potential cosmo-chronometer), 134Cs (to reproduce adequately the observed abundance ratio of the two s-only isotopes, 134Ba and 136Ba) and 94Nb (to solve the puzzle about the exact contribution of s-processing to 94Mo).
+XIV.4.3 Sandia Z
+The Sandia National Laboratories Z machine is the world’s largest pulsed power accelerator [595].
+It charges 22 MJ of electrical energy into a capacitor and then discharges it all at once (over approximately 100 ns) into targets ranging from millimeters to centimeters in size.
+The peak power reached is 80 terawatts, 15 times the steady-state electrical power generated by all of the world’s power plants.
+By concentrating this massive power into small targets, the Z machine can convert them into a state of high energy density (HED), which refers to an extreme state of matter with a pressure exceeding times atmospheric pressure.
+The Z machine has been used for a wide range of HED science, including inertial confinement fusion, dynamic material properties, and laboratory astrophysics.
+In particular, the Sandia Z machine can convert this electrical power into X-ray power using a scheme called the Z-pinch dynamic hohlraum [549].
+This is the most energetic X-ray source on earth and has been used by academic collaborations such as the Wooton Center for Astrophysical Plasma Properties and the Center for Laboratory Astrophysics.
+For the stellar opacity project (see Sect.
+XIII.4), they heat iron or oxygen to the conditions relevant to the base of the convection zone and measure their frequency-resolved opacities to test hypotheses for the solar opacity-abundance problem [97, 84, 82, 502].
+For the accretion disk projects, they heat silicon, iron, and neon photoionized plasmas to conditions similar to those at black hole accretion disks to experimentally test the validity of calculated atomic data, kinetics, and spectral formation [440].
+This is important for accurately interpreting high-resolution data from upcoming X-ray space telescopes such as XRISM.
+For the white dwarf photosphere project, they reproduce its photosphere conditions and validate spectroscopic methods used for understanding the age of the universe as well as constraining its evolutionary paths and supernova remnants [263, 484, 572, 571].
+These experiments reach temperatures ranging from to K and densities ranging from to electrons cm-3 and are performed simultaneously by placing their samples at different locations from the Z-pinch dynamic hohlraum.
+XIV.5 Storage rings for nuclear astrophysics studies
+XIV.5.1 CRYRING
+A main limitation of experimental studies involving unstable nuclei arises from the difficulty of producing radioactive ion beams of adequate intensity and purity at the energies of interest for astrophysical applications.
+Storage rings offer a key advantage over traditional ISOL and in-flight approaches as they allow for storing and re-circulating radioactive ions over and over again, thus allowing for multiple interactions of un-reacted beam particles with an in-ring target [609].
+Beam re-circulation significantly improves the quality of the beam as it results in orders-of-magnitude increased intensity (typical boosting factors of ), limited only by the duty cycle of the measurement [154], and in improved purity of the stored beam, as only beam particles with the right mass-to-charge ratio will survive in the beam orbit.
+One of the main technical challenges of storage rings comes from the requirement of ultra high vacuum conditions ( mbar) necessary to guarantee that ions can be recirculated with minimal losses.
+This, in turn, translates into a requirement for sufficiently thin targets (i.e.
+typically atoms cm-2) to minimise beam ion losses through scattering or electron capture (see [154] for further details).
+A dedicated low-energy storage ring, CRYRING [433], inherited from Stockholm University, was installed at GSI in 2016, as a Swedish in-kind contribution to the FAIR facility.
+The facility allows nuclear astrophysics studies to be done with radioactive nuclei.
+After its initial recommissioning [343], CRYRING now serves as a low-energy extension for experimental storage ring beams, as well as a standalone machine with a local ion source [296].
+After in-flight production in the Fragment Recoil Separator at relativistic energies, rare ions can now be cooled, post-decelerated and stored in the full range down to about 100 keV/.
+A study is underway to reach lower energy by means of a transverse low-energy beamline from the the FISIC project [582, 302].
+CRYRING is now equipped with an in-ring micro-droplet gas target, whose low density ( atoms cm-2) ensures minimal beam energy loss and straggling through the target, a unique advantage for charged particle spectroscopy at storage rings compared to standard techniques.
+In addition, a new detection chamber (Fig.
+15), CARME (CRYRING Array for Reaction MEasurements), was recently installed at CRYRING and fully commissioned in early 2022.
+Specifically designed to fulfil requirements of stored beam experiments and operating at a pressure of mbar, CARME can host up to eight double-sided-silicon-strip detectors with excellent angular and energy resolution.
+Finally, the upcoming installation of the FISIC transverse beam line [581] will allow to intercept the beam stored in CRYRING, thus opening up unprecedented opportunities for crossed-beam experiments.
+An exciting application may be the study of nuclear reactions between ion beams, i.e.
+unaffected by the electron screening, directly at the energies of astrophysical interest.
+Initial proposals for the study of nuclear astrophysics reactions have already been approved.
+Closing remarks
+XV.1 Summary
+This review summarizes the significant progress that has been made over the last decade in understanding the nuclear reactions that govern energy production in solar and stellar hydrogen burning.
+It also describes some of the plasma and atomic physics that effects the solar environment in which these reactions take place, as well as the diagnostics tools – solar neutrinos and helioseismolgy – by which we can probe that environment.
+For and , cross sections have been derived from first principles, using both potential theory and effective field theory.
+In the case of , new experimental data agree well with the theoretical predictions.
+, which governs a minor branch of solar fusion, must also be taken from theory.
+The S-factors for other pp-chain reactions are based on laboratory measurements.
+In some cases, including notably , most of the studies reviewed here are relatively far from the solar Gamow peak energies, underlining the need for theoretical and experimental tools to bridge the gap.
+The nuclear physics of the pp chain remains a significant uncertainty in SSM predictions of the individual neutrino fluxes - pp, pep, hep, 7Be, 8B, and CNO.
+The nuclear errors are comparable to, and in the case of the hep neutrinos greatly exceed, the “environmental” errors that are generated by other uncertainties in the SSM [519].
+For several reactions including and , the SF III recommended values are less precise than those of SF II.
+This fact should not be taken as a sign of regression, but instead as indicative of a more cautious treatment of uncertainties.
+In some of the error estimates made in this review, we have been mindful of past lessons on the potential impact of unidentified systematics.
+The status of our understanding of radiative opacities has been reviewed here, the first time the subject has been included in the Solar Fusion series.
+The debate generated by the solar composition problem has brought renewed attention to the complicated interplay between opacities and composition in the SSM.
+Given new experimental opportunities, especially luminous X-ray sources coupled with plasma targets, there is the expectation of rapid progress in this field over the next decade.
+There has been only very limited progress on electron screening since SF II.
+This field has important practical implications not only for the Sun but also for laboratory experiments, and may merit additional attention in the coming decade.
+Progress may depend on achieving a better understanding of low-energy stopping powers.
+XV.2 Recommended values for S-factors and their derivatives
+The SF III recommendations for nuclear reaction S-factors and their derivatives are presented in Table 1 of the Introduction.
+In the sections that follow the Introduction, we presented the details leading to these results and identified opportunities for future work, which are summarized directly below.
+XV.3 General recommendations for future work
+For all of the solar fusion reactions, obtaining high-precision experimental data in Gamow window remains an elusive goal, one that may not be reached in the foreseeable future.
+However, such direct measures are just one of many avenues for improving our understanding of the nuclear physics of the Sun.
+The recommendations below are in most cases summaries of more detailed discussions presented in this review; see the relevant sections for additional details.
+1.
+For all of the experimentally accessible pp-chain reactions and for which determines the CNO cycle rate, there is a need for at least one experimental data set spanning the entire energy range, encompassing the more limited data sets that currently exist.
+This will provide an important crosscheck on the normalizations of those data sets.
+Experimentalists can now tackle this challenge due to the new generation of underground accelerators that have the dynamic range needed for such studies.
+2.
+We recommended continued effort on reaction theory relevant to pp-chain reactions.
+The needs go beyond rates to include both astrophysical (e.g., precise neutrino spectra) and laboratory (e.g., -ray angular distributions) observables.
+3.
+We urge the community to set an ambitious goal for the precision of future S-factor measurements: reduce the nuclear physics uncertainties to a factor of two below current SSM “environmental” uncertainties, so that they are no longer a significant contributor to SSM neutrino flux uncertainties.
+With the exception of the hep neutrinos, this goal could be reached by a future reduction of a factor two, typically, in the individual S-factor uncertainties [519].
+As solar neutrino data are an important input into global neutrino oscillation analyses, nuclear uncertainties will continue to feed into that program until this goal is achieved.
+Improvements in and other CNO cycle S-factors are needed, as the current 8.4% uncertainty would be a limiting factor in extracting the solar core’s metallicity from a future large-volume CN neutrino experiment.
+The progress reported here on , with the uncertainty reduced significantly from SF II to 3.4%, should be continued, given the current and future 8B neutrino programs of Super-K and Hyper-K: these neutrinos are our best solar core thermometer.
+Finally, ongoing Super-K efforts to detect hep neutrinos provide strong motivation for improving estimates of and the associated neutrino spectrum.
+(Uncertainties in the high-energy tail of the 8B spectrum must also be reduced, as discussed in Section II.)
+4.
+New experimental methods are becoming available that could help us better understand electron screening in terrestrial and solar reactions.
+Plasma conditions resembling those of the Sun can now be produced in the laboratory.
+Reactions can be studied in rings, using ions stripped of their atomic shells.
+The level of theory activity in this field has declined over the past decade, so we hope experimental progress will help renew interest.
+5.
+In order to break the degeneracy between solar abundances and solar opacities, efforts should be made to improve our understanding of radiative opacities for the Sun’s principal metals.
+A high priority is the resolution of the current discrepancy between measured and modeled iron opacities.
+Solar physicists need access to state-of-the-art, open-source opacity codes.
+6.
+We recommend that resources be available to extend the use of current and next-generation dark matter detectors to solar neutrino detection.
+This includes designing new detectors so that they are highly capable for both applications.
+Some of the high-priority goals of solar neutrino physics – such as a 1% measurement of the pp neutrinos to check the luminosity constraint – might be achieved with a dual-purpose detector.
+7.
+The Solar Fusion program to periodically review the nuclear physics of the Sun and other hydrogen-burning stars should be continued, with the separation between studies being no more than ten years.
+The format should continue to be open, giving all researchers working in the field an opportunity to contribute, and broadly international.
+8.
+Future Solar Fusion studies should continue to strengthen the connections between this community and others interested in main sequence stars.
+SF III has placed increased emphasis on hydrogen burning at higher temperatures, plasma physics, opacities, and asteroseimology.
+However, with observations now being made of solar-like stars in their formation stages, we anticipate additional connections developing, well beyond the limits of this study.
+XV.4 Outlook
+Six decades after the initial observations of solar neutrinos by Ray Davis [224], significant advances have been made in our understanding of the SSM and indeed of hydrogen-burning stars in general.
+The driver of this progress has been experiment: the observations made of all of the principal solar neutrino sources; our improved understanding of the flavor physics of those neutrinos; the advances made in helioseismology, a second quantitative probe of the solar interior; and our improved knowledge of the input microphysics of the SSM, including nuclear cross sections and opacities.
+Despite this progress, improvements are still needed.
+In several cases the dominant uncertainty in our solar neutrino flux predictions stems from the limited precision of the nuclear S-factors.
+Similarly, uncertainties remain in our understanding of radiative transport in the Sun, including the continuing debate about the Sun’s primordial composition.
+As changes in composition often can be mimicked by changes in radiative opacities, our imperfect knowledge of the latter has slowed resolution of the solar composition problem.
+The prospects for substantial progress in the next decade are bright.
+Whereas at the time of SF II, there was only one underground ion accelerator worldwide, now there are five, opening up possibilities for independently cross-checking important nuclear data.
+The low background that can be achieved underground has led to the first measurements of reactions in the Gamow window.
+Other new experimental capabilities have come on line that are advancing our understanding of solar atomic physics, including low-energy storage rings and new plasma facilities, the latter made possible in part by advances in high-power lasers.
+The field is approaching the point where opacity measurements can be made under conditions closely approximating those of the solar interior.
+Theory has progressed as well, including advances in neutrino flavor physics that have made neutrinos once again a precise probe of solar physics.
+Techniques with a firmer footing in first principles, such as effective field theory and lattice QCD, are being employed in the extraction of pp chain S-factors.
+Conventional approaches like the R-matrix can now be applied to complex nuclear systems where the inclusion of multiple reaction channels is important.
+On the observational side, several very large neutrino detectors are under construction, including Hyper-Kamiokande, JUNO, and DUNE.
+Certain direct-detection dark matter experiments have reached a sensitivity where nuclear recoils due to the coherent solar neutrino scattering have become a background, leading to discussions of how such detectors might be optimized for solar neutrino detection [261].
+This underscores the remarkable radiopurity of these underground detectors.
+Finally, the helioseismic studies that proved so important in establishing the credibility of the SSM are now being extended to other stars – a major reason SF III was undertaken at this time.
+Asteroseismic studies have the potential to tell us what is special and what is common about the Sun as a star.
+Given current excitement about the emerging field of multi-messenger astrophysics, it is perhaps fitting to remember that the Sun was the prototype for this field.
+An array of physical measurements of the Sun – its mass, radius, luminosity, abundances – were combined with neutrino flux and helioseismic measurements to constrain and then test the SSM.
+The result was the discovery of neutrino mass and flavor mixing.
+This success was a multi-disciplinary effort that has not yet run its course.
+We still have uncertainties in our characterization of the Sun’s nuclear, atomic, and weak interaction physics that, if improved, would make the Sun an even more powerful laboratory.
+As our best known and nearest star, the Sun is the test of our understanding of the structure and evolution of other main-sequence stars.
+If the past is any indicator of the future, the efforts we make over the next decade to make the Sun an even more precisely calibrated reference will yield a rich return in our understanding of physics.
+List of Symbols and Abbreviations
+- ANC
+asymptotic normalization coefficient.
+- BBN
+Big Bang nucleosynthesis.
+- CME
+common-mode error.
+- EFT
+effective field theory, and variants: chiral (EFT), pion-less (EFT), hybrid (EFT*), etc.
+- LO, NLO, NNLO
+leading order, next to (next to) leading order, etc.
+- SSM
+Standard Solar Model.
+Acknowledgements
+We gratefully acknowledge discussions with many colleagues. The help they provided contributed significantly to this review.
+We thank the following institutions for their support of the SF III effort:
+Physics Frontier Center N3AS (Network for Neutrinos, Nuclear Astrophysics, and Symmetries), UC Berkeley, host of the SF III workshop.
+N3AS is supported by the U.S.
+National Science Foundation under cooperative agreement 2020275.
+Institute for Nuclear Theory, University of Washington. The INT is supported by the US Department of Energy under grant DE-FG02-00ER41132.
+European Union, “Chemical Elements as Tracers of the Evolution of the Cosmos – Infrastructures for Nuclear Astrophysics” (ChETEC-INFRA), HORIZON2020 project number 101008324.
+The Heising Simons Foundation, grant 2017-228.
+The authors of SF III acknowledge the following support:
+U.S. National Science Foundation, through grants
+PHY-1430152 (Joint Institute for Astrophysics) and OISE-1927130 (IReNA), Michigan State University
+PHY-1913620 and PHY-2209184, Mississippi State University
+PHY-2011890 and PHY-2310059, Notre Dame University
+PHY-2111426, University of Tennessee
+PHY-2108339, University of Wisconsin
+U.S. Department of Energy, through contracts
+DE-AC02-07CH11359, Fermi National Accelerator Laboratory. Fermilab is operated by Fermi Research Alliance LLC for the Office of Science.
+DE-AC02-05CH11231, Lawrence Berkeley National Laboratory. LBNL is operated by the University of California for the Office of Science.
+DE-AC52-07NA27344, Lawrence Livermore National Laboratory. LLNL is operated by the Lawrence Livermore National Security LLC for NNSA.
+89233218NCA000001, Los Alamos National Laboratory. LANL is operated by Triad National Security LLC for NNSA.
+DE-NA0003525, Sandia National Laboratories. Sandia National Laboratories are managed and operated by NTESS for the NNSA.
+DE-AC05-00OR22725, Oak Ridge National Laboratory. ORNL is operated by UT-Battelle, LLC for the Office of Science.W
+U.S. Department of Energy, through awards
+DE-FG02-00ER41138 and DE-SC0004658, University of California, Berkeley
+DE-FG02-93ER40789, Colorado School of Mines
+DE-FG02-97ER41033 (TUNL), Duke University
+DE-SC0013617 (FRIB Theory Alliance), Michigan State University
+DE-FG02-97ER41041, University of North Carolina
+DE-FG02-88ER40387, DE-NA0003883, and DE-NA0004065, Ohio University
+DE-FG02-08ER41533, Texas A&M University, Commerce
+DE-FG02-97ER41020 (CENPA), University of Washington
+DE-SC0019465, University of Wisconsin
+DE-SC0023692, Iowa State University
+DE-SC0019257, San Diego State University
+Helmholtz Research Academy Hesse for FAIR, Germany
+Hungarian Scientific Research Fund grant FK134845, National Research, Development, and Innovation Office
+Israel Science Foundation grant 889/23
+Italian Ministero dell’Università e della Ricerca (MUR) grants 2022E2J4RK (program PRIN 2022) and 2017W4HA7S (program PRIN 2017)
+The Natural Sciences and Engineering Research Council of Canada (NSERC)
+Spanish Unidad de Excelencia María de Maeztu, grant CEX2020-001058-M, Generalitat de Catalunya, grant 2021-SGR-1526, and MICINN, grant PID2023-149918NB-I00
+U.K. Science and Technology Facilities Council, grant ST/V001051/1
+This paper describes objective technical results and analysis.
+Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S.
+Department of Energy, United States Government, or any other agency listed above.
+For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
+Appendix A Gallium Neutrino Source Cross Sections
+Here we describe in more detail the constraint that the electron capture rate for 71Ge places on the 51Cr and 37Ar neutrino source cross sections for 71Ga, expanding on the discussion of Section II.1.2.
+We discuss both the early treatment of Bahcall and the more recent analysis of Elliott et al.
+[256, 255], performed in support of the BEST experiment
+Following Bahcall [68], the electron (EC) rate can be written in terms of a dimensionless phase space factor :
+where is the energy available to the neutrino and we have omitted atomic shell effects for brevity.
+The last factor is the electron density at the nucleus within the radius .
+Multiplying this by the experimentally determined half-life yields the traditional value needed to calculate the neutrino cross section.
+Bahcall [68] defines a characteristic scale for neutrino cross sections on 71Ga in the form
+where .
+Bahcall [71] reports the number given above and states that “it is about 0.5% less than the value I have used since 1984.” Since these numbers differ by 2.2%, there is clearly an error in one of the 3 numbers.
+Recently Elliott et al.
+[256, 255] have recast the derivations with fundamental constants replacing numerical values and correction terms explicitly stated.
+They conclude that the ground-state cross section is 2.5% lower, and the excited-state contributions somewhat larger, than values used in the anomaly analysis.
+The adjusted net anomaly is minimally reduced.
+They find the following expression for the neutrino cross section on 71Ga leading to the ground state (gs) of 71Ge:
+where and eV is a small correction introduced by Bahcall for the energy lost to electronic rearrangement in the charge-changing reaction.
+The Gamow-Teller matrix element is determined from the measured half-life of the EC decay of 71Ge:
+In this expression, is a correction for overlap and exchange to account for the differences in the wave functions of the initial and final atomic states, is a radiative correction, and is a correction for the weak magnetism contribution to the transition probability.
+The factor 2 with is a spin statistical factor.
+Hence,
+The last line extracts the Bahcall cross-section scale and inserts the dimensionful parameters needed to scale the energy and momentum.
+The atomic factors of Elliott et al.
+[256] are equivalent to those of Bahcall:
+Unfortunately, none of the three most advanced calculations referenced in [71] of the densities has been published so the numerical factor 0.01443 is taken from the ratio in Eq.
+36.
+(The computer code GRASP [252] may still be available.) Rather than use the theoretical electron density of each subshell at the nucleus, Elliott et al.
+use the experimentally measured subshell ratios, which has the advantage that only a single overlap and exchange correction and a single energy are needed, for the state.
+As Bahcall has stated, overlap and exchange have little effect on the total EC rate, so Elliott et al.
+write Eq.
+41 as follows, with the inclusion of a small Q-value update,
+The resulting expression is,
+which is only 0.3% larger than the value given in Bahcall [71].
+The above cross-section expressions give a ground-state cross section 2.6% lower than in the cross section table, Table II in Bahcall [71].
+Whether this extends to the solar neutrino rates has not yet been evaluated, but in any scenario some reconsideration of the low-energy solar neutrino data may be required once a resolution to the Ga anomaly has emerged.
+Appendix B Bayesian Methods
+The treatment of uncertainties when dealing with multiple data sets was addressed in the appendix of SF II, including the Inflation Factor Method used by the Particle Data Group that was adopted by most of the SF II working groups.
+Readers are directed there.
+We now extend that discussion to include the Bayesian methods used by several of the SF III work groups.
+For erxample, in the analyisis of the 2H(,)3He reaction presented in Section IV.4, both Bayesian analysis and Bayes model averaging were utilized.
+Consider an optimization that utilizes Gaussian distributed priors in a constrained Bayesian analysis. With such priors, maximizing the likelihood amounts to minimizing an augmented
+The first term in the above equation is the standard and the second term derives from the Bayesian constraint on the parameters () with prior mean () and width (), which are chosen with some prior knowledge or can be optimized as described below.
+Of course, care must be taken when choosing values for the priors ( and ).
+In the absence of prior information, it is common to set .
+Similarly, one should not make too small, unless one has prior knowledge to do so, else this will bias the determination of .
+In the first term in Eq.
+(46), the double sum runs over the datasets () and individual results from each dataset (), with the mean value and stochastic uncertainty of each data point given by and , respectively.
+The theoretical model describing the S-factor data, , is a function of the energy () and a set of parameters () that must be determined.
+The quoted systematic uncertainties are parameterized by the normalization factors , with a prior of unit normalization and a width characterized by the quoted systematic uncertainty.
+It is straightforward within a Bayesian framework to utilize distribution functions for that are not Gaussian, such as a log-normal or other distributions.
+The parameters represent a normalization of the model function for a given dataset, which, viewed from a non-Bayesian perspective, can also be interpreted as normalization factors that must be applied to the data to match the “true” underlying distribution.
+Finally, are unknown extrinsic uncertainties [229].
+If one assumes that a smooth function of energy can accurately describe the data, then the data in a given experimental set may scatter about this presumed “true” value by more than is reflected by the quoted statistical uncertainties.
+It was suggested in de Souza et al.
+[229] that this extra scatter might be explained by some additional source of statistical uncertainty unbeknownst to the experimenter, or by some unknown systematic uncertainty that is different for each data point in the same data set, as opposed to a correlated systematic that affects all data points similarly.
+In either case, this extrinsic uncertainty can be accommodated in a Bayesian analysis framework by adding the additional uncertainty as a normal-distributed source of noise with a width that is constrained by the data.
+In de Souza et al.
+[229] and Moscoso et al.
+[488], was added as an absolute uncertainty, independent of the energy.
+For some data, Odell et al.
+[515] suggested adopting instead a relative uncertainty, such that the scale of the extrinsic fluctuations is proportional to the mean value of .
+This strategy of adding extra extrinsic uncertainty to the data sets can be viewed as an alternative to that of inflating the quoted statistical uncertainties by , which is often used for seemingly incompatible data sets [666], where is the per degree of freedom.
+An advantage of the extrinsic uncertainty method is that it uses the observed scatter within a given data set as a measure of the possible size of unreported uncertainties, as opposed to uniformly increasing the uncertainty in all data sets by the same relative amount.
+One can further perform a Bayes model averaging, which we implement and describe here.
+For a given model, after optimizing the posterior parameter distributions, the Bayes factor (BF) is proportional to the probability of the model given the data [351].
+Therefore, for a fixed data set, the BF can be used as a relative probability of each model, thus enabling a weighted model-averaging procedure.
+If we assume a uniform likelihood for each model, the expectation value and variance of a quantity is given by
+where denotes the expectation of given the model and denotes the probability of the model given the data (), which is given by
+with the Bayes factor of model .
+Similarly, for a given model, in the absence of prior information on the size of an unknown parameter, the optimal width of its prior can be estimated by finding the value of that maximizes the BF, which typically provides a reasonable approximation to marginalizing over the prior width.